A Complete Bibliography of Publications in Biometrics: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 September 2024 Version 2.00

Title word cross-reference

2^{M} [2084]. 2 × 2 [1901, 1518, 1155, 33, 570, 829, 1377, 1895, 587, 26, 274, 275, 332, 621]. 2 × 2 × K [1483]. 2 × C [829]. 2 × K [1579, 1777]. _{2ε} [1444]. χ^{2} [1904]. D [538]. F [31, 32, 2003, 2004, 1371, 1372]. F 2 [875]. F_{2} [1417]. H_{0} [1547]. k [1901]. k(2 × 2) [1902]. k × 2 [1904, 1905]. Λ [656]. M_{t} [1927]. N [446]. P [2053, 1455, 1456, 60, 61, 719, 562]. Pr(X < Y) − Pr(X > Y) [1106, 1107]. Q [933]. R^{2} [656]. R × C [1344]. t [31, 32, 629, 1371, 1372, 128]. T^{2} [656]. τ_{b} [1225]. × [1505]. x^{2} [1905].

-estimation [446]. -Optimal [538]. -Value [1455, 60, 1456, 61]. -Values [719, 562].

/under [1220].

1 [1932, 1736]. 10.1111/j.0006 [2116]. 10.1111/j.0006-341X.1999.00603.x [2116]. 1750 [369]. 1930 [1787]. 1961 [431]. 1963 [509]. 1965 [553]. 1967 [553]. 1969 [954]. 1982 [202]. 1983 [1065]. 1984 [209]. 1989 [211]. 1990 [432]. 1991 [598, 381, 598]. 1992 [558, 794, 954, 794]. 1993 [741, 1020, 1085]. 1994 [908, 1306, 1143, 1141]. 1995 [1641]. 1996 [1919]. 1997. [2115]. 1999 [2114, 2116, 2117].

2 [1541, 1914, 1785]. 2000 [1746, 1975, 1541]. 20th [1143]. 2b [1443]. 2nd [153, 1745, 1855, 1742, 150, 148].

3 [2033]. 3/4 [1141]. 341X.1999.00603.x [2116]. 3rd [1743].

4 [1084]. 44 [213, 271]. 45 [157]. 46 [270, 556]. 47 [775].

5 [1845]. 50 [1306, 2051]. 53 [2115]. 55 [2114, 2116, 2117]. 5th [377].

70th [1541, 1746, 1975].

89d [271]. 89h [213].

90k [157]. 92f [775]. ’93 [1084].

A-Bomb [496]. Aalen [1524, 1764]. Aberrant [825]. Ability [1097]. Abramson [96]. Abrupt [2053, 2054]. Absence [1388, 531]. Absolute [127, 1038, 1364, 1533, 2115]. Abundance [1433, 1809, 873, 1213, 186, 617, 1754, 1261, 947, 300]. Accelerated [238, 2003, 2004, 540, 1941, 1768]. accept [1547]. Access [1999]. Accommodating [182]. Accommodation [1828]. Accompany [198]. Account [1460, 1826]. Accounting [21, 1427]. Accuracy [1357, 1559]. Accurate [1610]. Acknowledgement [630]. Acknowledgements [2113]. Act [1498]. Action [609, 1119]. Active [1942, 408]. Activity [1705, 408, 1445]. Activity-Based [1445]. Actuaries [962]. Ad [1025]. Ad-hoc [1025]. Adaptive [977, 1001, 1885, 2093, 1560, 1570, 1233, 345, 1126, 354, 578, 920, 1723, 1937, 1306]. Added [1650]. Addition [1119]. Additional [1424]. Additive [739, 318, 1360, 1764, 1719, 2061, 456, 21, 2018]. Additivity [1475, 1208, 722]. Adequacy [1779, 253]. Adjust [1869]. Adjusted [912, 649, 1436, 1616, 1938, 1555, 562, 2062, 1948]. Adjusting [1214, 1037, 719, 1199]. Adjustment [1336, 584, 1564, 2027, 248, 1731, 1000]. Administration [529]. Admixture [1149]. Adolescent [347]. Adult [347]. Advanced [1443, 723, 896, 1243, 639]. Advances [1084, 1640, 844, 1740, 1916, 1536, 325, 1851]. Adverse [651, 1666]. Aerial [1511, 353, 1095]. Affected [73, 1804, 1700, 1807, 2005, 1805, 1808, 808]. Affected-Sibling-Pair [1700]. Affects [1611]. Afifi [1540, 423]. Aflatoxin [1728]. after [1863, 1384]. Against [18, 1748, 518, 934, 1694, 1461]. Age [1832, 862, 1028, 584, 1660, 1616, 2066, 111, 1477, 1767, 1318, 1609, 1764, 2005]. Age- [1318]. Age-Adjusted [1616]. Age-at-Onset [2066, 2005]. Age-Dependent [1764]. Age-Specific [862, 1028, 584, 1660, 111]. Age-Standardized [584]. Age-Time [1609]. Agents [1375, 1781]. Ages [1716]. Aggregate [1828, 1330]. Aggregated [1369]. Aggregation [342, 1103, 1210, 405, 387]. Aggregations [1754]. Aging [1059]. Agreement [612, 59, 1452, 1534, 392, 1336, 881, 1708, 705, 1526, 1709]. Agresti [250]. Agricultural [400, 165, 925, 1302]. Aid [1008]. Aided [1540, 423]. AIDS [343, 181, 1356, 697, 1099, 1276, 816, 2014, 602, 1043, 497, 995, 1169, 2019, 1512, 446, 1932, 1011]. Aircraft [1809]. Aitken [1912]. Aitkin [192]. Akaike [1977]. Albert [2116, 466]. Algebra [1798, 1971]. Algorithm [187, 1240, 693, 1485, 1686, 1224, 1825, 620, 1575, 1735, 1521, 585, 525, 534, 1939, 666, 1419, 1153, 1368, 1369]. Algorithmic [1494]. Algorithms [458, 1953, 362, 1073]. Alignments [1032]. Alimentaire [193]. All-Pairwise [1156]. All-Way [2084]. Alldredge [889]. Allele [356, 1378]. Alleles [477]. Allocation [1576, 1214, 1465]. Allocations [2098]. Allometric [2077]. Allow [649]. Allowing [2024, 1425]. Alpha [1770]. Alternating [341]. Alternative [1455, 1897, 1969, 1047, 1456]. Alternatives [33, 518, 1992]. Altman [501]. Always [408]. Alzheimer [1103]. Ambulatory [227]. Ames [658]. AMS [1633]. AMS-LaTEX [1633]. Analyse [771, 760, 1848, 315, 91, 1491, 90]. Analyses [977, 190, 1155, 1483, 649, 1626, 1176, 1758, 1440, 1376, 698, 1120, 1239, 479, 288, 1432, 1723, 679]. Analysing [2039, 910, 317, 1100, 1187]. Analysis [171, 344, 1875, 968, 1112, 979, 103, 329, 377, 552, 507, 511, 684, 686, 792, 843, 900, 1016, 848, 902, 852, 1076, 1138, 1399, 1300, 1447, 1451, 1542, 1637, 1497, 1543, 1541, 1746, 1974, 2042, 2108, 1975, 1976, 194, 319, 647, 929, 1157, 1293, 1458, 1113, 339, 1654, 1328, 1940, 1392, 1548, 780, 1408, 458, 1707, 1675, 1146, 1960, 280, 836, 1151, 1338, 371, 459, 1762, 1614, 1647, 930, 1523, 349, 1040, 912, 586, 1342, 227, 1511, 17, 401, 1645, 1186, 974, 1168, 1705, 360, 299, 337, 1881, 368, 626, 487, 478, 1661, 1010]. Analysis [375, 1990, 521, 659, 760, 138, 625, 1894, 1608, 1413, 1276, 28, 1147, 841, 681, 1395, 166, 822, 914, 47, 309, 1964, 2073, 1345, 913, 1473, 1226, 466, 1388, 1346, 1804, 969, 1627, 1671, 972, 1089, 1596, 695, 1713, 1217, 1733, 644, 1058, 938, 205, 1699, 180, 1492, 889, 80, 728, 631, 729, 148, 195, 370, 970, 1160, 1600, 1553, 1667, 287, 1312, 926, 320, 1222, 1839, 886, 305, 1149, 1767, 1868, 515, 1429, 922, 799, 1421, 420, 1030, 480, 207, 756, 2058, 1652, 1393, 2079, 117, 282, 294, 1006, 358, 174, 185, 749, 825, 520, 1682, 50, 994, 1718]. Analysis [1607, 1807, 868, 1514, 1672, 141, 1623, 995, 973, 394, 38, 1220, 1655, 1045, 1965, 1357, 445, 492, 1062, 363, 2056, 1394, 1430, 224, 613, 758, 1824, 361, 652, 333, 1362, 1171, 924, 1837, 1505, 1942, 1529, 1956, 800, 1185, 1092, 1662, 1486, 1771, 1605, 1478, 1812, 1384, 1322, 564, 275, 454, 2016, 2019, 1512, 1907, 1846, 603, 925, 1161, 1805, 1808, 2104, 71, 976, 705, 1876, 563, 1153, 391, 2088, 390, 167, 981, 1104, 36, 1154, 1937, 1407, 234, 1216, 1306, 1377, 274, 1117, 621, 771, 1296, 251, 688, 731, 1540, 1496, 1640, 1744, 1791, 1745, 1915, 1977].

Analysis [2106, 2043, 1980, 201, 943, 1069, 1396, 423, 322, 44, 1182, 724, 1489, 250, 102, 506, 1015, 1539, 144, 261, 888, 1183, 550, 1131, 1910, 726, 424]. Analytical [958, 1973]. Analyzing [1255, 1901, 1902, 1567, 571, 362, 33, 1026, 1825, 602, 1371, 1389, 220, 1532, 1343, 1925, 179, 1926, 1372, 1978, 465]. Ancienne [898]. Anderson [1976, 192, 939, 947, 1138, 887, 53]. Anderson-Hauck [939]. Anemia [221]. Angewandte [1794]. Angioplasty [1940]. Angler [1072]. Anglers [1479]. Angles [878]. Angular [514, 1997]. Anil [1400]. Animal [2042, 1433, 452, 299, 298, 625, 873, 1213, 1030, 1393, 67, 1810, 405, 129, 1280, 578, 1059]. Animals [404, 951]. Annotated [431, 509, 639, 954, 553]. Annual [211, 432, 598, 794, 1085, 1641, 1919]. Anonymized [1880]. Anova [1586, 31, 32, 880, 694, 1876, 1109, 1811]. Answers [1639]. Anthropology [1852]. Antibody [348, 1239]. Anticipation [2005]. Antihypertensive [1282]. Anwendung [1794]. Aoshima [1541, 1746, 1975]. Aperiodic [927]. Aplastic [221]. Aplysia [1705]. Apparent [129]. Applebaum [1538]. Applicable [1932]. Application [171, 612, 1504, 188, 717, 1103, 1716, 2026, 1661, 2096, 1099, 1879, 1725, 992, 1706, 1963, 1217, 1166, 1660, 2014, 175, 970, 602, 768, 1314, 1149, 125, 1868, 497, 670, 1947, 1810, 295, 1116, 1649, 1954, 2084, 1994, 519, 1480, 345, 221, 809, 1478, 1169, 524, 387, 1619, 1555, 918, 526, 1737, 889]. Applications [1112, 596, 637, 901, 1194, 1136, 1349, 1979, 1980, 1072, 1483, 942, 1949, 1130, 1659, 348, 604, 205, 1699, 1494, 1971, 1820, 1029, 932, 1564, 369, 617, 1851, 1064, 2019, 1907, 390, 807, 883, 1811, 1039, 446, 328, 635, 154, 507, 792, 1078, 1195, 1251, 1304, 1450, 1638, 1916, 2109, 2041, 591, 262, 86, 1492, 467, 316, 137, 554, 420, 43, 1685]. Applied [152, 470, 684, 1354, 1446, 1586, 1859, 1181, 321, 561, 627, 1910, 1399, 329, 265, 842, 899, 1542, 1744, 1791, 1915, 1182, 1348, 679, 88, 38, 422, 425, 39, 1493]. Appliquée [953]. Applying [906, 1936, 1101, 1842, 996]. Approach [393, 1112, 1572, 854, 441, 1551, 537, 1597, 1176, 910, 1933, 1172, 1991, 1028, 302, 695, 1660, 2014, 1049, 1759, 235, 1262, 411, 1108, 2009, 1334, 1833, 1840, 237, 2047, 448, 454, 1720, 1415, 1838, 234, 255, 264, 1011, 470, 789, 843, 1191, 887, 1003, 590, 1348, 1909, 1189, 1493]. Approaches [1940, 1762, 912, 2067, 1312, 1655, 1687]. Appropriate [790, 1488]. Approximate [571, 1985, 1778, 1036, 113, 270, 556, 1890, 267, 555]. Approximating [1374]. Approximation [1742, 2064, 1577, 567, 1068]. Approximations [488, 1386, 1884, 1029, 1547]. Aquatic [448, 1650, 1190, 1191]. Arabie [1588]. Arbitrarily [2024, 1320, 1660]. Arbitrary [1943]. Area [675, 579, 1284]. Areas [1486, 1702]. ARIMA [1345]. ARMA [1234]. Armed [857]. Arminger [1247]. Armitage [1495, 1536, 1784, 1836, 1209]. Arnold [1005]. Arrangements [402]. Arrival [130]. Art [588]. Arteries [516]. Arthur [737, 738]. Article [1806]. Artificial [844, 428]. Ary [465]. ascertained [446]. Ascertainment [2066]. Aspects [1614, 245, 302, 894, 1134]. Assay [1381, 1388, 69, 658, 2078, 495, 670, 122, 1874]. Assays [1704, 657, 1659, 1235, 1817, 825, 622, 266]. Assembled [222]. Assess [1534].

Assessing [1258, 183, 629, 1772, 487, 1728, 1748, 1391, 1165, 1557, 1761, 1609, 921, 1409, 2016, 222, 859, 1526, 2100]. Assessment [208, 211, 1692, 2117, 488, 1755, 1009, 1576, 1119, 651, 69, 175, 2112, 1779, 412, 769, 1585, 1487, 449, 1111, 1841, 1988, 1498]. Assessments [1452, 392]. Assignment [384]. Assignments [152]. Assisted [458, 1707]. Associated [651, 226]. Association [2046, 817, 52, 1365, 1717, 1389, 226, 916, 221, 1219, 1870, 1871, 1091]. Associations [67]. Assumption [341, 572]. Assurance [1198]. Aston [1450]. Astrostatistics [1449]. Asymmetry [1478]. Asymptotic [210, 1035, 1521, 12, 66, 391, 772, 328]. Asymptotics [1853]. Atkinson [1063, 1535]. Atlas [1089]. Atomic [838]. Attributable [169, 386, 709]. Attribute [661]. Attributes [793]. Audit [263]. Auditing [959]. Auditor [558, 741, 908, 381]. Audrain [898]. Auflage [1794]. Augmented [1330]. Auswahlverfahren [1854]. authors [2101]. Autocorrelation [1385]. Autologistic [1706]. Automation [960]. Autoregressive [1045]. Autoregressive-Type [1045]. aux [1116]. Average [1199, 227, 1224, 1037, 1116]. Ayton [1253].

B [552, 508, 470, 639, 687, 735, 732, 736, 792, 1015, 1016, 846, 957, 901, 902, 1077, 1083, 1082, 1354, 1446, 1497, 1855, 1792, 1695, 1793, 1796, 1790, 1913, 1976, 1914, 2115, 2114, 255, 1736, 551, 1003, 780, 1297, 1128, 951, 945, 2105, 591, 151, 204, 192, 197, 1396, 375, 546, 1785, 631, 148, 264, 1398, 1631, 2033, 1911, 195, 949, 320, 1489, 53, 93, 141, 1005, 1014, 45, 1846, 98, 312, 1150, 2097]. B-Splines [1150]. Babu [1449]. Baby [860]. Back [182]. Backcalculation [1108]. Backward [1678]. Bailar [893]. Bailey [1074]. Bain [683]. Baker [957]. Balakrishnan [1242, 503, 1005, 1740, 1916, 948, 1242, 1738, 1851]. Balance [241]. Balanced [574, 521]. Banded [765]. Bands [1035, 2061]. Bandwidth [1995]. Bandwidths [803, 1024]. Banks [1736, 1785, 2033]. Barbour [1068]. Bark [675, 1045]. Barnard [464]. Barndoff [1399]. Barndoff-Nielson [1399]. Barndorff [328, 894]. Barndorff-Nielsen [894]. Barndorff-Nielson [328]. Barnett [1302, 1692, 1066, 785, 1133, 2037]. Bart [2035]. Bartel [264]. Barthélemy [468]. Bartoszyński [1739]. Based [854, 343, 818, 169, 1038, 441, 1551, 2076, 2087, 79, 1324, 1923, 1673, 1170, 400, 1176, 910, 2067, 914, 293, 348, 180, 173, 65, 1049, 429, 525, 1262, 1340, 451, 874, 1897, 515, 648, 1156, 1700, 1235, 2051, 186, 540, 417, 1212, 12, 2080, 1668, 1722, 1666, 1992, 1827, 1153, 1204, 2018, 1594, 1603, 388, 704, 1547, 1613, 1445, 1000]. Baseline [494, 2000, 287, 1165]. Baselines [1903]. Bases [844, 953]. Basic [505]. Basics [1586]. Basilevsky [1492]. Basu [1400]. Batch [538, 350]. Batches [351]. Bauer [1306]. Bausell [639]. Bayes [153, 304, 1489, 2052, 27, 1108, 2049, 2050, 1516, 1489]. Bayes-Fisher [304]. Bayesian [2116, 1443, 1789, 852, 1451, 2097, 1113, 1422, 1989, 1381, 283, 919, 571, 1168, 1597, 337, 2068, 1823, 1959, 482, 1147, 1933, 822, 823, 1473, 561, 1991, 1213, 1667, 1759, 1895, 2058, 1861, 1817, 174, 2009, 750, 1171, 1505, 1942, 1956, 1275, 1290, 1668, 1812, 1599, 858, 821, 1153, 1117, 1941, 86, 1396, 1630, 1845, 2107]. Bayesienne [1004]. Be [474, 214, 272, 330, 380, 434, 512, 557, 599, 641, 690, 740, 796, 855, 907, 964, 1021]. Beck [790]. Becker [144]. Beetle [1045]. before [938, 369]. Begin [1832]. Beginners [896]. Beginning [408]. Behavior [1455, 1474, 129, 1456]. Behavioral [2035, 942, 465, 1247]. Bell [1792]. Belle [1587, 941]. Belsley [835]. Bennett [368]. Benzécri [681]. Beran [1139]. Berger [1845, 638]. Berlinsky [1590]. Bernado [1845]. Bernoulli [1369]. Berry [197, 1630, 1495]. Bertalanffy [406]. Best [5, 155, 323]. BETA [1398, 851, 1658, 1311, 1435, 1]. Beta-Binomial [851, 1311, 1435, 1]. Between [771, 629, 1718, 611, 1987, 663, 1199, 1534, 1425, 881, 1119, 1835, 1115, 1312, 1434, 1037, 1294, 817, 1131, 1365, 1724, 67, 1389, 226, 1430, 1148, 107, 108, 1613, 1091]. Between- [1718]. between-Population [1148]. Beyond [1586]. Bhattacharyya [1400]. Biadditive [833]. Bias [982, 1278, 1778, 863, 869, 2066, 1235, 29, 1756, 1095, 1952, 990, 310, 1874, 1284, 1702, 707]. Bias-Corrected [1874]. Biased [74, 237, 1233, 167]. Biases [132]. Bibliography [431, 509, 954, 553, 883]. Bickel [1001]. Bidirectional [1714]. Bijleveld [906]. Billard [1296]. Binary [1112, 1572, 102, 212, 433, 1452, 2114, 392, 1164, 1726, 746, 1523, 3, 911, 63, 910, 1050, 1100, 1936, 660, 2053, 2054, 891, 1984, 972, 1566, 604, 744, 671, 933, 1105, 1365, 1886, 812, 1389, 168, 745, 1954, 539, 492, 1994, 1877, 519, 2082, 573, 1382, 336, 1893, 1677, 1870, 1871, 387, 1509, 1439, 1615, 1593, 2085, 517, 388, 213]. Binding [926]. Binomial [1387, 1504, 267, 555, 851, 278, 2076, 172, 626, 113, 1271, 279, 1383, 114, 1311, 622, 133, 617, 532, 1435, 1, 124, 270, 556, 277]. Binomials [34]. Bioassay [15, 174, 484]. Bioassays [2052, 759]. Bioavailabilities [1487, 1111]. Bioavailability [1010]. Biodiversity [2096]. Bioeconomics [689]. Bioequivalence [1010, 694, 1487, 1111, 2098, 769]. Biographies [91]. Biography [464]. Biologen [1019]. Biological [1543, 1960, 837, 1973, 258, 947, 326, 632, 843, 259]. Biologiques [940]. Biology [596, 736, 844, 1018, 1688, 1693, 1129, 1074, 426, 546, 1132, 256, 1631, 593]. Biomass [628]. Biomathematics [955]. Biomedical [755, 50]. Biometric [158, 558, 741, 908, 381]. Biometrics [214, 272, 330, 380, 434, 474, 512, 557, 599, 641, 690, 740, 796, 855, 907, 964, 1021, 775, 213, 270, 556, 107, 157, 271, 1349, 2099, 2032]. Biométrie [892, 940]. Biometrisches [500]. Biometry [967, 1064, 1921, 1536, 1180, 1296]. Biopharmaceutical [1078]. Biostatistical [791, 727, 8, 1348]. Biostatistics [1300, 551, 1629, 1630, 1784, 377, 1076, 1077, 1135, 1587, 1852, 1858, 941]. Biotechnology [254]. Biplots [722]. Birds [453, 1649]. Birth [338]. Birthday [1541, 1975, 1746]. Bishop [1911]. Biswas [680, 1348]. Bivariate [393, 1638, 1567, 1114, 1313, 1484, 571, 747, 1872, 1329, 289, 698, 235, 316, 1776, 1343, 1219, 1431, 1512, 1870, 1871, 1838, 782]. Bland [49]. Blanks [1057]. Blinding [1233]. Block [1703, 1502, 333]. Blocked [516]. Blocks [1211]. Blom [1494]. Blood [347, 463, 227, 447, 1964, 294, 1622]. Blue [1649]. Boashash [792]. Bock [2039, 44]. Bofinger [956]. Bolasco [251]. Bole [675]. Bomb [838, 496]. Bone [1675, 1863]. Bonferroni [1685]. Bonferroni-Type [1685]. Bonferronio [82]. Book [92, 326, 632, 895, 1013, 251, 208, 1248, 1073, 102, 103, 152, 209, 55, 210, 153, 154, 104, 57, 101, 105, 211, 328, 329, 377, 378, 431, 266, 263, 430, 379, 432, 506, 552, 508, 509, 596, 597, 470, 472, 473, 471, 510, 507, 598, 639, 788, 687, 683, 688, 731, 793, 786, 684, 686, 689, 735, 789, 790, 640, 636, 732, 733, 734, 638, 685, 787, 637, 736, 791, 792, 794, 842, 844, 899, 1015, 843, 954, 845, 900, 1016, 955, 846, 956, 959, 957]. Book [960, 1017, 1018, 901, 961, 1019, 904, 905, 962, 898, 906, 847, 848, 849, 897, 902, 903, 953, 958, 963, 1020, 1134, 1190, 1191, 1076, 1077, 1078, 1135, 1140, 1194, 1195, 1249, 1250, 1079, 1136, 1080, 1192, 1083, 1137, 1193, 1196, 1251, 1197, 1252, 1074, 1081, 1082, 1084, 1138, 1139, 1198, 1253, 1085, 1141, 1302, 1399, 1445, 1449, 1349, 1303, 1301, 1350, 1353, 1400, 1354, 1300, 1304, 1446, 1305, 1450, 1401, 1447, 1448, 1351, 1539, 1542, 1586, 1587, 1540, 1588, 1637, 1589, 1497, 1543, 1544]. Book [1634, 1636, 1541, 1638, 1496, 1639, 1498, 1640, 1495, 1590, 1635, 1641, 1740, 1744, 1791, 1794, 1741, 1852, 1690, 1745, 1855, 1688, 1693, 1792, 1795, 1689, 1691, 1746, 1798, 1854, 1857, 1859, 1695, 1793, 1856, 1696, 1796, 1853, 1692, 1742, 1790, 1799, 1743, 1915, 1916, 1978, 2107, 1917, 2037, 1913, 1974, 2042, 2039, 2108, 1918, 2109, 1975, 1979, 1976, 1977, 2038, 2106, 2110, 1914, 2040, 2041, 2043, 2111, 1980, 1919, 194, 144, 319, 255, 940, 887, 1066, 1736, 201, 551, 1003, 1681, 426, 780, 1072, 1297, 594]. Book [1128, 951, 198, 1683, 41, 318, 590, 834, 1629, 1490, 87, 261, 890, 945, 836, 1536, 1737, 2105, 202, 1244, 149, 371, 252, 1684, 139, 469, 591, 151, 142, 97, 327, 592, 727, 783, 2035, 943, 1583, 147, 204, 325, 839, 262, 1009, 1186, 324, 85, 634, 944, 1246, 1298, 91, 1001, 150, 942, 1443, 193, 888, 1183, 368, 550, 838, 1012, 1397, 1786, 192, 1071, 1069, 547, 197, 86, 840, 1010, 1396, 1630, 376, 375, 138, 317, 725, 1188, 51, 46, 423, 322, 546, 1686, 206, 841, 941, 468, 681, 1130, 1395, 1785, 952, 47, 48, 466, 1181]. Book [465, 44, 321, 1182, 1491, 1849, 891, 42, 373, 1848, 199, 314, 1008, 1245, 1132, 1299, 254, 1584, 205, 1492, 889, 1002, 1004, 728, 631, 723, 140, 100, 148, 313, 257, 260, 256, 264, 323, 500, 553, 504, 467, 464, 589, 635, 633, 785, 680, 948, 896, 837, 893, 1180, 1247, 1243, 1129, 1133, 1348, 1296, 1347, 1444, 1398, 1633, 1494, 1538, 1631, 1582, 1784, 1787, 1739, 1850, 1973, 2036, 2033, 1911, 1971, 2102, 2112, 1908, 1972, 2034, 2103, 195, 370, 95, 429, 146, 894, 316, 724, 781, 949, 1242, 249, 196, 946, 137, 191, 1909, 1189, 143, 502, 320, 1063, 679]. Book [1489, 1537, 1845, 1789, 1131, 52, 372, 53, 554, 420, 998, 374, 1067, 207, 1000, 1065, 88, 90, 1007, 1187, 1687, 1006, 145, 1682, 49, 50, 1910, 503, 258, 315, 253, 93, 141, 782, 1070, 1005, 1738, 89, 1068, 1632, 38, 548, 369, 726, 730, 200, 421, 54, 505, 1062, 1851, 588, 835, 1064, 593, 950, 1394, 1585, 1788, 424, 784, 422, 1184, 96, 1075, 999, 1014, 1185, 1011, 40, 94, 549, 682, 425, 45, 1907, 428, 1846, 2104, 98, 947, 39, 136, 1844, 892, 250, 43, 501, 203, 99, 259, 1493, 1847, 427, 1685, 1912, 312]. Book [1254]. Bookstein [1691, 593]. Bootstrap [1638, 775, 236, 1565, 307, 1172, 1631, 543, 890, 945, 1737]. Border [773]. Borgan [2115]. Bortz [1848]. Both [698]. Bound [22]. Boundaries [398].

Boundary [1770]. Bounds [576, 2057, 580]. Bourque [732]. Bowerman [470]. Bowman [328]. Box [41]. Boyer [687]. Braak [47]. Bradford [508]. Brain [178, 1094]. Branching [1070]. Bratley [148]. Braun [1065]. Break [480]. Breakthroughs [1741]. Breast [1832, 1619]. Breckling [205]. Breeding [1349, 862, 147, 204, 111, 2015, 1280, 1811, 147]. Breslow [207, 986]. Bridge [1012]. Brillinger [1296]. British [52]. Brockwell [1795]. Bromek [685]. Brookmeyer [1011]. Bross [1248, 730]. Brown [790, 204, 1129, 1072]. Brownie [502]. Buckland [947]. Buckley [515]. Budgets [179]. Bugaj [1739]. Building [41, 257]. Built [1147]. Bulked [1430]. Bunke [139, 139]. Burnham [947]. Bus [1999]. Bus-Route [1999]. Business [329].

C [92, 103, 209, 506, 597, 471, 687, 684, 689, 843, 845, 959, 962, 898, 906, 847, 848, 958, 1194, 1136, 1192, 1083, 1399, 1543, 1744, 1745, 1855, 1859, 1915, 2042, 1918, 2043, 2111, 2114, 940, 1736, 201, 1072, 1683, 261, 1737, 1244, 142, 727, 839, 1186, 85, 1298, 1001, 1010, 1395, 1785, 47, 1182, 373, 1004, 728, 729, 100, 313, 323, 467, 635, 893, 1247, 1494, 2033, 1911, 95, 316, 502, 1063, 1789, 998, 90, 145, 503, 1005, 1632, 1394, 1184, 999, 1185, 45, 428, 136, 43, 1912, 312]. C.A.MAN [458]. Calculation [1012, 84, 1479]. Calculations [169, 1673, 669, 2093, 1558, 827, 396, 438, 532]. Calculus [1979, 1590]. Calibration [1272, 1240, 1614, 1646, 585, 411, 290, 1466, 1603, 374]. Cambridge [1305, 161]. Campbell [789, 255]. Cancer [171, 853, 1255, 1832, 1026, 1115, 625, 822, 1209, 207, 1393, 825, 9, 530, 1996, 345, 1027, 496, 1731, 524, 1093, 1619, 1515]. Cancers [346]. Canonical [1980]. Capsule [1478]. Capture [872, 112, 187, 763, 836, 1151, 452, 1989, 1900, 864, 302, 1929, 1049, 618, 502, 805, 997, 1331, 1648, 1756, 1337, 1460, 2084, 1362, 1927, 19, 352]. Capture-Mark-Recapture [1337]. Carcinogenesis [1209, 1678]. Carcinogenicity [703, 298, 1659, 1436, 1030, 67, 2090, 95]. Care [1248, 904, 1077, 181]. Caribou [1754]. Caries [936]. Carlin [1396, 1489]. Carlo [546, 1631, 1971, 283, 1511, 1825, 675, 866, 1706, 1957, 1699, 1890, 1928, 1861, 1812, 1117]. Carrier [1388]. Carroll [1584, 40]. Carter [94]. Case [2117, 1293, 1458, 980, 1258, 1038, 667, 1524, 218, 177, 386, 1580, 11, 1367, 1042, 1438, 311, 284, 80, 1863, 10, 1533, 2066, 1767, 671, 710, 2009, 479, 1714, 1225, 338, 224, 581, 285, 524, 391, 859, 167, 1988, 981, 306, 2115, 1917, 1296, 1789]. Case-Cohort [980, 581]. Case-Control [2117, 1293, 1258, 1038, 667, 1524, 386, 1580, 11, 1367, 311, 284, 80, 10, 1533, 1767, 710, 2009, 479, 1225, 224, 285, 524, 391, 859, 167, 1988, 981, 306, 2115]. Case-Crossover [1714]. Case-Sibling [2066]. Casebook [1296, 726]. Casella [638, 998]. Cases [970, 2095]. Casti [1188]. Castri [259]. Caswell [199]. Catalogue [773]. Catalyzed [538]. Catch [1255, 1433, 1026, 1471, 1479, 454]. Catch-at-Length [454]. Catch-Effort [1433, 1471]. Catch-Up [1255, 1026]. catchabilities [1034]. Catchability [872]. Categorical [2024, 187, 1293, 1829, 678, 983, 2012, 669, 1131, 457, 799, 1765, 1886, 971, 219, 141, 672, 1833, 2013, 1943, 1670, 1415, 250]. Categories [1567, 1579, 712]. Category [2024]. Caulcutt [849]. Causal [983, 943, 384]. Cause [127, 1209]. Cause-Specific [127]. Cautionary [996]. Caveat [1050]. CBDEA [845]. CD4 [975, 918]. CD4-Lymphocyte [975]. Celeux [1491]. Cell [463, 33, 1237, 294, 613, 26, 1411, 918, 1653]. Cell-Lineage [1411]. Cells [657, 1949]. Cellular [895]. Censored [1366, 1320, 1052, 2116, 1934, 1548, 1889, 1923, 487, 1823, 1358, 676, 1959, 1099, 1825, 1936, 816, 1028, 444, 1410, 1963, 1058, 1660, 2014, 602, 1600, 1820, 126, 1944, 1618, 1289, 481, 1203, 2060, 540, 122, 995, 1769, 804, 1044, 2081, 1945, 1956, 1092, 1169, 2019, 390, 1768, 1948]. Censoring [393, 6, 156, 268, 647, 929, 1780, 801, 803, 532, 1946, 1610, 2, 4, 1594, 157, 271]. Censorship [1659, 2063]. Census [1595, 202, 28, 429]. Census-Based [429]. Censuses [1906]. Centre [507]. Centres [1402, 483]. Certain [1120]. Cervical [524]. Chain [1511, 1825, 522, 1706, 1699, 1928, 1051, 1861, 2081, 615, 1812, 1946, 1117]. Chain-of-Events [2081, 1946]. Chains [179, 1382, 640]. Chan [960]. Change [1654, 1749, 1936, 1043, 287, 125, 239, 182, 882, 407, 1096]. Change-in-Ratio [1749, 407, 1096]. Change-Point [1654]. Changepoint [2053, 2054, 1391]. Changes [6, 156, 268, 1025, 2, 4, 523, 157, 271]. Channel [482]. Chaos [894]. Chaos-Statistical [894]. Character [761]. Characteristic [1549, 529, 1655, 1284]. Characteristics [982, 718, 921]. Characterization [949, 228]. Charles [1064]. Chatfield [1399, 103, 85, 1298]. Chatterjee [631, 1296]. Chaudhuri [1400, 143]. Cheese [1402, 483]. Chemical [1208]. Chemistry [958]. Chemometrics [57]. Chemoprevention [625, 1393]. Chémotaxie [760]. Chemotaxis [760]. Chen [597, 151]. Chi [1681, 409]. Chi-Square [409]. Chi-Squared [1681]. Child [2005]. Children [838]. China [597]. Choe [943]. Choice [881, 676, 305, 1118, 455, 16, 2095]. Choices [2024]. Choquet [1194]. Choquet-Deny [1194]. Chow [2042, 1186, 1010]. Christensen [1639, 1910]. Chromosome [1867]. Chronological [431, 509, 954, 553]. Chudnovsky [378]. Ciminera [95]. Circadian [408, 1326]. Circular [1997, 576, 1006, 706]. Clamp [482, 1553]. Clarice [2117]. Clark [689, 732, 1540, 97, 423]. Clarke [505]. Class [612, 659, 936, 1357, 2070, 440, 808]. Classes [1699, 539]. Classical [561, 723, 2027]. Classification [176, 1395, 1687, 1482, 1585, 1588, 44]. Classifications [660]. Classified [815]. Clegg [204]. Clément [955]. Cleveland [55, 944]. Clients [545]. Clinical [900, 853, 854, 1087, 1448, 1637, 1640, 1664, 441, 700, 1218, 1882, 1762, 930, 987, 227, 2008, 1821, 1177, 1009, 1531, 1176, 478, 857, 2069, 822, 244, 466, 439, 698, 889, 13, 1124, 886, 305, 497, 1842, 1187, 978, 9, 530, 1938, 310, 1883, 1505, 1942, 1668, 2029, 757, 1722, 603, 2074, 858, 1666, 2072, 1723, 523, 1932, 935, 36, 1039]. Clinicians [1793]. Clogg [1247]. Clonal [348]. Closed [1595, 491, 863, 864, 805, 1331, 1648, 613, 2074]. Closure [1927]. Cluster [1901, 1902, 1940, 1125, 322, 311, 1556, 1560, 971, 1718, 1570, 1126, 242, 354, 1107, 1106]. Clustered [1940, 1993, 1507, 1627, 1671, 1289, 932, 1319, 219, 168, 83, 745, 1718, 1514, 1954, 569, 1046, 492, 1994, 519, 2001, 387]. Clustering [292, 1463, 25, 360, 66, 462, 704, 1588]. Clusters [1653]. Coal [1575, 1735, 1969]. Coarse [755]. Cochran [1836]. Cochran-Armitage [1836]. Codon [979]. Coefficient [2089, 1291, 876, 413, 1438, 495, 1776, 1620, 246, 1409, 928, 1606]. Coefficients [1758, 10, 295, 2025, 22, 1814, 1091]. Cohen [45, 59, 426, 503]. Coherence [861]. Coherent [1513]. Cohort [968, 343, 980, 127, 429, 410, 701, 207, 1060, 361, 581, 222, 601]. Cokriging [291]. Collaboration [8]. Collected [209, 1065]. Collection [1494]. Collett [891, 1007]. Collinearity [835]. Colon [825]. Color [1478]. Column [1717]. Combination [608, 830, 1509]. Combinations [278, 2052]. Combinatorial [1851]. Combined [1119, 276, 60, 928, 61]. Combining [269, 1749, 1895, 2009, 939, 75]. comes [1994]. Comment [1534, 1440, 886, 677, 2098, 775]. Comments [1240, 83]. Common [1265, 1434, 1776, 12, 1501]. Commonly [1894, 769]. Commonsense [789, 255]. Community [47, 242, 882]. Companion [1444]. Comparability [184]. Comparative [1087, 1240, 1551, 1420, 1658, 585, 935, 2037]. Compare [1410]. Comparing [1052, 1461, 654, 1879, 287, 673, 1289, 1502, 1621, 2060, 1287, 1622, 1031, 1033, 1574, 1702, 1694]. Comparison [6, 156, 268, 340, 1940, 879, 218, 70, 1830, 317, 2096, 1565, 444, 1671, 284, 68, 1025, 1124, 587, 479, 1672, 30, 826, 443, 1708, 2072, 1723, 2, 523, 1966, 1967, 1709, 332, 4, 157, 271]. Comparisons [1387, 210, 1315, 1324, 857, 286, 1156, 1065, 1922, 1632, 220, 1064, 1992, 1547]. Compartmental [643, 623]. Competing [1676, 402, 1276, 1101, 1842]. Competition [1896]. Completely [815]. Completeness [584]. Complex [506, 736, 1639, 1664, 866, 1804, 993, 1983, 859]. Compliance [497, 310]. Complications [1940]. Comply [1752]. Component [511, 1338, 1149, 1961, 294, 358, 715, 219, 60, 61]. Components [1875, 1200, 693, 1286, 1881, 1285, 989, 1208, 1434, 1831, 605, 363, 1474, 991, 71, 43, 22, 547, 998]. Comportement [193]. Composite [1581, 1467, 1145]. Composition [882]. Compositional [1675]. Computation [1251, 635, 807, 260, 319]. Computational [1693, 1303]. Computationally [1485]. Computations [1836]. Computer [844, 1249, 458, 1707, 896, 1540, 87, 890, 423]. Computer-Aided [1540, 423]. Computer-Assisted [458, 1707]. Computer-Intensive [87]. Computers [378]. Computing [198, 28, 260, 1437, 1129, 1129, 505]. Concentration [1119, 1208, 1217, 1874]. Concentration-Response [1208]. Conception [633, 2047, 860, 1369]. Concepts [1793, 1751, 249, 1788]. Concerning [1050]. Concise [1856, 2036]. Concisely [896]. Concomitant [1295, 217, 1277]. Concordance [1534, 1291, 495, 1620]. Condition [1999]. Conditional [156, 1661, 1830, 28, 669, 302, 1890, 411, 1895, 1669, 1842, 2025, 1212, 485, 26, 652, 1722, 4, 1547, 157, 621, 1195]. Conditioning [835]. Conditions [1878]. Conference [98]. Confidence [1958, 1934, 169, 1604, 2076, 1315, 1035, 657, 303, 1426, 1386, 414, 1679, 1390, 604, 565, 1265, 2094, 1029, 239, 1144, 1274, 34, 76, 1520, 826, 12, 2061, 1896, 757, 1610, 22, 1874, 1550, 1107, 1417, 1106]. Configural [688, 784]. Configurations [615]. Conflict [1188]. Confounders [1523, 2100, 1761, 134, 394, 485]. Confounding [1940, 1563]. Congenital [867]. Conjointe [90]. Connectance [1729]. Connections [749]. Conover [2038]. Conquest [1296]. Conrad [152]. Consensus [60, 61]. Considerations [1218, 1342, 1302]. Constant [59, 299, 768]. Constrained [624, 707]. Constraints [649, 24]. Construct [1029]. Constructing [1051]. Construction [1770, 1417, 1816, 1585]. Consul [467]. Consultant [776, 545]. Contain [362, 1966, 1967]. Contemporary [198]. Contingency [393, 552, 739, 1392, 371, 1716, 659, 1894, 829, 721, 1344, 1717, 587, 455, 2084, 1161, 712, 456, 815, 332, 621]. Contingency-Table [393]. Continual [9, 1334]. Continued [26, 328]. Continuity [1264]. Continuity-Corrected [1264]. Continuous [340, 1770, 404, 1567, 881, 1358, 1464, 1839, 710, 1655, 334, 1994, 1343, 2018, 1594, 1242, 316, 1242]. Continuous-Time [1839, 2018]. Continuously [226]. Contoured [2091]. Contribution [656]. Contributions [597, 471, 1510, 163, 165, 162, 1924]. Control [955, 2117, 1293, 1258, 1038, 667, 1524, 834, 654, 2046, 386, 1580, 11, 1367, 311, 284, 80, 10, 1533, 2066, 1767, 1502, 1621, 710, 2009, 479, 1225, 2007, 484, 224, 285, 1942, 2090, 524, 391, 859, 167, 1988, 981, 306, 2115, 105, 793, 1692]. Controlling [286]. Controls [339, 31, 1186, 1436, 758, 32]. Controversy [163]. contubernalis [1478]. Conventional [920]. Convergence [1360]. Convex [1829, 1204]. Cook [1347, 726, 707]. Cooke [505]. Cooper [1245, 726]. Cooperation [1188]. Coordinate [120]. Coppi [251]. Copula [1219]. Copy [1551]. Cordray [726]. Cornelius [132]. Coronary [1940]. Correct [216, 1095]. Corrected [1264, 84, 1874]. Correcting [134]. Correction [6, 212, 267, 433, 555, 511, 739, 1087, 1306, 1402, 1581, 1858, 1797, 2115, 2114, 2116, 2117, 213, 113, 270, 556, 1879, 1172, 1756, 157, 271]. Corrections [155, 156, 267, 268, 269, 852, 850, 853, 854, 851, 1199, 1254, 1255, 1256, 1200, 1451, 1307, 1403, 1404, 1452, 1642, 1643, 1499, 1694, 451]. Correlated [1112, 1643, 1328, 1462, 1523, 1041, 1997, 1554, 1262, 1839, 1283, 1558, 1525, 971, 868, 1719, 539, 774, 2082, 1837, 1486, 1893, 1316, 563, 1310, 1593, 1702, 517, 1317]. Correlation [2114, 1534, 982, 814, 1291, 876, 629, 70, 1438, 230, 1359, 1189, 495, 820, 1311, 1292, 1783, 1776, 520, 1620, 246, 1370, 921, 1877, 580, 1987, 2021, 1814]. Correlations [1828, 664, 1225, 1047]. Correspondances [760]. Correspondence [368, 760, 166, 681]. Corrigendum [265, 595]. Corriger [216]. Cost [336, 433]. Cost-Efficient [336, 433]. Costs [1503, 548]. Coulton [1784]. Council [904]. Count [1689, 975, 1596, 577, 1262, 1508, 1525, 35, 1095, 1650, 1280, 115, 1730]. Counties [1115, 1294]. Counting [1140]. Counts [395, 714, 1043, 460, 405, 26, 1605]. Coupled [1313]. Coupling [486]. Courgeau [91]. Course [684, 1192, 1690, 594, 1296, 53, 1005, 469]. Courts [208]. Covariance [280, 1698, 692, 542, 1565, 1395, 2073, 1627, 1671, 1521, 287, 1360, 1385, 2027, 527, 363, 115, 1555, 663, 21, 1154]. Covariance-Adjusted [1555]. Covariances [334]. Covariate [433, 2003, 2004, 1936, 170, 1172, 1725, 1818, 1391, 701, 745, 1718, 225, 1094, 248, 336, 1501, 1943, 1309, 1407]. Covariate-Time [225]. Covariates [187, 929, 1336, 131, 1931, 566, 1227, 1819, 1957, 602, 1600, 817, 1765, 1669, 1886, 1267, 2030, 1270, 528, 226, 1044, 2001, 1236, 1670, 1159, 1615, 1625, 1948]. Coverage [805, 1034]. Cows [81]. Cox [102, 328, 1399, 193, 589, 1682, 1617, 1917, 1230, 986, 2064, 335, 1825, 1936, 14, 1577, 1818, 525, 1885, 125, 1283, 1765, 1101, 1947, 1955, 1824, 1236, 1827]. Cracking [771]. Craven [505]. CRC [2036]. Creighton [1192]. Cressie [847, 592]. Crisis [1248]. Criteria [1766, 754]. Criterion [185, 1520]. Critical [700, 384]. Cropping [1287]. Cross [746, 871, 176, 320, 617, 917, 815, 889, 889]. Cross-Classification [176]. Cross-Classified [815]. Cross-Over [746, 871, 320, 889, 889]. Cross-Sectional [917]. Cross-Tabulations [617]. Crosses [1211]. Crossover [1328, 574, 1155, 1342, 1440, 117, 1714, 1569]. Crossroads [1248]. Crowder [848, 424, 2106]. Crypt [825]. Culture [1622]. Cumulative [1663, 1412, 2061, 712]. Cure [582, 1124]. Cured [1203, 443]. Curtailing [1722]. Curve [473, 1113, 406, 1002, 1262, 229, 296, 2027, 1514, 1478, 1528, 1284, 1091, 1792]. Curves [1881, 121, 175, 297, 1231, 1769, 1655, 2061, 1024, 1599, 2021, 1174, 1937, 1702]. Cycles [2023]. Cyclic [1221, 1249]. Cycling [952]. Cytogenetics [376]. Cytonuclear [1380]. Cytotoxic [412].

D [1013, 1248, 102, 328, 378, 506, 508, 471, 789, 636, 845, 959, 960, 962, 1134, 1190, 1077, 1194, 1079, 1080, 1193, 1196, 1399, 1449, 1350, 1353, 1354, 1446, 1305, 1401, 1447, 1587, 1496, 1590, 1791, 1792, 1798, 1796, 1790, 1799, 1917, 1974, 1976, 2106, 1914, 2114, 2116, 255, 1736, 1297, 1240, 951, 590, 1490, 836, 1737, 149, 371, 97, 943, 1583, 204, 325, 839, 1246, 91, 150, 1397, 1786, 192, 1071, 197, 1396, 1630, 317, 941, 1130, 1785, 952, 465, 891, 1299, 254, 1584, 889, 1002, 728, 729, 100, 323, 500, 589]. D [1180, 1129, 1296, 1347, 1494, 1538, 1973, 2033, 1911, 370, 316, 946, 502, 1789, 554, 1007, 1682, 253, 93, 1068, 726, 730, 505, 835, 1585, 424, 422, 40, 94, 549, 947, 501]. Dagnelie [953]. Dagpunar [46]. Daily [2023, 2047]. Daly [1246, 896, 1444]. Daniel [377, 1076, 425]. Darling [877]. Data [344, 251, 1504, 393, 1366, 802, 1112, 1572, 2086, 1320, 865, 1828, 212, 263, 433, 507, 731, 684, 732, 844, 843, 1497, 2039, 2106, 2043, 2114, 2116, 194, 187, 647, 929, 1164, 1157, 1293, 1832, 1934, 1522, 746, 1258, 1272, 1654, 1829, 1940, 1548, 1889, 1259, 404, 1524, 1567, 1675, 1146, 18, 1151, 1882, 2071, 702, 1614, 1647, 678, 1518, 452, 1523, 1676, 1989, 494, 1813, 1381, 2003, 2004, 2000, 3, 911, 63, 281, 227, 1511, 1923, 17, 177, 748, 974, 1953, 1246, 1698, 299, 386, 1580, 362, 70, 1069, 487, 864, 216, 1823, 1358, 1396, 1115]. Data [375, 697, 1758, 1050, 1036, 1488, 1959, 521, 760, 317, 453, 2012, 1099, 1476, 828, 1208, 322, 1100, 1507, 1624, 1879, 653, 1061, 1825, 816, 1968, 1763, 823, 2073, 1281, 913, 1473, 44, 1656, 985, 1984, 1028, 289, 755, 885, 444, 1410, 669, 610, 1627, 1671, 972, 1376, 1089, 1721, 1596, 1862, 1713, 1929, 1963, 1383, 406, 1058, 938, 1166, 1660, 2014, 230, 231, 577, 1296, 970, 602, 1600, 1820, 126, 1944, 1667, 1618, 1262, 1178, 1189, 1533, 1289, 481, 1222, 1839, 2078, 1489, 305, 1767, 1051, 1165, 1557, 1503, 116, 1508, 457, 933, 932, 1283, 1344, 1657, 1267]. Data [53, 2092, 701, 1292, 756, 2059, 1203, 1319, 1318, 1609, 2079, 117, 812, 1700, 1007, 1187, 294, 1525, 971, 2030, 749, 531, 1389, 1922, 16, 355, 520, 2060, 1057, 1951, 415, 168, 83, 745, 1718, 868, 1514, 540, 2007, 338, 1672, 1373, 141, 122, 995, 437, 295, 394, 1769, 672, 1833, 1370, 804, 1044, 1532, 2065, 398, 996, 1954, 484, 2023, 569, 1046, 617, 445, 492, 1094, 835, 288, 2056, 593, 35, 416, 1877, 1060, 519, 2001, 990, 350, 758, 2013, 1602, 2081, 1945, 333, 616, 1219, 1431, 1774, 1732, 2070, 2082, 809, 1837, 1529, 1822, 1956, 1275, 573]. Data [1092, 1771, 336, 1411, 1478, 1812, 1946, 1322, 1169, 1341, 1266, 2019, 2090, 1512, 1846, 1893, 82, 1280, 1316, 1677, 1870, 1871, 115, 2104, 1943, 387, 1552, 1619, 1096, 250, 563, 821, 753, 601, 1415, 2031, 2017, 2088, 390, 1730, 486, 981, 663, 1439, 815, 1526, 1932, 1869, 1481, 1903, 1838, 1594, 1546, 918, 1268, 1593, 1407, 1948, 2115, 1039, 213, 1317, 1117, 1976, 102, 1015, 847, 848, 902, 1350, 1351, 1539, 1689, 1691, 1978, 1976, 144, 1066, 261, 836, 592, 1583, 944, 375, 465, 891, 42, 1008, 1582, 370, 1131, 1006, 50, 141, 96, 138, 47]. Data [1847]. Data-Bases [844]. Database [1982]. Datasets [845]. Dates [2065]. David [2117, 589, 1536, 795]. Davidian [1583]. Davis [1795]. Davison [1737]. Dawid [1845]. Day [100, 207]. Daykin [962]. Deactivation [538]. Dead [1140, 1222]. Dean [1974]. DeAngelis [952]. Death [145, 1384]. December [1306]. Décès [51]. Decision [1197, 1916, 1762, 1668, 2029, 1153]. Defined [1567, 761, 1044, 615]. Definition [876]. Definitions [1751]. Defoliation [1238]. Degradation [350]. Degree [303, 2059]. Degrees [1772, 135]. Delaney [317]. Delay [326, 1923, 995]. Deleterious [1928]. Delivered [1143]. Dellaert [147]. Delta [703, 119]. Demand [1493]. Dementia [1660]. Demétrio [2114]. Demographic [1751, 1049]. Démographique [91]. Demography [526]. Denominator [1772]. Densities [1386, 1995, 714, 76]. Density [1675, 946, 922, 178, 301, 921, 1126]. Deny [1194]. Departures [1208]. Dependence [2059, 1788]. Dependencies [1682]. Dependent [212, 3, 1711, 131, 1099, 1656, 1725, 1819, 2063, 223, 489, 1764, 868, 1236, 1309, 1924, 1526, 213]. Deriso [2105]. Derivatives [1024]. Derived [1038]. Dervin [940]. Descent [807]. describing [110]. Design [341, 266, 379, 507, 900, 1135, 1300, 1637, 1640, 1815, 201, 980, 780, 1218, 836, 1614, 1315, 1342, 1186, 476, 1471, 583, 2075, 254, 889, 121, 1049, 1160, 137, 886, 1502, 922, 480, 207, 2051, 1269, 1612, 1910, 9, 164, 1469, 1505, 1942, 1930, 1185, 275, 1599, 603, 222, 1665, 167, 1439, 1781, 274, 591, 888, 1062, 1974, 2042, 887, 202, 840, 1010, 320, 1062]. Designed [319, 17, 1703, 1212, 1846]. Designing [1999, 1883, 317]. Designs [1366, 433, 643, 773, 1328, 814, 574, 1291, 1114, 1425, 1327, 2022, 400, 402, 1830, 1679, 660, 1173, 1329, 1816, 2011, 410, 931, 607, 1379, 648, 538, 117, 15, 1817, 1714, 1263, 573, 1454, 336, 1501, 2029, 440, 1722, 124, 354, 1935, 1249, 138, 1063]. Desirable [608, 830]. Detailed [1564]. Detect [1463, 1208, 1530]. Detectability [920]. Detected [1754]. Detecting [529, 297, 875, 715, 242, 882, 1323]. Detection [294, 67, 1239, 1031, 1619, 1515]. Determination [1699, 701, 710, 479, 1145, 1668]. Determined [1165]. Determining [398]. Development [597, 626, 110, 849, 2042]. Developmental [1504, 1125, 1232, 702, 349, 1711, 1994, 449, 758]. Developments [1448, 1707, 1016]. Deviance [850, 713]. Deviation [1364]. Devlin [1792]. Devore [430]. Diabetic [992]. Diagnoses [1276]. Diagnosing [1413]. Diagnosis [936]. Diagnostic [1694, 1053, 340, 1461, 335, 747, 1227, 14, 1701, 1357, 2082, 1486, 496, 1559, 1702]. Diagnostics [2064, 1380, 173, 1601, 835, 984, 734, 1012]. Diagonals [712]. Diagonals-Parameter [712]. Diallel [1211]. Dialogue [57]. Diameter [823, 235]. Dichotomous [881, 661, 62, 220, 569]. Dictionary [313, 1180, 1913]. Dielman [329]. Dietary [1199, 1964, 1037]. Difference [267, 555, 2076, 1835, 113, 1733, 1657, 443, 668, 1056, 270, 556]. Differences [299, 243, 1131, 1925, 1926, 541]. Different [184, 1031]. Differential [1979, 177, 1748, 596, 1250]. Differentials [232]. Diffusion [178]. Diggle [1539, 791, 1879]. Digit [399]. Digital [1638, 1857]. Dilated [1952]. Diller [896]. Dilution [266, 1704, 1381, 657, 1646, 751, 2078, 1235, 1817, 622, 385, 826, 1874]. Dimensional [1408, 586, 402, 1345, 714, 178, 1992, 815]. Dimensions [401]. Dinosaur [2096]. Diploid [1606]. Directed [1777]. Direction [205]. Directional [205, 141]. Dirichlet [1427]. Dirschedl [1303]. discours [1844]. Discovery [1351, 1300]. Discrete [1196, 340, 1684, 884, 375, 1464, 1721, 370, 781, 782, 334, 445, 2006, 1343, 2001, 1480, 1774, 1946, 82, 1316, 1093, 352, 1710, 1317, 1738]. Discrete-Time [445, 1946]. Discrete-Valued [1684]. Discriminant [1458, 1881, 1413, 1182, 304, 1601, 185, 1153, 841]. Discriminante [1491]. Discriminants [1555]. Discrimination [304, 128, 486]. Discussion [1238, 1807, 1806, 107]. Disease [342, 344, 802, 1828, 144, 1157, 1727, 292, 1755, 1146, 1960, 1470, 1103, 1048, 2067, 293, 348, 2075, 992, 65, 756, 1609, 445, 652, 221, 66, 462, 222]. Diseases [529, 1884, 2066, 67]. Disequilibria [1380]. Disequilibrium [2089, 1373]. Disinfectant [1388]. Disorders [387]. Dispersed [1504]. Dispersion [1404, 279, 671, 1215, 1220, 133]. dissimilarity [1613]. Distance [659, 1773, 1434, 1205, 1153, 1613, 947]. Distances [2057, 107, 772, 108]. Distinct [531]. Distinguishing [1470]. Distortion [993]. Distributed [118, 772]. Distribution [851, 1402, 1638, 1335, 930, 1282, 876, 1923, 1206, 816, 823, 1226, 1237, 1659, 1089, 1706, 483, 1243, 235, 114, 534, 450, 1105, 1374, 670, 1482, 613, 12, 564, 1169, 1, 66, 1619, 1561, 772, 1753, 526, 2062, 1653, 1121, 1948, 1427, 1855, 948, 196, 1301]. Distribution-Free [876, 1482, 564, 1301]. Distributions [1743, 1567, 463, 884, 1772, 1207, 1410, 1465, 467, 235, 316, 2091, 1144, 294, 81, 1776, 995, 1394, 1092, 2016, 1710, 1242, 196, 781, 949, 1242, 782, 1738, 210]. Diversity [1924, 99]. Diversity-Dependent [1924]. Dixon [427]. DNA [1012, 146, 979, 1123, 1551, 915, 2057, 1982, 107, 108, 1613, 1104]. Do [189]. Dobson [252]. Document [896, 1444]. Documents [896]. Dodge [194]. Does [480]. DOI [2116]. Doll [1209]. Domain [1094, 1918]. Donev [1063]. Données [216, 760, 315]. Dose [1098, 853, 1307, 1573, 404, 702, 2083, 2087, 1711, 298, 120, 608, 830, 121, 175, 228, 716, 16, 530, 484, 496, 1599, 1260, 1666, 1509, 1159]. Dose-Dependent [1711]. Dose-Finding [1666]. Dose-Response [1098, 1307, 404, 702, 2087, 120, 121, 175, 228, 716, 16, 1599, 1159]. Doses [16]. Double [647, 1803, 628, 1834, 451, 867, 939, 1057, 1095, 1983]. Double-Count [1095]. Double-Platform [1803, 1983]. Doubling [1598]. Doubly [1392, 816, 1028, 602, 294, 1169, 2019]. Doubly-Censored [1028]. Dowdy [634]. down [79]. Drake [259]. Draper [41]. Driven [1847]. Drop [1421, 1677]. Drop-Out [1677]. Drop-Outs [1421]. Dropouts [720, 416, 544]. Drug [351, 623, 497, 178, 2016, 1509, 1009]. Drugs [1282, 412]. Dual [574, 529, 2010]. Dual-Balanced [574]. Dudewica [635]. Dudewicz [956, 151, 635, 54, 428, 897, 1857, 469]. Dudewiczo [1638]. Due [1940, 1755, 1756, 182]. Duffy [370]. d’un [771, 193, 760]. d’une [575, 1116]. Dunnett [218, 1064]. Duration [493, 2097, 1099, 13, 1160, 123, 1289]. Duration-Dependent [1099]. Durations [1891, 1892]. Durrett [733]. Duscha [1197]. Dykstra [140]. Dynamic [2107, 1214, 1546, 55]. Dynamics [895, 1380, 1932, 110, 2105, 952, 258].

E. [1411]. Early [1547, 1664, 1645, 1176, 13, 2072]. Ecological [1191, 1879, 99]. Ecologists [2035]. Ecology [1190, 1914, 1297, 1493, 1013, 47]. Econometric [1399]. Econometrics [1399, 1979]. Economic [1302]. Economics [264, 329]. Ecosystem [2112]. Ecosystems [1071]. Ed [2051]. ED50 [1679, 16]. Edge [1879]. Edge-Effect [1879]. Edition [153, 377, 786, 847, 1745, 1855, 1742, 1743, 150, 148]. Editor [1352, 1782, 2099, 2032, 1535]. Editorial [56, 1127, 778, 1179, 1256]. Edwards [786]. Effect [1272, 359, 1523, 1282, 770, 1772, 298, 308, 1459, 1879, 1763, 1646, 694, 1863, 827, 1519, 168, 921, 1454, 496, 300, 1606, 517]. Effective [2083, 1606, 1321, 100]. Effects [1504, 212, 1200, 1581, 1157, 1940, 1882, 1467, 1462, 494, 654, 1418, 3, 63, 606, 1597, 770, 1119, 1206, 402, 1036, 1488, 1990, 176, 1624, 653, 1122, 822, 1281, 170, 969, 1963, 498, 1359, 1562, 525, 1944, 1759, 1120, 1563, 1149, 1477, 1502, 1165, 1557, 116, 812, 2027, 1831, 745, 1718, 1840, 1996, 572, 1357, 1409, 1094, 288, 485, 1952, 384, 1332, 1824, 774, 1480, 1732, 1896, 991, 1322, 1873, 1280, 1316, 1677, 2049, 2050, 1870, 1871, 1876, 1236, 436, 1325, 859, 1368, 1516, 21, 1869, 1550, 1110, 213, 1034, 1317]. Efficacy [987, 1998, 1177, 1388, 1228, 698, 499, 624, 1666, 1992]. Efficiency [2117, 1815, 403, 1998, 2012, 1577, 410, 1319, 117, 745, 2013, 1126, 1988, 535]. Efficient [433, 773, 1224, 121, 810, 336, 222, 2062, 1948, 1001]. Effort [872, 1433, 1999, 1471, 1096]. Efron [945]. Eggs [1878]. Eiselt [901]. Eisen [98]. Either [1052]. Elandt [2043]. Elandt-Johnson [2043]. Electrophoretic [586]. Elementary [151, 361, 1494, 1136, 1353, 327]. Elements [613, 1911, 1745, 260]. Elementwise [1404, 1215]. Elevations [1991]. Élevées [1116]. Elimination [1173]. ELISA [69]. Elliptically [2091]. Elsevier [1180]. Embleton [902]. Embryonic [626]. Emigration [77]. Emphasis [163]. Emphasising [316]. Empirical [403, 488, 1940, 1567, 2052, 884, 1489, 1108, 1144, 1169, 2049, 2050, 1912, 153, 41, 257]. Employing [1430]. Encyclopedia [2036, 1858, 1736, 1785, 1784, 2033, 312]. End [408]. Ended [1579]. Enders [1399]. Endpoints [1766, 1158, 698, 1371, 286, 886, 1662, 603, 2074, 2072, 1720, 1372]. Engelhardt [683]. Engineer [105]. Engineering [430, 1398, 596, 472]. Engineers [100, 148, 473, 264]. English [367]. Enregistrement [51]. Entre [771]. Entropies [2096]. Entropy [472]. Entry [1342]. Environment [1692, 1645, 1608, 1133, 1532, 2065, 785]. Environmental [1302, 1498, 1917, 2039, 1914, 1297, 1244, 1751, 1184, 427, 2039]. Environmentally [1470]. Environments [1796, 366, 1796]. Envisioning [504]. Enzyme [69, 538, 415]. Enzyme-Catalyzed [538]. Enzyme-Kinetic [415]. Enzyme-Linked [69]. Epidemic [181, 575, 1350]. Epidemics [1356, 1048]. epidémie [575]. Epidemiologic [651, 309, 125, 288]. Epidemiology [1140, 1564, 1011, 788, 1080]. Epigenetic [736]. Epileptic [395]. Episodes [1725]. Epistatic [1925, 1926]. Equality [1385, 527, 350, 1814]. Equation [1572, 450, 515, 1615, 1625, 1838, 352]. Equations [1112, 596, 955, 1194, 1793, 717, 177, 1050, 1627, 1671, 970, 932, 824, 1365, 1765, 1292, 749, 1840, 334, 2006, 1943, 1552, 2031, 2017, 2088, 981, 1407, 1678]. Equicorrelation [928]. Equilibria [356]. Equilibrium [72, 1374, 1339]. Equivalence [1087, 1576, 64, 604, 2079, 1612, 711, 935]. Eradicate [189]. Erläuterungen [500]. Erläuterungen/Register [500]. Error [433, 880, 1238, 1614, 1866, 1313, 1821, 1475, 747, 1931, 1386, 309, 1172, 1646, 1818, 1584, 80, 287, 286, 1961, 499, 1564, 2057, 528, 89, 572, 2082, 336, 1501, 1310, 1109, 1481, 984, 535, 1786]. Errors [793, 1328, 1757, 1549, 1168, 1985, 1278, 362, 936, 118, 673, 868, 548, 385, 2015, 774]. Errors-in-Variables [1168]. Essays [1400]. Essential [1079, 150]. Essentials [845]. Establishing [1612]. Estimate [628, 172, 1999, 1438, 1809, 10, 1311, 1996, 484, 1205]. Estimated [1390, 1279, 2016, 1107, 1106]. Estimates [865, 211, 127, 169, 1038, 703, 236, 1813, 568, 764, 1041, 1278, 1778, 985, 1820, 932, 1319, 135, 921, 2080, 1148, 1323, 928, 77, 663]. Estimating [1112, 1572, 1199, 2114, 188, 1428, 1727, 1433, 1232, 1960, 452, 1749, 351, 717, 1998, 606, 1580, 747, 1050, 2067, 1763, 243, 447, 1330, 132, 1627, 1671, 873, 1089, 2075, 714, 694, 1818, 970, 240, 1037, 618, 805, 451, 497, 450, 515, 1503, 457, 932, 824, 1365, 1765, 1292, 2059, 749, 531, 1922, 530, 995, 1769, 1840, 301, 334, 2006, 1877, 1754, 485, 1887, 2081, 616, 1469, 221, 1126, 182, 1024, 1472, 947, 1943, 1552, 2031, 2017, 2088, 1090, 981, 1615, 1625, 1753, 1838, 352, 541, 1407, 853].

Estimation [59, 1098, 1617, 267, 555, 595, 957, 1307, 1404, 2115, 343, 1906, 980, 1272, 1889, 1335, 667, 1524, 693, 1467, 1551, 1462, 494, 2052, 1993, 871, 79, 862, 491, 1923, 1048, 1953, 646, 1698, 1001, 386, 692, 1170, 650, 1772, 1042, 521, 1099, 1773, 832, 113, 828, 389, 1995, 1224, 1122, 816, 1471, 823, 709, 675, 1679, 866, 660, 583, 885, 584, 1929, 1888, 1383, 406, 565, 231, 121, 1213, 1359, 483, 1049, 346, 525, 946, 1533, 2066, 582, 111, 411, 1314, 489, 1149, 1477, 867, 922, 1144, 1508, 1105, 1724, 997, 619, 81, 1215, 1776, 2051, 801, 16]. Estimation [803, 1519, 1270, 1221, 745, 503, 540, 916, 1986, 437, 133, 1116, 1648, 385, 1277, 1094, 2084, 20, 1430, 1150, 613, 1205, 581, 26, 917, 806, 1650, 1263, 809, 674, 1290, 1419, 1946, 1261, 1528, 668, 1169, 45, 1873, 1152, 2049, 2050, 1032, 1093, 300, 1096, 1378, 821, 1670, 408, 1310, 1987, 19, 2, 4, 157, 271, 1515, 1097, 2018, 2062, 1653, 1948, 575, 270, 556, 277, 446, 6, 156, 268, 1402, 1581, 1116, 1002, 372, 43]. Estimator [851, 1797, 359, 986, 676, 126, 1283, 276, 810, 12, 2048, 1435, 1, 1204, 1568, 1284]. Estimators [1803, 770, 11, 70, 863, 1661, 1510, 627, 932, 587, 186, 30, 417, 706, 990, 708, 407, 1047, 391, 1768, 1034, 332]. Ethical [245]. Ethnic [1731]. Étude [193, 760]. Eudyptula [1649]. Eusocial [1378]. Evaluate [1620]. Evaluating [1053, 1832, 1821, 1177, 649, 1357, 580, 1708, 1709, 1982]. Evaluation [266, 1306, 977, 702, 1042, 1228, 460, 1563, 1701, 1564, 391, 1912, 2040, 834]. Evans [889]. Event [1196, 2116, 1420, 1324, 1881, 1959, 1602, 1605]. Event-Related [1881]. Events [171, 1237, 2081, 1501, 1946, 446]. Everitt [255, 552, 1913]. Every [105]. Everyone [896]. Evidence [208, 1850, 1982, 1912]. Evoked [567, 1094]. Evolution [314, 1921, 1104, 915, 955]. Evolutionary [736, 1188, 1274, 1457]. Evolutionary-Symmetric [1457]. Evolving [208]. Exact [210, 1604, 2076, 1286, 692, 1285, 570, 176, 477, 1390, 669, 284, 972, 1265, 1371, 1372, 460, 1025, 397, 1962, 978, 1836, 605, 567, 2084, 26, 1896, 1382, 275, 391, 2021, 1377, 274, 621, 33]. Example [1402, 1238, 483, 141, 732, 631]. Examples [1012, 755, 1129, 1000, 733]. Excess [1524, 2085]. Exchangeability [522, 1161]. Exchangeable [1100, 1507]. Excluding [1606]. Exercise [1123]. Existence [608, 830]. Expanded [786]. Expectation [485]. Expected [31, 26, 32]. Expense [1295, 217]. Experiment [79, 298, 352]. Experimental [379, 1135, 643, 177, 888, 1786, 1063, 2051, 164, 2056, 1935, 266, 254, 1185]. Experiments [887, 1428, 977, 1766, 349, 1040, 1295, 17, 401, 400, 217, 1703, 625, 317, 476, 302, 579, 889, 1616, 137, 926, 937, 1030, 1118, 1393, 67, 1623, 1062, 826, 396, 448, 333, 1362, 800, 1059, 2018, 234, 517, 1306, 1974, 319, 502, 1846]. Experimentwise [286]. Expert [428]. Explanation [726]. Explanations [1850]. Explanatory [979, 911, 1432]. Exploration [1743]. Exploratory [938, 728, 729]. Explore [1206]. Exploring [170]. Exponential [27, 1819, 1890, 2094, 701]. Exponentiality [877]. Exposure [2117, 651, 1872, 2075, 1891, 1922, 288, 485, 1892, 361, 652, 222, 1988].

Exposure-Disease [2075, 652, 222]. Exposures [1714]. Extended [1167, 290]. Extension [267, 555, 2077, 1259, 1834, 218, 401, 113, 933, 1212, 270, 556, 110]. Extensions [1922, 1686]. Extent [2081]. External [524]. Extra [622, 277]. Extra-Binomial [622, 277]. extra-Poisson [277]. Extrapolation [818]. Eye [688].

F [506, 596, 786, 957, 905, 1134, 1135, 1084, 1305, 1544, 1691, 1692, 1918, 1914, 780, 1297, 1246, 1772, 546, 1395, 47, 373, 728, 729, 140, 313, 264, 633, 785, 893, 1133, 1444, 1494, 1631, 195, 1537, 1845, 50, 726, 1064, 593, 1394, 259, 718]. F-Distributions [1772]. Factor [1542, 1492]. Factorial [1366, 400, 1228, 1505, 800]. Factors [342, 1828, 1214, 1772, 710, 1516]. Failure [238, 1882, 2003, 2004, 2000, 281, 1725, 1863, 2059, 540, 1774, 668, 2019, 1152, 1316, 1941, 753, 1466, 1317]. Fairly [161]. Falconer [154, 1589]. Falguerolles [1256]. Falk [1075]. False [673]. Familial [342, 1103, 70, 1758, 1210, 664, 1370, 387, 1047]. Families [1463, 1359, 1890, 2094]. Family [2011, 1477, 1767, 232, 520, 60, 923, 809, 1321, 1814, 61]. Family-Specific [1477]. Family-Wide [60, 61]. Fan [2109]. Fang [597, 196]. Faulbaum [1084]. Fecundability [1658, 399]. Feeding [583, 1118]. Feeding-Preference [1118]. Feigelson [1449]. Feldman [636]. Ferguson [1690]. Fertility [925]. Fertilization [1054]. Festschrift [1400, 1064]. Few [1832]. Fibrillar [1402, 483]. Fidelity [1649]. Fiducial [1143]. Field [426, 401, 402, 29, 925, 234]. Field-Plot [234]. Fields [1299, 1399]. Fieller [119, 1679, 1550]. Fienberg [471, 1792, 208]. Fifty [1248]. Figures [1912]. Filter [583, 454]. Final [344, 1157]. Final-State [1157]. Finance [1399]. Financial [558, 741, 908, 381]. Finding [1573, 2087, 1260, 1666, 322]. Findley [203, 203]. Fingleton [141]. Finie [1116]. Finite [1571, 1834, 591, 620, 173, 232, 1939, 233, 1116, 705, 1606]. Finn [1976]. First [684, 1192, 594, 469, 1008, 1296, 53, 1005]. Firth [2114]. Fish [2105, 1716, 579, 1878]. Fisher [902, 1587, 1143, 1628, 941, 304, 379, 161, 33, 570, 166, 1006, 163, 164, 165, 162, 368]. Fisheries [1072, 846]. Fishing [1999]. Fissuration [771]. Fit [1726, 1258, 518, 1962, 1955, 1416, 1827, 1325, 1768, 388, 1110, 1927, 323]. Fitting [1423, 473, 1638, 702, 2003, 2004, 177, 884, 1281, 1318, 24, 178, 338, 2065, 35, 1945, 285]. Five [303]. Five-Degree-of-Freedom [303]. Fixed [303, 1772, 659, 176, 1562, 1332, 1236, 1509]. Fixed-Distance [659]. Fixed-Dose [1509]. Fixed-Effect [1772]. Fixed-Width [303]. Fleming [767, 820]. Fletcher [1914, 1297]. Flexible [1866, 1760, 988, 859, 1326]. Fligner [2035]. Florence [795]. Florida [1511]. Flow [376]. Follmann [2116]. Follow [1503, 288]. Follow-Up [1503, 288]. Following [853, 1485, 530, 1321]. follows [2101]. Folly [1248]. Food [1199, 952, 1037, 994, 1729]. Foodchain [94]. Foot [348]. Forage [2015].

Ford [1191]. Forecast [145]. Forecasting [2107, 181, 1795]. Forensic [1912]. Forest [1581, 1467, 1991, 235, 1116, 1014, 427]. Forested [1071]. Forestiers [1116]. Forestry [1029]. Form [1898, 1227, 1132]. Formation [1809]. Formula [1643, 1554]. Formulae [1264]. Formulas [1521]. Forschung [1848]. Forthofer [1300]. Fortuner [844]. Forum [776]. Found [1982]. Foundation [377, 1076]. Foundations [154, 1080]. Four [400, 1733, 1275]. Four-Level [400]. Four-Sequence [1733]. Four-Taxon [1275]. Fowler [426, 1859]. Fox [148, 636, 734, 2040]. Fractal [154, 1589, 1132]. Fractals [1299]. Fraction [1580, 709, 26]. Fractions [328]. Frailties [1617, 1965]. Frailty [283, 282, 619, 973, 1719, 1838, 1546]. Framework [1667, 1824]. Francis [1083, 192]. Frank [1086]. Franklin [839]. Fraud [1248, 730]. Fraudulent [189, 365, 418]. Frederick [471]. Free [876, 931, 1482, 564, 1301, 1002]. Freedom [303, 1772, 135]. Freeman [1134]. French [575, 771, 216, 760, 627, 1116, 367]. Frequencies [979, 571, 1378, 1613]. Frequency [688, 1199, 1689, 1918, 278, 657, 1037, 916, 1094, 792]. Frequent [633]. Front [7, 58, 106, 159, 215, 273, 331, 382, 435, 475, 513, 559, 600, 642, 691, 742, 797, 856, 909, 965, 1022, 1088, 1142, 1201, 1257, 1308, 1355, 1405, 1453, 1500, 1545, 1591, 1644, 1697, 1747, 1800, 1860, 1920, 1981, 2044]. Frontiers [956, 635, 428, 210]. FRS [589]. Fry [843]. Full [389, 1100, 1262, 232]. Full-Half [232]. Fuller [1399, 89]. Fully [1501]. Function [1272, 1770, 1889, 1663, 2022, 1986, 437, 1150, 1169, 1153, 1753, 526]. Functional [1194, 1240, 139, 1227, 585, 1478]. Functions [1073, 1581, 1232, 1467, 1313, 650, 1147, 1884, 68, 1764, 178, 711, 1743]. Fundamental [1143]. Fundamentals [1862, 729, 841, 1354, 551, 727, 728]. Further [936, 874]. Fuzzy [1130].

G [266, 510, 638, 955, 956, 1190, 1077, 1079, 1197, 1252, 1253, 1449, 1447, 1542, 1586, 1587, 1588, 1543, 1544, 1495, 1745, 2039, 1977, 2110, 2111, 2114, 144, 201, 775, 41, 590, 1490, 261, 1244, 149, 97, 1009, 150, 942, 1786, 138, 725, 1686, 841, 941, 1130, 47, 1491, 373, 1848, 100, 264, 464, 1247, 1633, 1494, 1909, 320, 372, 998, 1187, 315, 141, 200, 505, 588, 1184, 1014, 94, 549, 98, 501, 203, 427, 1685, 1912]. Gail [1011]. Galambos [1181, 1685]. Galéries [771]. Galleries [771]. Gametes [1090]. Gametic [1510]. Gamma [1617, 622, 973]. Gani [1181]. Garrott [261]. Gatsonis [1789]. Gaussian [433, 704, 1360, 927, 336]. GEE [1726, 1281]. Gehan [1448]. Gelman [1396]. Gels [586]. Gene [1933, 1314, 1928, 1477, 916, 1430, 807, 875]. General [1875, 968, 493, 1775, 1238, 177, 864, 1188, 1028, 1521, 229, 1474, 1824, 1362, 1528]. Generalization [767]. Generalizations [1209]. Generalized [344, 1875, 1112, 1572, 1403, 1003, 1335, 252, 717, 1035, 974, 1580, 1036, 1488, 1172, 1627, 1671, 620, 498, 970, 874, 1961, 824, 1292, 24, 219, 749, 415, 1840, 1229, 62, 1520, 759, 438, 934, 1419, 1943, 1555, 1325, 981, 1615, 1625, 984, 1162, 1427, 318, 467, 191]. Generated [1249, 1430]. Generation [1645, 1971, 46]. Genes [1792, 1598, 772]. Genetic [1499, 204, 325, 1867, 550, 838, 1012, 1397, 1442, 375, 1565, 1207, 1804, 1163, 1868, 831, 2057, 1700, 1205, 1263, 614, 809, 21, 788, 1497, 375]. Genetical [2108, 162]. Genetically [1470]. Geneticist [161]. Genetics [1856, 204, 163, 161, 97, 324, 98, 203, 682]. genetik [549]. Genome [476]. Genomes [1693]. Genomic [1551, 2046]. Genotype [1608, 109, 1532, 2065, 615]. Genotype-By-Environment [2065]. Genotype-Environment [1532]. Genotypes [366, 1598, 1378, 1924]. Genotypic [761]. Genotyping [1519]. GENSTAT [679]. Gentle [1971, 1971]. Geographic [1996]. Geometric [1335, 590, 1490, 1658, 922, 801]. Geometrical [1299]. Geometry [154, 1250, 593, 203, 1589]. Gerber [1976]. Gershenfeld [2102]. Getz [200]. Gianola [325]. Gibbons [1401, 1496]. Gibbs [575, 1676, 653]. Gifi [724]. Gijbels [2109]. Gilbert [1009]. Giller [1190]. Gillespie [1856]. Gillman [100]. Giltinan [1583]. Giri [1017]. Given [1724, 1384]. Glasbey [1543]. Glasserman [1196]. GLIM [1083, 192]. GLM [537]. Global [1578, 744, 1955, 259, 211]. GLS [1241]. Gnanadesikan [1445]. Godbole [1252]. Gold [1694, 1461, 1708, 1709]. Golden [846, 901]. Goldstein [596]. Goode [2039]. Goodman [98]. Goodness [1726, 323, 518, 1962, 1955, 1416, 1927, 1827, 1325, 388]. Goodness-of-Fit [1726, 1955, 1416, 1827, 1325, 388, 1927]. Goodwin [736]. Goosens [1444]. Gosset [464]. Govindarajulu [686]. GPSS [469]. Graft [1863]. Graphical [321]. Graphics [55, 1347]. Graphs [1013]. Grasses [2065]. Grätzer [1633]. Gray [376]. Grazing [201]. Green [1083, 1003]. Greenhouse [1296]. Greenwood [1681, 1643, 1554]. Gregoire [1014]. Gregor [163]. Grenander [1911]. Grice [95]. Grimm [942]. Groundwater [1401]. Group [172, 1381, 2008, 1176, 1485, 478, 1390, 1888, 698, 1160, 2011, 123, 1025, 1124, 2093, 305, 1897, 397, 1379, 1156, 767, 978, 1840, 572, 396, 1290, 1472, 564, 2074]. Grouped [647, 929, 1567, 281, 1358, 913, 289, 610, 701, 294, 2001, 124, 753]. Grouping [754]. Groups [1548, 218, 754, 322, 767]. Groves [259, 548]. Growth [1113, 17, 1786, 1132, 1713, 406, 1262, 229, 297, 296, 2027, 295, 1337, 1528, 2017, 2021, 427, 1937, 1091, 950]. Guénoche [468]. Guide [101, 639, 790, 1300, 1976, 1681, 547, 896, 1444, 2104, 960, 1445, 1856, 85, 1298, 52, 148]. Guidelines [987, 307, 858, 775, 959]. Gutierrez [1493]. Guttorp [1582, 1070].

H [92, 431, 509, 471, 954, 901, 906, 1136, 1192, 1198, 1640, 1799, 1917, 2108, 1979, 1977, 1980, 1072, 594, 198, 1629, 1536, 139, 147, 204, 888, 368, 1396, 317, 1130, 1395, 47, 465, 44, 199, 1245, 1299, 313, 264, 553, 896, 1444, 1494, 1973, 95, 1189, 502, 679, 374, 1000, 1065, 1005, 726, 505, 1394, 1788, 96, 1014, 1011, 94, 39, 259, 1912]. H.-J [1198]. Haaland [254]. Habitat [1071]. Haemophilus [2097]. Haenszel [1087, 1412, 12, 753, 935, 1593]. Haenszel-Type [1087, 1412, 935]. Hahn [725, 896]. Haight [200]. Hald [1787, 369]. Half [232]. Hambleton [1849]. Hamilton [684]. Hammond [325]. Hand [2106, 1246, 1585, 424]. Handbook [1246, 1245, 1398, 679, 681, 1855, 1244, 142, 1397, 1849, 948, 1247]. Handbuch [549]. Handcock [1296]. Haphazard [1330]. Haplotype [73]. Harbour [1809]. Hard [1388]. Härdle [899]. Hardy [2089, 72, 477, 1057, 1373, 130]. Harley [94]. Harris [466]. Harrison [2107]. Harter [954, 553, 431, 509, 954, 553]. Hartl [97]. Hartmann [726]. Harvesting [200]. Harville [1798]. Hasard [1193]. Hastie [318]. Hastings [104]. Hatching [1878]. Hauck [939]. Haul [921]. Hawkins [963]. Hayakawa [1541, 1746, 1975]. Hazard [1889, 986, 14, 1884, 1618, 803, 1150, 1864, 526, 535]. Hazards [870, 1320, 238, 1578, 2071, 1663, 2000, 752, 697, 1147, 389, 1825, 1936, 170, 444, 798, 751, 68, 768, 1780, 125, 528, 1769, 1363, 1824, 1771, 696, 1826, 1827, 2018]. Head [1478]. Heading [2065]. Healey [264]. Healing [2026]. Health [1248, 211, 377, 432, 510, 598, 794, 904, 906, 1020, 1076, 1077, 1085, 1141, 1587, 1641, 1852, 1919, 255, 181, 941, 1008, 768, 253, 145]. Hedayat [591]. Hedges [726, 1245]. Heiberger [319]. Hellekalek [2110]. Heller [735]. Henderson [52]. Hengeveld [258]. Herbert [686]. Heredity [163]. Heritabilities [2015]. Heritability [1604, 1813, 1778]. Herrendörfer [549]. Heterogeneity [765, 1900, 1608, 23, 1173, 992, 439, 831, 1951, 1331, 677, 567, 923, 614]. Heterogeneous [850, 713, 1556, 1756, 242]. Heteroscedastic [644, 118, 1876]. Heteroscedasticity [1120]. Hidden [1553, 1952, 1684]. Hierarchical [2097, 2071, 27, 1473, 2055, 1895, 2009, 363, 1054]. High [1408]. High-Dimensional [1408]. Higher [661]. Higher-Order [661]. Highest [1116]. Highly [1727, 2071, 2002, 1830, 1166]. Hildebrand [1799]. Hildrew [1190]. Hill [508, 508]. Hinde [192]. Hines [502]. Hinkelmann [1062]. Hinkley [1399, 589, 1737]. Hirayama [429]. Hirsch [1008]. Histograms [879]. Historical [339, 1813, 1436, 1537, 1165, 484, 758, 2090]. History [687, 1143, 1787, 355, 145, 369, 901]. Hit [759]. HIV [1872, 1880, 1963, 180, 1108, 917, 182, 1932, 918]. HIV-1 [1932]. Hjorth [890]. HLA [73, 72, 993, 1166, 1057, 221]. HLA/Disease [221]. Hoaglin [471, 728, 729]. hoc [1025]. Hocking [1907]. Hodges [1789]. Hoffmann [1494]. Hoffmann-Jorgensen [1494]. Holland [1793]. Holst [1494, 1068]. Holt [506]. Home [1204]. Home-Range [1204]. Homogeneity [1499, 911, 1273, 1442, 1163, 1441, 1657, 1899, 1373, 306]. Homoscedasticity [533, 779, 1482]. Honor [686, 1541, 1740, 1746, 1975, 1064]. Honour [1016, 1400, 1181, 589]. Hoppe [1064]. Hoppensteadt [1918]. Horgan [1543]. Hormonal [1217]. Horn [1854]. Horvitz [1803]. Horvitz-Thompson [1803]. Hosmer [253, 422]. Hotelling [656]. Household [1998]. Hsu [1632]. Hubert [1588]. Huberty [1182]. Hughes [1796, 1796]. Hull [1829, 1204]. Hull-Based [1204]. Human [1813, 550, 1397, 1748, 976, 1326, 1090]. Hutchinson [845, 316, 845]. Huxley [950]. Hybrid [1457]. Hybridization [1551]. Hydropsyche [1478]. Hymenoptera [1378]. Hypergeometric [534]. Hypotheses [1380, 114, 934, 1635, 87]. Hypothesis [307, 1455, 664, 1868, 1363, 1883, 1275, 674, 543, 775, 1456].

I/II [1165, 1666]. Ideas [1251]. Identifiability [949]. Identification [2083, 109, 927]. Identifying [1880]. Identity [807]. Ignorability [755]. Ignorable [1164]. Ignoring [774]. II [1638, 1494, 207, 1218, 1114, 1425, 84, 1165, 1668, 2029, 440, 1666, 956, 1497, 206]. II. [862]. IIB [858]. III [431, 1741, 893, 1857]. Ill [450, 145]. Ill-Posed [450]. Illustration [1575, 1735]. Image [1566, 1089, 1543]. Images [1850]. Immigration [77]. Immune [1203]. Immunity [2097, 206]. Immunoassay [1272, 1646, 1603, 960]. Immunodeficiency [1748, 976]. Immunogold [670]. Immunosorbent [69]. Impact [1580, 1964, 439]. Implants [1711]. Implementable [1817]. Implementation [123]. Implications [881, 1711, 384]. Implicitly [1044]. Import [1097]. Importance [938]. Impossible [1233]. Improve [2117, 1225, 1988]. Improved [1271, 873, 2057]. Improvement [325, 769, 1530, 1185]. Improvements [1721]. Improving [666]. Imputation [1664, 1058, 390, 1625]. In-Vitro [412]. Inbreeding [1606]. Incidence [493, 1960, 1663, 649, 747, 1042, 2067, 584, 1660, 1616, 1318, 1609, 995, 445, 917, 2080, 2061, 182, 1731, 1472]. Including [763, 1316, 1317]. Inclusion [1880]. Incomplete [929, 2003, 2004, 1703, 910, 2073, 293, 1376, 1383, 1503, 1765, 1886, 1267, 2030, 1211, 1454]. Inconsistency [986]. Incorporating [339, 1218, 1813, 716]. Increase [31, 32]. Increased [703]. Increasing [1295, 217]. Incubation [343]. Independence [393, 1518, 1050, 1344, 1431, 621]. Independent [558, 741, 908, 1993, 1119, 381, 829, 1780, 489, 34, 1864, 1966, 1574, 1967]. Index [645, 1749, 233, 405, 66]. Index-Removal [1749]. Indicators [2006]. Indices [954]. Individual [1258, 452, 2089, 1998, 1208, 2079, 1672, 656, 1487, 1111]. Individuals [1463, 1880, 1203]. Induced [1470, 2063]. Industrial [635]. Inequalities [1685]. Infant [976, 1937]. Infected [1463]. Infection [343, 816, 1880, 1108, 1724, 756, 917, 182]. Infectious [344, 144, 1157, 1727, 1048, 65, 756, 445]. Infectiousness [1998]. Infectivity [180]. Inference [1366, 595, 685, 956, 1743, 1272, 1038, 1517, 1422, 1989, 283, 591, 368, 1143, 2068, 131, 1990, 1656, 1646, 723, 1562, 1890, 502, 1314, 2078, 397, 1669, 1412, 1861, 1700, 1055, 750, 662, 1651, 503, 93, 1070, 384, 924, 1883, 1457, 2021, 562, 1603, 1288, 1091, 65, 379, 638, 905, 1135, 1192, 2037, 194, 727, 1443, 314, 140, 1739, 1394, 999]. Inferences [293, 1219]. Inferential [490, 1932]. Inferred [228]. Inferring [1123]. Infestante [760]. Infesting [760]. Inflation [1121]. Influence [342, 1815, 1578, 487, 1438, 173, 1712, 81, 984]. Influence-Based [173]. Influential [1413, 297, 2095]. influenzae [2097]. Information [160, 383, 560, 743, 966, 1202, 1304, 1406, 1592, 1801, 2045, 1813, 1295, 217, 1703, 1224, 240, 1160, 1424, 75, 1096, 504, 1538]. Informative [6, 156, 268, 647, 720, 1036, 1488, 801, 1807, 544, 1677, 1805, 1808, 2, 4, 1869, 157, 271]. Informatively [1092]. Infromation [269]. Inhomogeneity [129].

Inhomogeneous [360]. Initiation [1226, 1945]. Input [65]. Insect [355, 759]. Inspection [793]. Institute [904]. Institutional [822]. Instrument [673]. Insults [1174]. Intake [1199, 1964, 1037, 1987]. Integral [1517, 450]. Integrals [188]. Integrated [1647, 579]. Intelligence [844, 428, 1792]. Intensity [1227, 1277]. Intensive [890, 87]. Intent [1667, 1421]. Intent-to-Treat [1667, 1421]. Intention [1376]. Intention-to-Treat [1376]. Intentionally [1439]. Interaction [25, 1115, 2069, 1725, 1391, 1950, 1294, 225, 2084, 611, 1162, 1811]. Interactions [1468, 2055, 655]. Interactive [198]. Interblock [1703]. Intercept [766]. Interclass [70, 1370]. Intercourse [860]. Interest [455, 1389]. Interference [1327]. Intergradient [1703]. Interim [977, 930, 1155, 1342, 1420, 1176, 1440, 698, 1723, 36, 1306]. Interindividual [1206]. Interlaboratory [1876]. Intermédiaires [216]. Intermediate [216]. Intermittent [1953, 651]. International [524, 98]. Internet [1695]. Interobserver [881]. Interphase [1123]. Interpolation [903]. Interpretation [1758, 413, 1008, 1682]. Interpreting [1658, 120, 1056, 53]. Interrater [1336, 1526]. Interrelations [1853]. Interspecific [1925, 1926]. Interval [1098, 267, 555, 1889, 459, 79, 491, 1358, 1099, 832, 113, 1825, 1936, 1679, 660, 1600, 1618, 411, 1437, 1311, 1724, 1221, 1651, 540, 122, 1769, 804, 2084, 826, 2081, 1945, 1896, 1956, 1605, 1946, 1169, 391, 22, 1594, 1107, 270, 556, 1106, 1307]. Interval-Censored [1889, 1099, 1825, 1600, 1618, 540, 122, 1769, 804, 2081, 1945, 1956, 1169]. Intervals [1958, 1934, 169, 1604, 2076, 1315, 657, 303, 1426, 1386, 414, 1390, 604, 565, 1265, 1029, 239, 1144, 34, 76, 1520, 1874, 1550, 1417, 725]. Intervariety [1703]. Intervention [490]. Interventions [1557]. Intra [921]. Intra-Haul [921]. Intraclass [2114, 814, 876, 1438, 1359, 820, 1311, 1877, 2021, 1814]. Intracluster [1783]. Intraindividual [606]. Introduction [731, 1918, 198, 945, 252, 322, 256, 633, 1347, 1633, 1399, 791, 1399, 1537, 104, 683, 688, 1017, 1300, 1544, 1693, 1795, 198, 149, 262, 49, 892]. Introductory [1018, 1137, 1688]. Intuitive [1629]. Invalid [587]. Invariant [1358]. Invariants [662]. Invasions [259, 258]. Inventaires [1116]. Inventories [1116]. Inventory [1014]. Inverse [1098, 1307, 568, 1650, 21]. Inversion [71]. Investigation [836, 574]. Investigations [1728, 255]. Invited [107]. ion [2006]. Irreducibility [615]. Irreducible [1051]. Isotonic [2026]. Issues [788, 1964, 1133]. Item [1849, 957]. Iterative [974]. IV [509, 1975].

J [102, 153, 104, 101, 430, 506, 596, 472, 471, 788, 683, 789, 790, 640, 734, 791, 1015, 843, 900, 956, 906, 848, 897, 902, 1135, 1249, 1250, 1192, 1197, 1074, 1081, 1139, 1198, 1445, 1449, 1349, 1301, 1304, 1450, 1351, 1539, 1588, 1637, 1638, 1496, 1498, 1855, 1795, 1689, 1857, 1856, 1853, 1742, 1978, 2107, 2042, 2039, 2108, 2109, 1979, 1976, 2038, 1914, 2043, 1980, 255, 940, 887, 1003, 426, 780, 1297, 594, 318, 590, 834, 1490, 890, 945, 836, 2105, 252, 469, 151, 2035, 147, 1186, 1246]. J [1001, 1183, 368, 550, 838, 1012, 1397, 192, 547, 86, 1010, 1396, 376, 725, 1188, 46, 322, 546, 1686, 468, 681, 1395, 47, 1181, 321, 1182, 1491, 1849, 373, 1848, 1132, 1584, 205, 889, 728, 729, 723, 148, 257, 256, 264, 323, 589, 635, 896, 893, 1243, 1296, 1494, 1631, 1973, 1971, 2112, 195, 370, 95, 894, 191, 143, 502, 679, 1845, 1789, 372, 53, 1067, 1687, 145, 253, 141, 1632, 38, 726, 421, 54, 950, 1394, 1585, 424, 784, 96, 40, 94, 682, 45, 428, 2104, 98, 947, 39, 1844, 43, 259, 1685]. J.-M [1844]. J.-P [940, 1186, 1010, 468, 681, 1491]. j.0006-341X.1999.00603.x [2116]. Jackknife [932, 1283, 276]. Jackson [547]. Jacobson [707]. Jain [682]. James [515]. Jansen [1349]. Janson [1068]. January [558, 741, 908, 381]. Japan [1097, 429]. Jaworski [2112]. Jenks [378]. Jensen [894]. Jeu [1193]. JMP [2104, 2104]. Joe [1788]. John [209, 1249, 1065]. Johnson [793, 1634, 1740, 1791, 2043, 781, 1242, 1738, 312, 1741, 2043]. Johnstone [896]. Joint [1522, 609, 1105, 1607, 448, 859, 1481]. Joint-Regression [1607]. Jointly [455]. Jolicoeur [892]. Jolly [1034]. Jones [1072, 320, 1683, 1189]. Jongman [47]. Jöreskog [1542]. Jorgensen [1494, 1128]. Jump [1812]. Jumping [1429]. Jumps [925]. June [2115, 2116]. Jurecková [1853].

k-Sample [1842]. K.-T [196]. K.-Y [1539]. Kaandorp [1132]. Kahler [204]. Kalman [454]. Kaplan [708]. Kappa [59, 1273, 832, 660, 10, 661]. Kappa-Type [661]. Kappel [596]. Kapur [472, 1304]. Karian [469]. Karlqvist [1188]. Karr [905]. Kass [1789]. Kaufman [322]. Kavalieris [1914]. Kearsey [2108]. Keller [264]. Kemp [781]. Kempthorne [1062]. Kempton [2040]. Kendall [894, 1225, 1443, 1395, 723, 1243, 1394]. Kendler [788]. Kenward [320]. Keppel [138]. Kernel [1418, 1170, 1995, 1715, 1683]. Kernel-Based [1170]. Kernels [1429, 803]. Key [1943]. Kidney [1725]. Killed [348]. Kimber [848]. Kimura [1240]. Kinetic [538, 415]. Kinetics [623, 632]. Kitagawa [1977]. Klaassen [1001]. Klar [253]. Kleffe [43]. Kleinbaum [1447]. kleiner [1848]. klinische [1848]. Know [105]. Knowledge [1454, 1841, 1786]. Known [161, 754, 997, 652, 2048, 19]. Knox [1950]. Knudson [1209]. Kocherlakota [782, 782]. Kohne [1306]. Kolassa [1742]. Koopmans [198, 198]. Kopka [896, 1444]. Kotz [793, 1741, 1736, 1785, 2033, 781, 1242, 196, 1738, 312, 1634, 1740]. Krewski [839]. Kriging [1986, 291]. Krishnaiah [142]. Krishnan [1686]. Kruger [259]. Kruskal [471]. Krzanowski [1395, 1394]. Kumar [1400]. Kurzgefasste [1848]. Kutner [39]. Kutzbach [100].

L [632, 209, 328, 431, 266, 430, 509, 470, 473, 683, 793, 684, 732, 638, 954, 846, 901, 906, 848, 1077, 1445, 1539, 1587, 1588, 1634, 1635, 1740, 1794, 1741, 1696, 1799, 1917, 1914, 887, 1736, 426, 1072, 951, 198, 1684, 97, 327, 147, 204, 262, 942, 1188, 322, 941, 1785, 952, 48, 1245, 1584, 140, 100, 148, 264, 553, 464, 1180, 1296, 1444, 1398, 1494, 2033, 95, 894, 781, 949, 1242, 196, 554, 1738, 1068, 726, 593, 947, 203, 312]. Laake [947]. Laboratory [959]. Lack [1768]. Lack-of-Fit [1768]. Lad [1537]. Lai [845, 316]. Lajos [1181]. Lamport [1444]. L’Analyse [1004]. Land [787]. Landmark [1691, 593]. Landscape [47]. Lange [1296]. Langholz [2115]. Larcher [2110]. Large [1690, 880, 866, 1473, 429, 1509, 1067]. Large-Scale [429]. Larva [760]. Larve [760]. Latency [816]. Latent [612, 1048, 2028, 659, 936, 1525, 569, 1046, 1357, 2006, 1332, 2082, 1893, 705, 1559, 1369]. Latin [1269]. Lattice [1703]. Lavit [90]. Law [1057, 130, 1912]. Lawler [1544]. Layout [654]. Layouts [1621, 1183]. LD [79]. Lead [1472]. Leading [1634]. Learning [1447]. Least [1364, 974, 799, 117, 749, 1837]. Lecture [1143]. Lectures [1628]. Lee [1251, 1300, 1069]. Leemis [1696]. Left [1052, 1923, 1769]. Left- [1052]. Left-Censored [1923]. Left-Truncated [1769]. Legait [637]. Legay [1844]. Lehman [2104]. Lehmann [1635]. Leitfaden [1848]. Lelièvre [91]. Lemak [1448]. Lemeshow [591, 253, 2041, 422]. Lemons [1498, 1498]. Length [1572, 454, 1039]. Leslie [110]. Lesquoy [898]. Lesquoy-de [898]. Lethal [121, 1314]. Letter [1782, 2032, 1535]. Leukemia [1863]. L’Évaluation [51]. Level [400, 1624, 1165]. Levels [1728, 1963, 497, 539, 361]. Levy [2041, 591]. Lewis [956, 902, 148, 1066]. L’expérience [1844]. Liaison [771]. Liang [1539]. Library [1395, 1394]. Lienert [1848]. Life [351, 48, 355, 2060, 1719, 45, 1768, 429]. Life-Style [429]. Lifetime [650, 298, 2092, 1948]. Ligand [926]. Ligand-Binding [926]. Light [726]. Likelihood [59, 1875, 511, 1743, 1551, 1462, 1040, 119, 1626, 1778, 989, 910, 521, 389, 1100, 1510, 709, 302, 1167, 620, 1660, 2014, 1049, 970, 1562, 1262, 460, 1701, 489, 937, 867, 1886, 1700, 81, 1235, 358, 715, 2051, 1831, 16, 662, 1334, 540, 133, 1994, 1223, 581, 26, 1983, 438, 1027, 1822, 1982, 1378, 1876, 928, 1670, 859, 1415, 1284, 786]. Likelihood-Based [1551, 910, 1049, 1700, 1235, 2051]. Likelihoods [1898, 1895]. Lilienfeld [1080]. Limit [1752]. Limited [1906]. Limiting [266, 1381, 657, 1359, 1235, 1817]. Limits [1239, 12, 1610]. Linden [1849]. Linder [737, 738]. Lindgren [1082, 197]. Lindley [1305, 1134]. Lindman [888]. Lindsey [1689]. Line [1803, 1802, 536, 1418, 1757, 1993, 1750, 147, 1995, 714, 1213, 896, 1715, 766, 29, 76, 353, 237, 774, 1983, 1864, 300]. Lineage [1411]. Linear [326, 1875, 156, 1403, 1639, 1230, 1458, 1003, 594, 1128, 1524, 1604, 252, 1614, 1462, 1278, 1206, 1285, 1036, 1488, 1579, 1224, 1281, 561, 1172, 444, 1506, 620, 498, 1971, 1600, 1820, 191, 768, 1712, 874, 515, 1556, 1961, 457, 1564, 24, 1962, 1831, 662, 415, 605, 1229, 62, 1223, 438, 934, 1322, 1528, 1323, 1907, 39, 1054, 1561, 4, 1550, 984, 1162, 157, 470, 1636, 42, 1909]. Lines [1403, 1120, 1229]. Link [1581, 1467, 771, 1594]. Link-Based [1594]. Linkage [1499, 1442, 1804, 1312, 1163, 1868, 1429, 831, 1807, 1430, 614, 1805, 1808, 808, 550, 1397]. Linked [69]. Literature [2079]. Litigation [1912]. Litters [278]. Little [1649]. Liu [2042, 1186, 1010]. Live [1522, 1222]. Live-Recapture [1522]. Lives [2016]. Livestock [325]. Living [52]. Local [1578, 1889, 389, 1995, 1820, 1712, 2109]. Locally [1948]. Location [865, 1519, 1417, 901]. Locations [269, 75]. Loci [459, 1207, 1521, 1312, 1651, 1417]. Lock [1002]. Locus [109, 1166, 1519, 356, 1812]. Log [1675, 930, 2004, 1281, 1659, 1600, 1340, 1962, 345, 357, 1561, 711, 1966, 1574, 1967, 535, 2003, 1261]. Log- [2004, 2003]. Log-Linear [1281, 1600, 1561]. Log-Normal [1574, 1967]. Log-Rank [930, 1659, 1340, 345, 711, 535]. Logistic [1504, 112, 979, 155, 488, 1258, 169, 183, 1484, 1483, 386, 692, 1661, 1758, 1386, 1624, 709, 1984, 1329, 5, 948, 1669, 1374, 1783, 83, 1102, 422, 285, 1454, 1677, 2049, 2050, 387, 391, 1935]. Logistic-Bivariate [1484]. Logistic-Normal [1984]. Logit [1958, 1900, 226]. Logit-Rank [226]. Loglinear [1230, 455]. Lognormal [294, 186, 417]. Lognormal-Based [186, 417]. Logofet [1013]. Logrank [1160, 123, 1843, 1530]. London [161, 1143]. Long [1139, 1676, 1826]. Long-Term [1139, 1676, 1826]. Longitudinal [171, 2086, 507, 1200, 1539, 2106, 18, 862, 1420, 720, 1698, 2068, 2026, 910, 521, 653, 1061, 1763, 1984, 1713, 230, 231, 970, 1557, 1886, 1421, 749, 520, 1270, 750, 868, 394, 519, 544, 991, 1512, 1677, 1870, 1871, 115, 1481, 918, 1189]. Lorenzen [887]. Loss [84]. Lotka [1973]. Louis [726, 1489]. Louwes [147]. Low [529]. Lower [22]. Lübbert [1197]. Lumer [955]. Lunn [1246]. Lwanga [253]. Lwin [153]. Lymphocyte [975]. Lyons [788].

M [895, 55, 210, 101, 506, 473, 471, 788, 683, 789, 636, 787, 962, 848, 1134, 1195, 1250, 1083, 1445, 1541, 1693, 1746, 1854, 1696, 1743, 2107, 2108, 1975, 2106, 2114, 2117, 319, 255, 1681, 1072, 1683, 871, 2035, 943, 1583, 147, 204, 192, 51, 841, 1130, 48, 889, 1002, 1180, 1247, 1296, 1444, 1494, 1739, 1911, 146, 249, 320, 1845, 420, 1067, 1187, 49, 548, 200, 505, 1064, 424, 1011, 94, 98, 39, 1844, 203, 259]. M-Estimation [871]. Macdonald [1684, 326]. Machin [255, 789]. MACSYMA [735]. Mager [57, 840]. Magnus [1979]. Magurran [99]. Mahalanobis [1613]. Main [1501]. Major [109, 1477, 772, 875]. Making [96]. Male [665]. Malformations [867]. Mallows [209]. Management [689, 846, 264, 1072, 2112]. Managers [1799]. Manatee [1511]. Mandel [834, 1183]. Manly [951, 1914, 780, 1297, 546, 1631, 195]. Mann [78, 1137, 1259]. Mantel [1087, 1412, 12, 753, 935, 1593]. Manton [1130]. Manual [1852, 198, 1444, 1083]. Many [2024]. Maple [1251]. Mapped [269, 75]. Mapping [459, 476, 1207, 1521, 1651]. Maps [1693, 1089]. Marazzi [1073]. March [2114]. Marden [1978]. Mardia [588]. Marginal [1164, 746, 1567, 2000, 2026, 1386, 1281, 539]. Marginally [1984]. Margins [1968]. Marine [301, 921]. Maritz [153, 1301]. Mark [1802, 236, 491, 764, 985, 579, 1337, 806, 77]. Mark-Recapture [1802, 236, 491, 764, 985, 579, 806, 77]. Marker [2086, 1312, 1165, 1700, 1231, 917, 2080, 772]. Marker-Based [2080]. Markers [1821, 1933, 1807, 1430, 1805, 1808, 1349]. Markov [640, 1684, 395, 802, 1484, 2020, 1511, 1953, 1099, 1825, 522, 1706, 1699, 1553, 460, 1839, 1928, 1779, 489, 1051, 1861, 355, 179, 1945, 615, 1382, 1812, 1946, 1117, 1268]. Markov-Dependent [489]. Markovian [763, 915]. Marriott [1395, 313, 1394]. Marrow [1863]. Mars [2058]. Marten [201]. Martens [374]. Martingale [1929, 1827, 352]. Marubini [1187]. Masson [940]. Matched [1258, 1367, 284, 80, 480, 1612, 224, 275, 391, 306, 274]. Matched-Pairs [480, 1612, 275, 274]. Matches [480]. Matching [2002, 11, 1279, 810]. Mate [1649]. Math [1633]. Mathai [210]. Mathematical [1013, 209, 154, 683, 256, 1787, 2102, 1537, 54, 689, 1251, 1859, 146, 554]. Mathematics [104, 507, 636, 1450, 1918, 100, 1129, 2036, 378, 687, 1398]. Mathematik [1019]. Matheson [1185]. Mating [1166, 498]. Matings [1263, 1606]. Matrices [1565, 1385, 527, 1013]. Matrix [1404, 1521, 1215, 1922, 71, 21, 110, 1798, 1979, 199]. Matter [7, 58, 106, 159, 215, 273, 331, 382, 435, 475, 513, 559, 600, 642, 691, 742, 797, 856, 909, 965, 1022, 1088, 1142, 1201, 1257, 1308, 1355, 1405, 1453, 1500, 1545, 1591, 1644, 1697, 1747, 1800, 1860, 1920, 1981, 2044]. Maurice [1016]. Maximally [1904, 1905, 409, 442]. Maximum [59, 1875, 1752, 1462, 1778, 521, 1510, 709, 620, 970, 1562, 1160, 123, 489, 867, 81, 1831, 16, 662, 540, 133, 1223, 581, 26, 1378, 1876, 928, 859, 1284, 472]. Maximum-Entropy [472]. Maximum-Penalized-Likelihood [489]. Maxwell [317]. Mayo [1786]. Mazumdar [256]. McCabe [149]. McClave [264]. McConway [1246]. McCullagh [191]. McCulloch [1789, 998]. McDonald [951]. McGill [55]. McGlynn [203]. McLachlan [1686, 841]. McNemar [1382]. Mead [137]. Mean [494, 1760, 2022, 650, 303, 1485, 1459, 309, 1265, 80, 534]. Means [878, 694, 451, 1029, 334, 1966, 1574, 1967]. Measure [716, 1613]. Measured [1821, 1931, 1818, 1262, 226, 1481]. Measurement [433, 1336, 1238, 1614, 1866, 1757, 1040, 1291, 1549, 1583, 1475, 936, 1386, 1172, 607, 1961, 499, 1564, 528, 813, 1623, 437, 416, 774, 336, 1501, 1310, 984, 99, 1584, 89]. Measurements [983, 1997, 542, 2091, 799, 297, 1231, 1840, 2056, 333, 834, 1395]. Measures [507, 1195, 1578, 748, 1376, 695, 366, 116, 1365, 499, 1416, 396, 1343, 696, 563, 436, 1987, 1516, 1869, 424, 1304]. Measuring [1118]. Mécanistes [216]. Mechanism [1383, 672, 384]. Mechanisms [929]. Mechanistic [216]. Medhi [1081]. Median [121, 668]. Medical [508, 1448, 1495, 255, 8, 1701, 52, 1503, 1007, 49, 501, 255, 789, 790, 893]. Medication [497]. Medicine [904]. Meeker [725]. Meier [958, 708]. Meldahl [427]. Member [1359]. Members [808]. Memorial [1143, 1628]. Memory [1575, 1735, 1969]. Mendel [163]. Mendenhall [327, 264, 554]. Mengersen [956]. Merzbach [687]. Meta [339, 1483, 974, 1661, 1990, 1149, 2079, 1672, 1322, 726]. Meta-Analyses [1483]. Meta-Analysis [339, 974, 1661, 1990, 1149, 2079, 1672, 1322, 726]. Metaregression [1964]. Metastasis [346, 1515, 1753]. Methadone [1602]. Method [968, 1255, 340, 703, 536, 1418, 283, 119, 571, 1278, 1026, 1823, 1206, 1997, 1880, 132, 873, 1950, 618, 1025, 937, 1609, 766, 939, 1674, 9, 1334, 1622, 492, 26, 1457, 182, 1771, 1708, 1528, 1873, 82, 1361, 1653, 1709]. Méthode [898, 1844]. Methoden [1794]. Méthodes [51]. Methodology [639, 1587, 1638, 1857, 201, 941, 148, 1362, 883, 680]. Methods [210, 379, 787, 792, 844, 958, 1074, 1198, 1401, 1638, 1496, 1495, 1852, 1743, 2040, 2041, 1815, 188, 255, 1038, 1433, 1072, 189, 365, 590, 1490, 87, 236, 890, 1737, 702, 591, 2035, 1750, 325, 606, 1177, 1069, 375, 1476, 546, 1748, 1872, 14, 675, 44, 1577, 1089, 1299, 1929, 1492, 1631, 146, 249, 1124, 1489, 1537, 457, 1105, 1886, 420, 1564, 1067, 207, 1715, 1000, 1861, 1235, 1389, 1922, 50, 2060, 1225, 2065, 1851, 1031, 826, 1095, 1754, 519, 361, 418, 1983, 1822, 1014, 564, 859, 1723, 1912, 1625, 1574, 388, 277, 845, 198]. Methods [1971, 153, 842, 1301, 1696, 198, 139, 197, 88, 1910, 503, 1632, 1847, 265, 595, 1907]. Methods/Comparisons/Research [210]. Metropolis [1429]. Meulen [635, 428]. Meyerthole [1197]. Mezei [473]. Microbiological [1388]. Microbiology [1191]. Microcomputers [790]. Microsome [658]. Mid [284]. Mid-P [284]. Midge [130]. Migration [616, 1469]. Miller [898, 1586, 421]. Milton [594]. Minimally [2053, 2054]. Minimum [2083, 1773]. minor [1649]. Minoru [1541, 1746, 1975]. Misclassification [362, 288, 496, 1731]. Misclassified [2075, 1922, 821]. Mishra [54]. Mismeasurement [1872]. Missing [929, 18, 1614, 1953, 1464, 1036, 1488, 885, 1383, 1957, 970, 1667, 1891, 1344, 1669, 1776, 672, 990, 1892, 1732, 1943, 1552, 2031, 1439, 1615, 1625, 1407]. Missing-Data [929]. Missingness [1943, 1869]. Misspecification [770, 279, 535]. Misspecified [869]. Mitochondrial [644]. Mittag [1198]. Mittelbach [1444]. Mixed [854, 1200, 1704, 441, 1604, 1462, 2000, 2020, 1898, 1900, 606, 692, 1772, 1206, 1285, 1464, 176, 1224, 1281, 866, 561, 694, 1713, 1963, 1909, 1712, 1556, 1961, 116, 2027, 1831, 605, 1996, 1532, 2065, 1223, 241, 1092, 1419, 991, 1322, 1873, 1280, 1316, 1677, 2049, 2050, 1870, 1871, 436, 1325, 1309, 486, 1104, 1550, 1317]. Mixed-Effect [694]. Mixed-Effects [1831, 1322, 436, 1325, 1550]. Mixed-Form [1898]. Mixing [1144, 367]. Mixture [395, 511, 404, 1571, 989, 1226, 620, 173, 460, 489, 1868, 1267, 358, 2030, 1239, 1939, 338, 122, 35, 1474, 809, 674, 1152, 705]. Mixtures [458, 1707, 1422, 537, 1208, 869, 1596, 294, 715, 1924]. MLE [1331]. MLEs [1521]. Model [344, 1504, 59, 395, 802, 2086, 1320, 1617, 238, 156, 212, 471, 854, 1200, 643, 2097, 1230, 818, 1157, 704, 1113, 819, 1258, 339, 1704, 441, 1524, 183, 41, 890, 1517, 1604, 1151, 986, 1882, 1333, 2064, 172, 335, 1760, 1663, 2003, 2004, 568, 2000, 3, 63, 1484, 1483, 2020, 17, 1177, 646, 692, 1278, 626, 623, 1358, 697, 1036, 1488, 317, 1099, 176, 1386, 389, 1825, 1936, 2073, 1281, 561, 170, 583, 969, 348, 694, 1818, 751, 1167, 1506, 620, 2014, 577, 257, 518, 665, 525, 1944, 1759, 582, 931, 460, 1885, 229, 1839, 1780, 1779, 1149, 125, 1895, 457]. Model [1765, 297, 1209, 701, 296, 2058, 1318, 1764, 1374, 1525, 971, 412, 1831, 1239, 355, 528, 83, 1607, 622, 540, 122, 1996, 1719, 461, 1044, 2065, 353, 1363, 445, 2006, 356, 35, 759, 1969, 1865, 241, 1925, 1680, 1480, 2061, 2082, 809, 1956, 1092, 1435, 1927, 991, 1322, 1512, 1893, 1873, 1280, 1316, 696, 712, 1826, 1059, 705, 1827, 1941, 1054, 1326, 2095, 1090, 1516, 4, 1926, 1481, 1615, 1104, 1768, 1546, 1268, 535, 1091, 1110, 213, 707, 915, 110, 1317, 157, 1002]. Model-Based [818, 704]. Model-Building [41]. Model-Free [931, 1002]. Modèle [51, 1844, 193]. Modèles [216]. Modeling [2024, 870, 1504, 1572, 6, 901, 1978, 1918, 1522, 1726, 746, 765, 2028, 864, 298, 1991, 1468, 1247, 1582, 1944, 235, 658, 1262, 801, 1939, 1951, 1810, 2023, 2001, 2047, 1474, 1774, 1561, 427, 2, 1903, 1559, 1369, 575, 271, 635, 2102]. Modelisation [940, 575]. Modelling [342, 268, 2109, 347, 181, 278, 73, 975, 192, 521, 453, 542, 1028, 1725, 610, 1132, 1575, 1735, 589, 1701, 1970, 399, 485, 1689, 891, 837, 1007]. Models [612, 872, 1875, 1423, 112, 979, 470, 472, 739, 1192, 1399, 1350, 1451, 1403, 1581, 1859, 1696, 2043, 403, 1578, 609, 169, 1003, 404, 1802, 762, 1571, 693, 1684, 1467, 2071, 1123, 702, 1614, 1866, 1333, 678, 1462, 1749, 1381, 283, 862, 983, 1900, 537, 943, 1035, 752, 1583, 1867, 606, 1168, 1953, 833, 1597, 1626, 2028, 1985, 1475, 362, 1772, 1778, 1234, 216, 1206, 1285, 86, 402, 676, 514, 1464, 659, 936, 1147, 1624, 653, 1224, 522, 1227, 14, 709, 293, 866, 321, 1656, 42, 1172, 1984, 444, 992, 1706, 1963, 1167, 1957, 173, 498, 721]. Models [1465, 1863, 1129, 1600, 1553, 949, 68, 1909, 111, 805, 1779, 489, 1712, 874, 1868, 1556, 1961, 116, 1267, 499, 2092, 1360, 24, 812, 1962, 619, 455, 2030, 2027, 219, 1682, 1729, 415, 168, 1331, 605, 338, 295, 1955, 1229, 2025, 62, 1532, 2065, 1460, 1045, 539, 569, 1046, 1357, 1416, 1409, 1994, 1223, 1788, 1060, 1033, 1602, 1332, 1945, 290, 1171, 285, 438, 1219, 934, 1529, 1864, 1501, 1419, 1946, 1528, 1907, 1152, 1677, 2049, 2050, 1870, 1871, 115, 357, 456, 1841, 1325, 1368, 601, 1309, 1310, 2017, 860, 1516, 1159, 663, 815, 1932, 2018, 1594, 1550, 918, 1288]. Models [984, 1678, 1935, 388, 1110, 1162, 1811, 1039, 1194, 1636, 1639, 2107, 594, 1128, 318, 252, 1001, 199, 1584, 191, 89, 45, 39, 1129]. Modern [956, 1849, 633, 2027, 684, 686, 1446, 1638, 1857, 469, 151, 54]. Modes [384]. Modification [2008, 1554, 1643]. Modifications [1721]. Modified [718, 1518, 1435, 1419, 82]. Modifying [31, 32, 629]. Molecular [1349, 1275, 1032]. Molenberghs [1909]. Mollison [1350]. Moment [851, 1, 277]. Moments [1980, 1374, 772]. Monitoring [1914, 2116, 987, 1420, 64, 2026, 857, 1959, 1178, 1379, 572, 1666, 523, 1401, 1297]. Monotherapies [1228]. Monotone [1147, 175, 1769, 1196]. Monotonic [696]. Monte [546, 1631, 1971, 283, 1511, 1825, 675, 866, 1706, 1957, 1699, 1890, 1928, 1861, 1812, 1117]. Moolgavkar [1209]. Mooney [259]. Moore [149]. Moose [20]. Morgan [1015]. Morgenthaler [784]. Morphometric [1691, 593]. Mortality [667, 1115, 2067, 610, 429, 1810, 496, 1216]. Most [633]. Mosteller [471, 728, 729, 893, 726]. Mother [976]. Motility [1949, 895]. Motion [1705, 1949]. Motulsky [1629]. Mouth [348]. Movement [1810, 300]. Mover [2086]. Moyenne [1116]. MR0931633 [271]. MR0963914 [213]. MR1029611 [157]. MR1132543 [775]. MSI [1746, 1975, 1541]. MSI-2000 [1746, 1975, 1541]. MTD [757]. Mukherjee [1400]. Multi [171, 1645, 857]. Multi-Armed [857]. Multi-Environment [1645]. Multi-Type [171]. Multicenter [1882, 1177, 822, 1938]. Multichannel [1553, 1094]. Multidimensional [120, 1763, 961].

Multienvironment [1122]. Multifactor [655]. Multilevel [969]. Multinomial [281, 518, 821, 1561, 1935]. Multiple [267, 555, 1053, 1039, 1428, 1573, 1664, 763, 699, 218, 1315, 974, 1041, 11, 863, 131, 2096, 113, 28, 1872, 166, 477, 1721, 1816, 1058, 65, 1616, 661, 1371, 744, 286, 886, 1895, 1717, 1030, 710, 820, 1118, 1000, 1065, 1231, 83, 246, 220, 539, 569, 2006, 519, 810, 1454, 2029, 1323, 1260, 603, 2074, 1361, 390, 1439, 1625, 1603, 270, 556, 1372, 1237, 1854, 1632, 1064]. Multiple-Objective [1816]. Multiple-Recapture [863, 28]. Multiple-Sacrifice [1616]. Multiple-Stage [1454]. Multiple-Step [699]. Multiplicative [739, 1122, 1227, 1149, 1564, 1318, 1532, 1925, 456, 1926, 1162, 1811]. Multiplicative-Epistatic [1926]. Multiplicity [2028, 719]. Multiply [306]. Multiply-Matched [306]. Multipoint [1429, 1805, 1807, 1808]. Multiresource [1014]. Multiresponse [568, 583]. Multisite [854, 441]. Multispecies [882]. Multistage [700, 1960, 1209, 233, 135]. Multistratum [563]. Multitype [2055]. Multivariate [1540, 1541, 1791, 1745, 1746, 1975, 819, 1408, 1882, 693, 1834, 2052, 2000, 752, 1050, 2069, 423, 1395, 522, 913, 1473, 321, 1173, 1506, 1944, 724, 196, 2091, 933, 1267, 296, 2059, 812, 2027, 1389, 973, 1738, 516, 398, 334, 1965, 1394, 519, 448, 1680, 1774, 1822, 564, 1161, 1316, 1838, 1546, 1937, 1091, 1039, 1317, 57, 946, 374, 1682, 1788, 1184]. Multiway [251, 371]. Murase [895]. Murray [1250]. Museum [1143]. Must [189]. Mutation [1209, 21]. Mutational [1773]. Mutations [1209]. Myers [594]. Myth [1264].

N [326, 472, 793, 1017, 847, 902, 1077, 1194, 1197, 1074, 1300, 1304, 1446, 1448, 1634, 1740, 1741, 1916, 2040, 2043, 144, 41, 592, 589, 948, 1296, 2102, 2112, 781, 1242, 249, 1063, 1789, 207, 1006, 1682, 503, 1005, 1738, 54, 1851, 312]. Naes [374]. Nagaraja [1005]. Nakache [1491]. Name [1248, 887]. Namkoong [98]. Narain [324]. Narayana [510]. Narrative [1850]. National [904, 1498]. Natural [1542, 1498, 1143, 1143]. Nature [170, 2102]. Nearest [772]. Needs [181]. Neel [838]. Negative [1271, 279, 673, 133, 617, 1561]. Neglected [23, 677]. Neighbour [773, 1815, 403]. Neighbour-Designs [773]. Neighbouring [1115, 1294]. Nelder [191]. Nelson [635, 428, 635]. Nephropathy [992]. Nest [762, 132, 1951, 1649]. Nested [667, 1863, 1533, 232, 1033, 524, 2115]. Nesting [1028, 539]. Nestlings [453]. Neter [39]. Networks [1911, 1911, 894]. Neudecker [1979]. Neural [1911, 1687, 1911]. Neuroimaging [1112]. Neuron [1654]. Neurons [1918]. Neurophysiological [674]. Neurosciences [92]. Neutrality [1380]. Neuts [1494]. Nevill [1535]. Ng [196]. Nichols [502]. Nielsen [894]. Nielson [328, 1399]. Niewiadomska [1739]. Niewiadomska-Bugaj [1739]. Nightingale [795]. Nikulin [1681]. No [887, 1057, 775, 213, 270, 556, 777, 157, 271]. No-Name [887]. Non [210, 1164, 763, 1206, 829, 722, 927, 1033, 496, 1154, 704]. Non-Additivity [722]. Non-Cancer [496]. Non-Gaussian [927, 704]. Non-Ignorable [1164]. Non-Independent [829]. Non-Linear [1206]. Non-Markovian [763]. Non-Nested [1033]. Non-Null [210]. Non-Parametric [1154]. Non-Response [1164]. Nonadditive [459]. Noncentral [534]. Noncommunicating [1699]. Nonidentified [798]. Nonignorable [929, 1293, 1383, 1833]. Nonignorably [1552, 2031]. Nonlinear [609, 1272, 678, 1035, 606, 1597, 583, 1584, 118, 1262, 116, 1607, 295, 2065, 290, 436, 1325, 1368, 1550, 1450, 139, 1583, 373, 724, 372]. Nonmonotonic [2087]. Nonnegativity [24]. Nonnormal [437]. Nonparametric [1366, 1423, 1052, 265, 842, 899, 2038, 343, 1775, 879, 1392, 667, 359, 654, 2083, 1923, 1698, 1991, 366, 565, 1820, 346, 287, 582, 1502, 88, 1270, 1331, 226, 1514, 1986, 237, 706, 1946, 425, 668, 811, 1341, 976, 2072, 1720, 1515, 1284, 1496, 1003]. Nonproportional [238, 1771]. Nonrandomly [18]. Nonresponse [1293, 983, 799, 1833]. Nonsense [999]. Nonzero [1903]. Noreen [87]. Normal [511, 1855, 2064, 1484, 303, 1485, 989, 1226, 1984, 604, 1265, 1944, 196, 1434, 827, 358, 715, 1776, 30, 532, 1092, 1966, 1574, 1967]. Normally [118]. Northeast [2112]. Note [56, 739, 1352, 1573, 718, 1285, 542, 829, 1281, 1733, 1058, 10, 411, 1292, 766, 1620, 996, 990, 1241, 1896, 708, 456, 1377]. Notz [2035]. Nouveaux [898]. Nuclear [809]. Nuclei [1123]. Nucleoli [1402, 483]. Null [210]. Number [1906, 880, 1551, 1711, 989, 1971, 531, 801, 16, 2048, 1321, 1653]. Numbers [918]. Numerical [177, 260, 356, 640, 1971]. Nurseries [2015]. Nutrient [952, 1987]. Nutrition [1852, 1614, 1841, 510]. Nutritive [1268]. Nychka [1917].

O [1013, 328, 1399, 139, 193, 47, 894, 1845, 38, 1062]. Obituary [737, 37, 738, 1086, 795]. Object [2002]. Objective [1816]. O’Brien [767, 820]. O’Brien-Fleming [820]. O’Brien’s [1241]. Observable [408]. Observation [2020, 465]. Observational [2100, 1761, 1187, 861, 1952]. Observations [1052, 1452, 392, 359, 1099, 1413, 293, 1105, 1558, 1776, 801, 868, 564, 128, 1966, 1967, 1653]. Observed [455, 1057, 1277]. Observer [1452, 392, 1993]. Observers [1864]. Obtained [1199, 1170, 1037]. Obtaining [1521]. Occasional [1174]. Occupancy [1237]. Occupational [768, 1564]. O’Connell [470]. October [1143]. Odds [1958, 183, 11, 243, 1656, 2075, 1412, 587, 1922, 461, 276, 810, 12, 652, 357, 541, 332]. Off [1425]. Offspring [1906, 1145, 1378]. O’Hagan [1443]. Oligonucleotides [476]. Oliver [787]. OLS [1241]. Omitted [1523]. Omitting [185]. Oncology [1992]. One [654, 1621, 1229, 247, 62, 1482, 759, 1883, 934, 128, 1876, 1403]. One-Hit [759]. One-Sided [1229, 247, 62, 1883, 934, 1403]. One-Way [654, 1621, 1482, 1876]. Only [1384]. Onset [2028, 2066, 1477, 1767, 2005]. Ooijen [1349]. Open [1579, 1929, 1756, 1362]. Open-Ended [1579]. Operating [982, 718, 1549, 1655, 1284]. Operational [1537]. Operations [104, 148]. Optimal [689, 1762, 919, 2022, 529, 166, 583, 1329, 1816, 538, 2098, 124, 1665, 1924, 1375, 1781, 1197]. Optimality [69]. Optimization [700, 1093, 1054]. Optimum [643, 1576, 301, 1935, 1063]. Option [813]. Option-3 [813]. Orav [148]. Ord [723, 1243]. Order [1387, 431, 509, 736, 954, 360, 553, 661, 1777, 1651, 1424, 1005, 1611, 595, 140, 503]. Order-Directed [1777]. Order-Restricted [1387, 1651]. Ordered [1829, 1392, 678, 650, 2012, 1579, 669, 457, 1777, 1840, 923, 2013, 712, 1415]. Ordered-Heterogeneity [923]. Ordering [1933, 804]. Ordinal [802, 183, 1207, 1346, 969, 970, 817, 1131, 1412, 1292, 1046, 1316, 696, 1943, 611, 1903, 1317]. Ordinary [117]. Ordination [1394]. Organization [211, 432, 598, 794, 1641, 1919, 129, 211, 432, 598, 794, 1020, 1085, 1141, 1641, 1919]. Oriented [1348, 1909]. Original [1346]. Orr [1354]. Ostermann [1303]. Ostrowski [1246]. Other [1399, 1673, 2053, 2054, 1950, 234, 1684]. Ott [264, 1799, 327, 550, 1397]. Outcome [1164, 1821, 1872, 1343, 1439, 1369]. Outcomes [744, 799, 1389, 1939, 2029, 440, 1666, 1361, 1509, 1615, 517]. Outcrossing [1925, 1926]. Outlier [813, 1947]. Outliers [1323, 925, 1066]. Outlines [963]. Output [65]. Outs [1421]. over- [1220]. over-/under-Dispersion [1220]. Over-Dispersed [1504]. Overdispersed [1451, 1484, 1626, 1596, 518, 1171, 1280]. Overdispersion [626, 610, 1878, 115, 1415]. Overlap [1896, 1435]. Overviews [439]. Öztürk [635, 428].

P [57, 736, 791, 842, 845, 955, 959, 953, 958, 1134, 1190, 1080, 1137, 1196, 1252, 1253, 1303, 1400, 1450, 1306, 1539, 1588, 1640, 1495, 1792, 1795, 1853, 2042, 2110, 2040, 2041, 2116, 940, 775, 1003, 1681, 1683, 41, 1536, 1244, 149, 591, 151, 142, 1186, 324, 1001, 1012, 840, 1010, 322, 468, 681, 1491, 284, 1008, 254, 100, 148, 467, 635, 896, 1129, 1444, 1582, 1784, 1973, 316, 949, 191, 1489, 1845, 1789, 1067, 88, 1070, 1184, 1411, 682, 428, 947, 892, 203, 1493, 1847, 2054, 1000]. p-Value [1000]. Packages [790]. Pair [80, 1700]. Pair-Matched [80]. Paired [972, 1217, 1131, 1382]. Pairs [1804, 480, 1612, 1807, 2005, 275, 1805, 1808, 274]. Pairwise [1315, 857, 1156]. Palaeobiology [961]. Panchapakesan [1916]. Panel [341, 2020, 23, 1839, 677]. Papastavridis [1252]. Paper [107, 100]. Papers [214, 272, 330, 380, 434, 474, 512, 557, 599, 641, 690, 740, 796, 855, 907, 964, 1021, 1181, 1252]. Parallel [1403, 1313, 120, 1120, 1229]. Parameter [957, 488, 1048, 1224, 823, 985, 279, 1383, 932, 24, 133, 490, 1219, 1454, 712, 1874, 663, 45]. Parameters [267, 555, 1335, 1041, 1773, 113, 1471, 1818, 406, 1049, 118, 457, 2059, 81, 34, 916, 334, 1263, 1148, 2049, 2050, 1870, 1871, 2017, 1550, 270, 556]. Parametric [1423, 1866, 912, 131, 1957, 126, 971, 520, 437, 1044, 809, 1501, 1152, 1154, 1653]. Parasite [760]. Parent [1145, 2005]. Parent-Offspring [1145]. Parental [147]. Parsimony [314, 1457]. Part [1517, 1395, 1394]. Partial [1053, 1040, 1880, 1211, 1670]. Partially [1507, 126, 1807, 1277, 1805, 1808, 815]. Particle [963]. Partitioning [829]. Partly [2025]. Partner [180, 756]. Parts [206]. Passive [1146]. Past [314]. Patch [482, 1553]. Patchy [628]. Patel [1855]. Patient [1342, 2079, 412, 1672]. Patients [443, 1384, 524]. Patil [1244, 1184]. Pattern [1190, 654, 841, 1911, 1502, 1267, 2030, 1478, 408, 1911, 1687]. Pattern-Mixture [1267, 2030]. Patterns [1252, 347, 2055, 1217, 130, 1943]. Patton [1071]. Pavel [841]. Pay [480]. Payne [1302, 1083]. PCR [1430]. PCR-Generated [1430]. Peace [1078]. Peak [1282]. Peanuts [1728]. Pearson [464]. Pedigree [1699, 1051, 71, 808]. Pedigrees [866]. Penalized [1660, 2014, 489, 1599]. Penalty [1003]. Penetrance [1867]. Penguins [1649]. Pentikäinen [962]. Percent [535]. Percentile [566]. Perelson [206]. Performance [824, 820, 81]. Performances [1879]. Performing [477]. Perfusion [450]. Period [343, 1569]. Periodic [1832]. Périodique [193]. Periods [1048, 408]. Permitting [872]. Permutation [1364, 1997, 567, 2070, 1266]. Permutation-Type [1997]. Permutational [1659, 978, 1361]. Personalities [1634]. Perspective [1555, 259, 1798, 317]. Perspectives [1077]. Pesonen [962]. Pest [1752, 1097]. Petersen [2048]. Pfaffenberger [1695]. Ph [637]. Pharmaceutical [2042, 919, 1186]. Pharmacochemistry [840]. Pharmacodynamic [883]. Pharmacokinetic [1024, 883]. Pharmacokinetic/Pharmacodynamic [883]. Pharmacokinetics [188]. Phase [2068, 451, 9, 20, 853, 1218, 1114, 1425, 1531, 84, 1165, 530, 1668, 2029, 440, 757, 858, 1666]. Phenomenes [940]. Phenotype [447]. Phenotypic [366, 761]. Philosophical [1537]. Phylogenetic [1861, 662]. Phylogenies [1457]. Physics [596]. Physiology [256]. Piecewise [566, 1819, 68, 973]. Piegorsch [1917]. Pigs [278]. Pilot [648, 1375, 1781]. Plackett [464]. Plan [20]. Plane [754, 613, 1653, 1639]. Planifiées [760]. Planned [1197, 1263, 760]. Planning [64]. Plans [135]. Plant [1349, 2040, 1755, 1470, 1048, 1510, 1706, 618, 1097, 1811, 204]. Plant-Capture [618]. Plants [1314, 1928]. Platform [1803, 1983]. Pleiotropic [1652]. Pleszczynska [685]. Plot [706, 234]. Plots [773, 335, 120, 1227, 357]. Plotting [1175]. Plus [1116, 1446]. Point [2110, 1654, 2055, 1299, 1029, 125, 1715, 587, 332, 905]. Points [1043, 235]. Poisson [467, 341, 850, 1451, 1626, 1237, 583, 1991, 1596, 713, 1891, 670, 277, 1277, 1892, 1171, 490, 1309, 1121, 1068]. Policy [1498, 471, 1790, 2039]. Pollock [1072, 132, 502]. Polution [1692]. Polygenic [1360]. Polyhazard [2092]. Polymorphic [1166]. Polymorphisms [772]. Polynomial [2109, 1288]. Polynomials [566]. Polysampling [784]. Polytomous [1627, 1671, 749]. Pons [193]. Pooled [2117, 985, 1988]. Pooling [1960, 350, 708]. Pooni [2108]. Population [1038, 452, 1749, 591, 97, 204, 863, 864, 375, 2067, 828, 1995, 1471, 985, 199, 1929, 1348, 618, 805, 1029, 1508, 921, 1116, 1648, 1337, 2084, 20, 1887, 1487, 1111, 806, 1148, 1096, 1493, 1932, 1966, 2018, 883, 1967, 1039, 875, 110, 1856, 200]. Population-Based [1038, 2067]. Populations [493, 2041, 1595, 1993, 862, 591, 491, 360, 51, 1271, 579, 1265, 837, 195, 1434, 111, 232, 1055, 233, 1373, 301, 1756, 2080, 1362, 549, 128, 578, 407, 1606, 1417, 1140, 1973, 682, 947]. Populations-genetik [549]. Porpoise [1809]. Portmanteau [1234]. Posed [450]. Positive [673, 1965, 1622]. Possible [1760, 440]. Possibly [1966, 1967]. Post [1093]. Post-Treatment [1093]. Postcalving [1754]. Posterior [1743, 1822]. Posttest [308, 1529]. Potato [147]. Potency [645, 646, 1173, 240, 716, 385, 1090]. Potential [1881, 567, 1094]. Potentially [1464]. pour [216]. Pouvant [51]. Powder [963]. Powder-Particle [963]. Powell [1411]. Power [511, 748, 1367, 669, 644, 1312, 1836, 358, 1674, 1899, 1212, 1031, 310, 438, 753, 1547]. Powerful [715, 78]. Prabhakaran [682]. Practical [843, 960, 1422, 302, 137, 824, 9, 784, 384, 858, 473, 962, 2038, 2106, 426, 1846, 501]. Practice [901, 1740, 1977, 149, 245, 946, 1909, 1279, 685, 1077]. Praktischer [1848]. Precision [236, 1295, 217, 1148]. Predict [2097]. Predicting [1238, 27]. Prediction [347, 1238, 1663, 216, 676, 561, 627, 444, 2025, 774, 216]. Predictions [1513]. Predictive [59, 930, 1968, 244, 1506, 750, 1516]. Preference [1118, 399]. Preferred [699]. Pregnancy [296, 1480]. Preparation [896, 1444]. Preparations [1173]. Presence [6, 156, 268, 1755, 1238, 494, 1414, 1459, 1388, 1410, 126, 673, 820, 538, 1203, 531, 801, 1220, 925, 2072, 1734, 19, 2, 4, 1284, 1702, 157, 271]. Presence/Absence [1388]. Presentation [1346]. Presenting [1276, 1515]. Press [86]. Pressure [347, 227, 1964]. Pressures [347]. Pretest [308, 1529]. Pretest-Posttest [308]. Pretest/Posttest [1529]. Prevalence [493, 1960, 2068, 2067, 529, 1318, 1609, 1341]. Prevalent [343, 1060, 601]. Prevention [171, 345, 1097, 906]. Prevosti [1205]. Price [631]. Priestley [1016]. Primary [1724, 1570, 354]. Primer [1490]. Principal [1338, 1881, 547, 1434]. Principles [508, 86, 137, 263, 97, 46]. Prior [1813, 240, 1043, 81]. Priority [630]. Privacy [1695]. Probabilistic [1696, 894]. Probabilités [637, 315]. Probabilities [1934, 452, 862, 28, 447, 1330, 576, 111, 1379, 1842, 2047, 712]. Probability [59, 683, 1017, 853, 1136, 1192, 1252, 1400, 1980, 151, 262, 244, 1181, 1271, 1494, 949, 827, 74, 530, 369, 1756, 1570, 1851, 1487, 652, 1111, 1075, 1722, 860, 807, 1515, 1494, 897, 1253, 1353, 430, 636, 733, 1793, 264, 1494, 1538, 1739]. Probability-of-Exposure [652]. Proband [387]. Probands [1103]. Probit [1408, 1462, 1931, 1624]. Probit-Linear [1462]. Problem [511, 1402, 1775, 2063, 483, 768, 411, 450, 358, 2060, 225, 290, 85, 1298]. Problems [1013, 1728, 1237, 1699, 1093, 1494, 1075, 950]. Procedure [699, 218, 919, 478, 1158, 1100, 1721, 1156, 767, 820, 1287, 490, 291, 578, 486]. Procedures [956, 1853, 1573, 654, 27, 1410, 938, 286, 1897, 527, 226, 1260, 2074, 775, 1197]. Proceedings [507, 98]. Process [341, 6, 268, 1123, 771, 1186, 2026, 131, 714, 65, 1525, 2, 575, 271, 1190, 427]. Processes [92, 905, 850, 1484, 2020, 629, 747, 1345, 1596, 1348, 713, 1081, 1139, 1544, 1188, 93, 1070]. Processing [896, 732]. Processus [771, 575]. Production [1906]. Products [280, 447]. Professional [1498].

Professionals [255]. Professor [686, 1400, 1541, 1746, 1975, 161]. Progenies [1510]. Progeny [109]. Progeny-Tested [109]. Prognostic [1214, 1270]. Program [1472]. Progression [975]. Progressive [747, 1884, 2014]. Projection [1416, 182]. Projections [211]. Promotion [825]. Propelled [1949]. Propensity [770, 1279]. Proper [580]. Properties [643, 1834, 11, 467, 2072, 1047]. Proportion [172, 477, 867, 443]. Proportional [870, 1320, 1578, 183, 2071, 1663, 752, 697, 1147, 389, 1825, 1936, 170, 444, 798, 751, 68, 1780, 125, 528, 461, 1570, 1363, 1824, 696, 1826, 1827]. Proportionality [183]. Proportions [1387, 1571, 2076, 1835, 1390, 1888, 673, 1220]. Prospective [2009, 1674, 1899]. Protect [1695]. Protecting [18]. Protection [699, 1748]. Prove [1912]. Proximity [468]. Psychosocial [788]. Public [471, 906, 1790]. Publication [982]. Publish [100]. Published [214, 272, 330, 380, 434, 474, 512, 557, 599, 641, 690, 740, 796, 855, 907, 964, 1021, 1322]. Pyzdek [105].

Q [725]. QSAR [57]. QTL [1652, 1031]. Quadratic [1413]. Quadrats [269, 1269, 75, 242]. Quadrature [188]. Qualitative [225, 655]. Qualitatives [1491]. Quality [105, 793, 959, 1198, 912, 2007, 2062, 1948]. Quality-Adjusted [912, 2062, 1948]. Quantal [1015, 818, 703, 1125, 702, 537, 121, 16, 122, 484, 1930, 674]. Quantal-Response [702]. Quantifying [236, 1034]. Quantile [1170, 30]. Quantiles [1934, 1567, 1952, 1610]. Quantitatifs [90]. Quantitative [1496, 2108, 459, 1773, 1207, 465, 1388, 1521, 1312, 1519, 1651, 449, 1263, 1011, 1812, 98, 71, 1417, 2105]. Quantities [1850]. Quarantine [1097]. Quarterly [598, 794, 1020, 1141, 211, 432, 1641, 1919]. Quasi [612, 2110, 1167, 937, 1837, 1415]. Quasi-Least [1837]. Quasi-Likelihood [1167, 937, 1415]. Quasi-Random [2110]. Quasi-Symmetric [612]. Quasilikelihood [1310]. Questionnaire [1199, 1037]. Questions [1639]. Quinn [2105].

R [251, 102, 328, 378, 379, 597, 470, 507, 639, 790, 733, 638, 787, 844, 959, 961, 898, 848, 849, 958, 1134, 1077, 1194, 1249, 1302, 1399, 1445, 1303, 1300, 1401, 1542, 1586, 1636, 1639, 1791, 1795, 1854, 1799, 2039, 1979, 2040, 2043, 2117, 319, 940, 775, 594, 41, 318, 590, 1490, 261, 945, 836, 2105, 1244, 142, 327, 943, 262, 888, 368, 1071, 47, 1849, 42, 1008, 1584, 889, 140, 100, 257, 260, 264, 504, 464, 635, 1347, 1739, 1850, 137, 1189, 679, 1789, 554, 998, 1682, 50, 1910, 258, 548, 726, 200]. R [1184, 1075, 1185, 1011, 40, 1907, 428, 947, 136, 43, 259, 427]. R. [161, 164, 162]. Rade [1398]. Radhaiah [510]. Radiation [717, 524]. Radio [261, 1754]. Radio-Tracking [261]. Radioimmunoassay [385]. Radionuclides [94]. Raffaelli [1190]. Random [393, 1572, 212, 1581, 1796, 2110, 1157, 1882, 1467, 1462, 351, 3, 63, 1597, 720, 1531, 1206, 1285, 1036, 1488, 1990, 176, 46, 1624, 653, 969, 1659, 1299, 1166, 498, 970, 525, 1944, 1759, 607, 1131, 1149, 1477, 812, 801, 803, 295, 1357, 1409, 416, 613, 1824, 1480, 544, 1610, 2049, 2050, 1876, 1368, 1516, 1869, 234, 1110, 1039, 213, 1796, 1971, 1796, 2110]. Random-Effects [1157, 653, 969]. Randomization [1998, 1177, 1830, 1426, 1233, 624, 546, 1631]. Randomized [1766, 930, 1964, 287, 886, 1502, 333, 603, 1432]. Randomly [1259, 126]. Randomness [269, 878, 1879, 75]. Range [1156, 1204]. Rank [854, 1978, 982, 441, 1470, 930, 2087, 1358, 1659, 442, 1340, 1897, 226, 1513, 448, 345, 1841, 711, 535]. Rank-Based [854, 441, 2087, 1897]. Ranked [1797, 1335, 1465, 74, 1568]. Ranking [1880]. Rankings [1217]. Ranks [880]. Rao [510, 1194, 43, 1016, 1195, 1852, 1244, 142, 949, 1434, 1184, 136]. Rapid [26]. Rasch [1180, 549, 500]. Rasmussen [731]. Rat [2023]. Ratchet [576]. Rate [1828, 1727, 986, 1048, 298, 1479, 1616, 2066, 1222, 286, 803, 1337, 572, 1148, 976]. Rater [612]. Rates [2116, 1428, 765, 747, 1959, 1468, 584, 1124, 1108, 1962, 1996, 1622, 616, 1731, 524, 1309]. Rating [705]. Ratings [661]. Ratio [1958, 511, 703, 1675, 1604, 1749, 1282, 1278, 11, 989, 1656, 1271, 2075, 1733, 761, 2066, 1437, 1724, 997, 587, 358, 715, 174, 1831, 34, 276, 810, 12, 581, 652, 438, 1027, 668, 1722, 357, 407, 1096, 19, 535, 332]. Ratio-Defined [761]. Ratios [243, 1701, 1412, 1922, 1982, 1550, 541]. Ratkowski [889]. Rawlings [38]. Rayner [323]. Re [77]. Re-Immigration [77]. Reactions [651, 538]. Read [1736, 1785, 2033, 1855, 312]. Reader [906]. Readers [1716]. Readings [1498, 1498]. Real [398]. Reassessment [9, 1334]. Recapture [872, 112, 187, 1522, 1802, 763, 236, 1151, 1647, 452, 1989, 491, 764, 1900, 863, 864, 28, 985, 1929, 406, 579, 1049, 502, 1222, 805, 997, 1331, 1648, 1756, 1337, 1460, 2084, 806, 1362, 1927, 2017, 77, 19, 352, 2018]. Recaptures [862, 1330, 111]. Receiver [1549, 1655, 1284]. Recognition [1911, 841, 841, 1911]. Recommended [853, 530]. Reconstructing [314]. Recordings [482, 1553]. Recovery [1595, 1522, 1333, 1647, 1703, 453, 1222, 145, 616, 1117]. Recovery/Recapture [1647]. Recreational [1999]. Recruitment [1337]. Rectangle [1703]. Recurrent [171, 1420, 1324, 1602, 1605]. Red [463, 294]. Reduced [1867, 1952, 448, 1841]. Reduced-Rank [448, 1841]. Reducing [1295, 217, 309, 80]. Reduction [1778, 863, 1235, 232, 813]. Rees [1079, 150]. Reexamination [1411]. Referee [1527]. Reference [1170, 529, 896, 1444]. Reflect [1900]. Regards [898]. Regional [1054]. Regions [2094, 1274]. Register [500]. Registration [867]. Registries [2067]. Registry [1146]. Regression [870, 1423, 112, 1617, 238, 155, 329, 1403, 647, 929, 1164, 1578, 488, 819, 1258, 1003, 1524, 183, 1364, 2071, 139, 1834, 335, 494, 1760, 1295, 1035, 73, 1168, 386, 2002, 217, 487, 2026, 1661, 676, 514, 1758, 1464, 1931, 413, 1386, 1459, 1624, 14, 1281, 913, 1656, 1346, 969, 1329, 289, 444, 2100, 1577, 1468, 5, 1706, 751, 1383, 1957, 761, 644, 1058, 1347, 1600, 1820, 118, 1944, 1618, 1120, 1885, 1701, 1779, 515, 1761, 1891, 1283, 1669, 499, 1101, 1564, 2059, 1319, 282, 1525, 1783, 1239, 1947, 168, 1910, 1607, 1102, 1145, 995, 1955, 38, 1229, 1655, 2025]. Regression [516, 2065, 1954, 1416, 1938, 1363, 288, 1033, 448, 1892, 1680, 652, 774, 285, 1529, 1454, 1501, 1605, 291, 2019, 1323, 1907, 1677, 2049, 2050, 39, 387, 357, 1236, 436, 1309, 1466, 390, 1903, 1568, 1268, 2085, 1407, 388, 1110, 898, 193, 1797, 899, 373, 421, 835, 422, 40, 734, 631, 420]. Regression-to-the-Mean [1760]. Regressions [1170]. Regressive [1758]. Regularly [1832]. Reid [589]. Reinsel [1745]. Rejection [1725]. Rejmánek [259]. Rejoinder [1535, 1808]. Relapsing [802]. Relapsing-Remitting [802]. Related [236, 1338, 1867, 1881, 1507, 44, 1492, 1133, 196]. Relating [2027, 1279, 860]. Relation [59, 1350]. Relations [139]. Relationship [1199, 1048, 975, 1119, 1964, 346, 1037, 723]. Relationships [1071, 2075, 228, 222]. Relative [1230, 667, 31, 1551, 657, 626, 1042, 184, 1173, 240, 481, 1700, 1674, 1899, 1807, 385, 950, 1148, 1805, 1808, 306, 32]. Release [1083]. Reliability [848, 1148, 45, 1032, 1696]. Remarks [177]. Remission [1384]. Remitting [802]. REML [693, 692, 1224, 1122, 644, 619]. Removal [1749, 873, 2088]. Renewable [689]. Renewal [131]. Rennolls [202]. Renovated [1175]. Renshaw [837]. Repeated [507, 1164, 1704, 359, 1040, 1291, 983, 1583, 748, 988, 1931, 542, 1395, 1880, 1376, 695, 604, 607, 223, 239, 2091, 116, 799, 1105, 1365, 499, 1623, 437, 672, 1840, 1416, 2056, 416, 396, 1343, 424, 333, 696, 563, 436, 1516, 1869]. Repeated-Measurements [333]. Repeated-Measures [396]. Repeatedly [2080]. Replicate [1452, 392]. Replicates [1310]. Replication [1123, 777]. replied [2101]. reply [775]. Report [558, 741, 908, 381]. Reporting [995]. Reporting-Delay [995]. Reports [1880]. Reposer [51]. Represent [754]. Representation [26]. Representations [468]. Reproducibility [1620, 1409]. Reproductive [349, 81, 1650]. Requirements [1612]. Resampling [1410, 1000]. Resampling-Based [1000]. Research [210, 263, 639, 904, 849, 1077, 1448, 201, 780, 8, 138, 1008, 1245, 207, 38, 889, 264, 104, 1495, 634, 1007, 501]. Residual [850, 650, 713, 1783, 357]. Residuals [1947, 1827]. Resight [1522]. Resightings [862, 111]. Resistance [1470]. Resistant [2002]. Resnick [1792]. Resolution [1566, 1312]. Resource [787, 951]. Resources [1498, 689, 204]. Respond [1792]. Responding [657]. Response [1098, 433, 1015, 957, 1307, 1643, 818, 1164, 703, 1125, 404, 41, 702, 2087, 1425, 537, 400, 514, 120, 1208, 1849, 698, 121, 1554, 175, 13, 228, 1878, 1412, 716, 219, 1239, 16, 122, 62, 220, 484, 1930, 1024, 336, 496, 1599, 1316, 1677, 1870, 1871, 696, 1552, 2031, 1159, 523, 1317]. Responses [1726, 349, 987, 1420, 1324, 910, 1464, 1173, 348, 798, 604, 1365, 1886, 1717, 1954, 334, 539, 1650, 1316, 1943, 611, 427, 1317]. Responsible [1527]. Responsive [300]. Restoration [482, 1566]. Restricted [1387, 1468, 140, 1562, 1831, 1651, 1223, 1987]. Restrictions [1321]. Results [2100, 439, 410, 874, 1761, 1621, 1843, 1655]. Retaining [827]. Retherford [943]. Retrospective [168]. retrospectively [446]. Reuse [1476]. Reveal [1227]. Reverse [1722]. Reversible [1812]. Review [92, 326, 632, 895, 1013, 251, 208, 1248, 1073, 102, 103, 152, 209, 55, 210, 153, 154, 104, 57, 101, 105, 329, 377, 378, 431, 266, 263, 430, 379, 506, 552, 508, 509, 596, 597, 470, 472, 473, 471, 510, 507, 639, 788, 687, 683, 688, 731, 793, 786, 684, 686, 689, 735, 789, 790, 640, 636, 732, 733, 734, 638, 685, 787, 637, 736, 791, 792, 842, 844, 899, 1015, 843, 900, 1016, 955, 846, 956, 959, 957, 960, 1017, 1018, 901, 961, 1019, 904]. Review [905, 962, 898, 906, 847, 848, 849, 897, 902, 903, 953, 958, 963, 1020, 1134, 1190, 1191, 1076, 1077, 1078, 1135, 1140, 1194, 1195, 1249, 1250, 1079, 1136, 1080, 1192, 1083, 1137, 1193, 1196, 1251, 1197, 1252, 1074, 1081, 1082, 1084, 1138, 1139, 1198, 1253, 1085, 1141, 1302, 1445, 1449, 1349, 1303, 1301, 1350, 1353, 1400, 1354, 1300, 1304, 1446, 1305, 1450, 1401, 1447, 1448, 1351, 1539, 1542, 1586, 1587, 1540, 1588, 1637, 1589, 1497, 1543, 1544, 1634, 1636, 1541, 1638, 1496, 1639, 1640, 1495]. Review [1590, 1635, 1740, 1744, 1791, 1794, 1741, 1852, 1690, 1745, 1855, 1688, 1693, 1792, 1795, 1689, 1691, 1746, 1798, 1854, 1857, 1859, 1695, 1793, 1856, 1696, 1853, 1692, 1742, 1790, 1799, 1743, 1915, 1916, 1978, 2107, 1917, 2037, 1913, 1974, 2042, 2039, 2108, 1918, 2109, 1975, 1979, 1977, 2038, 2106, 2110, 1914, 2040, 2041, 2043, 2111, 1980, 194, 144, 319, 940, 887, 1066, 1736, 201, 551, 1003, 1681, 426, 780, 1072, 1297, 594, 1128, 951, 1683, 41, 318, 590, 834, 1629, 1490, 87, 261, 890, 945, 836, 1536]. Review [1737, 2105, 202, 1244, 149, 371, 252, 1684, 139, 469, 151, 142, 97, 327, 592, 727, 783, 2035, 943, 1583, 147, 204, 325, 839, 262, 1009, 1186, 324, 85, 634, 944, 1246, 1298, 91, 1001, 150, 942, 1443, 193, 888, 1183, 368, 550, 838, 1012, 1397, 1786, 192, 1071, 1069, 547, 197, 86, 840, 1010, 1396, 1630, 376, 375, 138, 317, 725, 1188, 51, 46, 423, 322, 546, 1686, 206, 941, 468, 681, 1130, 1395, 1785, 952, 47, 48, 466, 1181, 465, 44, 321, 1182, 1491, 1849, 891, 42, 373, 1848, 199, 314, 1008, 1245, 1132, 1299, 254, 1584, 205]. Review [1492, 1002, 1004, 728, 631, 729, 723, 140, 313, 257, 260, 256, 323, 500, 504, 467, 464, 589, 633, 785, 680, 948, 837, 893, 1180, 1247, 1243, 1133, 1348, 1347, 1398, 1633, 1538, 1631, 1582, 1784, 1787, 1739, 1850, 1973, 2036, 2033, 2102, 2112, 195, 370, 95, 429, 146, 894, 316, 724, 781, 949, 249, 946, 137, 191, 1909, 1189, 143, 502, 320, 1063, 679, 1489, 1537, 1845, 1789, 1131, 52, 372, 53, 554, 420, 998, 374, 1067, 207, 1000, 1065, 88, 90, 1007, 1187, 1687, 1006, 145, 1682, 49, 50, 1910, 503, 258, 315, 253, 93, 141, 782, 1070]. Review [1005, 1738, 89, 1068, 1632, 38, 548, 369, 726, 730, 200, 164, 421, 54, 505, 1062, 1851, 588, 835, 1064, 593, 950, 1394, 1585, 1788, 424, 784, 422, 1184, 96, 1075, 999, 1014, 1185, 1011, 40, 94, 549, 682, 425, 45, 1907, 428, 1846, 2104, 98, 947, 39, 136, 1844, 892, 250, 43, 501, 203, 99, 259, 1493, 1847, 427, 1685, 312]. Reviews [211, 328, 432, 598, 794, 954, 845, 1254, 1399, 1498, 1641, 1796, 1976, 1919, 255, 198, 591, 841, 889, 100, 148, 264, 553, 635, 896, 1129, 1296, 1444, 1494, 1911, 1971, 1908, 1972, 2034, 2103, 1242, 196, 1912]. Revised [847]. Revisited [498]. Reyment [1542, 961]. Rhythms [1326]. Rice [1250]. Ridout [2114]. Riede [1019]. Riegelman [1008]. Right [1052, 6, 156, 268, 1923, 1410, 801, 2, 4, 1948, 157, 271]. Right-Censored [1052, 1410, 1948]. Right-Truncated [1923]. Rijckevorsel [906]. Riley [145]. Ring [1333, 1117]. Ring-Recovery [1333, 1117]. Ringed [453]. Rinne [1198]. Ripley [1911, 1446, 93]. Risk [342, 1828, 962, 1498, 127, 169, 1038, 1524, 299, 386, 243, 1533, 481, 52, 1657, 710, 1724, 1764, 1700, 1674, 449, 581, 1650, 2061, 1599, 541, 306, 2115]. Risks [1230, 1676, 1012, 651, 1276, 1101, 1842, 1899, 2085]. Ritov [1001]. RNA [1963]. Roads [1023]. Robbins [686]. Robert [1004]. Robertson [140]. Robinson [100]. Robson [836]. Robust [865, 1073, 980, 536, 139, 1324, 2012, 1471, 1679, 695, 644, 1049, 1215, 277, 527, 916, 1623, 1954, 1223, 2013, 1469, 573, 1323, 1404, 1853]. Robustness [880, 279, 1888, 186, 417, 241, 1893, 784]. ROC [1563, 1231, 1514, 1520, 1486, 811, 1702]. Roeder [1792]. Role [208, 384]. Root [1226]. Rosner [551, 83]. Ross [372]. Rossert [1973]. Rossi [1789]. Rossman [1351]. Rothery [1129, 1129]. Roughness [1003]. Rousseuw [322]. Route [1999, 784]. Routines [1073]. Roving [1479]. Ruark [427]. Rubin [1396]. Rule [1531, 303, 13]. Rules [439, 1585]. Running [227]. Runs [1603, 1252]. Ruppert [1584, 40]. Ryan [1296].

S [251, 208, 1073, 152, 55, 153, 552, 471, 507, 788, 731, 793, 790, 637, 898, 1190, 1137, 1252, 1301, 1400, 1300, 1539, 1497, 1634, 1636, 1741, 1690, 1693, 1792, 1916, 1913, 2042, 2108, 1975, 1976, 2041, 2114, 2116, 255, 1736, 775, 1681, 594, 890, 836, 149, 591, 783, 204, 1009, 1186, 634, 944, 86, 1010, 1396, 375, 138, 317, 206, 1785, 42, 889, 631, 464, 680, 1348, 1296, 1347, 2033, 146, 894, 781, 949, 1242, 196, 1789, 372, 420, 998, 1000, 253, 782, 1738, 1068, 726, 54, 588, 950, 784, 422, 1846, 98]. S [947, 203, 427, 312, 1446]. S-Plus [1446]. S.-C [1186, 1010]. Sachs [1794]. Sacrifice [649, 1616]. Saddle [1029]. Saddle-Point [1029]. Saddlepoint [488, 1386, 1884]. Safeties [184]. Safety [1009, 1042, 698, 1666, 1992]. Salamander [498]. Salhi [51]. Sall [2104]. Salmon [1810, 806]. Salmonella [658]. Salmonella/Microsome [658]. Samarin [1444]. Same [1461, 1694]. Sample [511, 1690, 1694, 1595, 1775, 1461, 1834, 1989, 281, 919, 2008, 748, 1673, 11, 1367, 570, 1476, 1264, 414, 669, 2063, 565, 1562, 2094, 618, 805, 286, 2093, 1558, 701, 710, 827, 1842, 1067, 1843, 1836, 358, 2060, 479, 1612, 1674, 1899, 750, 253, 1145, 804, 396, 810, 1598, 532, 1668, 757, 821, 1509, 1768, 1375, 1034]. Sample-Size [1836, 1145]. Sample-Size-Optimal [919]. Sampled [1373, 242]. Samples [1958, 1052, 212, 819, 3, 1879, 1330, 364, 1167, 1289, 1557, 1680, 1175, 1966, 1574, 1967, 1107, 213, 1106]. Sampling [872, 1797, 647, 1901, 1902, 1335, 1125, 814, 1752, 628, 1676, 1418, 1995, 653, 311, 579, 1213, 1465, 195, 10, 410, 451, 1311, 1560, 1715, 1269, 74, 766, 168, 135, 29, 30, 301, 1570, 76, 237, 20, 1205, 615, 1126, 354, 578, 947, 920, 300, 859, 1987, 167, 107, 108, 1568, 591, 142, 327, 143, 2041, 591, 783, 2035, 1014]. Samuel [1740]. Samuels [48]. Sandell [1494]. Sandwich [1661]. Santner [370]. Saporta [315]. SAS [1909]. SAS-Oriented [1909]. Saunders [736]. Saved [2060]. Saville [590, 1490]. Scale [1190, 1346, 429, 239, 1323, 1768]. Scales [269, 75]. Scan [576]. Schappacher [596]. Scheaffer [1445, 327, 264, 262, 554]. Scheduled [1832]. Scheffé [1550]. Scheme [1051, 813, 615]. Schizophrenic [1474]. Schmidt [94]. Schmitz [1197]. School [1418]. Schrage [148]. Schreuder [1014]. Schull [838]. Schürmann [1687]. Schwartz [1193]. Science [1248, 472, 471, 1193, 189, 1186, 1398, 588, 846]. Sciences [1736, 1785, 2033, 312, 377, 430, 1076, 1542, 1587, 1543, 1634, 1859, 371, 942, 941, 48, 1247]. Sciences-Supplement [312]. Scientific [379, 100, 635, 1582, 730, 730]. Scientists [473, 1792, 1912, 100]. Sclerosis [65]. SCOPE [259]. Score [1333, 1103, 770, 284, 1344, 1777, 1674, 1370, 1110, 1121]. Scores [166, 996, 1279]. Scott [1305, 946]. Screened [2080, 1472]. Screening [1255, 1832, 1026, 308, 529, 2010, 223, 1947, 345, 1027, 1472, 1665]. Screenings [1832]. Sealy [464]. Search [1982]. Searle [998, 1636, 42]. Seasonality [1221, 927]. Seber [373, 1034]. Second [360, 451, 98]. Second-Order [360]. Second-Phase [451]. Secondary [1724, 354]. Section [1653]. Sectional [917]. Sections [613]. Segel [632]. Segmented [1288]. Segregation [993, 981]. Seizure [395, 460]. Selected [1252, 1904, 1905, 368, 2053, 2054, 409, 442]. Selecting [222]. Selection [819, 951, 890, 1517, 1151, 1333, 699, 147, 1143, 1823, 1995, 348, 1167, 1506, 229, 1165, 1557, 648, 232, 2030, 74, 233, 1225, 1424, 1570, 421, 1064, 356, 1680, 1339, 1956, 1321, 1516]. Selections [607]. Selective [1519]. Selectivity [1118]. Self [1447, 1949, 1883]. Self-Designing [1883]. Self-Learning [1447]. Self-Propelled [1949]. Semi [355, 1945, 1501, 1946, 1152]. Semi-Markov [355, 1945, 1946]. Semi-Parametric [1501, 1152]. Semiparametric [1259, 1001, 1985, 1099, 1872, 2073, 525, 1759, 1477, 1605, 1873, 1941, 918]. Sen [420, 1067, 1853]. Senn [889]. Sense [96, 999]. Sensitivity [717, 2100, 1761, 412, 224, 1952, 1472]. Separable [1345]. September [2117]. Sequence [2053, 2054, 1733, 1275, 1032, 1104, 1613]. Sequences [802, 979, 1693, 2057, 107, 108, 146]. Sequential [686, 900, 1078, 1637, 1548, 1762, 699, 1114, 1313, 919, 2008, 1176, 1485, 478, 1271, 1816, 698, 604, 1160, 2011, 123, 1178, 1025, 1124, 2093, 305, 1897, 397, 239, 1379, 1156, 767, 15, 978, 2060, 572, 385, 1430, 1771, 1261, 564, 1722, 2074, 124, 523]. Sequentially [1197]. Serial [1342, 1646, 230, 577, 1189, 2078, 1609]. Serially [1837]. Series [395, 791, 1016, 1399, 1745, 1795, 1977, 1705, 1042, 205, 1043, 1262, 460, 1779, 874, 927, 826, 1670, 1375, 1781, 1399, 103, 1138, 1684, 1742]. Seroconverters [918]. Serum [497]. Set [1797, 1335, 1465, 74, 1568]. Sets [1258, 659, 757, 2110, 1246, 1130]. Setting [798, 1824]. Settings [2087]. Several [269, 340, 1940, 218, 1894, 1265, 587, 1373, 361, 75, 1966, 1967, 1377, 110, 332, 621]. Sewall [37]. Sex [278, 997, 860, 19]. Shamoo [263]. Shanbhag [1194]. Shape [963, 1338, 716, 2111]. Shape-Related [1338]. Shaped [1764]. Shapes [1299]. Shapiro [877]. Shared [1617]. Sharing [73]. Shedding [1428]. Shelf [351, 2112, 2016]. Shelf-Life [351]. Shenton [328]. Sherman [2112]. Shift [359]. Shimizu [1975, 1541, 1746]. Short [476]. Shott [255]. Should [105]. Shrinkage [627]. Shukla [1077]. Sib [232, 1606]. Sibling [1804, 2066, 1700]. Siblings [73]. Sibship [1047]. Sickness [145]. Sided [1403, 1229, 247, 62, 1883, 934]. Siegmund [1806]. Sign [1414, 1611, 1970, 1734]. Signal [792]. Signals [567]. Signed [1513]. Significance [988, 366, 1221, 247, 62, 60, 1205, 61]. Significant [1208, 228]. Silini [94]. Silverman [1003]. SIMEX [1961]. Similar [1487, 1111].

Simonelli [1685]. Simonoff [1296]. Simple [872, 187, 606, 1997, 27, 23, 931, 1780, 767, 1269, 1057, 677, 492, 1107, 1106]. Simpler [1485]. Simplified [1287]. Simpson [679]. Simulated [1983]. Simulating [930]. Simulation [1638, 1857, 469, 635, 1117, 148, 148, 195]. Simultaneous [1315, 1222, 1340, 562, 1550]. Singer [959, 1067]. Single [1342, 649, 482, 1510, 2053, 2054, 1359, 1312, 618, 587, 1454, 332]. Single- [1454]. Single-Channel [482]. Single-Member [1359]. Single-Plant [1510]. Single-Sacrifice [649]. Singpurwalla [1789]. singularities [1427]. Sinha [591]. Siotani [1541, 1746, 1975]. Sir [589]. Sires [109]. Site [1999]. Situation [211]. Situations [824]. Size [344, 511, 1402, 1338, 452, 1418, 1749, 919, 2008, 748, 1673, 1367, 863, 828, 1264, 669, 1929, 483, 346, 618, 805, 2093, 1558, 701, 710, 827, 1836, 358, 74, 479, 1612, 1674, 1899, 253, 1145, 29, 301, 1648, 1570, 237, 2084, 20, 396, 1887, 613, 1598, 532, 1290, 1668, 1619, 1047, 1606, 1204, 1515, 1753, 1375, 2018, 1653]. Size- [1338]. Size-Biased [74, 237]. Size-Metastasis [346]. Sizes [33, 570, 414, 1814]. Skalski [836]. Skeletal [1140]. Skew [1465]. Skewness [267, 555, 113, 270, 556]. Skin [171]. Skinner [506]. Sleep [2023, 387]. Slope [1278, 174, 801]. Slopes [1673, 350]. Small [1958, 269, 819, 1989, 571, 1246, 11, 33, 1879, 1167, 1562, 2094, 1843, 810, 1680, 1233, 75, 757, 821, 2111]. Small-Sample [1989, 11, 2094, 810, 757, 821]. Smayda [2112]. Smith [848, 1973, 506, 1134, 1845]. Smolt [806]. Smooth [323]. Smoothed [1881]. Smoothers [1820]. Smoothing [566, 14, 577, 175, 1108, 388, 1683]. Snapshots [1494]. Snell [679, 102, 589]. Sobel [1247]. Sober [314]. Social [371, 264, 1247]. Society [158, 558, 741, 908, 381]. Sodium [1964]. Soete [1588]. SoftStat [1084]. Software [1084, 2104]. Soil [787, 771]. Sol [771]. Solid [346, 1515]. Solution [640, 356]. Solutions [177, 198, 897]. Solving [85, 1298]. Some [404, 702, 177, 606, 308, 936, 14, 755, 302, 410, 1340, 874, 1621, 83, 774, 934, 1266, 115, 1768, 1209]. Sophisticated [938]. Sorting [673]. Source [756]. Sources [639]. Space [577, 837, 1950, 1189, 1810, 615]. Spaced [1061, 230, 1262, 868]. Spain [1115]. Span [45]. Sparse [828, 1656, 1657]. Spatial [269, 847, 850, 878, 1518, 592, 1750, 401, 1645, 360, 629, 1234, 1115, 1476, 1879, 2055, 713, 1294, 1269, 93, 75, 234, 1216, 903, 141]. Spatial-Temporal [1216]. Spatially [1112, 1525, 563]. spatio [575, 575]. Spatio-temporal [575]. spatio-temporelle [575]. Special [1458, 529, 163]. Species [1706, 531, 1925, 1435, 1926, 110]. Specific [127, 862, 1028, 584, 1660, 111, 1477, 1318, 2085]. Specification [752]. Specified [1835, 1984]. Specimens [1960]. Speed [205]. Spencer [1790]. Spending [1770, 1313]. Sperm [1949]. Spermarche [231]. Spermaturia [231]. Spherical [902]. Spike [1654]. Spline [914]. Spline-Based [914]. Splines [1468, 1885, 1150, 1937, 1216]. Spreadsheet [473]. Sprent [842, 88, 1847]. SPRT [279]. SPSS [1976]. Square [409, 1269, 712]. Squared [1681, 309, 80]. Squares [974, 799, 117, 749, 1837]. Srivastava [1077, 420]. Stability [326, 1013, 366, 1287, 350, 924, 2016, 1186]. Stabilized [851, 1]. Stable [493, 1965]. Stables [51]. Stage [1906, 1218, 1158, 1608, 1560, 1930, 1454, 1528, 1569, 1781, 195]. Stage-Structured [1906, 195]. stages [110]. Standard [1694, 1461, 1985, 2015, 1708, 1709]. Standardized [243, 584, 1056, 541]. Stangl [1630]. Start [2104]. Starting [538]. State [395, 1157, 1123, 1484, 2020, 1099, 293, 2014, 577, 1189, 1810, 1339]. State-Space [577, 1189, 1810]. Statements [558, 741, 908, 381]. Stationary [1055]. Statistic [930, 989, 1659, 576, 1178, 1156, 1955, 1569]. Statistical [208, 328, 266, 265, 470, 471, 686, 790, 842, 845, 956, 905, 1078, 1135, 1192, 1084, 1138, 1302, 1399, 1301, 1305, 1634, 1541, 1635, 1791, 1852, 1746, 1696, 1853, 1743, 1916, 2037, 1975, 1976, 2111, 1066, 1736, 609, 340, 594, 458, 198, 181, 1146, 189, 365, 890, 469, 1282, 2035, 325, 1380, 881, 245, 625, 1748, 841, 109, 1785, 1964, 1089, 180, 140, 313, 260, 464, 635, 680, 1247, 1739, 2033, 370, 894, 1312, 137, 926, 2078, 1537, 397, 831, 1393, 88, 1687, 412, 825, 994, 1055, 1910, 2007, 445, 505, 588, 1060]. Statistical [1095, 1754, 384, 418, 924, 1883, 614, 999, 601, 1847, 526, 312, 379, 638, 685, 787, 848, 902, 958, 1074, 1082, 1198, 1401, 1495, 1799, 2040, 194, 255, 590, 1490, 943, 1186, 324, 942, 368, 192, 1069, 725, 1130, 1008, 1492, 589, 249, 502, 207, 1006, 50, 93, 1070]. Statistically [1208, 228]. Statistician [1798, 85, 1298]. Statisticians [189, 148, 545, 735]. Statistics [210, 211, 431, 430, 432, 509, 597, 473, 471, 598, 595, 731, 789, 790, 794, 954, 1018, 906, 897, 1020, 1085, 1141, 1445, 1400, 1354, 1446, 1351, 1641, 1740, 1688, 2039, 1979, 1980, 1919, 255, 1904, 1905, 426, 1356, 1244, 142, 1673, 1273, 1443, 1772, 86, 840, 1395, 723, 264, 553, 1243, 1129, 1296, 1787, 409, 1391, 661, 442, 1340, 679, 1845, 554, 503, 369, 1394, 1513, 1075, 165, 2104, 162, 501, 1593, 255, 198, 264, 1912, 1073, 152, 55, 328, 508, 683, 684, 1017, 1250, 1079, 1137, 1303, 1586, 1741]. Statistics [1742, 1917, 1913, 2038, 1244, 149, 151, 150, 321, 1299, 633, 893, 1494, 1971, 1789, 53, 1067, 49, 1005, 54, 1851, 1184, 425, 101, 510, 847, 849, 1139, 1448, 1692, 1790, 1914, 1297, 592, 839, 634, 197, 48, 785, 1133, 136]. Statistik [1794, 1848]. Statistique [1004, 898, 315, 953]. Statistiques [637]. statistischer [1794]. Stayer [2086]. Stefanski [1584]. Stein [903]. Steiner [1302, 633]. Step [699, 1041, 1621, 1361]. Step-Up [1041, 1621, 1361]. Stepwise [1430]. Stereology [1402, 483]. Sterile [665]. Stern [1396]. Stewart [640]. Stichproben [1848]. Stirzaker [1353]. Stochastic [344, 1194, 1544, 1953, 1348, 665, 949, 1374, 804, 1512, 1722, 1670, 2017, 1678, 92, 1081, 1582]. Stochasticity [1751]. Stochasticity-Concepts [1751]. Stolley [1080]. Stopping [1770, 1664, 1176, 439, 13, 1547]. Stoyan [1299, 1299]. Straight [774]. Strata [763, 451, 1233]. Strategies [100, 143, 1093, 107, 108]. Strategy [1151, 453, 1666]. Stratification [1760, 811]. Stratified [1387, 1595, 2071, 628, 1830, 309, 68, 1412, 479, 1674, 1648, 996, 2048, 285, 1732, 2070]. Stratum [398, 611]. Stress [1755, 427]. Strijbosch [266]. Strip [1570]. Strongly [81]. Structural [1863]. Structure [1196, 1350, 1698, 542, 1928, 232, 2021]. Structured [870, 1504, 1906, 623, 195]. Structures [1395, 1627, 1671, 1292, 1360, 520]. Structuring [25]. Stuart [723, 1243]. Studden [1980]. Student [1445, 464].

Students [1018, 1688]. Studies [171, 1828, 507, 1452, 1917, 2042, 2117, 818, 392, 127, 1038, 1125, 404, 667, 1232, 1960, 18, 763, 1614, 1549, 1103, 1420, 491, 1900, 1658, 2046, 720, 649, 881, 64, 11, 1367, 2068, 184, 1158, 311, 309, 23, 1028, 2100, 284, 2075, 694, 180, 80, 1296, 287, 2066, 1371, 1563, 1789, 1767, 671, 1761, 1421, 648, 1558, 701, 710, 756, 538, 825, 2009, 479, 1221, 1674, 1899, 1225, 677, 1756, 1212, 399, 224, 861, 1952, 350, 655, 581, 285, 1650, 2098, 544, 1930, 1708, 811, 1169, 2019, 2090, 524, 387, 1432, 1876, 391, 859, 1509, 2095, 167, 1988, 1439]. Studies [523, 1710, 306, 1372, 1709, 780, 1010, 466, 207, 1187, 253, 1181]. Study [341, 433, 717, 177, 760, 660, 1863, 970, 429, 13, 1928, 497, 239, 1878, 225, 338, 1337, 288, 810, 221, 336, 1501, 524, 1720, 838]. Study/Validation [1501]. Studying [1418]. Style [429]. Sub [278]. Sub-Binomial [278]. Subba [1016]. Subclasses [519, 407]. Subclusters [569]. Subdaily [2023]. Subgroup [190]. Subject [1164, 1293, 287, 1165, 1557, 799, 528, 1409, 1946, 2085]. Subject-Specific [2085]. Subjected [1174]. Subjective [1537, 1253]. Subpixel [1566]. Subset [852, 337, 1424, 421]. Subsets [155, 5]. Substances [1119]. Success [1792]. Sucking [1268]. Suen [465]. Sufficient [1209]. Summarizing [1949]. Summary [1673, 1869]. Sumpf [1180]. Superstitions [134]. Supplement [954, 1129, 312]. Supply [1493]. Supply-Demand [1493]. Surface [400, 120, 675, 1388]. Surfaces [1878, 41]. Surprising [410]. Surrogate [1766, 1821, 1158, 2072, 1720]. Surveillance [1146, 995, 1093]. Survey [732, 1072, 327, 1511, 1968, 143, 548, 787, 904, 548]. Surveys [1803, 1802, 1757, 1999, 1993, 1750, 1479, 750, 921, 353, 706, 1095, 917, 1983, 1864, 506, 1354]. Survival [1744, 1915, 647, 929, 1934, 1548, 762, 765, 1676, 912, 975, 974, 649, 1069, 1823, 1158, 2012, 27, 914, 913, 1473, 132, 1028, 444, 1410, 1577, 1862, 1819, 1884, 13, 1600, 126, 68, 1178, 817, 481, 1222, 1897, 239, 1379, 932, 1283, 648, 1203, 2058, 282, 619, 1007, 1187, 1270, 1951, 226, 225, 973, 1769, 1044, 1460, 1965, 1363, 2001, 2013, 1602, 1824, 1480, 1219, 1431, 2070, 1822, 1956, 800, 1662, 1771, 1266, 1720, 2095, 1090, 1481, 1838, 1594, 1546, 1174, 2062, 1447, 2043]. Survival-Adjusted [649]. Survivor [711]. Survivors [838, 496, 1826]. Survivorship [1530, 466]. Susceptibility [1998, 992]. Suspect [1982]. Sustainability [1302, 2112]. Svensson [1131]. Sweeting [848]. Swiss [1402, 483]. Symmetric [612, 1934, 1217, 1457, 196]. Symmetry [1716, 712]. Synergism [175]. Synergy [931]. Synthesis [1245]. Synthetic [476]. System [1083, 72, 120, 1444]. Systematic [844, 264, 1131]. Systems [326, 469, 867, 1287, 1622, 428, 736, 1196].

T [153, 105, 329, 506, 597, 470, 788, 685, 1015, 845, 1016, 1018, 962, 848, 902, 1191, 1140, 1251, 1197, 1074, 1138, 1541, 1744, 1690, 1688, 1746, 1915, 1975, 1976, 887, 1066, 1072, 318, 2105, 371, 727, 204, 1069, 1686, 140, 264, 1784, 2112, 370, 429, 316, 196, 1489, 374, 726, 1014, 682, 947, 644, 939]. t-REML [644]. t-Test [939]. T4 [1043]. Table [393, 570, 587, 533, 2084, 26, 332].

Tableaux [90]. Tables [267, 555, 739, 1901, 1902, 1392, 1904, 1905, 814, 371, 1518, 1483, 25, 33, 1716, 659, 1894, 570, 113, 1579, 829, 721, 1895, 1344, 1777, 1717, 777, 455, 1161, 712, 456, 815, 270, 556, 1377, 621, 552, 1305]. Tabulations [617]. Tag [1595, 1522, 1428, 616, 2017]. Tag-Recapture [2017]. Tag-Recovery [1595, 1522, 616]. Tag-Resight [1522]. Tag-Shedding [1428]. Tags [2048]. Tail [576]. Takács [1181]. Taking [1460, 1826]. Tandem [1809]. Taneja [1975, 151]. Tanner [1743]. Tanur [101, 471]. Target [1430]. Tarone [1721]. Tarter [1002]. Tassi [637]. Taxon [1275]. Technique [304, 665, 1095]. Techniques [1423, 328, 1125, 896, 2027, 957, 836]. Telemetry [865, 1754]. Temporal [1217, 1216, 575]. Temporary [1469]. temporelle [575]. Teng [1806]. Tensors [178]. Teratological [937]. Teratology [1125]. Term [1139, 1676, 1826]. Terms [979, 1532, 313]. Terwilliger [1397]. Test [1052, 210, 269, 511, 1458, 1775, 340, 982, 1829, 1259, 718, 1470, 2076, 930, 172, 2089, 1381, 1414, 876, 911, 1103, 752, 1380, 1576, 629, 988, 33, 1234, 989, 308, 2069, 570, 1579, 477, 23, 1388, 1721, 664, 2011, 1616, 768, 1289, 1780, 1441, 1897, 1344, 1777, 767, 1836, 358, 715, 877, 939, 1057, 677, 29, 672, 1655, 804, 1938, 1611, 1970, 567, 2084, 60, 345, 1027, 75, 1896, 1382, 1927, 2090, 1260, 611, 2072, 1827, 753, 1569, 711, 1734, 1814, 1593, 535, 388, 1121, 915, 61, 78]. Test-Based [2076]. Tested [109]. Testing [1087, 1200, 878, 1573, 488, 700, 87, 172, 1273, 1041, 72, 1716, 1835, 1997, 1459, 1507, 307, 1390, 1210, 608, 830, 1888, 604, 993, 518, 118, 582, 1868, 1717, 820, 1203, 831, 1000, 1385, 1831, 533, 1221, 1674, 1899, 527, 1373, 1840, 417, 779, 1416, 1363, 2005, 363, 567, 416, 350, 1205, 1732, 614, 419, 1275, 1486, 991, 2074, 543, 1361, 1059, 1730, 935, 517, 1110, 775, 1681, 1635]. Tests [1387, 393, 1087, 1403, 1854, 1694, 1230, 1053, 1934, 1726, 340, 1328, 703, 1461, 1463, 31, 1364, 1333, 1518, 2087, 1286, 281, 1313, 1324, 748, 1442, 1358, 1285, 1830, 2012, 1894, 176, 1426, 529, 1879, 109, 914, 2053, 2054, 1271, 669, 284, 366, 1436, 698, 1166, 323, 1160, 114, 228, 1950, 123, 931, 223, 1371, 229, 1340, 744, 1701, 1163, 1502, 1961, 1657, 1621, 1842, 1962, 1843, 978, 769, 745, 226, 605, 1370, 1229, 247, 62, 1482, 1649, 1357, 60, 923, 2013, 1241, 438, 1431, 2070, 934, 443, 1341, 1266, 1722, 130, 1530, 1509, 808, 935, 1768, 306, 1702, 1110]. tests [32, 575, 1372, 61, 78, 1499]. Tetraploids [1339]. Text [1447]. Thall [1640]. Their [905, 1350, 1304, 347, 1072, 1737, 447, 1820, 1947, 369, 545, 613, 1966, 1967]. Theorem [119, 1143, 1550]. Theoretic [2029]. Theoretical [736, 206]. Théorie [637]. Théorique [953]. Théoriques [953]. Theory [733, 685, 955, 957, 901, 962, 1639, 1740, 1916, 2111, 1980, 594, 1128, 1443, 864, 197, 1492, 723, 589, 635, 1243, 1973, 946, 143, 420, 1632, 1279, 1205, 1668, 526, 1494, 1911, 326, 1136, 1082, 1690, 1849, 1243]. Therapeutic [624]. Thermograms [1865]. Thinking [1799]. Thisted [260]. Thomas [951]. Thompson [1803, 783, 257]. Three [1548, 814, 1099, 1624, 823, 714, 2014, 1312, 81, 178, 1655, 440, 815, 1811]. Three-Dimensional [714, 178, 815]. Three-Level [1624].

Three-Parameter [823]. Three-State [1099, 2014]. Three-Way [1811]. Threshold [1778, 419]. Thresholds [2085]. Thumb [303]. Thursday [1143]. Tibshirani [318, 945]. Tied [1577]. Ties [1414, 1970, 1734]. Tiku [1180]. Time [395, 238, 103, 1016, 1138, 1199, 1255, 1399, 1745, 1795, 1977, 818, 292, 1882, 1684, 452, 2003, 2004, 2000, 281, 975, 1705, 1026, 131, 1936, 816, 1725, 1468, 610, 1819, 1436, 205, 1863, 235, 1950, 1037, 460, 817, 1839, 1779, 874, 2059, 1318, 1609, 1714, 927, 868, 226, 225, 540, 122, 2025, 445, 2001, 179, 1480, 1774, 1946, 1472, 1384, 2019, 462, 1152, 1236, 1670, 1941, 753, 1466, 352, 2018, 1594, 2062, 1399, 837, 791, 792]. Time-Adjusted [1436]. Time-Dependent [131, 1725, 1819, 868, 1236]. Time-Frequency [792]. Time-Specific [1318]. Time-to-Response [122]. Time-Varying [1936, 1725, 2025]. Timecourse [348]. Times [1643, 343, 700, 974, 1577, 1554, 1262, 305, 1719, 1945, 668, 1316, 1317]. Timing [1123, 860]. Tips [896]. Tissue [178]. Tits [1575, 1735, 1969]. Tolley [1130]. Tomassone [898, 940]. Tomato [1226]. Tome [953]. Tong [196]. Tongeren [47]. Tool [264, 38]. Tools [1691, 593, 1932, 1743]. Topics [680]. Topologies [1275]. Total [1705, 2067, 1479, 531]. Tour [1590]. Toussaint [147]. Toxicity [853, 1125, 1232, 1218, 349, 1425, 987, 1711, 530, 449, 448, 758]. Toxicological [610]. Toxicology [702, 1994, 1650, 839]. Toxins [609]. Tracking [261, 2081]. Tracks [1045]. Trade [1425]. Trade-Off [1425]. Trails [1122]. Train [1654]. Trait [459, 881, 1207, 1521, 1312, 1519, 1651, 1145, 1812, 705, 1417]. Traits [1773, 1207, 1804, 71, 2108]. Trajectories [1508, 1274]. Transect [1803, 1802, 536, 1418, 1757, 1993, 1750, 1995, 1213, 1715, 29, 76, 353, 237, 1983, 1864, 300]. Transformations [1675, 1475, 2073, 40]. Transient [1887, 1530]. Transients [1460]. Transit [2006]. Transition [2028, 1330]. Transitions [763]. Transmission [1727, 976]. Transplant [1725]. Transplantation [1863, 1384]. Transplants [1384]. Transport [178]. Transversal [111]. Transversion [1457]. Trapping [922]. Travelled [1023]. Treat [1376, 1667, 1421]. Treatment [1548, 404, 31, 494, 654, 1324, 1177, 770, 1214, 1459, 1507, 1763, 694, 1733, 1391, 1502, 1165, 767, 363, 1952, 1732, 182, 1992, 1093, 611, 517, 32]. Treatment-Covariate [1391]. Treatments [880, 574, 699, 1576, 1214, 1156, 1621, 1612, 2072]. Tree [870, 1504, 823, 1185]. Tree-Structured [870, 1504]. Trees [202, 1819, 481, 468]. Trend [1483, 1579, 1507, 1436, 228, 1616, 768, 1777, 1836, 1938, 2090, 1730]. Trends [1199, 1276, 1468, 1037, 1221, 1714, 484]. Trial [171, 853, 1640, 1664, 227, 1176, 1485, 822, 123, 497, 1842, 530, 2007, 1938, 757, 275, 624, 1722, 124, 274]. Trials [854, 1255, 1087, 441, 700, 1218, 1882, 1762, 930, 871, 1114, 1425, 987, 1155, 1342, 919, 2008, 1645, 1177, 1531, 1026, 478, 857, 2069, 1872, 1964, 244, 1440, 1345, 84, 439, 1228, 698, 889, 13, 1160, 1124, 886, 2093, 305, 1897, 397, 1165, 1187, 978, 9, 310, 345, 1027, 1883, 1505, 1942, 1668, 2029, 603, 2074, 925, 858, 1666, 1992, 2072, 1569, 1665, 1723, 1932, 935, 1375, 1781, 36, 1039, 900, 1448, 1637, 1009, 320]. Tribute [1134, 588]. Trinomial [349]. Trough [1282]. Trough-to-Peak [1282]. True [1199, 1455, 1037, 1456]. Truncated [1320, 1259, 1923, 697, 1028, 1660, 2014, 670, 294, 1783, 1102, 995, 1769, 1602, 1169]. Truncation [988, 565, 1946]. Truncation-Flexible [988]. Truth [730, 136]. Tsuang [788]. Tube [714]. Tubes [1289]. Tuckwell [1494, 92, 1136]. Tufte [504, 1850]. Tukey [209, 1065, 728, 729, 784]. Tumor [2028, 298, 1616, 1030, 1341, 1619, 1753]. Tumorigenic [645, 646, 716]. Tumorigenicity [299]. Turckheim [898, 193]. Turkman [1692, 785, 1133]. Tutorial [896]. Twice [161]. Twin [1813, 663, 1710]. Two [395, 1052, 511, 739, 1451, 1694, 1534, 403, 1595, 1775, 775, 879, 1392, 1461, 814, 1218, 574, 2076, 1315, 586, 281, 1484, 2020, 25, 401, 748, 1214, 629, 650, 1183, 2068, 402, 1158, 659, 1608, 2067, 176, 1345, 660, 307, 1173, 1410, 1671, 2063, 366, 1228, 1733, 565, 721, 235, 1289, 286, 1209, 767, 1560, 777, 812, 294, 358, 715, 533, 2060, 1612, 1331, 34, 247, 804, 1482, 1622, 356, 20, 396, 1171, 1505, 1930, 1528, 128, 1992, 456, 407, 1569, 1723, 711, 1768, 1781, 1574, 1702, 1263]. Two-Allele [356]. Two-by-Two [814, 1228]. Two-Component [511, 294, 358]. Two-Dimensional [586, 402, 1345, 1992]. Two-Group [396]. Two-Locus [356]. Two-Mutation [1209]. Two-Period [1569]. Two-Phase [2068, 20]. Two-Sample [1595, 1775, 281, 748, 2063, 565, 286, 2060, 804, 1768]. Two-Sided [247]. Two-Stage [1218, 1158, 1608, 1560, 1930, 1528, 1569, 1781]. Two-State [395, 1484, 2020]. Two-Treatment [1733]. Two-Way [739, 1392, 1315, 25, 1183, 659, 176, 1173, 721, 777, 533, 1482, 456, 1263]. Type [171, 1194, 1087, 880, 1999, 1997, 661, 1412, 1045, 935, 2097, 572, 1685]. Types [1428, 1748, 1030]. Typesetting [896]. Typicality [1458].

U [687, 890, 1911, 1764]. U-Shaped [1764]. Umbach [2117]. Umbrella [654, 1502]. UMVUE [1485]. Unaffected [404]. Unbalanced [693, 1315, 1772, 1285, 176, 561, 42, 1621, 605, 1263, 436]. Unbounded [526]. Uncertain [1662]. Uncertainty [1134, 636]. Unconditional [1830, 1894, 275, 274]. under-Dispersion [1220]. Under-or [626]. Understanding [264, 1075]. Unequal [872, 989, 1465, 1109, 1814, 1034]. Unequally [1041, 1061, 230, 1262, 868]. Unevenly [772]. Unified [1687, 143]. Uniformity [1997]. uniformly [78]. Unifying [2011]. Unique [1339]. Unit [1040, 1623, 301]. Units [1570, 354]. Univariate [1638, 752, 1242, 2027, 1870, 1871, 781]. Universal [291]. Unknown [343, 756, 1945, 101]. Unmeasured [2100, 1761, 394]. Unplanned [1176]. Unread [2048]. Unstratified [479]. Up-and-down [79]. Update [1736, 1785, 2033]. Upon [1120]. Upper [576, 1952]. Upper-Tail [576]. Upton [141, 959]. Urn [1233]. Use [1423, 979, 328, 790, 1852, 1534, 494, 1900, 2069, 570, 1436, 938, 1340, 937, 1165, 484, 1185, 2090, 524, 1569, 1869, 1710, 1678]. Used [1894, 769]. User [547, 1444]. Users [896]. Uses [893]. Using [865, 1581, 2117, 188, 392, 1832, 1335, 1240, 1524, 1675, 236, 693, 1467, 1422, 1333, 172, 1418, 1295, 1511, 1380, 1881, 2028, 217, 476, 653, 566, 1130, 1345, 1679, 1226, 1804, 1028, 1468, 1809, 1627, 1089, 1929, 1884, 620, 1521, 1863, 896, 240, 1600, 585, 525, 1839, 1928, 1895, 867, 495, 1557, 922, 1669, 2058, 1609, 2079, 531, 1831, 1269, 1270, 1939, 1674, 1807, 1225, 540, 1373, 973, 1996, 2025, 1416, 1277, 567, 1095, 1754, 1150, 1487, 1279, 1474, 758, 652, 917, 1111, 616, 1469, 1837, 1275, 1435, 1812, 1322, 1323, 1805, 1808, 2104, 300, 1096, 1378, 1054, 1988, 1516, 808]. Using [981, 1937, 1407, 1216, 1671, 1452, 216]. Utilisation [216]. Utility [1898]. Utilization [1337]. Utilizing [1208].

V [378, 510, 732, 1134, 1251, 1302, 1399, 1305, 1540, 1852, 1692, 2037, 1975, 887, 1066, 1737, 838, 423, 1245, 553, 589, 785, 1133, 726, 588, 682]. Vaccine [1998, 1042, 1748, 1872, 348, 1239]. Valencia [1115]. Valeurs [1116]. Valid [1990]. Validating [2080]. Validation [890, 1766, 1186, 364, 2075, 495, 1501]. Validity [1815, 403, 293, 1577]. Valsecchi [1187]. Value [1455, 761, 633, 1000, 60, 1456, 61]. Valued [1684]. Values [700, 1170, 1464, 287, 1116, 652, 419, 1454, 719, 562]. Variability [31, 606, 32]. Variable [872, 2024, 1295, 217, 1823, 1991, 1506, 2066, 801, 1225, 921, 569, 1046, 706, 1332, 1893, 1047, 523, 1559]. Variables [979, 1534, 911, 1168, 2028, 1278, 1491, 1596, 1818, 81, 455, 185, 219, 1277, 656, 1343, 1369, 1427]. Variance [1875, 1617, 980, 1272, 1328, 169, 1232, 280, 1604, 693, 2089, 1286, 2022, 1278, 888, 1285, 521, 1608, 1224, 1521, 1437, 534, 1961, 932, 1283, 219, 1831, 1910, 135, 605, 1672, 1986, 437, 991, 71, 43, 22, 21, 277, 1200, 729, 728, 1907, 998]. Variance-Covariance [21]. Variances [1040, 764, 989, 308, 1556, 1623, 1109]. Variate [46]. Variates [827, 532]. Variation [278, 626, 827, 622, 1996, 1409, 277]. Varieties [1470]. Variety [2040, 1645, 1122, 1345, 2007]. Vary [452]. Varying [1936, 1725, 803, 2025, 1024]. Vector [1572]. Vegetation [628]. Venables [1446]. Ventilating [1289]. Ventilation [450]. Ventilation-Perfusion [450]. Venzon [1209]. Verbeke [1909]. Verification [1053, 1284, 1702]. Verma [1077]. Versatile [1340]. Versions [1414, 1894, 1734]. Versus [1998, 362, 1160, 1025, 1672, 1863]. Vertebrate [862, 111]. Very [1176]. VI [209, 553]. Via [2073, 805, 1110, 846, 386, 1281, 1034, 1429, 1861, 1117, 1369]. View [1494, 1687]. Viewing [1716]. VII [954]. VIII [954, 1065]. Violations [572]. Virological [1932]. Virus [1748, 976]. Viruses [759]. Visibility [1095, 707]. Vision [844]. Visual [1850]. Visualization [946]. Visualizing [944]. Visweswara [510]. Vital [1468]. Vitro [412, 717]. Vivo [1090, 1932]. Voelkl [1976]. Vol [1141, 1641, 1741, 1746, 1975, 1919, 1244, 142, 1443, 1243, 1242, 207, 141]. Vollandt [1854]. Vols [1541]. Volume [209, 160, 431, 383, 509, 560, 743, 954, 966, 1202, 1406, 1592, 1740, 1796, 1801, 2045, 1736, 463, 1785, 675, 553, 2033, 1065, 294, 312]. Volumes [1494]. Vos [143]. Voss [1974]. vs [730, 247]. Vue [637].

W [209, 55, 377, 266, 596, 471, 786, 640, 899, 960, 1076, 1250, 1082, 1138, 1399, 1349, 1446, 1305, 1586, 1543, 1791, 1917, 1976, 2038, 1980, 1003, 87, 1684, 327, 2035, 147, 944, 838, 197, 376, 725, 1395, 1849, 728, 729, 148, 264, 323, 896, 1444, 2036, 894, 781, 196, 946, 143, 52, 554, 1065, 253, 89, 200, 1064, 1394, 784, 422, 94, 425, 39, 427]. Wackerly [554]. Wald [1772, 1437]. Waldron [1140]. Walk [1531]. Walks [1796, 1796]. Wand [1683]. Wang [999]. Warrack [264]. Warren [427]. Wasil [846]. Wasserman [39]. Water [1133]. Waterfowl [765, 1126]. Waterman [1693, 146]. Watkins [1445]. Watson [588]. Watt [1018, 1688]. Way [739, 1392, 654, 1315, 25, 1183, 659, 176, 1173, 721, 1621, 777, 533, 1482, 2084, 1263, 456, 1876, 1811]. Weak [835]. Wearden [634]. Web [922]. Webs [994, 1729, 952]. Webster [787]. Weibull [238, 646, 823, 540]. Weibull-Based [540]. Weight [338]. Weighted [762, 1291, 1929, 114, 1340, 799, 1843, 749, 345, 1552, 1530, 535]. Weighting [40]. Weinberg [2117, 2089, 72, 477, 1057, 1373, 130]. Weir [98, 1497, 204, 375]. Weisberg [1347]. Weisstein [2036]. Welch [1576]. Well [161, 100]. Well-Known [161]. Wellner [1001]. Wermouth [1682]. West [2107]. Westergren [1398]. Westfall [1000]. Wheat [627]. Where [1052, 715]. White [261]. Whitehead [900, 1637]. Whitney [1259, 78]. Whittaker [321]. Whitten [45]. Who [443]. Wichern [1791]. Wickens [371]. Wide [60, 61]. Width [303]. Wilcoxon [1178, 78]. Wild [373]. Wildlife [865, 261, 836, 1071]. Wilk [877]. Wilks [656]. William [464]. Williams [1185, 1249]. Williamson [259]. Wilson [507, 775]. Wind [205]. Winter [627]. Wintering [1126]. Within [1040, 1718, 1409, 269, 1463, 1556, 178, 1623, 75, 1148]. within- [1148]. Within-Cluster [1718, 1556]. Within-Subject [1409]. Within-Unit [1040, 1623]. without [1793, 1879, 1472, 71]. Witmer [1445]. Wood [590, 1490, 1014]. Woodbury [1130]. Woodlands [202]. Woolson [50]. Woolston [100]. words [1613]. Worked [1012]. Working [1627, 1292]. Works [209, 1065]. Workshop [507, 1351]. World [211, 432, 598, 794, 1020, 1085, 1141, 1641, 1919, 967, 1494, 211, 432, 598, 794, 1641, 1919, 211, 432, 598, 794, 1020, 1085, 1141, 1641, 1919]. Wörterbuch [500]. Wright [1253, 140, 37]. Write [100]. Writing [100, 100]. Wu [793].

X [597, 793]. XVIIIth [1143].

Yandell [1846]. Yang [597]. Yao [1196]. Yates [1086]. Years [1248, 2060]. Yield [1581, 1467, 1345, 627]. Young [1000]. Yule [933].

Z [686, 469]. Zedeck [138]. Zeger [1539]. Zeisel [1912]. Zero [1966, 1967, 1121]. Zidovudine [497]. Zolman [1135]. Zucchini [1684]. Züchtungsmethodik [549]. Zünd [958].

References

Tamura:1987:SME

[1] Roy N. Tamura and S. Stanley Young. A stabilized moment estimator for the beta-binomial distribution. Biometrics, 43(4):813–824, December 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531535. See corrections [851].

Wu:1988:ECC

[2] Margaret C. Wu and Raymond J. Carroll. Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process. Biometrics, 44(1):175–188, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531905. See corrections [6, 271].

Coffey:1988:REM

[3] Mary Coffey. A random effects model for binary data from dependent samples. Biometrics, 44(3):787–801, September 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2531591. See corrections [212, 213].

Wu:1989:ECC

[4] Margaret C. Wu and Kent R. Bailey. Estimation and comparison of changes in the presence of informative right censoring: Conditional linear model. Biometrics, 45(3):939–955, September 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2531694. See corrections [156, 157].

Hosmer:1989:BSL

[5] David W. Hosmer, Borko Jovanovic, and Stanley Lemeshow. Best subsets logistic regression. Biometrics, 45(4):1265–1270, December 1989. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531779. See corrections [155].

Anonymous:1989:CEC

[6] Anonymous. Correction: Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring by modeling the censoring process. Biometrics, 45(4):1347, December 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531805. See [4].

Anonymous:1990:FMa

[7] Anonymous. Front matter. Biometrics, 46(1):i–iv, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531626.

Ellenberg:1990:BCM

[8] Jonas H. Ellenberg, Peter Armitage, Thomas C. Chalmers, Edmund A. Gehan, Judith R. O’Fallon, Stuart J. Pocock, and Marvin Zelen. Biostatistical collaboration in medical research. Biometrics, 46(1):1–32, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531627. With discussion and a

rejoinder by the author.

OQuigley:1990:CRM

[9] John O’Quigley, Margaret Pepe, and Lloyd Fisher. Continual reassessment method: a practical design for phase 1 clinical trials in cancer. Biometrics, 46(1):33–48, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531628.

Kraemer:1990:NCC

[10] Helena Chmura Kraemer and Daniel A. Bloch. A note on case-control sampling to estimate kappa coefficients. Biometrics, 46(1):49–59, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531629.

Ejigou:1990:SSP

[11] Ayenew Ejigou. Small-sample properties of odds ratio estimators under multiple matching in case-control studies. Biometrics, 46(1):61–69, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531630.

Sato:1990:CLC

[12] Tosiya Sato. Confidence limits for the common odds ratio based on the asymptotic distribution of the Mantel–Haenszel estimator. Biometrics, 46(1):71–80, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531631.

Kim:1990:SDC

[13] Kyungmann Kim and Anastasios A. Tsiatis. Study duration for clinical trials with survival response and early stopping rule. Biometrics, 46(1):

81–92, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531632.

Gray:1990:SDM

[14] Robert J. Gray. Some diagnostic methods for Cox regression models through hazard smoothing. Biometrics, 46(1):93–102, March 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531633.

McLeish:1990:SDB

[15] D. L. McLeish and D. Tosh. Sequential designs in bioassay. Biometrics, 46(1):103–116, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531634.

Muller:1990:CND

[16] Hans-Georg Müller and Thomas Schmitt. Choice of number of doses for maximum likelihood estimation of the ED50 for quantal dose-response data. Biometrics, 46(1):117–129, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531635.

Cullis:1990:MAG

[17] B. R. Cullis and C. A. McGilchrist. A model for the analysis of growth data from designed experiments. Biometrics, 46(1):131–142, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531636.

Brown:1990:PAN

[18] C. Hendricks Brown. Protecting against nonrandomly missing data in longitudinal studies. Biometrics, 46(1):143–155, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531637.

Wolter:1990:CRE

[19] Kirk M. Wolter. Capture–recapture estimation in the presence of a known sex ratio. Biometrics, 46(1):157–162, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531638.

Rivest:1990:TPS

[20] Louis-Paul Rivest, Hélène Crépeau, and Michel Crête. A two-phase sampling plan for the estimation of the size of a moose population. Biometrics, 46(1):163–176, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531639.

Wray:1990:AME

[21] Naomi R. Wray. Accounting for mutation effects in the additive genetic variance-covariance matrix and its inverse. Biometrics, 46(1):177–186, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531640.

Wang:1990:LBC

[22] C. Ming Wang. On the lower bound of confidence coefficients for a confidence interval on variance components. Biometrics, 46(1):187–192, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531641.

Hamerle:1990:STN

[23] Alfred Hamerle. On a simple test for neglected heterogeneity in panel studies. Biometrics, 46(1):193–199, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531642. See comment [677].

McDonald:1990:FGL

[24] John W. McDonald and Ian D. Diamond. On the fitting of generalized linear models with nonnegativity parameter constraints. Biometrics, 46(1):201–206, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531643.

Corsten:1990:SIT

[25] L. C. A. Corsten and J. B. Denis. Structuring interaction in two-way tables by clustering. Biometrics, 46(1):207–215, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531644.

Satten:1990:CFR

[26] Glen A. Satten and Lawrence L. Kupper. Continued fraction representation for expected cell counts of a 2 × 2 table: a rapid and exact method for conditional maximum likelihood estimation. Biometrics, 46(1):217– 223, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531645.

Geisser:1990:HBP

[27] Seymour Geisser. On hierarchical Bayes procedures for predicting simple exponential survival. Biometrics, 46(1):225–230, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531646.

Garthwaite:1990:AMR

[28] Paul H. Garthwaite and Stephen T. Buckland. Analysis of a multiplerecapture census by computing conditional probabilities. Biometrics, 46(1):231–238, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531647.

Otto:1990:SBL

[29] Mark C. Otto and Kenneth H. Pollock. Size bias in line transect sampling: a field test. Biometrics, 46(1):239–245, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531648.

Parrish:1990:CQE

[30] Rudolph S. Parrish. Comparison of quantile estimators in normal sampling. Biometrics, 46(1):247–257, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531649.

Brownie:1990:MATa

[31] Cavell Brownie, Dennis D. Boos, and Jacqueline Hughes-Oliver. Modifying the t and ANOVA F tests when treatment is expected to increase variability relative to controls. Biometrics, 46(1):259–266, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531650.

Brownie:1990:MATb

[32] Cavell Brownie, Dennis D. Boos, and Jacqueline Hughes-Oliver. Modifying the t and ANOVA F tests when treatment is expected to increase variability relative to controls. Biometrics, 46(1):259–266, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Engeman:1990:AFE

[33] Richard Engeman, George D. Swanson, and William R. Rice. Alternatives to Fisher’s “Exact test” for analyzing 2 × 2 tables with small cell sizes. Biometrics, 46(1):267–269, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531651.

Nurminen:1990:CIR

[34] Markku Nurminen, Olli Miettinen, John J. Gart, and Jun mo Nam. Confidence intervals for the ratio of the parameters of two independent binomials. Biometrics, 46(1):269–272, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531652.

Richardson:1990:FMM

[35] S. C. Richardson, V. T. Farewell, and D. A. Sprott. Fitting a mixture model to count data. Biometrics, 46(1):273, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531653.

Yee:1990:IAC

[36] K. F. Yee, P. R. Freeman, Sung C. Choi, and Patricia A. Pepple. Interim analysis of clinical trials. Biometrics, 46(1):274–275, March 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531654.

Anonymous:1990:OSW

[37] Anonymous. Obituary: Sewall Wright 1889–1988. Biometrics, 46(1):277– 279, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531655.

Paterson:1990:BRA

[38] L. J. Paterson. Book review: Applied Regression Analysis: a Research Tool, by J. O. Rawlings. Biometrics, 46(1):281–282, March 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531656.

Thoni:1990:BRA

[39] H. Thöni. Book review: Applied Linear Regression Models, by J. Neter, W. Wasserman, M. H. Kutner. Biometrics, 46(1):282–283, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531657.

Sprent:1990:BRT

[40] P. Sprent. Book review: Transformations and Weighting in Regression, by R. J. Carroll, D. Ruppert. Biometrics, 46(1):283, March 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531658.

Brown:1990:BRE

[41] R. A. Brown. Book review: Empirical Model-Building and Response Surfaces, by G. E. P. Box, N. R. Draper. Biometrics, 46(1):283–284, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531659.

Hasted:1990:BRL

[42] A. Hasted. Book review: Linear Models for Unbalanced Data, by S. R. Searle. Biometrics, 46(1):284, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531660.

Venables:1990:BRE

[43] W. N. Venables. Book review: Estimation of Variance Components and Applications, by C. R. Rao, J. Kleffe. Biometrics, 46(1):284–285, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531661.

Hand:1990:BRC

[44] D. J. Hand. Book review: Classification and Related Methods of Data Analysis, by H. H. Bock. Biometrics, 46(1):286, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531662.

Sundberg:1990:BRP

[45] R. Sundberg. Book review: Parameter Estimation in Reliability and Life Span Models, by A. C. Cohen, B. J. Whitten. Biometrics, 46(1):286–287, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531663.

Gentle:1990:BRP

[46] J. E. Gentle. Book review: Principles of Random Variate Generation, by J. Dagpunar. Biometrics, 46(1):287, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531664.

Green:1990:BRD

[47] R. H. Green. Book review: Data Analysis in Community and Landscape Ecology, by R. H. G. Jongman, C. J. F. Braak, O. F. R. Van Tongeren.

Biometrics, 46(1):287–288, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531665.

Habbema:1990:BRS

[48] J. D. F. Habbema. Book review: Statistics for the Life Sciences, by M. L. Samuels. Biometrics, 46(1):288, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531666.

Murray:1990:BRI

[49] G. D. Murray. Book review: Introduction to Medical Statistics, by M. Bland. Biometrics, 46(1):289, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531667.

Murray:1990:BRS

[50] G. D. Murray. Book review: Statistical Methods for the Analysis of Biomedical Data, by R. F. Woolson. Biometrics, 46(1):289, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531668.

Garrido:1990:BRL

[51] J. Garrido. Book review: L’Évaluation de l’Enregistrement des Décès par les Méthodes Pouvant Reposer sur le Modèle des Populations Stables, by M. Salhi. Biometrics, 46(1):289–290, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531669.

Liddell:1990:BRL

[52] F. D. K. Liddell. Book review: Living with Risk: The British Medical Association Guide, by W. Henderson. Biometrics, 46(1):290–291, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531670.

Loukas:1990:BRI

[53] S. Loukas. Book review: Interpreting Data: a First Course in Statistics, by A. J. B. Anderson. Biometrics, 46(1):291, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531672.

Rasch:1990:BRM

[54] D. Rasch. Book review: Modern Mathematical Statistics, by E. J. Dudewicz, S. N. Mishra. Biometrics, 46(1):291, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531671.

Anonymous:1990:BRDa

[55] Anonymous. Book review: Dynamic Graphics for Statistics, by W. S. Cleveland, M. E. McGill. Biometrics, 46(1):292, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531673.

Anonymous:1990:EN

[56] Anonymous. Editorial note. Biometrics, 46(1):292, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531675.

Anonymous:1990:BRM

[57] Anonymous. Book review: Multivariate Chemometrics in QSAR: a Dialogue, by P. P. Mager. Biometrics, 46(1):292, March 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531674.

Anonymous:1990:FMb

[58] Anonymous. Front matter. Biometrics, 46(2):i–iv, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531433.

Aickin:1990:MLE

[59] Mikel Aickin. Maximum likelihood estimation of agreement in the constant predictive probability model, and its relation to Cohen’s kappa. Biometrics, 46(2):293–302, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531434.

Rice:1990:CCVa

[60] William R. Rice. A consensus combined P -value test and the family-wide significance of component tests. Biometrics, 46(2):303–308, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531435.

Rice:1990:CCVb

[61] William R. Rice. A consensus combined P -value test and the family-wide significance of component tests. Biometrics, 46(2):303–308, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piegorsch:1990:OSS

[62] Walter W. Piegorsch. One-sided significance tests for generalized linear models under dichotomous response. Biometrics, 46(2):309–316, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531436.

Conaway:1990:REM

[63] Mark R. Conaway. A random effects model for binary data. Biometrics, 46(2):317–328, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531437.

Durrleman:1990:PME

[64] Sylvain Durrleman and Richard Simon. Planning and monitoring of equivalence studies. Biometrics, 46(2):329–336, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531438.

Joseph:1990:MSI

[65] Lawrence Joseph, Christina Wolfson, and David B. Wolfson. Is multiple sclerosis an infectious disease? Inference about an input process based on the output. Biometrics, 46(2):337–349, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531439.

Tango:1990:ADI

[66] Toshiro Tango. Asymptotic distribution of an index for disease clustering. Biometrics, 46(2):351–357, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531440.

Mitchell:1990:DAB

[67] Toby J. Mitchell and Bruce W. Turnbull. Detection of associations between diseases in animal carcinogenicity experiments. Biometrics, 46(2): 359–374, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531441.

Kronborg:1990:PCS

[68] Dorte Kronborg and Peter Aaby. Piecewise comparison of survival functions in stratified proportional hazards models. Biometrics, 46(2):375– 380, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531442.

Karpinski:1990:OAE

[69] K. F. Karpinski. Optimality assessment in the enzyme-linked immunosorbent assay (ELISA). Biometrics, 46(2):381–390, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531443.

Eliasziw:1990:CRE

[70] Michael Eliasziw and Allan Donner. Comparison of recent estimators of interclass correlation from familial data. Biometrics, 46(2):391–398, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531444.

Thompson:1990:PAQ

[71] E. A. Thompson and R. G. Shaw. Pedigree analysis for quantitative traits: Variance components without matrix inversion. Biometrics, 46(2): 399–413, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2531445.

Eguchi:1990:THW

[72] Shinto Eguchi and Masaaki Matsuura. Testing the Hardy–Weinberg equilibrium in the HLA system. Biometrics, 46(2):415–426, June 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531446.

Darlington:1990:RMH

[73] G. A. Darlington and V. T. Farewell. Regression modelling of HLA haplotype sharing in affected siblings. Biometrics, 46(2):427–434, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531447.

Muttlak:1990:RSS

[74] Hassen A. Muttlak and Lyman L. McDonald. Ranked set sampling with size-biased probability of selection. Biometrics, 46(2):435–445, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531448.

Shaw:1990:TSR

[75] M. W. Shaw. A test of spatial randomness on small scales, combining information from mapped locations within several quadrats. Biometrics, 46(2):447–458, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2531449. See corrections [269].

Quang:1990:CID

[76] Pham Xuan Quang. Confidence intervals for densities in line transect sampling. Biometrics, 46(2):459–472, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531450.

Whitehead:1990:MRE

[77] Hal Whitehead. Mark-recapture estimates with emigration and reimmigration. Biometrics, 46(2):473–479, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531451.

Streitberg:1990:TUM

[78] Bernd Streitberg and Joachim Roehmel. On tests that are uniformly more powerful than the Wilcoxon–Mann–Whitney test. Biometrics, 46(2):481–484, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531452.

Choi:1990:IEL

[79] Sung C. Choi. Interval estimation of the LD 50 based on an up-and-down experiment. Biometrics, 46(2):485–492, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531453.

Kalish:1990:RMS

[80] Leslie A. Kalish. Reducing mean squared error in the analysis of pairmatched case-control studies. Biometrics, 46(2):493–499, June 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531454.

McNally:1990:MLE

[81] Richard J. Q. McNally. Maximum likelihood estimation of the parameters of the prior distributions of three variables that strongly influence reproductive performance in cows. Biometrics, 46(2):501–514, June 1990.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531455.

Tarone:1990:MBM

[82] R. E. Tarone. A modified Bonferronio method for discrete data. Biometrics, 46(2):515–522, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531456.

Neuhaus:1990:SCR

[83] John M. Neuhaus, Nicholas P. Jewell, Bernard Rosner, and Tor D. Tosteson. Some comments on Rosner’s multiple logistic model for clustered data. Biometrics, 46(2):523–534, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531457.

Hilden:1990:CLC

[84] J. Hilden and R. Sylvester. Corrected loss calculation for Phase II trials. Biometrics, 46(2):535–538, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531458.

Derr:1990:BRP

[85] J. A. Derr. Book review: Problem Solving: a Statistician’s Guide, by C. Chatfield. Biometrics, 46(2):539–540, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531459.

Fearn:1990:BRB

[86] T. Fearn. Book review: Bayesian Statistics: Principles, Models, and Applications, by S. J. Press. Biometrics, 46(2):540, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531460.

Buckland:1990:BRC

[87] S. T. Buckland. Book review: Computer-Intensive Methods for Testing Hypotheses, by E. W. Noreen. Biometrics, 46(2):540–541, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531461.

McCulloch:1990:BRA

[88] C. E. McCulloch. Book review: Applied Nonparametric Statistical Methods, by P. Sprent. Biometrics, 46(2):541–542, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531462. See corrigendum [265].

Papaioannou:1990:BRM

[89] T. Papaioannou. Book review: Measurement Error Models, by W. A. Fuller. Biometrics, 46(2):542, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531463.

McHale:1990:BRA

[90] D. McHale. Book review: Analyse Conjointe de Tableaux Quantitatifs, by C. Lavit. Biometrics, 46(2):542–543, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531464.

Diamond:1990:BRA

[91] I. D. Diamond. Book review: Analyse Démographique des Biographies, by D. Courgeau, É. Lelièvre. Biometrics, 46(2):543–544, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531465.

Abakuks:1990:BRS

[92] A. Abakuks. Book review: Stochastic Processes in the Neurosciences, by H. C. Tuckwell. Biometrics, 46(2):544, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531467.

Ord:1990:BRSa

[93] J. K. Ord. Book review: Statistical Inference for Spatial Processes, by B. D. Ripley. Biometrics, 46(2):544, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531466.

Stather:1990:BRR

[94] J. W. Stather. Book review: Radionuclides in the Foodchain, by M. W. Carter, J. H. Harley, G. D Schmidt, G. Silini. Biometrics, 46(2): 545, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2531468.

Kilpatrick:1990:BRC

[95] S. J. Kilpatrick. Book review: Carcinogenicity, by H. C. Grice, J. L. Ciminera. Biometrics, 46(2):545–546, June 1990. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2531469.

Shotter:1990:BRM

[96] M. V. Shotter. Book review: Making Sense of Data, by J. H. Abramson. Biometrics, 46(2):546, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531470.

Cormack:1990:BRP

[97] R. M. Cormack. Book review: Principles of Population Genetics, by D. L. Hartl, A. G. Clark. Biometrics, 46(2):546–547, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531471.

Thompson:1990:BRP

[98] R. Thompson. Book review: Proceedings of the Second International Conference on Quantitative Genetics, by B. S. Weir, E. J. Eisen, M. M. Goodman, G. Namkoong. Biometrics, 46(2):547, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531472.

Woiwod:1990:BRE

[99] I. P. Woiwod. Book review: Ecological Diversity and Its Measurement, by A. E. Magurran. Biometrics, 46(2):547–548, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531473.

Kemp:1990:BRE

[100] A. W. Kemp. Book reviews: Effective Writing Strategies for Engineers and Scientists, by D. C. Woolston, P. A. Robinson, G. Kutzbach; How to Write and Publish a Scientific Paper, by R. A. Day; Writing Mathematics Well, by L. Gillman. Biometrics, 46(2):548–549, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531474.

Anonymous:1990:BRS

[101] Anonymous. Book review: Statistics: a Guide to the Unknown, by J. M. Tanur. Biometrics, 46(2):549–550, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531475.

Anonymous:1990:BRAa

[102] Anonymous. Book review: Analysis of Binary Data, by D. R. Cox, E. J. Snell. Biometrics, 46(2):550, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2531476.

Anonymous:1990:BRAb

[103] Anonymous. Book review: The Analysis of Time Series, by C. Chatfield. Biometrics, 46(2):550, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2531477.

Anonymous:1990:BRI

[104] Anonymous. Book review: Introduction to the Mathematics of Operations Research, by K. J. Hastings. Biometrics, 46(2):550, June 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531478.

Anonymous:1990:BRWa

[105] Anonymous. Book review: What Every Engineer Should Know about Quality Control, by T. Pyzdek. Biometrics, 46(2):550, June 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531479.

Anonymous:1990:FMc

[106] Anonymous. Front matter. Biometrics, 46(3):i–iv, September 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532078.

Weir:1990:BIP

[107] B. S. Weir and C. J. Basten. A biometrics invited paper with discussion. sampling strategies for distances between DNA sequences. Biometrics, 46(3):551–582, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532079.

Weir:1990:SSD

[108] B. S. Weir and C. J. Basten. Sampling strategies for distances between DNA sequences. Biometrics, 46(3):551–582, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goffinet:1990:STI

[109] B. Goffinet, J. M. Elsen, and P. Le Roy. Statistical tests for identification of the genotype at a major locus of progeny-tested sires. Biometrics, 46(3):583–594, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532080.

Sondgerath:1990:ELM

[110] Dagmar Söndgerath and Otto Richter. An extension of the Leslie matrix model for describing population dynamics of species with several development stages. Biometrics, 46(3):595–607, September 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532081.

Lebreton:1990:EAS

[111] Jean-Dominique Lebreton, Georges Hemery, Jean Clobert, and Hervé Coquillart. The estimation of age-specific breeding probabilities from recaptures or resightings in vertebrate populations. i. transversal models. Biometrics, 46(3):609–622, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532082.

Alho:1990:LRC

[112] Juha M. Alho. Logistic regression in capture–recapture models. Biometrics, 46(3):623–635, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532083.

Gart:1990:AIE

[113] John J. Gart and Jun mo Nam. Approximate interval estimation of the difference in binomial parameters: Correction for skewness and extension to multiple tables. Biometrics, 46(3):637–643, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532084. See corrections [267, 270, 556].

Kocherlakota:1990:THW

[114] S. Kocherlakota and K. Kocherlakota. Tests of hypotheses for the weighted binomial distribution. Biometrics, 46(3):645–656, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532085.

Thall:1990:SCM

[115] Peter F. Thall and Stephen C. Vail. Some covariance models for longitudinal count data with overdispersion. Biometrics, 46(3):657–671, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532086.

Lindstrom:1990:NME

[116] Mary J. Lindstrom and Douglas M. Bates. Nonlinear mixed effects models for repeated measures data. Biometrics, 46(3):673–687, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532087.

Matthews:1990:ADC

[117] J. N. S. Matthews. The analysis of data from crossover designs: The efficiency of ordinary least squares. Biometrics, 46(3):689–696, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532088.

Kimura:1990:TNR

[118] Daniel K. Kimura. Testing nonlinear regression parameters under heteroscedastic, normally distributed errors. Biometrics, 46(3):697–708, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532089.

Cox:1990:FTL

[119] Christopher Cox. Fieller’s theorem, the likelihood and the delta method. Biometrics, 46(3):709–718, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532090.

Gennings:1990:IPM

[120] Chris Gennings, Kathryn S. Dawson, Walter H. Carter, Jr., and Raymond H. Myers. Interpreting plots of a multidimensional dose-response surface in a parallel coordinate system. Biometrics, 46(3):719–735, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532091.

Kalish:1990:EDE

[121] Leslie A. Kalish. Efficient design for estimation of median lethal dose and quantal dose-response curves. Biometrics, 46(3):737–748, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532092.

Pack:1990:MMI

[122] Simon E. Pack and Byron J. T. Morgan. A mixture model for intervalcensored time-to-response quantal assay data. Biometrics, 46(3):749– 757, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532093.

Lan:1990:IGS

[123] K. K. Gordon Lan and John M. Lachin. Implementation of group sequential logrank tests in a maximum duration trial. Biometrics, 46(3):759– 770, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532094.

Therneau:1990:ODG

[124] Terry M. Therneau, H. S. Wieand, and Myron Chang. Optimal designs for a grouped sequential binomial trial. Biometrics, 46(3):771–781, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532095.

Liang:1990:CPH

[125] Kung-Yee Liang, Steven G. Self, and Xinhua Liu. The Cox proportional hazards model with change point: an epidemiologic application. Biometrics, 46(3):783–793, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532096.

Klein:1990:PPE

[126] John P. Klein, Shih-Chang Chang Lee, and M. L. Moeschberger. A partially parametric estimator of survival in the presence of randomly censored data. Biometrics, 46(3):795–811, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532097.

Benichou:1990:EAC

[127] J. Benichou and Mitchell H. Gail. Estimates of absolute cause-specific risk in cohort studies. Biometrics, 46(3):813–826, September 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532098.

Sutradhar:1990:DOO

[128] Brajendra C. Sutradhar. Discrimination of observations into one of two t populations. Biometrics, 46(3):827–835, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532099.

Solow:1990:IAO

[129] Andrew Solow and Peter Tyack. Inhomogeneity and apparent organization in animal behavior. Biometrics, 46(3):837–839, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532100.

Theobald:1990:MAP

[130] C. M. Theobald. Midge arrival patterns and tests of the Hardy–Weinberg law. Biometrics, 46(3):841–847, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532101.

Fang:1990:PIM

[131] Ji-Qian Fang, Zhong-Lu Shi, Yi Wang, Xia Zhang, Dong-Lu Zeng, and Jian-Na Zhang. Parametric inference in a multiple renewal process with time-dependent covariates. Biometrics, 46(3):849–854, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532102.

Heisey:1990:BPC

[132] Dennis M. Heisey and Erik V. Nordheim. Biases in the pollock and cornelius method of estimating nest survival. Biometrics, 46(3):855–862, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532103.

Piegorsch:1990:MLE

[133] Walter W. Piegorsch. Maximum likelihood estimation for the negative binomial dispersion parameter. Biometrics, 46(3):863–867, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532104.

Mantel:1990:CCS

[134] Nathan Mantel, Priya Wickramaratne, and Theodore Holford. Confounders: Correcting superstitions. Biometrics, 46(3):869–872, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532105.

Notz:1990:DFE

[135] William Notz and Jonathan Bart. Degrees of freedom for estimates of the variance in multistage sampling plans. Biometrics, 46(3):873–874, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532106.

Tomassone:1990:BRS

[136] R. Tomassone. Book review: Statistics and Truth, by C. R. Rao. Biometrics, 46(3):875–876, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532107.

Kuehl:1990:BRD

[137] R. O. Kuehl. Book review: The Design of Experiments: Statistical Principles for Practical Applications, by R. Mead. Biometrics, 46(3):876–877, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532108.

Frank:1990:BRD

[138] O. Frank. Book review: Data Analysis for Research Designs, by G. Keppel, S. Zedeck. Biometrics, 46(3):877, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532109.

Carroll:1990:BRN

[139] R. J. Carroll. Book review: Nonlinear Regression, Functional Relations and Robust Methods, by H. Bunke, O. Bunke. Biometrics, 46(3):877–878, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532110.

Kelly:1990:BRO

[140] R. E. Kelly. Book review: Order Restricted Statistical Inference, by T. Robertson, F. T. Wright, R. L. Dykstra. Biometrics, 46(3):878–879, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532111.

Ord:1990:BRSb

[141] J. K. Ord. Book review: Spatial Data Analysis by Example, Vol. 2: Categorical and Directional Data, by G. J. G. Upton, B. Fingleton. Biometrics, 46(3):879, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532112.

Cormack:1990:BRH

[142] R. M. Cormack. Book review: Handbook of Statistics, Vol. 6: Sampling, by P. R. Krishnaiah, C. R. Rao. Biometrics, 46(3):879–880, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532113.

Laurence:1990:BRU

[143] A. F. Laurence. Book review: Unified Theory and Strategies of Survey Sampling, by A. Chaudhuri, J. W. E. Vos. Biometrics, 46(3):880, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532114.

Bailey:1990:BRA

[144] N. T. J. Bailey. Book review: Analysis of Infectious Disease Data, by N. G. Becker. Biometrics, 46(3):880–881, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532115.

Murphy:1990:BRS

[145] M. Murphy. Book review: Sickness, Recovery and Death: a History and Forecast of Ill Health, by J. C. Riley. Biometrics, 46(3):881–882, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532116.

Kirkwood:1990:BRM

[146] T. B. L. Kirkwood. Book review: Mathematical Methods for DNA Sequences, by M. S. Waterman. Biometrics, 46(3):882, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532117.

Curnow:1990:BRPa

[147] R. N. Curnow. Book review: Parental Line Breeding and Selection in Potato Breeding, by K. M. Louwes, H. A. J. M. Toussaint, L. M. W. Dellaert. Biometrics, 46(3):882–883, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532118.

Kemp:1990:BRG

[148] C. D. Kemp. Book reviews: Guide to Simulation, 2nd Edition, by P. Bratley, B. L. Fox, L. E. A. Schrage; Simulation Methodology for Statisticians, Operations Analysis, and Engineers, by P. A. W. Lewis, E. J. Orav. Biometrics, 46(3):883–884, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532119.

Calinski:1990:BRI

[149] T. Calinski. Book review: Introduction to the Practice of Statistics, by D. S. Moore, G. P. McCabe. Biometrics, 46(3):884–885, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532120.

Dugard:1990:BRE

[150] P. I. Dugard. Book review: Essential Statistics, 2nd Edition, by D. G. Rees. Biometrics, 46(3):885, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532121.

Cooke:1990:BRM

[151] D. Cooke. Book review: Modern Elementary Probability and Statistics, by E. J. Dudewicz, P. Chen, B. K. Taneja. Biometrics, 46(3):885–886, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532122.

Anonymous:1990:BRAc

[152] Anonymous. Book review: Assignments in Applied Statistics, by S. Conrad. Biometrics, 46(3):886, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532123.

Anonymous:1990:BRE

[153] Anonymous. Book review: Empirical Bayes Methods, 2nd Edition, by J. S. Maritz, T. Lwin. Biometrics, 46(3):886, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532124.

Anonymous:1990:BRF

[154] Anonymous. Book review: Fractal Geometry: Mathematical Foundations and Applications, by K. Falconer. Biometrics, 46(3):886–887, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532125.

Anonymous:1990:CBS

[155] Anonymous. Corrections: Best subsets logistic regression. Biometrics, 46(3):889, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532127. See [5].

Anonymous:1990:CEC

[156] Anonymous. Corrections: Estimation and comparison of changes in the presence of informative right censoring: Conditional linear model. Biometrics, 46(3):889, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532126. See [4].

Wu:1990:CEC

[157] M. C. Wu and K. R. Bailey. Correction to: “Estimation and comparison of changes in the presence of informative right censoring: conditional linear model” [Biometrics 45 (1989), no. 3, 939–955; MR1029611 (90k:62151)]. Biometrics, 46(3):889, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [4].

Anonymous:1990:BS

[158] Anonymous. The Biometric Society. Biometrics, 46(3):891–893, September 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532128.

Anonymous:1990:FMd

[159] Anonymous. Front matter. Biometrics, 46(4):i–iv, December 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532434.

Anonymous:1990:VI

[160] Anonymous. Volume information. Biometrics, 46(4):xiii–xxxii, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532433.

Edwards:1990:RFT

[161] A. W. F. Edwards. R. A. Fisher twice Professor of Genetics: London and Cambridge or “A fairly well-known geneticist. Biometrics, 46(4):897–904, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532435.

Thompson:1990:RFC

[162] E. A. Thompson. R. A. Fisher’s contributions to genetical statistics. Biometrics, 46(4):905–914, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532436.

Piegorsch:1990:FCG

[163] Walter W. Piegorsch. Fisher’s contributions to genetics and heredity, with special emphasis on the Gregor Mendel controversy. Biometrics, 46(4):915–924, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532437.

Preece:1990:RFE

[164] D. A. Preece. R. A. Fisher and experimental design: a review. Biometrics, 46(4):925–935, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532438.

Street:1990:FCA

[165] Deborah J. Street. Fisher’s contributions to agricultural statistics. Biometrics, 46(4):937–945, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532439.

Gower:1990:FOS

[166] J. C. Gower. Fisher’s optimal scores and multiple correspondence analysis. Biometrics, 46(4):947–961, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532440.

Weinberg:1990:DAC

[167] Clarice R. Weinberg and Sholom Wacholder. The design and analysis of case-control studies with biased sampling. Biometrics, 46(4):963–975, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532441.

Neuhaus:1990:ERS

[168] John M. Neuhaus and Nicholas P. Jewell. The effect of retrospective sampling on binary regression models for clustered data. Biometrics, 46(4):977–990, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532442.

Benichou:1990:VCC

[169] Jacques Benichou and Mitchell H. Gail. Variance calculations and confidence intervals for estimates of the attributable risk based on logistic models. Biometrics, 46(4):991–1003, December 1990. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532443.

Hastie:1990:ENC

[170] Trevor Hastie and Robert Tibshirani. Exploring the nature of covariate effects in the proportional hazards model. Biometrics, 46(4):1005–1016, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532444.

Abu-Libdeh:1990:AMT

[171] Hasan Abu-Libdeh, Bruce W. Turnbull, and Larry C. Clark. Analysis of multi-type recurrent events in longitudinal studies; application to a skin cancer prevention trial. Biometrics, 46(4):1017–1034, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532445.

Chen:1990:UGT

[172] Chao L. Chen and William H. Swallow. Using group testing to estimate a proportion, and to test the binomial model. Biometrics, 46(4):1035–1046, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532446.

Jorgensen:1990:IBD

[173] Murray A. Jorgensen. Influence-based diagnostics for finite mixture models. Biometrics, 46(4):1047–1058, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532447.

Mendoza:1990:BAS

[174] M. Mendoza. A Bayesian analysis of the slope ratio bioassay. Biometrics, 46(4):1059–1069, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532448.

Kelly:1990:MSA

[175] Colleen Kelly and John Rice. Monotone smoothing with application to dose-response curves and the assessment of synergism. Biometrics, 46 (4):1071–1085, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532449.

Gallo:1990:ETR

[176] Jose Gallo and André I. Khuri. Exact tests for the random and fixed effects in an unbalanced mixed two-way cross-classification model. Biometrics, 46(4):1087–1095, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532450.

Dalgaard:1990:FNS

[177] Peter Dalgaard and Michael Larsen. Fitting numerical solutions of differential equations to experimental data: a case study and some general remarks. Biometrics, 46(4):1097–1109, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532451.

Nievergelt:1990:FDF

[178] Yves Nievergelt. Fitting density functions and diffusion tensors to threedimensional drug transport within brain tissue. Biometrics, 46(4):1111– 1121, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532452.

Rugg:1990:ATB

[179] David J. Rugg and Richard R. Buech. Analyzing time budgets with Markov chains. Biometrics, 46(4):1123–1131, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532453.

Jewell:1990:SAH

[180] Nicholas P. Jewell and Stephen C. Shiboski. Statistical analysis of HIV infectivity based on partner studies. Biometrics, 46(4):1133–1150, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532454.

Brookmeyer:1990:SMA

[181] Ron Brookmeyer and Jiangang Liao. Statistical modelling of the AIDS epidemic for forecasting health care needs. Biometrics, 46(4):1151–1163, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532455.

Solomon:1990:ACD

[182] P. J. Solomon and S. R. Wilson. Accommodating change due to treatment in the method of back projection for estimating HIV infection incidence. Biometrics, 46(4):1165–1170, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532456.

Brant:1990:APP

[183] Rollin Brant. Assessing proportionality in the proportional odds model for ordinal logistic regression. Biometrics, 46(4):1171–1178, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532457.

Finch:1990:CRS

[184] Peter D. Finch. The comparability of relative safeties from different studies. Biometrics, 46(4):1179–1180, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532458.

Meulepas:1990:COV

[185] E. Meulepas. On a criterion for omitting variables in discriminant analysis. Biometrics, 46(4):1181–1183, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532459.

Myers:1990:RLB

[186] R. A. Myers and P. Pepin. The robustness of lognormal-based estimators of abundance. Biometrics, 46(4):1185–1192, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532460.

Baker:1990:SEA

[187] S. G. Baker. A simple EM algorithm for capture–recapture data with categorical covariates. Biometrics, 46(4):1193–1200, December 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532461.

Bailer:1990:EIU

[188] A. John Bailer and Walter W. Piegorsch. Estimating integrals using quadrature methods with an application in pharmacokinetics. Biometrics, 46(4):1201–1211, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532462.

Bross:1990:HEF

[189] Irwin D. Bross. How to eradicate fraudulent statistical methods: Statisticians must do science. Biometrics, 46(4):1213–1225, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532463.

Berry:1990:SA

[190] Donald A. Berry. Subgroup analyses. Biometrics, 46(4):1227–1230, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532464.

Lachenbruch:1990:BRG

[191] P. A. Lachenbruch. Book review: Generalized Linear Models, by P. Mc-Cullagh, J. A. Nelder. Biometrics, 46(4):1231–1232, December 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532465.

Ekholm:1990:BRS

[192] A. Ekholm. Book review: Statistical Modelling in GLIM, by M. Aitkin, D. Anderson, B. Francis, J. Hinde. Biometrics, 46(4):1232–1233, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532466.

Dutilleul:1990:BRM

[193] P. Dutilleul. Book review: Modèle de Régression de Cox Périodique et Étude d’un Comportement Alimentaire, by O. Pons, E. De Turckheim. Biometrics, 46(4):1234, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532467.

Armitage:1990:BRS

[194] P. Armitage. Book review: Statistical Data Analysis and Inference, by Y. Dodge. Biometrics, 46(4):1234–1235, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532468.

Kempton:1990:BRS

[195] R. A. Kempton. Book review: Stage-Structured Populations: Sampling, Analysis and Simulation, by B. F. J. Manly. Biometrics, 46(4):1235, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532469.

Krzanowski:1990:BRM

[196] W. J. Krzanowski. Book reviews: The Multivariate Normal Distribution, by Y. L. Tong; Symmetric Multivariate and Related Distributions, by K.-T. Fang, S. Kotz, K. W. Ng. Biometrics, 46(4):1235–1236, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532470.

Fang:1990:BRS

[197] J. Q. Fang. Book review: Statistics: Theory and Methods, by D. A. Berry, B. W. Lindgren. Biometrics, 46(4):1236–1237, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532471.

Bowman:1990:BRI

[198] A. W. Bowman. Book reviews: Introduction to Contemporary Statistical Methods, by L. H. Koopmans; Interactive Computing in Statistics, by L. H. Koopmans; Solutions Manual to Accompany Introduction to Contemporary Statistical Methods, by L. H. Koopmans. Biometrics, 46 (4):1237–1238, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532472.

Holgate:1990:BRM

[199] P. Holgate. Book review: Matrix Population Models, by H. Caswell. Biometrics, 46(4):1238, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532473.

Pope:1990:BRP

[200] J. A. Pope. Book review: Population Harvesting, by W. M. Getz, R. G. Haight. Biometrics, 46(4):1238–1239, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532474.

Barath:1990:BRG

[201] E. Baráth. Book review: Grazing Research: Design, Methodology, and Analysis, by G. C. Marten. Biometrics, 46(4):1239–1240, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532475.

Burkhart:1990:BRD

[202] H. E. Burkhart. Book review: Design of the Census of Woodlands and Trees 1979–1982, by K. Rennolls. Biometrics, 46(4):1240, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532476.

Wilson:1990:BRG

[203] S. R. Wilson. Book review: The Geometry of Genetics, by A. M. Findley, S. P. McGlynn, G. L. Findley. Biometrics, 46(4):1240–1241, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532477.

Curnow:1990:BRPb

[204] R. N. Curnow. Book review: Plant Population Genetics, Breeding, and Genetic Resources, by A. H. D. Brown, M. T. Clegg, A. L. Kahler, B. S. Weir. Biometrics, 46(4):1241, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532478.

Jensen:1990:BRA

[205] J. L. Jensen. Book review: The Analysis of Directional Time Series: Applications to Wind Speed and Direction, by J. Breckling. Biometrics, 46(4):1241–1242, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532479.

Gilks:1990:BRT

[206] W. Gilks. Book review: Theoretical Immunity, Parts I and II, by A. S. Perelson. Biometrics, 46(4):1242–1243, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532480.

Machin:1990:BRS

[207] D. Machin. Book review: Statistical Methods in Cancer Research, Vol. II: The Design and Analysis of Cohort Studies, by N. E. Breslow, N. E. Day. Biometrics, 46(4):1243–1244, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532481.

Aitken:1990:BRE

[208] C. G. G. Aitken. Book review: The Evolving Role of Statistical Assessment as Evidence in the Courts, by S. E. Fienberg. Biometrics, 46(4): 1244, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532482.

Anonymous:1990:BRC

[209] Anonymous. Book review: The Collected Works of John W. Tukey, Volume VI, More Mathematical, 1938–1984, by C. L. Mallows. Biometrics, 46(4):1244–1245, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532483.

Anonymous:1990:BRDb

[210] Anonymous. Book review: Distributions of Test Statistics: Exact and Asymptotic, Null and Non-Null. Methods/Comparisons/Research Frontiers, by A. M. Mathai. Biometrics, 46(4):1245, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532484.

Anonymous:1990:BRWb

[211] Anonymous. Book reviews: World Health Statistics Annual 1989, by World Health Organization; World Health Statistics Quarterly, by World Health Organization; Global Estimates for Health Situation Assessment and Projections, by World Health Organization. Biometrics, 46(4):1245, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532485.

Anonymous:1990:CRE

[212] Anonymous. Correction: A random effects model for binary data from dependent samples. Biometrics, 46(4):1247, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532486. See [3].

Coffey:1990:CRE

[213] M. Coffey. Correction to: “A random effects model for binary data from dependent samples” [Biometrics 44 (1988), no. 3, 787–801; MR0963914 (89h:62166)]. Biometrics, 46(4):1247, December 1990. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [3].

Anonymous:1990:PPB

[214] Anonymous. Papers to be published in biometrics. Biometrics, 46(4): 1251, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532487.

Anonymous:1991:FMa

[215] Anonymous. Front matter. Biometrics, 47(1):i–iv, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532489.

Faivre:1991:UDI

[216] R. Faivre, B. Goffinet, and D. Wallach. Utilisation de données intermédiaires pour corriger la prédiction de modèles mécanistes. (French)

[Using intermediate data to correct the prediction of mechanistic models]. Biometrics, 47(1):1–12, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532490.

Engel:1991:IPR

[217] B. Engel and P. Walstra. Increasing precision or reducing expense in regression experiments by using information from a concomitant variable. Biometrics, 47(1):13–20, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532491.

Cheung:1991:EDM

[218] Siu Hung Cheung and Burt Holland. Extension of Dunnett’s multiple comparison procedure to the case of several groups. Biometrics, 47(1): 21–32, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532492.

Miller:1991:GVC

[219] Michael E. Miller and J. Richard Landis. Generalized variance component models for clustered categorical response variables. Biometrics, 47(1):33–44, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532493.

Piegorsch:1991:MCA

[220] Walter W. Piegorsch. Multiple comparisons for analyzing dichotomous response. Biometrics, 47(1):45–52, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532494.

Self:1991:EHD

[221] Steven G. Self, Gary Longton, Kenneth J. Kopecky, and Kung-Yee Liang. On estimating HLA/disease association with application to a study of aplastic anemia. Biometrics, 47(1):53–61, March 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532495.

Wacholder:1991:SED

[222] Sholom Wacholder, Mitchell Gail, and David Pee. Selecting an efficient design for assessing exposure-disease relationships in an assembled cohort. Biometrics, 47(1):63–76, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532496.

Lau:1991:DRS

[223] Tai-Shing Lau. On dependent repeated screening tests. Biometrics, 47(1):77–86, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532497.

Rosenbaum:1991:SAM

[224] Paul R. Rosenbaum. Sensitivity analysis for matched case-control studies. Biometrics, 47(1):87–100, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532498.

OQuigley:1991:PCT

[225] John O’Quigley and Fabienne Pessione. The problem of a covariate-time qualitative interaction in a survival study. Biometrics, 47(1):101–115, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532499.

OQuigley:1991:NTA

[226] John O’Quigley and Ross L. Prentice. Nonparametric tests of association between survival time and continuously measured covariates: The logitrank and associated procedures. Biometrics, 47(1):117–127, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532500.

Crager:1991:RAA

[227] Michael R. Crager and Martha A. Reitman. Running average analysis of clinical trial ambulatory blood pressure data. Biometrics, 47(1):129– 137, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532501.

Kodell:1991:CDR

[228] Ralph L. Kodell and James J. Chen. Characterization of dose-response relationships inferred by statistically significant trend tests. Biometrics, 47(1):139–146, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532502.

Lee:1991:TMS

[229] Jack C. Lee. Tests and model selection for the general growth curve model. Biometrics, 47(1):147–159, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532503.

Jones:1991:USL

[230] Richard H. Jones and Francis Boadi-Boateng. Unequally spaced longitudinal data with AR(1) serial correlation. Biometrics, 47(1):161–175, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532504.

Jorgensen:1991:ESL

[231] Merete Jørgensen, Niels Keiding, and Niels Erik Skakkebaek. Estimation of spermarche from longitudinal spermaturia data. Biometrics, 47(1): 177–193, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532505.

Meuwissen:1991:RSD

[232] T. H. E. Meuwissen. Reduction of selection differentials in finite populations with a nested full-half sib family structure. Biometrics, 47(1): 195–203, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532506.

Norell:1991:MIS

[233] Lennart Norell, Thorvaldur Arnason, and Kristinn Hugason. Multistage index selection in finite populations. Biometrics, 47(1):205–221, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532507.

Zimmerman:1991:RFA

[234] Dale L. Zimmerman and David A. Harville. A random field approach to the analysis of field-plot experiments and other spatial experiments. Biometrics, 47(1):223–239, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532508.

Knoebel:1991:BDA

[235] Bruce R. Knoebel and Harold E. Burkhart. A bivariate distribution approach to modeling forest diameter distributions at two points in time. Biometrics, 47(1):241–253, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532509.

Buckland:1991:QPM

[236] Stephen T. Buckland and Paul H. Garthwaite. Quantifying precision of mark-recapture estimates using the bootstrap and related methods. Biometrics, 47(1):255–268, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532510.

Quang:1991:NAS

[237] Pham Xuan Quang. A nonparametric approach to size-biased line transect sampling. Biometrics, 47(1):269–279, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532511.

Anderson:1991:NHW

[238] Keaven M. Anderson. A nonproportional hazards Weibull accelerated failure time regression model. Biometrics, 47(1):281–288, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532512.

Lin:1991:RCI

[239] D. Y. Lin and L. J. Wei. Repeated confidence intervals for a scale change in a sequential survival study. Biometrics, 47(1):289–294, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532513.

Kim:1991:ERP

[240] P. T. Kim, E. M. Carter, and J. J. Hubert. Estimating relative potency using prior information. Biometrics, 47(1):295–301, March 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532514.

Rocke:1991:RBM

[241] David M. Rocke. Robustness and balance in the mixed model. Biometrics, 47(1):303–309, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532515.

Solow:1991:DCH

[242] Andrew R. Solow and Woollcott Smith. Detecting cluster in a heterogeneous community sampled by quadrats. Biometrics, 47(1):311–317, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532516.

Greenland:1991:ESR

[243] Sander Greenland and Paul Holland. Estimating standardized risk differences from odds ratios. Biometrics, 47(1):319–322, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532517.

Grieve:1991:PPC

[244] A. P. Grieve, S. C. Choi, and P. A. Pepple. Predictive probability in clinical trials. Biometrics, 47(1):323–330, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532518.

Finney:1991:EAS

[245] David J. Finney. Ethical aspects of statistical practice. Biometrics, 47(1): 331–339, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532519.

Nijsse:1991:MCC

[246] M. Nijsse. Multiple correlation coefficient. Biometrics, 47(1):341, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532520.

Peritz:1991:OSV

[247] Eric Peritz and M. Zelen. One-sided vs two-sided significance tests. Biometrics, 47(1):342, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532521.

Robinson:1991:CA

[248] Laurence D. Robinson and Nicholas P. Jewell. Covariate adjustment. Biometrics, 47(1):342–343, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532522.

Kourouklis:1991:BRS

[249] S. Kourouklis. Book review: Statistical Methods and Concepts, by M. N. Das. Biometrics, 47(1):345–346, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532523.

Upton:1991:BRC

[250] G. J. G. Upton. Book review: Categorical Data Analysis, by A. Agresti. Biometrics, 47(1):346, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532524.

Agresti:1991:BRM

[251] A. Agresti. Book review: Multiway Data Analysis, by R. Coppi, S. Bolasco. Biometrics, 47(1):346–347, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532525.

Campbell:1991:BRI

[252] M. J. Campbell. Book review: An Introduction to Generalized Linear Models, by A. J. Dobson. Biometrics, 47(1):347, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532526.

Ogston:1991:BRA

[253] S. A. Ogston. Book review: Adequacy of Sample Size in Health Studies, by S. Lemeshow, D. W. Hosmer, J. Klar, S. K. Lwanga. Biometrics, 47(1):347–348, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532527.

Hunter:1991:BRE

[254] E. A. Hunter. Book review: Experimental Design in Biotechnology, by P. D. Haaland. Biometrics, 47(1):348, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532528.

Bailey:1991:BRS

[255] N. T. J. Bailey. Book reviews: Statistical Methods for Medical Investigations, by B. S. Everitt; Medical Statistics: a Commonsense Approach, by M. J. Campbell, D. Machin; Statistics for Health Professionals, by S. Shott. Biometrics, 47(1):348–349, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532529.

Kemp:1991:BRI

[256] A. W. Kemp. Book review: An Introduction to Mathematical Physiology and Biology, by J. Mazumdar. Biometrics, 47(1):349–350, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532530.

Kemp:1991:BREa

[257] A. W. Kemp. Book review: Empirical Model Building, by J. R. Thompson. Biometrics, 47(1):350–351, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532531.

North:1991:BRD

[258] P. M. North. Book review: Dynamics of Biological Invasions, by R. Hengeveld. Biometrics, 47(1):351–352, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532532.

Woiwod:1991:BRB

[259] I. P. Woiwod. Book review: Biological Invasions: a Global Perspective (SCOPE 37), by J. A. Drake, H. A. Mooney, F. Di Castri, R. H. Groves, F. J. Kruger, M. Rejmánek, M. Williamson. Biometrics, 47(1): 352, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532533.

Kemp:1991:BREb

[260] C. D. Kemp. Book review: Elements of Statistical Computing: Numerical Computation, by R. A. Thisted. Biometrics, 47(1):352–353, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532534.

Buckland:1991:BRA

[261] S. T. Buckland. Book review: Analysis of Wildlife Radio-Tracking Data, by G. C. White, R. A. Garrott. Biometrics, 47(1):353–354, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532535.

Dagpunar:1991:BRI

[262] J. S. Dagpunar. Book review: Introduction to Probability and Its Applications, by R. L. Scheaffer. Biometrics, 47(1):354–355, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532536.

Anonymous:1991:BRPa

[263] Anonymous. Book review: Principles of Research Data Audit, by A. E. Shamoo. Biometrics, 47(1):355, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532537.

Kemp:1991:BRP

[264] C. D. Kemp and A. W. Kemp. Book reviews: Probability and Statistics for Engineers, by R. L. Scheaffer, J. T. McClave; Understanding Statistics, by L. Ott, W. Mendenhall; Statistics for Management and Economics: a Systematic Approach, by G. Keller, B. Warrack, H. Bartel; Statistics: a Tool for Social Research, by J. F. Healey. Biometrics, 47(1): 355–356, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532538.

Anonymous:1991:CAN

[265] Anonymous. Corrigendum: Applied Nonparametric Statistical Methods. Biometrics, 47(1):356, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532540. See [88].

Anonymous:1991:BRE

[266] Anonymous. Book review: Experimental Design and Statistical Evaluation of Limiting Dilution Assays, by L. W. G. Strijbosch. Biometrics, 47 (1):356, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532539.

Anonymous:1991:CAI

[267] Anonymous. Corrections: Approximate interval estimation of the difference in binomial parameters: Correction for skewness and extension to multiple tables. Biometrics, 47(1):357, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532543. See [113, 555].

Anonymous:1991:CEC

[268] Anonymous. Corrections: Estimation and comparison of changes in the presence of informative right censoring by modelling the censoring process. Biometrics, 47(1):357, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532541. See [2].

Anonymous:1991:CTS

[269] Anonymous. Corrections: A test of spatial randomness on small scales, combining infromation from mapped locations within several quadrats. Biometrics, 47(1):357, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532542. See [75].

Gart:1991:CAI

[270] J. J. Gart and J. Nam. Correction to: “Approximate interval estimation of the difference in binomial parameters: correction for skewness and extension to multiple tables” [Biometrics 46 (1990), no. 3, 637–643]. Biometrics, 47(1):357, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [113].

Wu:1991:CEC

[271] M. C. Wu and R. J. Carroll. Correction to: “Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process” [Biometrics 44 (1988), no. 1, 175–188; MR0931633 (89d:62109)]. Biometrics, 47(1):357, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [2].

Anonymous:1991:PPBa

[272] Anonymous. Papers to be published in biometrics. Biometrics, 47(1): 359, March 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532544.

Anonymous:1991:FMb

[273] Anonymous. Front matter. Biometrics, 47(2):i–iv, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532130.

Suissa:1991:MPTa

[274] Samy Suissa and Jonathan J. Shuster. The 2 × 2 matched-pairs trial: exact unconditional design and analysis. Biometrics, 47(2):361–372, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Suissa:1991:MPTb

[275] Samy Suissa and Jonathan J. Shuster. The 2×2 matched-pairs trial: Exact unconditional design and analysis. Biometrics, 47(2):361–372, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532131.

Pigeot:1991:JEC

[276] Iris Pigeot. A jackknife estimator of a combined odds ratio. Biometrics, 47(2):373–381, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532132.

Moore:1991:REV

[277] D. F. Moore and A. Tsiatis. Robust estimation of the variance in moment methods for extra-binomial and extra-Poisson variation. Biometrics, 47(2):383–401, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532133.

Brooks:1991:MSB

[278] Rodney J. Brooks, William H. James, and Elmer Gray. Modelling subbinomial variation in the frequency of sex combinations in litters of pigs. Biometrics, 47(2):403–417, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532134.

Hubbard:1991:RSN

[279] Dean J. Hubbard and O. Brian Allen. Robustness of the SPRT for a negative binomial to misspecification of the dispersion parameter. Biometrics, 47(2):419–427, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532135.

Brown:1991:AVC

[280] David Brown and Neal Alexander. The analysis of the variance and covariance of products. Biometrics, 47(2):429–444, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532136.

Cook:1991:TST

[281] J. A. Cook and J. F. Lawless. Two-sample tests with multinomial or grouped failure time data. Biometrics, 47(2):445–459, June 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532137.

McGilchrist:1991:RFS

[282] C. A. McGilchrist and C. W. Aisbett. Regression with frailty in survival analysis. Biometrics, 47(2):461–466, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532138.

Clayton:1991:MCM

[283] David G. Clayton. A Monte Carlo method for Bayesian inference in frailty models. Biometrics, 47(2):467–485, June 1991. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532139.

Hirji:1991:CEM

[284] Karim F. Hirji. A comparison of exact, mid-p, and score tests for matched case-control studies. Biometrics, 47(2):487–496, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532140.

Scott:1991:FLR

[285] A. J. Scott and C. J. Wild. Fitting logistic regression models in stratified case-control studies. Biometrics, 47(2):497–510, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532141.

Lehmacher:1991:PTS

[286] Walter Lehmacher, Gernot Wassmer, and Peter Reitmeir. Procedures for two-sample comparisons with multiple endpoints controlling the experimentwise error rate. Biometrics, 47(2):511–521, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532142.

Knoke:1991:NAC

[287] James D. Knoke. Nonparametric analysis of covariance for comparing change in randomized studies with baseline values subject to error. Biometrics, 47(2):523–533, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532143.

Reade-Christopher:1991:EEM

[288] Susan J. Reade-Christopher and Lawrence L. Kupper. Effects of exposure misclassification on regression analyses of epidemiologic follow-up study data. Biometrics, 47(2):535–548, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532144.

Heitjan:1991:RBG

[289] Daniel F. Heitjan. Regression with bivariate grouped data. Biometrics, 47(2):549–562, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532145.

Schwenke:1991:CPE

[290] James R. Schwenke and George A. Milliken. On the calibration problem extended to nonlinear models. Biometrics, 47(2):563–574, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532146.

Stein:1991:UKC

[291] A. Stein and L. C. A. Corsten. Universal kriging and cokriging as a regression procedure. Biometrics, 47(2):575–587, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532147.

Best:1991:DCT

[292] D. J. Best and J. C. W. Rayner. Disease clustering in time. Biometrics, 47(2):589–593, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532148.

Gruger:1991:VIB

[293] Jens Grüger, Richard Kay, and Martin Schumacher. The validity of inferences based on incomplete observations in disease state models. Biometrics, 47(2):595–605, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532149.

McLaren:1991:DTC

[294] C. E. McLaren, M. Wagstaff, G. M. Brittenham, and A. Jacobs. Detection of two-component mixtures of lognormal distributions in grouped, doubly truncated data: Analysis of red blood cell volume distributions. Biometrics, 47(2):607–622, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532150.

Palmer:1991:ANM

[295] M. J. Palmer, B. F. Phillips, and G. T. Smith. Application of nonlinear models with random coefficients to growth data. Biometrics, 47(2):623– 635, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532151.

Lundbye-Christensen:1991:MGC

[296] Søren Lundbye-Christensen. A multivariate growth curve model for pregnancy. Biometrics, 47(2):637–657, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532152.

Liski:1991:DIM

[297] Erkki P. Liski. Detecting influential measurements in a growth curves model. Biometrics, 47(2):659–668, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532153.

Finkelstein:1991:MED

[298] Dianne M. Finkelstein. Modeling the effect of dose on the lifetime tumor rate from an animal carcinogenicity experiment. Biometrics, 47(2):669– 680, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532154.

Dinse:1991:CRD

[299] Gregg E. Dinse. Constant risk differences in the analysis of animal tumorigenicity data. Biometrics, 47(2):681–700, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532155.

Turnock:1991:ERM

[300] Benjamin J. Turnock and Terrance J. Quinn, II. The effect of responsive movement on abundance estimation using line transect sampling. Biometrics, 47(2):701–715, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532156.

Pennington:1991:OSS

[301] Michael Pennington and Jon H. Vølstad. Optimum size of sampling unit for estimating the density of marine populations. Biometrics, 47(2):717– 723, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532157.

Huggins:1991:SPA

[302] R. M. Huggins. Some practical aspects of a conditional likelihood approach to capture experiments. Biometrics, 47(2):725–732, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532158.

Edelman:1991:FDF

[303] David Edelman. The five-degree-of-freedom rule of thumb for fixed-width confidence intervals for a normal mean. Biometrics, 47(2):733–739, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532159.

Huang:1991:NDT

[304] Xiang-Ning Huang and Bai-Bing Li. A new discriminant technique: Bayes-fisher discrimination. Biometrics, 47(2):741–744, June 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532160.

Li:1991:CTD

[305] Zhaohai Li and Nancy L. Geller. On the choice of times for data analysis in group sequential clinical trials. Biometrics, 47(2):745–750, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532161.

Zelterman:1991:THR

[306] Daniel Zelterman and Chap T. Le. Tests of homogeneity for the relative risk in multiply-matched case-control studies. Biometrics, 47(2):751– 755, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532162.

Hall:1991:TGB

[307] Peter Hall and Susan R. Wilson. Two guidelines for bootstrap hypothesis testing. Biometrics, 47(2):757–762, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532163. See comment [775].

Follmann:1991:ESS

[308] Dean A. Follmann. The effect of screening on some pretest-posttest test variances. Biometrics, 47(2):763–771, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532164.

Greenland:1991:RMS

[309] Sander Greenland and Leslie A. Kalish. Reducing mean squared error in the analysis of stratified epidemiologic studies. Biometrics, 47(2):773– 776, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532165.

Seigel:1991:CBP

[310] Daniel G. Seigel, Marvin J. Podgor, and Marvin Zelen. Compliance, bias, and power in clinical trials. Biometrics, 47(2):777–779, June 1991.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532166.

Graubard:1991:CCS

[311] Barry I. Graubard, Mitchell H. Gail, and Thomas R. Fears. Case-control studies with cluster sampling. Biometrics, 47(2):779–780, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532167.

deOliviera:1991:BRE

[312] J. Tiago de Oliviera. Book review: Encyclopedia of Statistical Sciences-Supplement Volume, by S. Kotz, N. L. Johnson, C. B. Read. Biometrics, 47(2):781–782, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532168.

Kemp:1991:BRD

[313] A. W. Kemp. Book review: A Dictionary of Statistical Terms, by F. H. C. Marriott. Biometrics, 47(2):782, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532169.

Holgate:1991:BRR

[314] P. Holgate. Book review: Reconstructing the Past: Parsimony, Evolution, and Inference, by E. Sober. Biometrics, 47(2):782–783, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532170.

Oger:1991:BRP

[315] R. Oger. Book review: Probabilités, Analyse des Données et Statistique, by G. Saporta. Biometrics, 47(2):783, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532171.

Kocherlakota:1991:BRC

[316] S. Kocherlakota. Book review: Continuous Bivariate Distributions, Emphasising Applications, by T. P. Hutchinson, C. D. Lai. Biometrics, 47(2): 783–784, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532172.

Freeman:1991:BRD

[317] G. H. Freeman. Book review: Designing Experiments and Analysing Data: a Model Comparison Perspective, by S. E. Maxwell, H. D. Delaney.

Biometrics, 47(2):784–785, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532173.

Brown:1991:BRG

[318] R. A. Brown. Book review: Generalized Additive Models, by T. J. Hastie, R. J. Tibshirani. Biometrics, 47(2):785–786, June 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532174.

Bailey:1991:BRC

[319] R. A. Bailey. Book review: Computation for the Analysis of Designed Experiments, by R. M. Heiberger. Biometrics, 47(2):786–787, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532175.

Laycock:1991:BRD

[320] P. J. Laycock. Book review: Design and Analysis of Cross-Over Trials, by B. Jones, M. G. Kenward. Biometrics, 47(2):787, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532176.

Hand:1991:BRG

[321] D. J. Hand. Book review: Graphical Models in Applied Multivariate Statistics, by J. Whittaker. Biometrics, 47(2):787–788, June 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532177.

Gentle:1991:BRF

[322] J. E. Gentle. Book review: Finding Groups in Data: an Introduction to Cluster Analysis, by L. Kaufman, P. J. Rousseuw. Biometrics, 47(2): 788, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532178.

Kemp:1991:BRS

[323] A. W. Kemp. Book review: Smooth Tests of Goodness of Fit, by J. C. W. Rayner, D. J. Best. Biometrics, 47(2):788–789, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532179.

Dentine:1991:BRS

[324] M. R. Dentine. Book review: Statistical Genetics, by P. Narain. Biometrics, 47(2):789, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532180.

Curnow:1991:BRA

[325] R. N. Curnow. Book review: Advances in Statistical Methods for Genetic Improvement of Livestock, by D. Gianola, K. Hammond. Biometrics, 47(2):789–790, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532181.

Abakuks:1991:BRB

[326] A. Abakuks. Book review: Biological Delay Systems: Linear Stability Theory, by N. MacDonald. Biometrics, 47(2):790, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532182.

Cormack:1991:BRE

[327] R. M. Cormack. Book review: Elementary Survey Sampling, by R. L. Scheaffer, W. Mendenhall, L. Ott. Biometrics, 47(2):790–791, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532183.

Anonymous:1991:BRAa

[328] Anonymous. Book reviews: Asymptotic Techniques for Use in Statistics, by O. E. Barndorff-Nielson, D. R. Cox; Continued Fractions in Statistical Applications, by K. O. Bowman, L. R. Shenton. Biometrics, 47(2): 791, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532184.

Anonymous:1991:BRAb

[329] Anonymous. Book review: Applied Regression Analysis for Business and Economics, by T. E. Dielman. Biometrics, 47(2):791–792, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532185.

Anonymous:1991:PPBb

[330] Anonymous. Papers to be published in biometrics. Biometrics, 47(2): 793, June 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532186.

Anonymous:1991:FMc

[331] Anonymous. Front matter. Biometrics, 47(3):i–iv, September 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532639.

Walter:1991:CSP

[332] S. D. Walter and R. J. Cook. A comparison of several point estimators of the odds ratio in a single 2 × 2 contingency table. Biometrics, 47(3):795– 811, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532640.

Schaalje:1991:ARM

[333] Bruce Schaalje, Ji Zhang, Sastry G. Pantula, and Kenneth H. Pollock. Analysis of repeated-measurements data from randomized block experiments. Biometrics, 47(3):813–824, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532641.

Prentice:1991:EEP

[334] Ross L. Prentice and Lue Ping Zhao. Estimating equations for parameters in means and covariances of multivariate discrete and continuous responses. Biometrics, 47(3):825–839, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532642.

Chen:1991:DPC

[335] Chen-Hsin Chen and P. C. Wang. Diagnostic plots in Cox’s regression model. Biometrics, 47(3):841–850, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532643.

Spiegelman:1991:CES

[336] Donna Spiegelman and Robert Gray. Cost-efficient study designs for binary response data with Gaussian covariate measurement error. Biometrics, 47(3):851–869, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2532644. See correction [433].

Dixon:1991:BSA

[337] Dennis O. Dixon and Richard Simon. Bayesian subset analysis. Biometrics, 47(3):871–881, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2532645. See corrections [852].

Oja:1991:FMM

[338] Hannu Oja, Markku Koiranen, and Paula Rantakallio. Fitting mixture models to birth weight data: a case study. Biometrics, 47(3):883–897, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532646.

Begg:1991:MIH

[339] Colin B. Begg and Louise Pilote. A model for incorporating historical controls into a meta-analysis. Biometrics, 47(3):899–906, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532647.

Beam:1991:SMC

[340] Craig A. Beam and H. Samuel Wieand. A statistical method for the comparison of a discrete diagnostic test with several continuous diagnostic tests. Biometrics, 47(3):907–919, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532648.

Albert:1991:DPS

[341] Paul S. Albert and C. Hendricks Brown. The design of a panel study under an alternating Poisson process assumption. Biometrics, 47(3):921– 932, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532649.

Aalen:1991:MIR

[342] Odd O. Aalen. Modelling the influence of risk factors on familial aggregation of disease. Biometrics, 47(3):933–945, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532650.

Bacchetti:1991:NEI

[343] Peter Bacchetti and Nicholas P. Jewell. Nonparametric estimation of the incubation period of AIDS based on a prevalent cohort with unknown infection times. Biometrics, 47(3):947–960, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532651.

Addy:1991:GSM

[344] Cheryl L. Addy, Ira M. Longini, Jr., and Michael Haber. A generalized stochastic model for the analysis of infectious disease final size data. Biometrics, 47(3):961–974, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532652.

Self:1991:AWL

[345] Steven G. Self. An adaptive weighted log-rank test with application to cancer prevention and screening trials. Biometrics, 47(3):975–986, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532653.

Kimmel:1991:NES

[346] Marek Kimmel and Betty J. Flehinger. Nonparametric estimation of the size-metastasis relationship in solid cancers. Biometrics, 47(3):987–1004, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532654.

Berkey:1991:MAB

[347] Catherine S. Berkey, Nan M. Laird, Isabelle Valadian, and Jane Gardner. Modelling adolescent blood pressure patterns and their prediction of adult pressures. Biometrics, 47(3):1005–1018, September 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532655.

Hingley:1991:CSB

[348] Peter J. Hingley. A clonal selection based timecourse model for antibody responses to killed vaccine, with applications to foot and mouth disease. Biometrics, 47(3):1019–1047, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532656.

Chen:1991:ATR

[349] James J. Chen, Ralph L. Kodell, Richard B. Howe, and David W. Gaylor. Analysis of trinomial responses from reproductive and developmental toxicity experiments. Biometrics, 47(3):1049–1058, September 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532657.

Ruberg:1991:PDS

[350] Stephen J. Ruberg and James W. Stegeman. Pooling data for stability studies: Testing the equality of batch degradation slopes. Biometrics, 47(3):1059–1069, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532658.

Chow:1991:EDS

[351] Shein-Chung Chow and Jun Shao. Estimating drug shelf-life with random batches. Biometrics, 47(3):1071–1079, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532659.

Yip:1991:MEE

[352] Paul Yip. A martingale estimating equation for a capture–recapture experiment in discrete time. Biometrics, 47(3):1081–1088, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532660.

Quang:1991:LTM

[353] Pham Xuan Quang and Richard B. Lanctot. A line transect model for aerial surveys. Biometrics, 47(3):1089–1102, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532661.

Thompson:1991:ACS

[354] Steven K. Thompson. Adaptive cluster sampling: Designs with primary and secondary units. Biometrics, 47(3):1103–1115, September 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532662.

Munholland:1991:SMM

[355] Patricia L. Munholland and John D. Kalbfleisch. A semi-Markov model for insect life history data. Biometrics, 47(3):1117–1126, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532663.

Renaud:1991:NSE

[356] J-C. Renaud and J. R. Morton. A numerical solution to the equilibria of the two-locus two-allele selection model. Biometrics, 47(3):1127–1133, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532664.

Tsujitani:1991:RPL

[357] Masaaki Tsujitani and Gary G. Koch. Residual plots for log odds ratio regression models. Biometrics, 47(3):1135–1141, September 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532665.

Mendell:1991:LRT

[358] Nancy R. Mendell, Henry C. Thode, Jr., and Stephen J. Finch. The likelihood ratio test for the two-component normal mixture problem:

Power and sample size analysis. Biometrics, 47(3):1143–1148, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532666. See correction [511].

Brunner:1991:NES

[359] E. Brunner. A nonparametric estimator of the shift effect for repeated observations. Biometrics, 47(3):1149–1153, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532667.

Diggle:1991:SOA

[360] P. J. Diggle and A. G. Chetwynd. Second-order analysis of spatial clustering for inhomogeneous populations. Biometrics, 47(3):1155–1163, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532668.

Sato:1991:EMC

[361] Tosiya Sato. Elementary methods of cohort analysis with several exposure levels. Biometrics, 47(3):1165–1170, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532669.

Ekholm:1991:AVM

[362] Anders Ekholm, Mark A. Espeland, and Siu L. Hui. Algorithms versus models for analyzing data that contain misclassification errors. Biometrics, 47(3):1171–1182, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532670.

Rayner:1991:THT

[363] Arthur A. Rayner, Christopher Bingham, and Stephen E. Fienberg. Testing hierarchical treatment components in analysis of covariance. Biometrics, 47(3):1183–1191, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532671.

Hirsch:1991:VS

[364] Robert P. Hirsch. Validation samples. Biometrics, 47(3):1193–1194, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532672.

Bross:1991:FSM

[365] Irwin D. Bross. Fraudulent statistical methods. Biometrics, 47(3):1194– 1196, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532673.

Huhn:1991:PSG

[366] Manfred Hühn and Raja Nassar. Phenotypic stability of genotypes over environments: On tests of significance for two nonparametric measures. Biometrics, 47(3):1196–1197, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532674.

Tee:1991:MEF

[367] Garry J. Tee. Mixing English and French. Biometrics, 47(3):1197, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532675.

Edwards:1991:BRS

[368] A. W. F. Edwards. Book review: Statistical Inference and Analysis: Selected Correspondence of R. A. Fisher, by J. H. Bennett. Biometrics, 47(3):1199–1200, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532676.

Pearce:1991:BRH

[369] S. C. Pearce. Book review: A History of Probability and Statistics and Their Applications before 1750, by A. Hald. Biometrics, 47(3):1200, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532677.

Kenward:1991:BRS

[370] M. G. Kenward. Book review: The Statistical Analysis of Discrete Data, by T. J. Santner, D. E. Duffy. Biometrics, 47(3):1200–1201, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532678.

Calinski:1991:BRM

[371] T. Caliński and P. Krajewski. Book review: Multiway Contingency Tables Analysis for the Social Sciences, by T. D. Wickens. Biometrics, 47 (3):1201–1202, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532679.

Liese:1991:BRN

[372] F. Liese. Book review: Nonlinear Estimation, by G. J. S. Ross. Biometrics, 47(3):1202–1203, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532680.

Hasted:1991:BRN

[373] A. Hasted. Book review: Nonlinear Regression, by G. A. F. Seber, C. J. Wild. Biometrics, 47(3):1203, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532681.

Lwin:1991:BRM

[374] T. Lwin. Book review: Multivariate Calibration, by H. Martens, T. Naes. Biometrics, 47(3):1203–1205, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532682.

Feytmans:1991:BRG

[375] E. Feytmans. Book review: Genetic Data Analysis: Methods for Discrete Population Genetic Data, by B. S. Weir. Biometrics, 47(3):1205, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532683.

Feytmans:1991:BRF

[376] E. Feytmans. Book review: Flow Cytogenetics, by J. W. Gray. Biometrics, 47(3):1205–1206, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532684.

Anonymous:1991:BRB

[377] Anonymous. Book review: Biostatistics: a Foundation for Analysis in the Health Sciences, 5th Edition, by W. W. Daniel. Biometrics, 47(3): 1206, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532686.

Anonymous:1991:BRCa

[378] Anonymous. Book review: Computers in Mathematics, by D. V. Chudnovsky, R. D. Jenks. Biometrics, 47(3):1206, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532687.

Anonymous:1991:BRS

[379] Anonymous. Book review: Statistical Methods, Experimental Design, and Scientific Inference, by R. A. Fisher. Biometrics, 47(3):1206, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532685.

Anonymous:1991:PPBc

[380] Anonymous. Papers to be published in biometrics. Biometrics, 47(3): 1207, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532688.

Fedder:1991:BSF

[381] Reznick Fedder. The Biometric Society financial statements and independent auditors’ report January 31, 1991. Biometrics, 47(3):1209–1212, September 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532689.

Anonymous:1991:FMd

[382] Anonymous. Front matter. Biometrics, 47(4):i–iv, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532380.

Anonymous:1991:VI

[383] Anonymous. Volume information. Biometrics, 47(4):vii–xxvi, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532379.

Rubin:1991:PIM

[384] Donald B. Rubin. Practical implications of modes of statistical inference for causal effects and the critical role of the assignment mechanism. Biometrics, 47(4):1213–1234, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532381.

Racine-Poon:1991:ERP

[385] A. Racine-Poon, C. Weihs, and A. F. M. Smith. Estimation of relative potency with sequential dilution errors in radioimmunoassay. Biometrics, 47(4):1235–1246, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532382.

Drescher:1991:ARE

[386] Karsten Drescher and Walter Schill. Attributable risk estimation from case-control data via logistic regression. Biometrics, 47(4):1247–1256, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532383.

Tosteson:1991:LRC

[387] Tor D. Tosteson, Bernard Rosner, and Susan Redline. Logistic regression for clustered binary data in proband studies with application to familial aggregation of sleep disorders. Biometrics, 47(4):1257–1265, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532384.

leCessie:1991:GFT

[388] S. le Cessie and J. C. van Houwelingen. A goodness-of-fit test for binary regression models, based on smoothing methods. Biometrics, 47(4):1267– 1282, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532385.

Gentleman:1991:LFL

[389] R. Gentleman and J. Crowley. Local full likelihood estimation for the proportional hazards model. Biometrics, 47(4):1283–1296, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532386.

Wei:1991:AMI

[390] Greg C. G. Wei and Martin A. Tanner. Applications of multiple imputation to the analysis of censored regression data. Biometrics, 47 (4):1297–1309, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532387.

Vollset:1991:EEA

[391] Stein E. Vollset, Karim F. Hirji, and Abdelmonem A. Afifi. Evaluation of exact and asymptotic interval estimators in logistic analysis of matched case-control studies. Biometrics, 47(4):1311–1325, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532388.

Baker:1991:URO

[392] Stuart G. Baker, Laurence S. Freedman, and M. K. B. Parmar. Using replicate observations in observer agreement studies with binary assessments. Biometrics, 47(4):1327–1338, December 1991. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532389. See corrections [1452].

Akritas:1991:TIB

[393] Michael G. Akritas and Clifford C. Clogg. Tests of independence for bivariate data with random censoring: a contingency-table approach. Biometrics, 47(4):1339–1354, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532390.

Palta:1991:ALD

[394] Mari Palta and Tzy-Jyun Yao. Analysis of longitudinal data with unmeasured confounders. Biometrics, 47(4):1355–1369, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532391.

Albert:1991:TSM

[395] Paul S. Albert. A two-state Markov mixture model for a time series of epileptic seizure counts. Biometrics, 47(4):1371–1381, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532392.

Rochon:1991:SSC

[396] James Rochon. Sample size calculations for two-group repeated-measures experiments. Biometrics, 47(4):1383–1398, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532393.

Lin:1991:ESI

[397] D. Y. Lin, L. J. Wei, and D. L. DeMets. Exact statistical inference for group sequential trials. Biometrics, 47(4):1399–1408, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532394.

Pla:1991:DSB

[398] Laura Pla. Determining stratum boundaries with multivariate real data. Biometrics, 47(4):1409–1422, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532395.

Ridout:1991:MDP

[399] Martin S. Ridout and Byron J. T. Morgan. Modelling digit preference in fecundability studies. Biometrics, 47(4):1423–1433, December 1991.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532396.

Edmondson:1991:ARS

[400] R. N. Edmondson. Agricultural response surface experiments based on four-level factorial designs. Biometrics, 47(4):1435–1448, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532397.

Cullis:1991:SAF

[401] B. R. Cullis and A. C. Gleeson. Spatial analysis of field experiments — an extension to two dimensions. Biometrics, 47(4):1449–1460, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532398.

Federer:1991:CED

[402] Walter T. Federer and Kaye E. Basford. Competing effects designs and models for two-dimensional field arrangements. Biometrics, 47(4):1461– 1472, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532399.

Baird:1991:EEV

[403] David Baird and Roger Mead. The empirical efficiency and validity of two neighbour models. Biometrics, 47(4):1473–1487, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532400.

Boos:1991:MMC

[404] Dennis D. Boos and Cavell Brownie. Mixture models for continuous data in dose-response studies when some animals are unaffected by treatment. Biometrics, 47(4):1489–1504, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532401.

Perry:1991:NIA

[405] Joe N. Perry and Mark Hewitt. A new index of aggregation for animal counts. Biometrics, 47(4):1505–1518, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532402.

James:1991:EBG

[406] Ian R. James. Estimation of von Bertalanffy growth curve parameters from recapture data. Biometrics, 47(4):1519–1530, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532403.

Udevitz:1991:CRE

[407] Mark S. Udevitz and Kenneth H. Pollock. Change-in-ratio estimators for populations with more than two subclasses. Biometrics, 47(4):1531–1546, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532404.

Vivaas:1991:ECA

[408] Helga Jónsdóttir Vivås and Steinar Engen. Estimation of circadian activity pattern when the beginning and end of active periods are not always observable. Biometrics, 47(4):1547–1556, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532405.

Koziol:1991:MSC

[409] James A. Koziol. On maximally selected chi-square statistics. Biometrics, 47(4):1557–1561, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532406.

Langholz:1991:ECS

[410] Bryan Langholz and Duncan C. Thomas. Efficiency of cohort sampling designs: Some surprising results. Biometrics, 47(4):1563–1571, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532407.

Lee:1991:NCA

[411] J. Jack Lee. A note on the conditional approach to interval estimation in the calibration problem. Biometrics, 47(4):1573–1580, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532408.

Minkin:1991:SMV

[412] Salomon Minkin. A statistical model for in-vitro assessment of patient sensitivity to cytotoxic drugs. Biometrics, 47(4):1581–1591, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532409.

Galbraith:1991:IRC

[413] Jane I. Galbraith, Scott L. Zeger, Kung-Yee Liang, and Paul S. Albert. The interpretation of a regression coefficient. Biometrics, 47(4):1593– 1596, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532410.

Grieve:1991:CIS

[414] A. P. Grieve and Stuart L. Beal. Confidence intervals and sample sizes. Biometrics, 47(4):1597–1603, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532411.

Nelder:1991:GLM

[415] J. A. Nelder, D. Ruppert, N. Cressie, and R. J. Carroll. Generalized linear models for enzyme-kinetic data. Biometrics, 47(4):1605–1615, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532412.

Ridout:1991:TRD

[416] Martin S. Ridout and Peter J. Diggle. Testing for random dropouts in repeated measurement data. Biometrics, 47(4):1617–1621, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532413.

Pennington:1991:TRL

[417] Michael Pennington, R. A. Myers, and P. Pepin. On testing the robustness of lognormal-based estimators. Biometrics, 47(4):1623–1624, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532414.

Schneiderman:1991:FSM

[418] Marvin A. Schneiderman and Irwin D. Bross. Fraudulent statistical methods. Biometrics, 47(4):1624–1628, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532415.

Silvapulle:1991:TTV

[419] Mervyn J. Silvapulle. On testing for threshold values. Biometrics, 47 (4):1628–1629, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532416.

Loynes:1991:BRR

[420] R. M. Loynes. Book review: Regression Analysis: Theory, Methods, and Applications, by A. K. Sen, M. S. Srivastava. Biometrics, 47(4):1631– 1632, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532417.

Raab:1991:BRS

[421] G. M. Raab. Book review: Subset Selection in Regression, by A. J. Miller. Biometrics, 47(4):1632, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532418.

Scott:1991:BRA

[422] A. J. Scott. Book review: Applied Logistic Regression, by D. W. Hosmer, S. Lemeshow. Biometrics, 47(4):1632–1633, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532419.

Gentle:1991:BRC

[423] J. E. Gentle. Book review: Computer-Aided Multivariate Analysis, by A. A. Afifi, V. Clark. Biometrics, 47(4):1633–1634, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532420.

Rosenberger:1991:BRA

[424] J. L. Rosenberger. Book review: Analysis of Repeated Measures, by M. J. Crowder, D. J. Hand. Biometrics, 47(4):1634, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532421.

Streitberg:1991:BRA

[425] B. Streitberg. Book review: Applied Nonparametric Statistics, by W. W. Daniel. Biometrics, 47(4):1634–1635, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532422.

Bjornsson:1991:BRP

[426] H. Björnsson. Book review: Practical Statistics for Field Biology, by J. Fowler, L. Cohen. Biometrics, 47(4):1635–1636, December 1991. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532423.

Wright:1991:BRP

[427] H. L. Wright. Book review: Process Modeling of Forest Growth Responses to Environmental Stress, by R. K. Dixon, R. S. Meldahl, G. A. Ruark, W. G. Warren. Biometrics, 47(4):1636–1637, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532424.

Talbot:1991:BRF

[428] M. Talbot. Book review: The Frontiers of Expert Systems and Artificial Intelligence, by E. J. Dudewicz, P. R. Nelson, A. Öztürk, E. C. van der Meulen. Biometrics, 47(4):1637, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532425.

Kilpatrick:1991:BRL

[429] S. J. Kilpatrick. Book review: Life-Style and Mortality: a Large-Scale Census-Based Cohort Study in Japan, by T. Hirayama. Biometrics, 47 (4):1637–1638, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532426.

Anonymous:1991:BRPb

[430] Anonymous. Book review: Probability and Statistics for Engineering and the Sciences, by J. L. Devore. Biometrics, 47(4):1638, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532427.

Anonymous:1991:BRCb

[431] Anonymous. Book review: The Chronological Annotated Bibliography of Order Statistics, Volume III: 1960–1961, by H. L. Harter. Biometrics, 47(4):1638–1639, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532428.

Anonymous:1991:BRW

[432] Anonymous. Book reviews: World Health Statistics Annual 1990, by World Health Organization; World Health Statistics Quarterly, by World Health Organization. Biometrics, 47(4):1639, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532429.

Anonymous:1991:CCE

[433] Anonymous. Correction: Cost-efficient study designs for binary response data with Gaussian covariate measurement error. Biometrics, 47(4):1641, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532430. See [336].

Anonymous:1991:PPBd

[434] Anonymous. Papers to be published in biometrics. Biometrics, 47(4): 1643, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532431.

Anonymous:1992:FMa

[435] Anonymous. Front matter. Biometrics, 48(1):i–iv, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532733.

Vonesh:1992:MEN

[436] Edward F. Vonesh and Randy L. Carter. Mixed-effects nonlinear regression for unbalanced repeated measures. Biometrics, 48(1):1–17, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532734.

Paik:1992:PVF

[437] Myunghee C. Paik. Parametric variance function estimation for nonnormal repeated measurement data. Biometrics, 48(1):19–30, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532735.

Self:1992:PCL

[438] Steven G. Self, Robert H. Mauritsen, and Jill Ohara. Power calculations for likelihood ratio tests in generalized linear models. Biometrics, 48(1): 31–39, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532736.

Hughes:1992:ISR

[439] Michael D. Hughes, Laurence S. Freedman, and Stuart J. Pocock. The impact of stopping rules on heterogeneity of results in overviews of clinical trials. Biometrics, 48(1):41–53, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532737.

Storer:1992:CPI

[440] Barry E. Storer. A class of Phase II designs with three possible outcomes. Biometrics, 48(1):55–60, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532738.

Boos:1992:RBM

[441] Dennis D. Boos and Cavell Brownie. A rank-based mixed model approach to multisite clinical trials. Biometrics, 48(1):61–72, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532739. See corrections [854].

Lausen:1992:MSR

[442] Berthold Lausen and Martin Schumacher. Maximally selected rank statistics. Biometrics, 48(1):73–85, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532740.

Sposto:1992:CTD

[443] Richard Sposto, Harland N. Sather, and Sherryl A. Baker. A comparison of tests of the difference in the proportion of patients who are cured. Biometrics, 48(1):87–99, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532741.

Heller:1992:PCS

[444] Glenn Heller and Jeffrey S. Simonoff. Prediction in censored survival data: a comparison of the proportional hazards and linear regression models. Biometrics, 48(1):101–115, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532742.

Rampey:1992:DTM

[445] Alvin H. Rampey, Jr., Ira M. Longini, Jr., Michael Haber, and Arnold S. Monto. A discrete-time model for the statistical analysis of infectious disease incidence data. Biometrics, 48(1):117–128, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532743.

Wang:1992:ERA

[446] Mei-Cheng Wang and Lai-Chu See. N-estimation from retrospectively ascertained events with applications to AIDS. Biometrics, 48(1):129– 141, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532744.

Greenwood:1992:EBP

[447] Sina R. Greenwood and G. A. F. Seber. Estimating blood phenotype probabilities and their products. Biometrics, 48(1):143–154, March 1992.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532745.

Ryan:1992:RRM

[448] D. A. J. Ryan, J. J. Hubert, E. M. Carter, J. B. Sprague, and J. Parrott. A reduced-rank multivariate regression approach to aquatic joint toxicity experiments. Biometrics, 48(1):155–162, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532746.

Ryan:1992:QRA

[449] Louise Ryan. Quantitative risk assessment for developmental toxicity. Biometrics, 48(1):163–174, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532747.

Lim:1992:EVP

[450] Lynette L.-Y. Lim and John Whitehead. Estimating the ventilationperfusion distribution: an ill-posed integral equation problem. Biometrics, 48(1):175–187, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532748.

Li:1992:ESM

[451] H. G. Li, H. T. Schreuder, D. D. Van Hooser, and G. E. Brink. Estimating strata means in double sampling with corrections based on second-phase sampling. Biometrics, 48(1):189–199, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532749.

Chao:1992:EPS

[452] Anne Chao, S-M Lee, and S-L Jeng. Estimating population size for capture–recapture data when capture probabilities vary by time and individual animal. Biometrics, 48(1):201–216, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532750.

Freeman:1992:MSR

[453] Stephen N. Freeman and Byron J. T. Morgan. A modelling strategy for recovery data from birds ringed as nestlings. Biometrics, 48(1):217–235, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532751.

Sullivan:1992:KFA

[454] Patrick J. Sullivan. A Kalman filter approach to catch-at-length analysis. Biometrics, 48(1):237–257, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532752.

Mesbah:1992:CLM

[455] M. Mesbah, J. Lellouch, and C. Huber. The choice of loglinear models in contingency tables when the variables of interest are not jointly observed. Biometrics, 48(1):259–265, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532753.

Tsujitani:1992:NAM

[456] Masaaki Tsujitani. A note on the additive and multiplicative models in two-way contingency tables. Biometrics, 48(1):267–269, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532754. See correction [739].

Lipsitz:1992:MEP

[457] Stuart R. Lipsitz. Methods for estimating the parameters of a linear model for ordered categorical data. Biometrics, 48(1):271–281, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532755.

Bohning:1992:CAA

[458] Dankmar Böhning, Peter Schlattmann, and Bruce Lindsay. Computerassisted analysis of mixtures (C.A.MAN): Statistical algorithms. Biometrics, 48(1):283–303, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532756.

Carbonell:1992:IMA

[459] E. A. Carbonell, T. M. Gerig, E. Balansard, and M. J. Asins. Interval mapping in the analysis of nonadditive quantitative trait loci. Biometrics, 48(1):305–315, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532757.

Le:1992:ELE

[460] Nhu D. Le, Brian G. Leroux, Martin L. Puterman, and Paul S. Albert. Exact likelihood evaluation in a Markov mixture model for time series of seizure counts. Biometrics, 48(1):317–323, March 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532758.

Peterson:1992:POM

[461] Bercedis Peterson, Frank E. Harrell, Jr., and Rollin Brant. Proportional odds model. Biometrics, 48(1):325–326, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532759.

Tango:1992:DCT

[462] Toshiro Tango, D. J. Best, and J. C. W. Rayner. Disease clustering in time. Biometrics, 48(1):326–327, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532760.

Cavill:1992:RBC

[463] I. Cavill and Christine E. McLaren. Red blood cell volume distributions. Biometrics, 48(1):327–328, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532761.

Kemp:1992:BRSa

[464] C. D. Kemp. Book review: ’Student’: a Statistical Biography of William Sealy Gosset, by E. S. Pearson, R. L. Plackett, G. A. Barnard. Biometrics, 48(1):329–330, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532762.

Hamerle:1992:BRA

[465] A. Hamerle. Book review: Analyzing Quantitative Behavioral Observation Data, by H. K. Suen, D. Ary. Biometrics, 48(1):330, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532763.

Habbema:1992:BRS

[466] J. D. F. Habbema. Book review: Survivorship Analysis for Clinical Studies, by E. K. Harris, A. Albert. Biometrics, 48(1):331, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532764.

Kemp:1992:BRG

[467] A. W. Kemp. Book review: Generalized Poisson Distributions: Properties and Applications, by P. C. Consul. Biometrics, 48(1):331–332, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532765.

Gordon:1992:BRT

[468] A. D. Gordon. Book review: Trees and Proximity Representations, by J.-P. Barthélemy, A. Guénoche. Biometrics, 48(1):332, March 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532766.

Cheng:1992:BRM

[469] R. C. H. Cheng. Book review: Modern Statistical, Systems, and GPSS Simulation: The First Course, by Z. A. Karian, E. J. Dudewicz. Biometrics, 48(1):332–333, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532767.

Anonymous:1992:BRL

[470] Anonymous. Book review: Linear Statistical Models: an Applied Approach, by B. L. Bowerman, R. T. O’Connell. Biometrics, 48(1):333, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532769.

Anonymous:1992:BRSa

[471] Anonymous. Book review: A Statistical Model: Frederick Mosteller’s Contributions to Statistics, Science, and Public Policy, by S. E. Fienberg, D. C. Hoaglin, W. H. Kruskal, J. M. Tanur. Biometrics, 48(1):333, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532768.

Anonymous:1992:BRM

[472] Anonymous. Book review: Maximum-Entropy Models in Science and Engineering, by J. N. Kapur. Biometrics, 48(1):333–334, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532770.

Anonymous:1992:BRP

[473] Anonymous. Book review: Practical Spreadsheet Statistics and Curve Fitting for Scientists and Engineers, by L. M. Mezei. Biometrics, 48(1): 334, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532771.

Anonymous:1992:PPBa

[474] Anonymous. Papers to be published in biometrics. Biometrics, 48(1): 335, March 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532772.

Anonymous:1992:FMb

[475] Anonymous. Front matter. Biometrics, 48(2):i–iv, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532294.

Fu:1992:DGM

[476] Y.-X. Fu, W. E. Timberlake, and J. Arnold. On the design of genome mapping experiments using short synthetic oligonucleotides. Biometrics, 48(2):337–359, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532295.

Guo:1992:PET

[477] Sun Wei Guo and Elizabeth A. Thompson. Performing the exact test of Hardy–Weinberg proportion for multiple alleles. Biometrics, 48(2):361– 372, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532296.

Falissard:1992:NPG

[478] Bruno Falissard and Joseph Lellouch. A new procedure for group sequential analysis in clinical trials. Biometrics, 48(2):373–388, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532297.

Nam:1992:SSD

[479] Jun mo Nam. Sample size determination for case-control studies and the comparison of stratified and unstratified analyses. Biometrics, 48(2):389– 395, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532298.

Lynn:1992:WDI

[480] Henry S. Lynn and Charles E. McCulloch. When does it pay to break the matches for analysis of a matched-pairs design? Biometrics, 48(2):397– 409, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532299.

LeBlanc:1992:RRT

[481] Michael LeBlanc and John Crowley. Relative risk trees for censored survival data. Biometrics, 48(2):411–425, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532300.

Fredkin:1992:BRS

[482] Donald R. Fredkin and John A. Rice. Bayesian restoration of singlechannel patch clamp recordings. Biometrics, 48(2):427–448, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532301.

Keiding:1992:ESD

[483] Niels Keiding and Lars Andersen. Estimation of the size distribution of fibrillar centres in nucleoli — an example of the “Swiss Cheese” problem in stereology. Biometrics, 48(2):449–458, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532302. See correction [1402].

Prentice:1992:UHC

[484] Ross L. Prentice, Robert T. Smythe, Daniel Krewski, and Mark Mason. On the use of historical control data to estimate dose response trends in quantal bioassay. Biometrics, 48(2):459–478, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532303.

Robins:1992:EEE

[485] James M. Robins, Steven D. Mark, and Whitney K. Newey. Estimating exposure effects by modelling the expectation of exposure conditional on confounders. Biometrics, 48(2):479–495, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532304.

Wernecke:1992:CPD

[486] Klaus-D. Wernecke. A coupling procedure for the discrimination of mixed data. Biometrics, 48(2):497–506, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532305.

Escobar:1992:AIR

[487] Luis A. Escobar and William Q. Meeker, Jr. Assessing influence in regression analysis with censored data. Biometrics, 48(2):507–528, June 1992.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532306.

Bedrick:1992:EAS

[488] Edward J. Bedrick and Joe R. Hill. An empirical assessment of saddlepoint approximations for testing a logistic regression parameter. Biometrics, 48(2):529–544, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532307.

Leroux:1992:MPL

[489] Brian G. Leroux and Martin L. Puterman. Maximum-penalized-likelihood estimation for independent and Markov-dependent mixture models. Biometrics, 48(2):545–558, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532308.

Shanmugam:1992:IPP

[490] Ramalingam Shanmugam. An inferential procedure for the Poisson intervention parameter. Biometrics, 48(2):559–565, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532309.

Cormack:1992:IEM

[491] R. M. Cormack. Interval estimation for mark-recapture studies of closed populations. Biometrics, 48(2):567–576, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532310.

Rao:1992:SMA

[492] J. N. K. Rao and A. J. Scott. A simple method for the analysis of clustered binary data. Biometrics, 48(2):577–585, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532311.

Alho:1992:PID

[493] Juha M. Alho. On prevalence, incidence, and duration in general stable populations. Biometrics, 48(2):587–592, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532312.

Chen:1992:UBD

[494] Shande Chen and Christopher Cox. Use of baseline data for estimation of treatment effects in the presence of regression to the mean. Biometrics, 48(2):593–598, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532313.

Lin:1992:AVU

[495] Lawrence I-Kuei Lin. Assay validation using the concordance correlation coefficient. Biometrics, 48(2):599–604, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532314.

Sposto:1992:EDM

[496] Richard Sposto, Dale L. Preston, Yukiko Shimizu, and Kiyohiko Mabuchi. The effect of diagnostic misclassification on non-cancer and cancer mortality dose response in A-bomb survivors. Biometrics, 48(2): 605–617, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532315.

Lim:1992:ECS

[497] Lynette L-Y Lim. Estimating compliance to study medication from serum drug levels: Application to an AIDS clinical trial of zidovudine. Biometrics, 48(2):619–630, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532316.

Karim:1992:GLM

[498] M. Rezaul Karim and Scott L. Zeger. Generalized linear models with random effects; salamander mating revisited. Biometrics, 48(2):631–644, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532317.

Liu:1992:ERM

[499] Xinhua Liu and Kung-Yee Liang. Efficacy of repeated measures in regression models with measurement error. Biometrics, 48(2):645–654, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532318.

Kemp:1992:BRB

[500] A. W. Kemp. Book review: Biometrisches Wörterbuch: Erläuterungen/ Register, by D. Rasch. Biometrics, 48(2):655–656, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532319.

Walter:1992:BRP

[501] S. D. Walter. Book review: Practical Statistics for Medical Research, by D. G. Altman. Biometrics, 48(2):656, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532320.

Lawson:1992:BRS

[502] A. B. Lawson. Book review: Statistical Inference for Capture–Recapture Experiments, by K. H. Pollock, J. D. Nichols, C. Brownie, J. E. Hines. Biometrics, 48(2):656–657, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532321.

Newson:1992:BRO

[503] R. Newson. Book review: Order Statistics and Inference: Estimation Methods, by N. Balakrishnan, A. C. Cohen. Biometrics, 48(2):657, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532322. See corrigen-

dum [595].

Kemp:1992:BRE

[504] C. D. Kemp. Book review: Envisioning Information, by E. R. Tufte. Biometrics, 48(2):657–658, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532323.

Rasch:1992:BRB

[505] D. Rasch. Book review: Basic Statistical Computing, by D. Cooke, A. H. Craven, G. M. Clarke. Biometrics, 48(2):658–659, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532324.

Anonymous:1992:BRAa

[506] Anonymous. Book review: Analysis of Complex Surveys, by C. J. Skinner, D. Holt, T. M. F. Smith. Biometrics, 48(2):659, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532325.

Anonymous:1992:BRWa

[507] Anonymous. Book review: Workshop on Design of Longitudinal Studies and Analysis of Repeated Measures Data: Proceedings of the Centre for Mathematics and Its Applications, by S. R. Wilson. Biometrics, 48(2): 659, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532326.

Anonymous:1992:BRB

[508] Anonymous. Book review: Bradford Hill’s Principles of Medical Statistics, by A. B. Hill, I. D. Hill. Biometrics, 48(2):659–660, June 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532327.

Anonymous:1992:BRC

[509] Anonymous. Book review: The Chronological Annotated Bibliography of Order Statistics, Volume IV: 1962–1963, by H. L. Harter. Biometrics, 48(2):660, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532329.

Anonymous:1992:BRSb

[510] Anonymous. Book review: Statistics in Health and Nutrition, by Rao K. Visweswara, G. Radhaiah, V. Narayana. Biometrics, 48(2):660, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532328.

Anonymous:1992:CLR

[511] Anonymous. Correction: The likelihood ratio test for the two-component normal mixture problem: Power and sample size analysis. Biometrics, 48(2):661, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532330. See

[358].

Anonymous:1992:PPBb

[512] Anonymous. Papers to be published in biometrics. Biometrics, 48(2): 663, June 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532331.

Anonymous:1992:FMc

[513] Anonymous. Front matter. Biometrics, 48(3):i–iv, September 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532333.

Fisher:1992:RMA

[514] N. I. Fisher and A. J. Lee. Regression models for an angular response. Biometrics, 48(3):665–677, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532334.

Lin:1992:LRA

[515] J. S. Lin and L. J. Wei. Linear regression analysis based on Buckley– James estimating equation. Biometrics, 48(3):679–681, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532335.

Percy:1992:BAM

[516] David F. Percy. Blocked arteries and multivariate regression. Biometrics, 48(3):683–693, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532336.

Zucker:1992:TET

[517] David Zucker and Janet Wittes. Testing the effect of treatment in experiments with correlated binary outcomes. Biometrics, 48(3):695–709, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532337.

Kim:1992:TGF

[518] Byung Soo Kim and Barry H. Margolin. Testing goodness of fit of a multinomial model against overdispersed alternatives. Biometrics, 48(3):711– 719, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532338.

Rosner:1992:MMC

[519] B. Rosner. Multivariate methods for clustered binary data with multiple subclasses, with application to binary longitudinal data. Biometrics, 48(3):721–731, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532339.

Munoz:1992:PFC

[520] Alvaro Muñoz, Vincent Carey, Jan P. Schouten, Mark Segal, and Bernard Rosner. A parametric family of correlation structures for the analysis of longitudinal data. Biometrics, 48(3):733–742, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532340.

Forcina:1992:MBL

[521] A. Forcina. Modelling balanced longitudinal data: Maximum likelihood estimation and analysis of variance. Biometrics, 48(3):743–750, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532341.

Gottschau:1992:EMM

[522] Adam Gottschau. Exchangeability in multivariate Markov chain models. Biometrics, 48(3):751–763, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532342.

Wu:1992:SMC

[523] Margaret C. Wu and K. K. Gordon Lan. Sequential monitoring for comparison of changes in a response variable in clinical studies. Biometrics, 48(3):765–779, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532343.

Thomas:1992:UER

[524] Duncan C. Thomas, Maria Blettner, and Nicholas E. Day. Use of external rates in nested case-control studies with application to the international radiation study of cervical cancer patients. Biometrics, 48(3):781–794, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532344.

Klein:1992:SER

[525] John P. Klein. Semiparametric estimation of random effects using the Cox model based on the EM algorithm. Biometrics, 48(3):795–806, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532345.

Zelterman:1992:SDU

[526] Daniel Zelterman. A statistical distribution with an unbounded hazard function and its application to a theory from demography. Biometrics, 48(3):807–818, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532346.

OBrien:1992:RPT

[527] Peter C. O’Brien. Robust procedures for testing equality of covariance matrices. Biometrics, 48(3):819–827, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532347.

Nakamura:1992:PHM

[528] Tsuyoshi Nakamura. Proportional hazards model with covariates subject to measurement error. Biometrics, 48(3):829–838, September 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532348.

Geisser:1992:OAD

[529] Seymour Geisser and Wesley Johnson. Optimal administration of dual screening tests for detecting a characteristic with special reference to low prevalence diseases. Biometrics, 48(3):839–852, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532349.

OQuigley:1992:EPT

[530] John O’Quigley. Estimating the probability of toxicity at the recommended dose following a Phase I clinical trial in cancer. Biometrics, 48(3):853–862, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532350. See corrections [853].

Mingoti:1992:ETN

[531] Sueli A. Mingoti and Glen Meeden. Estimating the total number of distinct species using presence and absence data. Biometrics, 48(3):863– 875, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532351.

Skalski:1992:SSC

[532] John R. Skalski. Sample size calculations for normal variates under binomial censoring. Biometrics, 48(3):877–882, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532352.

Mudholkar:1992:THT

[533] Govind S. Mudholkar and Ila C. Sarkar. Testing homoscedasticity in a two-way table. Biometrics, 48(3):883–888, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532353.

Liao:1992:AMV

[534] Jiangang Liao. An algorithm for the mean and variance of the noncentral hypergeometric distribution. Biometrics, 48(3):889–892, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532354.

Zucker:1992:EWL

[535] David M. Zucker. The efficiency of a weighted log-rank test under a percent error misspecification model for the log hazard ratio. Biometrics, 48(3):893–899, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532355.

Buckland:1992:RLT

[536] Stephen T. Buckland and Benjamin J. Turnock. A robust line transect method. Biometrics, 48(3):901–909, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532356.

Cox:1992:GAQ

[537] Christopher Cox. A GLM approach to quantal response models for mixtures. Biometrics, 48(3):911–928, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532357.

Malcata:1992:SOD

[538] F. Xavier Malcata. Starting D-optimal designs for batch kinetic studies of enzyme-catalyzed reactions in the presence of enzyme deactivation. Biometrics, 48(3):929–938, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532358.

Qaqish:1992:MMC

[539] Bahjat F. Qaqish and Kung-Yee Liang. Marginal models for correlated binary responses with multiple classes and multiple levels of nesting. Biometrics, 48(3):939–950, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532359.

Odell:1992:MLE

[540] Patricia M. Odell, Keaven M. Anderson, and Ralph B. D’Agostino. Maximum likelihood estimation for interval-censored data using a Weibullbased accelerated failure time model. Biometrics, 48(3):951–959, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532360.

Yu:1992:ESR

[541] Kai Fun Yu, Sander Greenland, and Paul W. Holland. On estimating standardized risk differences from odds ratios. Biometrics, 48(3):961–964, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532361.

Gill:1992:NMC

[542] P. S. Gill. A note on modelling the covariance structure of repeated measurements. Biometrics, 48(3):965–968, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532362.

Tibshirani:1992:BHT

[543] Robert Tibshirani, Peter Hall, and Susan R. Wilson. Bootstrap hypothesis testing. Biometrics, 48(3):969–970, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532363.

Shih:1992:IRD

[544] Weichung Joseph Shih. On informative and random dropouts in longitudinal studies. Biometrics, 48(3):970–972, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532364.

Rayner:1992:CST

[545] Arthur A. Rayner. Consultant statisticians and their clients. Biometrics, 48(3):972, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532365.

Gentle:1992:BRR

[546] J. E. Gentle. Book review: Randomization and Monte Carlo Methods in Biology, by B. F. J. Manly. Biometrics, 48(3):973–974, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532366.

Everitt:1992:BRU

[547] B. S. Everitt. Book review: A User’s Guide to Principal Components, by J. E. Jackson. Biometrics, 48(3):974, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532367.

Paterson:1992:BRS

[548] L. Paterson. Book review: Survey Errors and Survey Costs, by R. M. Groves. Biometrics, 48(3):974–975, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532368.

Stene:1992:BRH

[549] J. Stene. Book review: Handbuch der Populations-genetik und Züchtungsmethodik, by D. Rasch, G. Herrendörfer. Biometrics, 48(3): 975, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532369.

Edwards:1992:BRA

[550] A. W. F. Edwards. Book review: Analysis of Human Genetic Linkage, by J. Ott. Biometrics, 48(3):975–976, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532370.

Barath:1992:BRF

[551] E. Baráth. Book review: Fundamentals of Biostatistics, by B. A. Rosner. Biometrics, 48(3):976, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532371.

Anonymous:1992:BRAb

[552] Anonymous. Book review: The Analysis of Contingency Tables, by B. S. Everitt. Biometrics, 48(3):977, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532373.

Kemp:1992:BRC

[553] C. D. Kemp and A. W. Kemp. Book reviews: The Chronological Annotated Bibliography of Order Statistics, Volume V: 1964–1965 by H. L. Harter; The Chronological Annotated Bibliography of Order Statistics, Volume VI: 1966–1967, by H. L. Harter. Biometrics, 48(3):977, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532374.

Loukas:1992:BRM

[554] S. Loukas. Book review: Mathematical Statistics with Applications, by W. Mendenhall, D. D. Wackerly, R. L. Scheaffer. Biometrics, 48(3):977, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532372.

Anonymous:1992:CAI

[555] Anonymous. Correction: Approximate interval estimation of the difference in binomial parameters: Correction for skewness and extension to multiple tables. Biometrics, 48(3):979, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532375. See [113].

Gart:1992:CAI

[556] J. J. Gart and J. Nam. Correction to: “Approximate interval estimation of the difference in binomial parameters: correction for skewness and extension to multiple tables” [Biometrics 46 (1990), no. 3, 637–643; 1085812]. Biometrics, 48(3):979, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [113].

Anonymous:1992:PPBc

[557] Anonymous. Papers to be published in biometrics. Biometrics, 48(3):981, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532376.

Anonymous:1992:BSF

[558] Anonymous. The Biometric Society financial statements and independent auditors’ report January 31, 1992. Biometrics, 48(3):983–986, September 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532377.

Anonymous:1992:FMd

[559] Anonymous. Front matter. Biometrics, 48(4):i–iv, December 1992. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532692.

Anonymous:1992:VI

[560] Anonymous. Volume information. Biometrics, 48(4):ix–xxvii, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532691.

Harville:1992:CBP

[561] David A. Harville and Alicia L. Carriquiry. Classical and Bayesian prediction as applied to an unbalanced mixed linear model. Biometrics, 48(4):987–1003, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532693.

Wright:1992:AVS

[562] S. Paul Wright. Adjusted P -values for simultaneous inference. Biometrics, 48(4):1005–1013, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532694.

Verbyla:1992:AMS

[563] A. P. Verbyla and B. R. Cullis. The analysis of multistratum and spatially correlated repeated measures data. Biometrics, 48(4):1015–1032, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532695.

Su:1992:GSD

[564] John Q. Su and John M. Lachin. Group sequential distribution-free methods for the analysis of multivariate observations. Biometrics, 48 (4):1033–1042, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532696.

Johnson:1992:TSN

[565] Richard A. Johnson, Christopher H. Morrell, and Anton Schick. Twosample nonparametric estimation and confidence intervals under truncation. Biometrics, 48(4):1043–1056, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532697.

Goldstein:1992:PSU

[566] Harvey Goldstein and Huiqi Pan. Percentile smoothing using piecewise polynomials, with covariates. Biometrics, 48(4):1057–1068, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532698.

Raz:1992:THE

[567] Jonathan Raz and George Fein. Testing for heterogeneity of evoked potential signals using an approximation to an exact permutation test.

Biometrics, 48(4):1069–1080, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532699.

Clarke:1992:IEM

[568] G. P. Y. Clarke. Inverse estimates from a multiresponse model. Biometrics, 48(4):1081–1094, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532700.

Qu:1992:LVM

[569] Yinsheng Qu, George W. Williams, Gerald J. Beck, and Sharon Vander-Brug Medendorp. Latent variable models for clustered dichotomous data with multiple subclusters. Biometrics, 48(4):1095–1102, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532701.

Fu:1992:TES

[570] Y. X. Fu and J. Arnold. A table of exact sample sizes for use with Fisher’s exact test for 2 × 2 tables. Biometrics, 48(4):1103–1112, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532702.

DeSouza:1992:ABB

[571] Cynthia M. DeSouza. An approximate bivariate Bayesian method for analyzing small frequencies. Biometrics, 48(4):1113–1130, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532703.

Proschan:1992:EAV

[572] Michael A. Proschan, Dean A. Follmann, and Myron A. Waclawiw. Effects of assumption violations on Type I error rate in group sequential monitoring. Biometrics, 48(4):1131–1143, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532704.

Sitter:1992:RDB

[573] R. R. Sitter. Robust designs for binary data. Biometrics, 48(4):1145– 1155, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532705.

Carriere:1992:IDB

[574] Keumhee Chough Carrière and Gregory C. Reinsel. Investigation of dualbalanced crossover designs for two treatments. Biometrics, 48(4):1157– 1164, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532706.

Chadoeuf:1992:MST

[575] J. Chadœuf, D. Nandris, J. P. Geiger, M. Nicole, and J. C. Pierrat. Modélisation spatio-temporelle d’une epidémie par un processus de Gibbs: estimation et tests. (French) [Spatio-temporal modeling of an epidemic by a Gibbs process: estimation and tests]. Biometrics, 48 (4):1165–1175, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532707.

Krauth:1992:BUT

[576] J. Krauth. Bounds for the upper-tail probabilities of the circular ratchet scan statistic. Biometrics, 48(4):1177–1185, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532708.

Kashiwagi:1992:SSC

[577] Nobuhisa Kashiwagi and Takemi Yanagimoto. Smoothing serial count data through a state-space model. Biometrics, 48(4):1187–1194, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532709.

Thompson:1992:APS

[578] Steven K. Thompson, Fred L. Ramsey, and George A. F. Seber. An adaptive procedure for sampling animal populations. Biometrics, 48 (4):1195–1199, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532710.

Jensen:1992:IAS

[579] A. L. Jensen. Integrated area sampling and mark-recapture experiments for sampling fish populations. Biometrics, 48(4):1201–1205, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532711.

Shih:1992:ECP

[580] Weichung Joseph Shih and Wei-Min Huang. Evaluating correlation with proper bounds. Biometrics, 48(4):1207–1213, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532712.

Sato:1992:MLE

[581] Tosiya Sato. Maximum likelihood estimation of the risk ratio in casecohort studies. Biometrics, 48(4):1215–1221, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532713.

Laska:1992:NET

[582] Eugene M. Laska and Morris J. Meisner. Nonparametric estimation and testing in a cure model. Biometrics, 48(4):1223–1234, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532714.

Hatzis:1992:ODN

[583] Christos Hatzis and Kinley Larntz. Optimal design in nonlinear multiresponse estimation: Poisson model for filter feeding. Biometrics, 48 (4):1235–1248, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532715.

Hilsenbeck:1992:ECA

[584] Susan Galloway Hilsenbeck, Charles Kurucz, and Robert C. Duncan. Estimation of completeness and adjustment of age-specific and agestandardized incidence rates. Biometrics, 48(4):1249–1262, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532716.

Kimura:1992:FCC

[585] Daniel K. Kimura. Functional comparative calibration using an EM algorithm. Biometrics, 48(4):1263–1271, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532717. See comments [1240].

Conradsen:1992:ATD

[586] Knut Conradsen and Jan Pedersen. Analysis of two-dimensional electrophoretic gels. Biometrics, 48(4):1273–1287, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532718.

Mantel:1992:ICS

[587] Nathan Mantel, S. D. Walter, and R. J. Cook. An invalid comparison of several point estimators of the odds ratio in a single 2 × 2 contingency table. Biometrics, 48(4):1289–1295, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532719.

Read:1992:BRA

[588] C. B. Read. Book review: The Art of Statistical Science: a Tribute to G. S. Watson, by K. V. Mardia. Biometrics, 48(4):1297–1298, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532720.

Kemp:1992:BRSb

[589] A. W. Kemp. Book review: Statistical Theory and Modelling: In Honour of Sir David Cox, FRS, by D. V. Hinkley, N. Reid, E. J. Snell. Biometrics, 48(4):1298, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532721.

Brown:1992:BRS

[590] R. A. Brown. Book review: Statistical Methods: The Geometric Approach, by D. J. Saville, G. R. Wood. Biometrics, 48(4):1298–1299, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532722.

Cohen:1992:BRS

[591] G. R. Cohen. Book reviews: Sampling of Populations: Methods and Applications, by P. S. Levy, S. Lemeshow; Design and Inference in Finite Population Sampling, by A. S. Hedayat, B. K. Sinha. Biometrics, 48 (4):1299–1300, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532723.

Cormack:1992:BRS

[592] R. M. Cormack. Book review: Statistics for Spatial Data, by N. Cressie. Biometrics, 48(4):1300–1301, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532724.

Reyment:1992:BRM

[593] R. A. Reyment. Book review: Morphometric Tools for Landmark Data: Geometry and Biology, by F. L. Bookstein. Biometrics, 48(4):1301–1302, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532725.

Bohning:1992:BRF

[594] D. Böhning. Book review: A First Course in the Theory of Linear Statistical Models, by R. H. Myers, J. S. Milton. Biometrics, 48(4):1302–1303, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532726.

Anonymous:1992:COS

[595] Anonymous. Corrigendum: Order Statistics and Inference: Estimation Methods. Biometrics, 48(4):1303, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532730. See [503].

Anonymous:1992:BRDa

[596] Anonymous. Book review: Differential Equations with Applications in Biology, Physics, and Engineering, by J. A. Goldstein, F. Kappel, W. Schappacher. Biometrics, 48(4):1303, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532728.

Anonymous:1992:BRDb

[597] Anonymous. Book review: The Development of Statistics: Recent Contributions from China, by X. R. Chen, K. T. Fang, C. C. Yang. Biometrics, 48(4):1303, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532727.

Anonymous:1992:BRWb

[598] Anonymous. Book reviews: World Health Statistics Annual 1991, by World Health Organization; World Health Statistics Quarterly 1991, by World Health Organization. Biometrics, 48(4):1303, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532729.

Anonymous:1992:PPBd

[599] Anonymous. Papers to be published in biometrics. Biometrics, 48(4): 1305, December 1992. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532731.

Anonymous:1993:FMa

[600] Anonymous. Front matter. Biometrics, 49(1):i–iv, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532596.

Wang:1993:SMP

[601] Mei-Cheng Wang, Ron Brookmeyer, and Nicholas P. Jewell. Statistical models for prevalent cohort data. Biometrics, 49(1):1–11, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532597.

Kim:1993:ADC

[602] Mimi Y. Kim, Victor G. De Gruttola, and Stephen W. Lagakos. Analyzing doubly censored data with covariates, with application to AIDS. Biometrics, 49(1):13–22, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532598.

Tang:1993:DAR

[603] Dei-In Tang, Nancy L. Geller, and Stuart J. Pocock. On the design and analysis of randomized clinical trials with multiple endpoints. Biometrics, 49(1):23–30, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2532599. See comment [886].

Jennison:1993:SET

[604] Christopher Jennison and Bruce W. Turnbull. Sequential equivalence testing and repeated confidence intervals, with applications to normal and binary responses. Biometrics, 49(1):31–43, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532600.

Ofversten:1993:ETV

[605] J. Öfversten. Exact tests for variance components in unbalanced mixed linear models. Biometrics, 49(1):45–57, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532601.

Davidian:1993:SSM

[606] Marie Davidian and David M. Giltinan. Some simple methods for estimating intraindividual variability in nonlinear mixed effects models. Biometrics, 49(1):59–73, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532602.

Laster:1993:RMD

[607] Larry L. Laster and James Pickands III. Repeated measurement designs with random selections. Biometrics, 49(1):75–83, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532603.

Hung:1993:TED

[608] H. M. James Hung, G. Y. H. Chi, and R. J. Lipicky. Testing for the existence of a desirable dose combination. Biometrics, 49(1):85–94, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532604.

Barton:1993:NSM

[609] Curtis N. Barton, Robert C. Braunberg, and Leonard Friedman. Nonlinear statistical models for the joint action of toxins. Biometrics, 49(1): 95–105, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532605.

Hines:1993:MOT

[610] R. J. O’Hara Hines and J. F. Lawless. Modelling overdispersion in toxicological mortality data grouped over time. Biometrics, 49(1):107–121, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532606.

Uesaka:1993:TIB

[611] Hiroyuki Uesaka. Test for interaction between treatment and stratum with ordinal responses. Biometrics, 49(1):123–129, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532607.

Agresti:1993:QSL

[612] Alan Agresti and Joseph B. Lang. Quasi-symmetric latent class models, with application to rater agreement. Biometrics, 49(1):131–139, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532608.

Rusakov:1993:ESD

[613] D. A. Rusakov. Estimation of the size distribution of closed cell elements from analysis of their random plane sections. Biometrics, 49(1):141–149, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532609.

Shoukri:1993:STG

[614] M. M. Shoukri. Statistical testing of genetic linkage under heterogeneity. Biometrics, 49(1):151–161, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532610.

Sheehan:1993:IMC

[615] Nuala Sheehan and Alum Thomas. On the irreducibility of a Markov chain defined on a space of genotype configurations by a sampling scheme. Biometrics, 49(1):163–175, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532611.

Schwarz:1993:EMR

[616] Carl J. Schwarz, Jake F. Schweigert, and A. Neil Arnason. Estimating migration rates using tag-recovery data. Biometrics, 49(1):177–193, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532612.

Ramakrishnan:1993:NBC

[617] V. Ramakrishnan and D. Meeter. Negative binomial cross-tabulations, with applications to abundance data. Biometrics, 49(1):195–207, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532613.

Laska:1993:PCM

[618] Eugene M. Laska and Morris Meisner. A plant-capture method for estimating the size of a population from a single sample. Biometrics, 49(1): 209–220, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532614.

McGilchrist:1993:RES

[619] C. A. McGilchrist. REML estimation for survival models with frailty. Biometrics, 49(1):221–225, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532615.

Jansen:1993:MLG

[620] R. C. Jansen. Maximum likelihood in a generalized linear finite mixture model by using the EM algorithm. Biometrics, 49(1):227–231, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532616.

Yao:1993:EAC

[621] Qing Yao and David Tritchler. An exact analysis of conditional independence in several 2 × 2 contingency tables. Biometrics, 49(1):233–236, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532617.

ONeill:1993:GME

[622] Terence J. O’Neill and Helen C. O’Neill. A gamma model for extrabinomial variation in dilution assays. Biometrics, 49(1):237–242, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532618.

Faddy:1993:SCM

[623] M. J. Faddy. A structured compartmental model for drug kinetics. Biometrics, 49(1):243–248, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532619.

Tamura:1993:CRT

[624] Roy N. Tamura, Bradley J. Mills, and Jeffrey K. Lovelace. Constrained randomization in a therapeutic efficacy trial. Biometrics, 49(1):249–258, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532620.

Freedman:1993:SAA

[625] Laurence S. Freedman, Douglas N. Midthune, Charles C. Brown, Vernon Steele, and Gary J. Kelloff. Statistical analysis of animal cancer chemoprevention experiments. Biometrics, 49(1):259–268, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532621.

Engel:1993:AED

[626] Bas Engel and Joop te Brake. Analysis of embryonic development with a model for under-or overdispersion relative to binomial variation. Biometrics, 49(1):269–279, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532622.

Hebel:1993:SEA

[627] Pascale Hebel, Robert Faivre, Bruno Goffinet, and Daniel Wallach. Shrinkage estimators applied to prediction of French winter wheat yield.

Biometrics, 49(1):281–293, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532623.

Catchpole:1993:SDS

[628] W. R. Catchpole and E. A. Catchpole. Stratified double sampling of patchy vegetation to estimate biomass. Biometrics, 49(1):295–303, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532624.

Dutilleul:1993:MTA

[629] Pierre Dutilleul, Peter Clifford, Sylvia Richardson, and Denis Hemon. Modifying the t test for assessing the correlation between two spatial processes. Biometrics, 49(1):305–314, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532625.

Hook:1993:AP

[630] Ernest B. Hook, Ronald R. Regal, and Richard Cormack. Acknowledgement of priority. Biometrics, 49(1):315, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532626.

Kanji:1993:BRR

[631] G. K. Kanji. Book review: Regression Analysis, by Example, by S. Chatterjee, B. Price. Biometrics, 49(1):317, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532627.

Abakuks:1993:BRB

[632] A. Abakuks. Book review: Biological Kinetics, by L. A. Segel. Biometrics, 49(1):317–318, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532628.

Kemp:1993:BRM

[633] A. W. Kemp. Book review: The Most Frequent Value: Introduction to a Modern Conception of Statistics, by F. Steiner. Biometrics, 49(1):318, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532629.

Derr:1993:BRS

[634] J. Derr. Book review: Statistics for Research, by S. Dowdy, S. Wearden. Biometrics, 49(1):318–319, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532630.

Kemp:1993:BRF

[635] C. D. Kemp and A. W. Kemp. Book reviews: The Frontiers of Statistical Computation, Simulation, and Modeling, by P. R. Nelson, E. J. Dudewicz, A. Öztürk, E. C. Meulen; The Frontiers of Statistical Scientific Theory and Industrial Applications, by A. Öztürk, E. C. Van der Meulen, E. J. Dudewica, P. R. Nelson. Biometrics, 49(1):319, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532631.

Anonymous:1993:BRPa

[636] Anonymous. Book review: Probability: The Mathematics of Uncertainty, by D. Feldman, M. Fox. Biometrics, 49(1):320, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532633.

Anonymous:1993:BRTa

[637] Anonymous. Book review: Théorie des Probabilités en Vue des Applications Statistiques, by Ph. Tassi, S. Legait. Biometrics, 49(1):320, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532632.

Anonymous:1993:BRSa

[638] Anonymous. Book review: Statistical Inference, by G. Casella, R. L. Berger. Biometrics, 49(1):320–321, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532634.

Anonymous:1993:BRA

[639] Anonymous. Book review: Advanced Research Methodology: an Annotated Guide to Sources, by R. B. Bausell. Biometrics, 49(1):321, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532636.

Anonymous:1993:BRN

[640] Anonymous. Book review: Numerical Solution of Markov Chains, by W. J. Stewart. Biometrics, 49(1):321, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532635.

Anonymous:1993:PPBa

[641] Anonymous. Papers to be published in biometrics. Biometrics, 49(1): 323, March 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532637.

Anonymous:1993:FMb

[642] Anonymous. Front matter. Biometrics, 49(2):i–iv, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532546.

Atkinson:1993:OED

[643] Anthony C. Atkinson, Kathryn Chaloner, Agnes M. Herzberg, and June Juritz. Optimum experimental designs for properties of a compartmental model. Biometrics, 49(2):325–337, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532547.

James:1993:RRH

[644] Alan T. James, Joseph T. Wiskich, and Robert A. J. Conyers. t-REML for robust heteroscedastic regression analysis of mitochondrial power. Biometrics, 49(2):339–356, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532548.

Bailer:1993:ITP

[645] A. John Bailer and Christopher J. Portier. An index of tumorigenic potency. Biometrics, 49(2):357–365, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532549.

Dewanji:1993:WME

[646] A. Dewanji, D. Krewski, and M. J. Goddard. A Weibull model for the estimation of tumorigenic potency. Biometrics, 49(2):367–377, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532550.

Baker:1993:RAG

[647] Stuart G. Baker, Yohanan Wax, and Blossom H. Patterson. Regression analysis of grouped survival data: Informative censoring and double sampling. Biometrics, 49(2):379–389, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532551.

Liu:1993:SDP

[648] P. Y. Liu, Steve Dahlberg, and John Crowley. Selection designs for pilot studies based on survival. Biometrics, 49(2):391–398, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532552.

Dinse:1993:ECA

[649] Gregg E. Dinse. Evaluating constraints that allow survival-adjusted incidence analyses in single-sacrifice studies. Biometrics, 49(2):399–407, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532553.

Ebrahimi:1993:ETO

[650] Nader Ebrahimi. Estimation of two ordered mean residual lifetime functions. Biometrics, 49(2):409–417, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532554.

Feldmann:1993:EAR

[651] Uwe Feldmann. Epidemiologic assessment of risks of adverse reactions associated with intermittent exposure. Biometrics, 49(2):419–428, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532555.

Satten:1993:CRA

[652] Glen A. Satten and Lawrence L. Kupper. Conditional regression analysis of the exposure-disease odds ratio using known probability-of-exposure values. Biometrics, 49(2):429–440, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532556.

Gilks:1993:REM

[653] W. R. Gilks, C. C. Wang, B. Yvonnet, and P. Coursaget. Random-effects models for longitudinal data using Gibbs sampling. Biometrics, 49(2): 441–453, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532557.

Chen:1993:NPC

[654] Yuh-Ing Chen and Douglas A. Wolfe. Nonparametric procedures for comparing umbrella pattern treatment effects with a control in a one-way layout. Biometrics, 49(2):455–465, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532558.

Russek-Cohen:1993:QIM

[655] Estelle Russek-Cohen and Richard M. Simon. Qualitative interactions in multifactor studies. Biometrics, 49(2):467–477, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532559.

Rencher:1993:CIV

[656] Alvin C. Rencher. The contribution of individual variables to Hotelling’s T^{2}, Wilks’ Λ, and R^{2}. Biometrics, 49(2):479–489, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532560.

Cyr:1993:CIR

[657] Louis Cyr, Philip F. Rust, John R. Peters, Marcia K. Schmehl, and Harvey L. Bank. Confidence intervals for the relative frequency of responding cells in limiting dilution assays. Biometrics, 49(2):491–498, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532561.

Krewski:1993:MAS

[658] D. Krewski, B. G. Leroux, S. R. Bleuer, and L. H. Broekhoven. Modeling the Ames salmonella/microsome assay. Biometrics, 49(2):499–510, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532562.

Formann:1993:FDL

[659] Anton K. Formann. Fixed-distance latent class models for the analysis of sets of two-way contingency tables. Biometrics, 49(2):511–521, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532563.

Hale:1993:IEU

[660] Cecilia A. Hale and Joseph L. Fleiss. Interval estimation under two study designs for kappa with binary classifications. Biometrics, 49(2): 523–534, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532564.

Lau:1993:HOK

[661] Tai-Shing Lau. Higher-order kappa-type statistics for a dichotomous attribute in multiple ratings. Biometrics, 49(2):535–542, June 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532565.

Navidi:1993:PIL

[662] W. C. Navidi, G. A. Churchill, and A. von Haeseler. Phylogenetic inference: Linear invariants and maximum likelihood. Biometrics, 49(2): 543–555, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532566.

Williams:1993:CBP

[663] Christopher J. Williams. On the covariance between parameter estimates in models of twin data. Biometrics, 49(2):557–568, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532567.

Katapa:1993:THF

[664] R. S. Katapa. A test of hypothesis on familial correlations. Biometrics, 49(2):569–576, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532568.

Kimanani:1993:SMS

[665] Ebi Kalahi Kimanani and John Wycliffe Odhiambo. A stochastic model for the sterile male technique. Biometrics, 49(2):577–585, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532569.

Rai:1993:IEA

[666] S. N. Rai and D. E. Matthews. Improving the EM algorithm. Biometrics, 49(2):587–591, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532570.

Borgan:1993:NER

[667] Ørnulf Borgan and Bryan Langholz. Nonparametric estimation of relative mortality from nested case-control studies. Biometrics, 49(2):593– 602, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532571.

Su:1993:NED

[668] John Q. Su and L. J. Wei. Nonparametric estimation for the difference or ratio of median failure times. Biometrics, 49(2):603–607, June 1993.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532572.

Hilton:1993:PSS

[669] Joan F. Hilton and Cyrus R. Mehta. Power and sample size calculations for exact conditional tests with ordered categorical data. Biometrics, 49(2):609–616, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532573.

Matthews:1993:ATP

[670] J. N. S. Matthews and D. R. Appleton. An application of the truncated Poisson distribution to immunogold assay. Biometrics, 49(2):617–621, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532574.

Liang:1993:CSB

[671] Kung-Yee Liang and Peter McCullagh. Case studies in binary dispersion. Biometrics, 49(2):623–630, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532575.

Park:1993:TMD

[672] Taesung Park and Charles S. Davis. A test of the missing data mechanism for repeated categorical data. Biometrics, 49(2):631–638, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532576.

Lakshmi:1993:CPP

[673] Damaraju V. Lakshmi and Woollcott K. Smith. Comparing proportions in the presence of false positive and false negative instrument sorting errors. Biometrics, 49(2):639–641, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532577.

Smith:1993:MEN

[674] Bruce R. Smith. Mixture estimation and the neurophysiological quantal hypothesis. Biometrics, 49(2):643–651, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532578.

Gregoire:1993:EBS

[675] T. G. Gregoire, H. T. Valentine, and G. M. Furnival. Estimation of bole surface area and bark volume with Monte Carlo methods. Biometrics, 49(2):653–660, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532579.

Feingold:1993:CPE

[676] Marcia Feingold, Glenn Heller, and Jeffrey S. Simonoff. Choice of prediction estimator in censored regression models. Biometrics, 49(2):661–664, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532580.

Orme:1993:CST

[677] Chris Orme. A comment on “A simple test for neglected heterogeneity in panel studies”. Biometrics, 49(2):665–667, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532581. See [23].

Caudill:1993:NMO

[678] Steven B. Caudill. Nonlinear models for ordered categorical data. Biometrics, 49(2):669–670, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532582.

Lefkovitch:1993:BRA

[679] L. P. Lefkovitch. Book review: Applied Statistics: a Handbook of GEN-STAT Analyses, by E. J. Snell, H. R. Simpson. Biometrics, 49(2):671, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532583.

Kemp:1993:BRT

[680] A. W. Kemp. Book review: Topics in Statistical Methodology, by S. Biswas. Biometrics, 49(2):671–672, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532584.

Gordon:1993:BRC

[681] A. D. Gordon. Book review: Correspondence Analysis Handbook, by J.- P. Benzécri. Biometrics, 49(2):672, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532585.

Stene:1993:BRG

[682] J. Stene. Book review: Genetics of Populations, by J. P. Jain, V. T. Prabhakaran. Biometrics, 49(2):672–673, June 1993. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532586.

Anonymous:1993:BRIa

[683] Anonymous. Book review: Introduction to Probability and Mathematical Statistics, by L. J. Bain, M. Engelhardt. Biometrics, 49(2):673, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532587.

Anonymous:1993:BRMa

[684] Anonymous. Book review: Modern Data Analysis: a First Course in Applied Statistics, by L. C. Hamilton. Biometrics, 49(2):673, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532588.

Anonymous:1993:BRSb

[685] Anonymous. Book review: Statistical Inference: Theory and Practice, by T. Bromek, E. Pleszczynska. Biometrics, 49(2):673, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532589.

Anonymous:1993:BRMb

[686] Anonymous. Book review: Modern Sequential Statistical Analysis in Honor of Professor Herbert Robbins, by Z. Govindarajulu. Biometrics, 49(2):673–674, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532590.

Anonymous:1993:BRH

[687] Anonymous. Book review: A History of Mathematics, by C. B. Boyer, U. C. Merzbach. Biometrics, 49(2):674, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532593.

Anonymous:1993:BRIb

[688] Anonymous. Book review: Introduction to Configural Frequency Analysis, by A. von Eye. Biometrics, 49(2):674, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532592.

Anonymous:1993:BRMc

[689] Anonymous. Book review: Mathematical Bioeconomics: The Optimal Management of Renewable Resources, by C. Clark. Biometrics, 49(2):

674, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532591.

Anonymous:1993:PPBb

[690] Anonymous. Papers to be published in biometrics. Biometrics, 49(2): 675, June 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532594.

Anonymous:1993:FMc

[691] Anonymous. Front matter. Biometrics, 49(3):i–iv, September 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532188.

Drum:1993:REE

[692] Melinda L. Drum and Peter McCullagh. REML estimation with exact covariance in the logistic mixed model. Biometrics, 49(3):677–689, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532189.

Calvin:1993:REU

[693] James A. Calvin. REML estimation in unbalanced multivariate variance components models using an EM algorithm. Biometrics, 49(3):691–701, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532190.

Hsuan:1993:ETM

[694] Francis C. Hsuan. Estimating treatment means in a mixed-effect ANOVA model for bioequivalence studies. Biometrics, 49(3):703–713, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532191.

Huggins:1993:RAA

[695] R. M. Huggins. A robust approach to the analysis of repeated measures. Biometrics, 49(3):715–720, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532192.

Thornquist:1993:PHM

[696] Mark D. Thornquist. Proportional hazards model for repeated measures with monotonic ordinal response. Biometrics, 49(3):721–730, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532193.

Finkelstein:1993:PHM

[697] Dianne M. Finkelstein, Dirk F. Moore, and David A. Schoenfeld. A proportional hazards model for truncated AIDS data. Biometrics, 49(3):731– 740, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532194.

Jennison:1993:GST

[698] Christopher Jennison and Bruce W. Turnbull. Group sequential tests for bivariate response: Interim analyses of clinical trials with both efficacy and safety endpoints. Biometrics, 49(3):741–752, September 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532195.

Chen:1993:MSS

[699] T. Timothy Chen and Richard Simon. A multiple-step selection procedure with sequential protection of preferred treatments. Biometrics, 49(3):753–761, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532196.

Brittain:1993:OMT

[700] Erica H. Brittain and Kent R. Bailey. Optimization of multistage testing times and critical values in clinical trials. Biometrics, 49(3):763–772, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532197.

Lui:1993:SSDa

[701] Kung-Jong Lui. Sample size determination for cohort studies under an exponential covariate model with grouped data. Biometrics, 49(3):773– 778, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532198.

Carr:1993:ESM

[702] Gregory J. Carr and Christopher J. Portier. An evaluation of some methods for fitting dose-response models to quantal-response developmental toxicology data. Biometrics, 49(3):779–791, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532199.

Bieler:1993:RED

[703] Gayle S. Bieler and Rick L. Williams. Ratio estimates, the delta method, and quantal response tests for increased carcinogenicity. Biometrics, 49(3):793–801, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532200.

Banfield:1993:MBG

[704] Jeffrey D. Banfield and Adrian E. Raftery. Model-based Gaussian and non-Gaussian clustering. Biometrics, 49(3):803–821, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532201.

Uebersax:1993:LTF

[705] John S. Uebersax and William M. Grove. A latent trait finite mixture model for the analysis of rating agreement. Biometrics, 49(3):823–835, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532202.

Quang:1993:NEV

[706] Pham Xuan Quang. Nonparametric estimators for variable circular plot surveys. Biometrics, 49(3):837–852, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532203.

Evans:1993:CCJ

[707] Marc A. Evans and Douglas G. Bonett. A constrained Cook–Jacobson model of visibility bias. Biometrics, 49(3):853–859, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532204.

Srinivasan:1993:NPK

[708] Cidambi Srinivasan and Mai Zhou. A note on pooling Kaplan–Meier estimators. Biometrics, 49(3):861–864, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532205.

Greenland:1993:MLE

[709] Sander Greenland and Karsten Drescher. Maximum likelihood estimation of the attributable fraction from logistic models. Biometrics, 49(3):865–872, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532206.

Lui:1993:SSDb

[710] Kung-Jong Lui. Sample size determination for multiple continuous risk factors in case-control studies. Biometrics, 49(3):873–876, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532207.

Wellek:1993:LRT

[711] Stefan Wellek. A log-rank test for equivalence of two survivor functions. Biometrics, 49(3):877–881, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532208.

Tomizawa:1993:DPS

[712] Sadao Tomizawa. Diagonals-parameter symmetry model for cumulative probabilities in square contingency tables with ordered categories. Biometrics, 49(3):883–887, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532209.

Lawson:1993:DRH

[713] Andrew B. Lawson. A deviance residual for heterogeneous spatial Poisson processes. Biometrics, 49(3):889–897, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532210. See corrections [850].

Horgan:1993:ETD

[714] G. W. Horgan, S. T. Buckland, and L. A. Mackie-Dawson. Estimating three-dimensional line process densities from tube counts. Biometrics, 49(3):899–906, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532211.

Mendell:1993:WLR

[715] Nancy R. Mendell, Stephen J. Finch, and Henry C. Thode, Jr. Where is the likelihood ratio test powerful for detecting two component normal mixtures? Biometrics, 49(3):907–915, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532212.

Meier:1993:MTP

[716] Kristen L. Meier, A. John Bailer, and Christopher J. Portier. A measure of tumorigenic potency incorporating dose-response shape. Biometrics, 49(3):917–926, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532213.

Cologne:1993:AGE

[717] John B. Cologne, Randy L. Carter, Shoichiro Fujita, and Sadayuki Ban. Application of generalized estimating equations to a study of in vitro radiation sensitivity. Biometrics, 49(3):927–934, September 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532214.

Blair:1993:NOC

[718] R. Clifford Blair, Shlomo Sawilowsky, C. Brownie, D. D. Boos, and J. Hughes-Oliver. A note on the operating characteristics of the modified F test. Biometrics, 49(3):935–939, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532215.

Westfall:1993:AVM

[719] P. H. Westfall, S. S. Young, and S. Paul Wright. On adjusting P -values for multiplicity. Biometrics, 49(3):941–945, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532216.

Diggle:1993:IRD

[720] Peter J. Diggle and Weichung J. Shih. On informative and random dropouts in longitudinal studies. Biometrics, 49(3):947–949, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532217.

Kateri:1993:MTW

[721] Maria Kateri and M. Tsujitani. Models in two-way contingency tables. Biometrics, 49(3):950–951, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532218.

Mandel:1993:NAB

[722] John Mandel and Hugh G. Gauch, Jr. Non-additivity and biplots. Biometrics, 49(3):952–954, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532219.

Keiding:1993:BRK

[723] N. Keiding. Book review: Kendall’s Advanced Theory of Statistics, 2: Classical Inference and Relationship, by A. Stuart, J. K. Ord. Biometrics, 49(3):955–956, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532220.

Kocherlakota:1993:BRN

[724] S. Kocherlakota. Book review: Nonlinear Multivariate Analysis, by A. Gifi. Biometrics, 49(3):956, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532221.

Freeman:1993:BRS

[725] G. H. Freeman. Book review: Statistical Intervals, by G. J. Hahn, W. Q. Meeker. Biometrics, 49(3):956–957, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532222.

Pearce:1993:BRM

[726] S. C. Pearce. Book review: Meta-Analysis for Explanation: a Casebook, by T. D. Cook, H. Cooper, D. S. Cordray, H. Hartmann, L. V. Hedges, R. J. Light, T. A. Louis, F. Mosteller. Biometrics, 49(3):957, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532223.

Cormack:1993:BRF

[727] R. M. Cormack. Book review: Fundamentals of Biostatistical Inference, by C. T. Le. Biometrics, 49(3):957–958, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532224.

Kanji:1993:BRF

[728] G. K. Kanji. Book review: Fundamentals of Exploratory Analysis of Variance, by D. C. Hoaglin, F. Mosteller, J. W. Tukey. Biometrics, 49(3): 958, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532225.

Kanji:1993:RFE

[729] G. K. Kanji. Review: Fundamentals of Exploratory Analysis of Variance, by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 49(3):958, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL http://www.jstor.org/stable/2532225.

Pearce:1993:BRS

[730] S. C. Pearce. Book review: Scientific Fraud vs. Scientific Truth, by I. D. Bross. Biometrics, 49(3):958, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532226.

Anonymous:1993:BRIc

[731] Anonymous. Book review: An Introduction to Statistics with Data Analysis, by S. Rasmussen. Biometrics, 49(3):959, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532228.

Anonymous:1993:BRPb

[732] Anonymous. Book review: Processing Data: The Survey Example, by L. B. Bourque, V. A. Clark. Biometrics, 49(3):959, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532229.

Anonymous:1993:BRPc

[733] Anonymous. Book review: Probability: Theory and Examples, by R. Durrett. Biometrics, 49(3):959, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532227.

Anonymous:1993:BRR

[734] Anonymous. Book review: Regression Diagnostics, by J. Fox. Biometrics, 49(3):959, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532230.

Anonymous:1993:BRMd

[735] Anonymous. Book review: MACSYMA for Statisticians, by B. Heller. Biometrics, 49(3):959–960, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532231.

Anonymous:1993:BRTb

[736] Anonymous. Book review: Theoretical Biology: Epigenetic and Evolutionary Order from Complex Systems, by B. Goodwin, P. Saunders. Biometrics, 49(3):960, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532232.

Abt:1993:OAL

[737] Klaus Abt. Obituary: Arthur Linder 1904–1993. Biometrics, 49(3): 961–966, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1993:OAL

[738] Anonymous. Obituary: Arthur Linder 1904–1993. Biometrics, 49 (3):961–966, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532233.

Anonymous:1993:CNA

[739] Anonymous. Correction: A note on the additive and multiplicative models in two-way contingency tables. Biometrics, 49(3):967, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532234. See [456].

Anonymous:1993:PPBc

[740] Anonymous. Papers to be published in biometrics. Biometrics, 49(3):969, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532235.

Anonymous:1993:BSF

[741] Anonymous. The Biometric Society financial statements and independent auditors’ report January 31, 1993. Biometrics, 49(3):971–974, September 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532236.

Anonymous:1993:FMd

[742] Anonymous. Front matter. Biometrics, 49(4):i–iv, December 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532239.

Anonymous:1993:VI

[743] Anonymous. Volume information. Biometrics, 49(4):vii–xxvi, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532238.

Lefkopoulou:1993:GTM

[744] Myrto Lefkopoulou and Louise Ryan. Global tests for multiple binary outcomes. Biometrics, 49(4):975–988, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532240.

Neuhaus:1993:EET

[745] John M. Neuhaus. Estimation efficiency and tests of covariate effects with clustered binary data. Biometrics, 49(4):989–996, December 1993.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532241.

Becker:1993:MMB

[746] Mark P. Becker and Cecile C. Balagtas. Marginal modeling of binary cross-over data. Biometrics, 49(4):997–1009, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532242.

Espeland:1993:EID

[747] Mark A. Espeland, Julia T. Rushing, and Arthur DeVault. Estimating incidence and diagnostic error rates for bivariate progressive processes. Biometrics, 49(4):1010–1021, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532243.

Dawson:1993:SPT

[748] Jeffrey D. Dawson and Stephen W. Lagakos. Size and power of twosample tests of repeated measures data. Biometrics, 49(4):1022–1032, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532244.

Miller:1993:ALP

[749] Michael E. Miller, Charles S. Davis, and J. Richard Landis. The analysis of longitudinal polytomous data: Generalized estimating equations and connections with weighted least squares. Biometrics, 49(4):1033–1044, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532245.

Nandram:1993:BPI

[750] B. Nandram and J. Sedransk. Bayesian predictive inference for longitudinal sample surveys. Biometrics, 49(4):1045–1055, December 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532246.

Hughes:1993:RDP

[751] Michael D. Hughes. Regression dilution in the proportional hazards model. Biometrics, 49(4):1056–1066, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532247.

Crouchley:1993:STU

[752] Robert Crouchley and Andrew Pickles. A specification test for univariate and multivariate proportional hazards models. Biometrics, 49(4):1067– 1076, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532248.

Wallenstein:1993:PMH

[753] Sylvan Wallenstein and Janet Wittes. The power of the Mantel–Haenszel test for grouped failure time data. Biometrics, 49(4):1077–1087, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532249.

Eslava-Gomez:1993:CRG

[754] G. Eslava-Gomez and F. H. C. Marriott. Criteria to represent groups in the plane when the grouping is known. Biometrics, 49(4):1088–1098, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532250.

Heitjan:1993:ICD

[755] Daniel F. Heitjan. Ignorability and coarse data: Some biomedical examples. Biometrics, 49(4):1099–1109, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532251.

Magder:1993:AID

[756] Laurence Magder and Ron Brookmeyer. Analysis of infectious disease data from partner studies with unknown source of infection. Biometrics, 49(4):1110–1116, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532252.

Storer:1993:SSC

[757] Barry E. Storer. Small-sample confidence sets for the MTD in a Phase I clinical trial. Biometrics, 49(4):1117–1125, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532253.

Ryan:1993:UHC

[758] Louise Ryan. Using historical controls in the analysis of developmental toxicity data. Biometrics, 49(4):1126–1135, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532254.

Ridout:1993:GOH

[759] M. S. Ridout, J. S. Fenlon, and P. R. Hughes. A generalized one-hit model for bioassays of insect viruses. Biometrics, 49(4):1136–1141, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532255.

Fraile:1993:ACD

[760] Laurent Fraile, Yves Escoufier, and André Raibaut. Analyse des correspondances de données planifiées: Étude de la chémotaxie de la larve infestante d’un parasite. (French) [Planned data correspondence analysis: Study of the chemotaxis of the infesting larva of a parasite]. Biometrics, 49(4):1142–1153, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532256.

Iwaisaki:1993:RGP

[761] Hiroaki Iwaisaki and James W. Wilton. Regression of genotypic on phenotypic value of a ratio-defined character. Biometrics, 49(4):1154–1163, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532257.

Bromaghin:1993:WNS

[762] Jeffrey F. Bromaghin and Lyman L. McDonald. Weighted nest survival models. Biometrics, 49(4):1164–1172, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532258.

Brownie:1993:CRS

[763] C. Brownie, J. E. Hines, J. D. Nichols, K. H. Pollock, and J. B. Hestbeck. Capture–recapture studies for multiple strata including non-Markovian transitions. Biometrics, 49(4):1173–1187, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532259.

Cormack:1993:VMR

[764] R. M. Cormack. Variances of mark-recapture estimates. Biometrics, 49(4):1188–1193, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532260.

Burnham:1993:MHS

[765] Kenneth P. Burnham and Eric A. Rexstad. Modeling heterogeneity in survival rates of banded waterfowl. Biometrics, 49(4):1194–1208, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532261.

Muttlak:1993:NLI

[766] Hassen A. Muttlak and S. M. Sadooghi-Alvandi. A note on the line intercept sampling method. Biometrics, 49(4):1209–1215, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532262.

Lui:1993:SGO

[767] Kung-Jong Lui. A simple generalization of the O’Brien and Fleming group sequential test procedure to more than two treatment groups. Biometrics, 49(4):1216–1219, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532263.

Le:1993:TLT

[768] Chap T. Le. A test for linear trend in constant hazards and its application to a problem in occupational health. Biometrics, 49(4):1220–1224, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532264.

Munk:1993:ICU

[769] Axel Munk. An improvement on commonly used tests in bioequivalence assessment. Biometrics, 49(4):1225–1230, December 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532265.

Drake:1993:EMP

[770] Christiana Drake. Effects of misspecification of the propensity score on estimators of treatment effect. Biometrics, 49(4):1231–1236, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532266.

Chadoeuf:1993:ALP

[771] Joel Chadoeuf, Andre Kretzschmar, Michel Goulard, and Keith Smettem. Analyse de la liaison entre un processus de galéries et la fissuration d’un sol. (French) [Analysis of the link between a galleries process and soil cracking]. Biometrics, 49(4):1237–1244, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532267.

Woolliams:1993:AMD

[772] J. A. Woolliams, C. S. Haley, and K. Lange. Asymptotic moments and distribution of nearest marker distances when major genes and marker polymorphisms are unevenly distributed. Biometrics, 49(4):1245–1251, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532268.

Azais:1993:CEN

[773] J.-M. Azaïs, R. A. Bailey, and H. Monod. A catalogue of efficient neighbour-designs with border plots. Biometrics, 49(4):1252–1261, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532269.

Schaalje:1993:SEI

[774] G. Bruce Schaalje and Richard A. Butts. Some effects of ignoring correlated measurement errors in straight line regression and prediction. Biometrics, 49(4):1262–1267, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532270.

Becher:1993:BHT

[775] Heiko Becher, Peter Hall, and Susan R. Wilson. Bootstrap hypothesis testing procedures. Comment on: “Two guidelines for bootstrap hypothesis testing” [Biometrics 47 (1991), no. 2, 757–762; MR1132543 (92f:62061)] by P. G. Hall and S. R. Wilson with a reply by Hall and Wilson. Biometrics, 49(4):1268–1272, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532271. See [307].

Anonymous:1993:CF

[776] Anonymous. The consultant’s forum. Biometrics, 49(4):1273, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532272.

Mandel:1993:TWT

[777] John Mandel. Two-way tables with no replication. Biometrics, 49 (4):1275–1276, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532273.

Snee:1993:E

[778] Ronald D. Snee, Govind S. Mudholkar, and I. C. Sarkar. [Editorial]. Biometrics, 49(4):1276–1278, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532274.

Piepho:1993:TH

[779] Hans-Peter Piepho, Govind S. Mudholkar, and I. C. Sarkar. Testing for homoscedasticity. Biometrics, 49(4):1279–1280, December 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532275.

Bjornsson:1993:BRD

[780] H. Björnsson. Book review: The Design and Analysis of Research Studies, by B. F. J. Manly. Biometrics, 49(4):1281–1282, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532276.

Kocherlakota:1993:BRU

[781] S. Kocherlakota. Book review: Univariate Discrete Distributions, by N. L. Johnson, S. Kotz, A. W. Kemp. Biometrics, 49(4):1282, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532277.

Ord:1993:BRB

[782] J. K. Ord. Book review: Bivariate Discrete Distributions, by S. Kocherlakota, K. Kocherlakota. Biometrics, 49(4):1282–1283, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532278.

Cormack:1993:BRS

[783] R. M. Cormack. Book review: Sampling, by S. K. Thompson. Biometrics, 49(4):1283–1284, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532279.

Rosenberger:1993:BRC

[784] J. L. Rosenberger. Book review: Configural Polysampling: a Route to Practical Robustness, by S. Morgenthaler, J. W. Tukey. Biometrics, 49(4):1284, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532280.

Kemp:1993:BRS

[785] A. W. Kemp. Book review: Statistics for the Environment, by V. Barnett, K. F. Turkman. Biometrics, 49(4):1284–1285, December 1993. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532281.

Anonymous:1993:BRL

[786] Anonymous. Book review: Likelihood, Expanded Edition, by A. W. F. Edwards. Biometrics, 49(4):1285, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532282.

Anonymous:1993:BRSc

[787] Anonymous. Book review: Statistical Methods in Soil and Land Resource Survey, by R. Webster, M. A. Oliver. Biometrics, 49(4):1285–1286, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532283.

Anonymous:1993:BRG

[788] Anonymous. Book review: Genetic Issues in Psychosocial Epidemiology, by M. T. Tsuang, K. S. Kendler, M. J. Lyons. Biometrics, 49(4):1286, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532286.

Anonymous:1993:BRMe

[789] Anonymous. Book review: Medical Statistics: a Commonsense Approach, by M. J. Campbell, D. Machin. Biometrics, 49(4):1286, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532284.

Anonymous:1993:BRMf

[790] Anonymous. Book review: Medical Statistics on Microcomputers: a Guide to the Appropriate Use of Statistical Packages, by R. A. Brown, J. S. Beck. Biometrics, 49(4):1286, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532285.

Anonymous:1993:BRTc

[791] Anonymous. Book review: Time Series: a Biostatistical Introduction, by P. J. Diggle. Biometrics, 49(4):1286, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532287.

Anonymous:1993:BRTd

[792] Anonymous. Book review: Time-Frequency Signal Analysis: Methods and Applications, by B. Boashash. Biometrics, 49(4):1286–1287, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532288.

Anonymous:1993:BRId

[793] Anonymous. Book review: Inspection Errors for Attributes in Quality Control, by N. L. Johnson, S. Kotz, X. Wu. Biometrics, 49(4):1287, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532289.

Anonymous:1993:BRW

[794] Anonymous. Book reviews: World Health Statistics Annual 1992, by World Health Organization; World Health Statistics Quarterly 1992, by World Health Organization. Biometrics, 49(4):1287, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532290.

Utts:1993:OFN

[795] Jessica Utts. Obituary: Florence Nightingale David (1909–1993). Biometrics, 49(4):1289–1291, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532291.

Anonymous:1993:PPBd

[796] Anonymous. Papers to be published in biometrics. Biometrics, 49(4): 1293, December 1993. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532292.

Anonymous:1994:FMa

[797] Anonymous. Front matter. Biometrics, 50(1):i–iv, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533191.

Hoover:1994:NRP

[798] Donald R. Hoover and Yanhua He. Nonidentified responses in a proportional hazards setting. Biometrics, 50(1):1–10, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533192.

Lipsitz:1994:WLS

[799] Stuart R. Lipsitz, Nan M. Laird, and David P. Harrington. Weighted least squares analysis of repeated categorical measurements with outcomes subject to nonresponse. Biometrics, 50(1):11–24, March 1994.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533193.

Slud:1994:AFS

[800] Eric V. Slud. Analysis of factorial survival experiments. Biometrics, 50(1):25–38, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533194.

Mori:1994:SEP

[801] Motomi Mori, Robert F. Woolson, and George G. Woodworth. Slope estimation in the presence of informative right censoring: Modeling the number of observations as a geometric random variable. Biometrics, 50(1): 39–50, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533195.

Albert:1994:MMS

[802] Paul S. Albert. A Markov model for sequences of ordinal data from a relapsing-remitting disease. Biometrics, 50(1):51–60, March 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533196.

Muller:1994:HRE

[803] Hans-Georg Müller and Jane-Ling Wang. Hazard rate estimation under random censoring with varying kernels and bandwidths. Biometrics, 50(1):61–76, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533197.

Petroni:1994:TST

[804] Gina R. Petroni and Robert A. Wolfe. A two-sample test for stochastic ordering with interval-censored data. Biometrics, 50(1):77–87, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533198.

Lee:1994:EPS

[805] Shen-Ming Lee and Anne Chao. Estimating population size via sample coverage for closed capture–recapture models. Biometrics, 50(1):88–97, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533199.

Schwarz:1994:MRE

[806] Carl James Schwarz and J. Brian Dempson. Mark-recapture estimation of a salmon smolt population. Biometrics, 50(1):98–108, March 1994.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533200.

Whittemore:1994:PGI

[807] Alice S. Whittemore and Jerry Halpern. Probability of gene identity by descent: Computation and applications. Biometrics, 50(1):109–117, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533201.

Whittemore:1994:CTL

[808] Alice S. Whittemore and Jerry Halpern. A class of tests for linkage using affected pedigree members. Biometrics, 50(1):118–127, March 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533202.

Shoukri:1994:PEG

[809] M. M. Shoukri and G. J. McLachlan. Parametric estimation in a genetic mixture model with application to nuclear family data. Biometrics, 50(1): 128–139, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533203.

Sasieni:1994:SSS

[810] Peter Sasieni. Small-sample study of an efficient estimator of the odds ratio under multiple matching. Biometrics, 50(1):140–148, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533204.

Sukhatme:1994:SNR

[811] Shashikala Sukhatme and C. A. Beam. Stratification in nonparametric ROC studies. Biometrics, 50(1):149–163, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533205.

McDonald:1994:TRE

[812] Barry W. McDonald. Two random effects models for multivariate binary data. Biometrics, 50(1):164–172, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533206.

Namgung:1994:ORO

[813] Yearnok Y. Namgung and Mark C. K. Yang. Outlier reduction by an option-3 measurement scheme. Biometrics, 50(1):173–182, March 1994.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533207.

Bodian:1994:ICT

[814] Carol A. Bodian. Intraclass correlation for two-by-two tables under three sampling designs. Biometrics, 50(1):183–193, March 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533208.

Williamson:1994:MTD

[815] G. David Williamson and Michael Haber. Models for three-dimensional contingency tables with completely and partially cross-classified data. Biometrics, 50(1):194–203, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533209.

Gomez:1994:EIT

[816] Guadalupe Gómez and Stephen W. Lagakos. Estimation of the infection time and latency distribution of AIDS with doubly censored data. Biometrics, 50(1):204–212, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533210.

Le:1994:ABS

[817] Chap T. Le, Patricia M. Grambsch, and Thomas A. Louis. Association between survival time and ordinal covariates. Biometrics, 50(1):213–219, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533211.

Bailer:1994:MBT

[818] A. John Bailer and Randall J. Smith. Model-based time extrapolation for quantal response studies. Biometrics, 50(1):220–225, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533212.

Bedrick:1994:MSM

[819] Edward J. Bedrick and Chih-Ling Tsai. Model selection for multivariate regression in small samples. Biometrics, 50(1):226–231, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533213.

Lui:1994:POF

[820] Kung-Jong Lui. The performance of the O’Brien-Fleming multiple testing procedure in the presence of intraclass correlation. Biometrics, 50(1):

232–236, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533214.

Viana:1994:BSS

[821] M. A. G. Viana. Bayesian small-sample estimation of misclassified multinomial data. Biometrics, 50(1):237–243, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533215.

Gray:1994:BAI

[822] Robert J. Gray. A Bayesian analysis of institutional effects in a multicenter cancer clinical trial. Biometrics, 50(1):244–253, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533216.

Green:1994:BET

[823] Edwin J. Green, Francis A. Roesch, Jr., Adrian F. M. Smith, and William E. Strawderman. Bayesian estimation for the three-parameter Weibull distribution with tree diameter data. Biometrics, 50(1):254–269, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533217.

Lipsitz:1994:PGE

[824] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Endel J. Orav, and Nan M. Laird. Performance of generalized estimating equations in practical situations. Biometrics, 50(1):270–278, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533218.

Minkin:1994:SAA

[825] Salomon Minkin. Statistical analysis of aberrant crypt assays for colon cancer promotion studies. Biometrics, 50(1):279–288, March 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533219.

Ridout:1994:CCI

[826] M. S. Ridout. A comparison of confidence interval methods for dilution series experiments. Biometrics, 50(1):289–296, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533220.

Lui:1994:ERP

[827] Kung-Jong Lui. The effect of retaining probability variation on sample size calculations for normal variates. Biometrics, 50(1):297–300, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533221.

Gazey:1994:PSE

[828] W. J. Gazey and Anne Chao. Population size estimation for sparse data. Biometrics, 50(1):301–305, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533222.

Golbeck:1994:NPC

[829] A. L. Golbeck and B. S. Everitt. A note on partitioning 2×C contingency tables into non-independent 2 × 2 tables. Biometrics, 50(1):305–307, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533223.

Hung:1994:TED

[830] H. M. James Hung. Testing for the existence of a desirable dose combination. Biometrics, 50(1):307–308, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533224.

Mangin:1994:STG

[831] B. Mangin, B. Goffinet, M. M. Shoukri, and G. M. Lathrop. Statistical testing in genetic linkage under heterogeneity. Biometrics, 50(1):308– 309, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533225.

Garner:1994:IEK

[832] Barry Garner, Cecilia A. Hale, and Joseph L. Fleiss. Interval estimation for kappa. Biometrics, 50(1):309–310, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533226.

Denis:1994:BM

[833] Jean-Baptiste Denis and John C. Gower. Biadditive models. Biometrics, 50(1):310–311, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533227.

Brown:1994:BRE

[834] R. A. Brown. Book review: Evaluation and Control of Measurements, by J. Mandel. Biometrics, 50(1):313–314, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533228.

Read:1994:BRC

[835] C. B. Read. Book review: Conditioning Diagnostics: Collinearity and Weak Data in Regression, by D. A. Belsley. Biometrics, 50(1):314–315, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533229.

Buckland:1994:BRT

[836] S. T. Buckland. Book review: Techniques for Wildlife Investigation: Design and Analysis of Capture Data, by J. R. Skalski, D. S. Robson. Biometrics, 50(1):315, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533230.

Kemp:1994:BRMa

[837] A. W. Kemp. Book review: Modelling Biological Populations in Space and Time, by E. Renshaw. Biometrics, 50(1):315–316, March 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533231.

Edwards:1994:BRCa

[838] A. W. F. Edwards. Book review: The Children of Atomic Bomb Survivors: a Genetic Study, by J. V. Neel, W. J. Schull. Biometrics, 50(1): 316, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533232.

Curnow:1994:BRS

[839] R. N. Curnow. Book review: Statistics in Toxicology, by D. Krewski, C. Franklin. Biometrics, 50(1):316–317, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533233.

Fearn:1994:BRDa

[840] T. Fearn. Book review: Design Statistics in Pharmacochemistry, by P. P. Mager. Biometrics, 50(1):317, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533234.

Glasbey:1994:BRD

[841] C. A. Glasbey. Book reviews: Discriminant Analysis and Statistical Pattern Recognition, by G. McLachlan; Fundamentals of Pattern Recognition, by M. Pavel. Biometrics, 50(1):317–318, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533235.

Anonymous:1994:BRAa

[842] Anonymous. Book review: Applied Nonparametric Statistical Methods, by P. Sprent. Biometrics, 50(1):318, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533237.

Anonymous:1994:BRB

[843] Anonymous. Book review: Biological Data Analysis: a Practical Approach, by J. C. Fry. Biometrics, 50(1):318, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533236.

Anonymous:1994:BRAb

[844] Anonymous. Book review: Advances in Computer Methods for Systematic Biology: Artificial Intelligence, Data-Bases, Computer Vision, by R. Fortuner. Biometrics, 50(1):319, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533239.

Anonymous:1994:BRDa

[845] Anonymous. Book reviews: The Datasets from CBDEA, by T. P. Hutchinson, C. D. Lai; Essentials of Statistical Methods, by T. P. Hutchinson. Biometrics, 50(1):319, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533240.

Anonymous:1994:BRFa

[846] Anonymous. Book review: Fisheries Management via Management Science, by B. L. Golden, E. A. Wasil. Biometrics, 50(1):319, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533241.

Anonymous:1994:BRSa

[847] Anonymous. Book review: Statistics for Spatial Data, Revised Edition, by N. A. C. Cressie. Biometrics, 50(1):319, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533238.

Anonymous:1994:BRSb

[848] Anonymous. Book review: Statistical Analysis of Reliability Data, by M. J. Crowder, A. C. Kimber, R. L. Smith, T. J. Sweeting. Biometrics, 50(1):319–320, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533242.

Anonymous:1994:BRSc

[849] Anonymous. Book review: Statistics in Research and Development, by R. Caulcutt. Biometrics, 50(1):320, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533243.

Anonymous:1994:CDR

[850] Anonymous. Corrections: A deviance residual for heterogeneous spatial Poisson processes. Biometrics, 50(1):321, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533244. See [713].

Anonymous:1994:CSM

[851] Anonymous. Corrections: A stabilized moment estimator for the betabinomial distribution. Biometrics, 50(1):321, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533245. See [1].

Anonymous:1994:CBS

[852] Anonymous. Corrections: Bayesian subset analysis. Biometrics, 50(1): 322, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533247. See [337].

Anonymous:1994:CEP

[853] Anonymous. Corrections: Estimating the probability of toxicity at the recommended dose following a Phase I clinical trial in cancer. Biometrics, 50(1):322, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533248. See [530].

Anonymous:1994:CRB

[854] Anonymous. Corrections: A rank-based mixed model approach to multisite clinical trials. Biometrics, 50(1):322, March 1994. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533246. See [441].

Anonymous:1994:PPBa

[855] Anonymous. Papers to be published in biometrics. Biometrics, 50(1): 323, March 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533249.

Anonymous:1994:FMb

[856] Anonymous. Front matter. Biometrics, 50(2):i–iv, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533375.

Follmann:1994:MPC

[857] Dean A. Follmann, Michael A. Proschan, and Nancy L. Geller. Monitoring pairwise comparisons in multi-armed clinical trials. Biometrics, 50(2): 325–336, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533376.

Thall:1994:PBG

[858] Peter F. Thall and Richard Simon. Practical Bayesian guidelines for Phase IIB clinical trials. Biometrics, 50(2):337–349, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533377.

Wacholder:1994:FML

[859] S. Wacholder and C. R. Weinberg. Flexible maximum likelihood methods for assessing joint effects in case-control studies with complex sampling. Biometrics, 50(2):350–357, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533378.

Weinberg:1994:MRT

[860] Clarice R. Weinberg, Beth C. Gladen, and Allen J. Wilcox. Models relating the timing of intercourse to the probability of conception and the sex of the baby. Biometrics, 50(2):358–367, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533379.

Rosenbaum:1994:COS

[861] Paul R. Rosenbaum. Coherence in observational studies. Biometrics, 50(2):368–374, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533380.

Clobert:1994:EAS

[862] Jean Clobert, Jean-Dominique Lebreton, Dominique Allaine, and J. M. Gaillard. The estimation of age-specific breeding probabilities from recaptures or resightings in vertebrate populations: II. longitudinal models. Biometrics, 50(2):375–387, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533381.

Evans:1994:BRM

[863] Marc A. Evans and Douglas G. Bonett. Bias reduction for multiplerecapture estimators of closed population size. Biometrics, 50(2):388– 395, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533382.

Evans:1994:GTM

[864] Marc A. Evans, Douglas G. Bonett, and Lyman L. McDonald. A general theory for modeling capture–recapture data from a closed population. Biometrics, 50(2):396–405, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533383.

Anderson-Sprecher:1994:REW

[865] Richard Anderson-Sprecher. Robust estimates of wildlife location using telemetry data. Biometrics, 50(2):406–416, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533384.

Guo:1994:MCE

[866] Sun Wei Guo and Elizabeth A. Thompson. Monte Carlo estimation of mixed models for large complex pedigrees. Biometrics, 50(2):417– 432, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533385.

Lie:1994:MLE

[867] Rolv Terje Lie, Ivar Heuch, and Lorentz M. Irgens. Maximum likelihood estimation of the proportion of congenital malformations using double registration systems. Biometrics, 50(2):433–444, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533386.

Nunez-Anton:1994:ALD

[868] Vicente Núñez-Antón and George G. Woodworth. Analysis of longitudinal data with unequally spaced observations and time-dependent correlated errors. Biometrics, 50(2):445–456, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533387.

Gray:1994:BMM

[869] Gerry Gray. Bias in misspecified mixtures. Biometrics, 50(2):457–470, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533388.

Ahn:1994:TSP

[870] Hongshik Ahn and Wei-Yin Loh. Tree-structured proportional hazards regression modeling. Biometrics, 50(2):471–485, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533389.

Chi:1994:MEC

[871] Eric M. Chi. M-estimation in cross-over trials. Biometrics, 50(2):486– 493, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533390.

Agresti:1994:SCR

[872] Alan Agresti. Simple capture–recapture models permitting unequal catchability and variable sampling effort. Biometrics, 50(2):494–500, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533391.

Hirst:1994:IRM

[873] David Hirst. An improved removal method for estimating animal abundance. Biometrics, 50(2):501–505, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533392.

Li:1994:TSM

[874] W. K. Li. Time series models based on generalized linear models: Some further results. Biometrics, 50(2):506–511, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533393.

Loisel:1994:DMG

[875] Patrice Loisel, Bruno Goffinet, Hervé Monod, and Gustavo Montes De Oca. Detecting a major gene in an F 2 population. Biometrics, 50(2): 512–516, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533394.

Commenges:1994:ICC

[876] Daniel Commenges and Hélène Jacqmin. The intraclass correlation coefficient: Distribution-free definition and test. Biometrics, 50(2):517–526, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533395.

Metz:1994:SWT

[877] Johan A. J. Metz, Patsy Haccou, and Evert Meelis. On the Shapiro– Wilk test and Darling’s test for exponentiality. Biometrics, 50(2):527– 530, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533396.

Assuncao:1994:TSR

[878] Renato Assunção. Testing spatial randomness by means of angles. Biometrics, 50(2):531–537, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533397.

Best:1994:NCT

[879] D. J. Best. Nonparametric comparison of two histograms. Biometrics, 50(2):538–541, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533398.

Brownie:1994:TER

[880] Cavell Brownie and Dennis D. Boos. Type I error robustness of ANOVA and ANOVA on ranks when the number of treatments is large. Biometrics, 50(2):542–549, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533399.

Donner:1994:SIC

[881] Allan Donner and Michael Eliasziw. Statistical implications of the choice between a dichotomous or continuous trait in studies of interobserver agreement. Biometrics, 50(2):550–555, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533400.

Solow:1994:DCC

[882] Andrew R. Solow. Detecting change in the composition of a multispecies community. Biometrics, 50(2):556–565, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533401.

Yuh:1994:PPP

[883] Lianng Yuh, Stuart Beal, Marie Davidian, Ferrin Harrison, Allen Hester, Kenneth Kowalski, Edward Vonesh, and Russell Wolfinger. Population pharmacokinetic/pharmacodynamic methodology and applications: a bibliography. Biometrics, 50(2):566–575, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533402.

Douglas:1994:EFD

[884] J. B. Douglas, Brian Leroux, and Martin L. Puterman. Empirical fitting of discrete distributions. Biometrics, 50(2):576–579, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533403.

Heitjan:1994:EMD

[885] Daniel F. Heitjan and Peter J. Diggle. Estimation with missing data. Biometrics, 50(2):580, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533404.

Lehmacher:1994:CDA

[886] Walter Lehmacher, Gernot Wassmer, and Peter Reitmeir. Comment on: On the design and analysis of randomized clinical trials with multiple endpoints. Biometrics, 50(2):581–583, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533405. See [603].

Bailey:1994:BRD

[887] R. A. Bailey. Book review: Design of Experiments: a No-Name Approach, by T. J. Lorenzen, V. L. Anderson. Biometrics, 50(2):584–585, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533406.

Edmondson:1994:BRA

[888] R. N. Edmondson. Book review: Analysis of Variance in Experimental Design, by H. R. Lindman. Biometrics, 50(2):585–586, June 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533407.

Jones:1994:BRC

[889] B. Jones. Book reviews: Cross-Over Experiments: Design, Analysis, and Application, by D. A. Ratkowski, M. A. Evans, J. R. Alldredge; Cross-Over Trials in Clinical Research, by S. Senn. Biometrics, 50(2): 586, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533408.

Buckland:1994:BRC

[890] S. T. Buckland. Book review: Computer Intensive Statistical Methods: Validation Model Selection and Bootstrap, by J. S. U. Hjorth. Biometrics, 50(2):586–587, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533409.

Harkness:1994:BRM

[891] W. L. Harkness. Book review: Modelling Binary Data, by D. Collett. Biometrics, 50(2):587–588, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533410.

Trenkel:1994:BRI

[892] V. Trenkel. Book review: Introduction à la Biométrie, by P. Jolicoeur. Biometrics, 50(2):588, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533411.

Kemp:1994:BRMb

[893] A. W. Kemp. Book review: Medical Uses of Statistics, by J. C. Bailar III, F. Mosteller. Biometrics, 50(2):588–589, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533412.

Kirkwood:1994:BRN

[894] T. B. L. Kirkwood. Book review: Networks and Chaos-Statistical and Probabilistic Aspects, by O. E. Barndorff-Nielsen, J. L. Jensen, W. S. Kendall. Biometrics, 50(2):589, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533413.

Abakuks:1994:BRD

[895] A. Abakuks. Book review: The Dynamics of Cellular Motility, by M. Murase. Biometrics, 50(2):589–590, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533414.

Kemp:1994:BRL

[896] C. D. Kemp. Book reviews: LaTEX for Everyone: a Reference Guide and Tutorial for Typesetting Documents Using a Computer, by J. Hahn; LaTEX, Concisely, by A. Johnstone; LaTEX Line by Line: Tips and Techniques for Document Processing, by A. Diller; A Guide to LaTEX: Document Preparation for Beginners and Advanced Users, by H. Kopka, P. W. Daly. Biometrics, 50(2):590–591, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533415.

Anonymous:1994:BRSd

[897] Anonymous. Book review: Solutions in Statistics and Probability, by E. J. Dudewicz. Biometrics, 50(2):591, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533416.

Anonymous:1994:BRRa

[898] Anonymous. Book review: La Régression: Nouveaux Regards sur une Ancienne Méthode Statistique, by R. Tomassone, S. Audrain, E. Lesquoyde Turckheim, C. Miller. Biometrics, 50(2):591–592, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533417.

Anonymous:1994:BRAc

[899] Anonymous. Book review: Applied Nonparametric Regression, by W. Härdle. Biometrics, 50(2):592, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533418.

Anonymous:1994:BRDb

[900] Anonymous. Book review: The Design and Analysis of Sequential Clinical Trials, by J. Whitehead. Biometrics, 50(2):592, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533419.

Anonymous:1994:BRL

[901] Anonymous. Book review: Location Modeling in Practice: Applications, Theory, and History, by B. L. Golden, H. A. Eiselt. Biometrics, 50(2): 592, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533422.

Anonymous:1994:BRSe

[902] Anonymous. Book review: Statistical Analysis of Spherical Data, by N. I. Fisher, T. Lewis, B. J. J. Embleton. Biometrics, 50(2):592, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533420.

Anonymous:1994:BRSf

[903] Anonymous. Book review: Spatial Interpolation, by A. Stein. Biometrics, 50(2):592, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533421.

Anonymous:1994:BRN

[904] Anonymous. Book review: A National Health Care Survey, by National Research Council Institute of Medicine. Biometrics, 50(2):592–593, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533423.

Anonymous:1994:BRPa

[905] Anonymous. Book review: Point Processes and Their Statistical Inference, by A. F. Karr. Biometrics, 50(2):593, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533425.

Anonymous:1994:BRRb

[906] Anonymous. Book review: A Reader on Applying Statistics in Public Health and Prevention, by J. L. A. Rijckevorsel, C. C. J. H. Bijleveld. Biometrics, 50(2):593, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533424.

Anonymous:1994:PPBb

[907] Anonymous. Papers to be published in biometrics. Biometrics, 50(2): 595, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533426.

Anonymous:1994:BSF

[908] Anonymous. The Biometric Society financial statements and independent auditors’ report January 31, 1994. Biometrics, 50(2):597–600, June 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533427.

Anonymous:1994:FMc

[909] Anonymous. Front matter. Biometrics, 50(3):i–iv, September 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532774.

Fitzmaurice:1994:AIL

[910] Garrett M. Fitzmaurice, Nan M. Laird, and Stuart R. Lipsitz. Analysing incomplete longitudinal binary responses: a likelihood-based approach. Biometrics, 50(3):601–612, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532775.

Commenges:1994:THB

[911] Daniel Commenges, Luc Letenneur, Hélène Jacqmin, Thierry Moreau, and Jean-François Dartigues. Test of homogeneity of binary data with explanatory variables. Biometrics, 50(3):613–620, September 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532776.

Cole:1994:PAQ

[912] Bernard F. Cole, Richard D. Gelber, and Keaven M. Anderson. Parametric approaches to quality-adjusted survival analysis. Biometrics, 50(3):621–631, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532777.

Guo:1994:RAM

[913] S. W. Guo and D. Y. Lin. Regression analysis of multivariate grouped survival data. Biometrics, 50(3):632–639, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532778.

Gray:1994:SBT

[914] Robert J. Gray. Spline-based tests in survival analysis. Biometrics, 50(3):640–652, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532779.

Kelly:1994:TMM

[915] Colleen Kelly. A test of the Markovian model of DNA evolution. Biometrics, 50(3):653–664, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532780.

Olson:1994:REG

[916] Jane M. Olson. Robust estimation of gene frequency and association parameters. Biometrics, 50(3):665–674, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532781.

Satten:1994:EIH

[917] Glen A. Satten and Ira M. Longini, Jr. Estimation of incidence of HIV infection using cross-sectional marker surveys. Biometrics, 50(3):675– 688, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532782.

Zeger:1994:SML

[918] Scott L. Zeger and Peter J. Diggle. Semiparametric models for longitudinal data with application to CD4 cell numbers in HIV seroconverters. Biometrics, 50(3):689–699, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532783.

Cressie:1994:SSO

[919] Noel Cressie and Jonathan Biele. A sample-size-optimal Bayesian procedure for sequential pharmaceutical trials. Biometrics, 50(3):700–711, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532784.

Thompson:1994:DCA

[920] Steven K. Thompson and George A. F. Seber. Detectability in conventional and adaptive sampling. Biometrics, 50(3):712–724, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532785.

Pennington:1994:AEI

[921] Michael Pennington and Jon Helge Vølstad. Assessing the effect of intrahaul correlation and variable density on estimates of population characteristics from marine surveys. Biometrics, 50(3):725–732, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532786.

Link:1994:DEU

[922] William A. Link and Richard J. Barker. Density estimation using the trapping web design: a geometric analysis. Biometrics, 50(3):733–745, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532787.

Rice:1994:OHF

[923] William R. Rice and Steven D. Gaines. The ordered-heterogeneity family of tests. Biometrics, 50(3):746–752, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532788.

Shao:1994:SIS

[924] Jun Shao and Shein-Chung Chow. Statistical inference in stability analysis. Biometrics, 50(3):753–763, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532789.

Taplin:1994:AAF

[925] Ross H. Taplin and Adrian E. Raftery. Analysis of agricultural field trials in the presence of outliers and fertility jumps. Biometrics, 50(3):764–781, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532790.

Lai:1994:SAL

[926] Tze Leung Lai and Limin Zhang. Statistical analysis of ligand-binding experiments. Biometrics, 50(3):782–797, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532791.

Normolle:1994:IAS

[927] Daniel P. Normolle and Morton B. Brown. Identification of aperiodic seasonality in non-Gaussian time series. Biometrics, 50(3):798–812, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532792.

Viana:1994:CML

[928] Marlos A. G. Viana. Combined maximum likelihood estimates for the equicorrelation coefficient. Biometrics, 50(3):813–820, September 1994.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532793.

Baker:1994:RAG

[929] Stuart G. Baker. Regression analysis of grouped survival data with incomplete covariates: Nonignorable missing-data and censoring mechanisms. Biometrics, 50(3):821–826, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532794.

Chang:1994:IAR

[930] Myron N. Chang and Jonathan J. Shuster. Interim analysis for randomized clinical trials: Simulating the predictive distribution of the logrank test statistic. Biometrics, 50(3):827–833, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532795.

Laska:1994:SDM

[931] Eugene M. Laska, Morris Meisner, and Carole Siegel. Simple designs and model-free tests for synergy. Biometrics, 50(3):834–841, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532796.

Lipsitz:1994:JEV

[932] Stuart R. Lipsitz, Keith B. G. Dear, and Lueping Zhao. Jackknife estimators of variance for parameter estimates from estimating equations with applications to clustered survival data. Biometrics, 50(3):842–846, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532797.

Lipsitz:1994:EYM

[933] Stuart R. Lipsitz and Garrett Fitzmaurice. An extension of Yule’s Q to multivariate binary data. Biometrics, 50(3):847–852, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532798.

Silvapulle:1994:TAO

[934] Mervyn J. Silvapulle. On tests against one-sided hypotheses in some generalized linear models. Biometrics, 50(3):853–858, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532799.

Yanagawa:1994:MHT

[935] T. Yanagawa, T. Tango, and Y. Hiejima. Mantel–Haenszel-type tests for testing equivalence or more than equivalence in comparative clinical trials. Biometrics, 50(3):859–864, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532800. See correction [1087].

Formann:1994:MEC

[936] Anton K. Formann. Measurement errors in caries diagnosis: Some further latent class models. Biometrics, 50(3):865–871, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532801.

Liang:1994:UQL

[937] Kung-Yee Liang and John Hanfelt. On the use of the quasi-likelihood method in teratological experiments. Biometrics, 50(3):872–880, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532802.

Jeffers:1994:IED

[938] J. N. R. Jeffers. The importance of exploratory data analysis before the use of sophisticated procedures. Biometrics, 50(3):881–883, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532803.

Nam:1994:MCD

[939] Yong W. Nam and Axel Munk. On a method of combining double t-test and anderson-hauck test. Biometrics, 50(3):884–886, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532804.

Bailey:1994:BRB

[940] N. T. J. Bailey. Book review: Biometrie: Modelisation de Phenomenes Biologiques, by R. Tomassone, C. Dervin, J.-P Masson. Biometrics, 50(3):887–888, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532805.

Gonin:1994:BRB

[941] R. Gonin. Book review: Biostatistics: a Methodology for the Health Sciences, by L. D. Fisher, G. van Belle. Biometrics, 50(3):888–889, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532806.

Dugard:1994:BRS

[942] P. I. Dugard. Book review: Statistical Applications for the Behavioral Sciences, by L. G. Grimm. Biometrics, 50(3):889, September 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532807.

Cox:1994:BRS

[943] D. R. Cox. Book review: Statistical Models for Causal Analysis, by R. D. Retherford, M. K. Choe. Biometrics, 50(3):889–890, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532808.

Derr:1994:BRV

[944] J. A. Derr. Book review: Visualizing Data, by W. S. Cleveland. Biometrics, 50(3):890, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532809.

Buckland:1994:BRI

[945] S. T. Buckland. Book review: An Introduction to the Bootstrap, by B. Efron, R. J. Tibshirani. Biometrics, 50(3):890–891, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532810.

Krzanowski:1994:BRM

[946] W. J. Krzanowski. Book review: Multivariate Density Estimation: Theory, Practice, and Visualization, by D. W. Scott. Biometrics, 50(3):891, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532811.

Thompson:1994:BRD

[947] S. K. Thompson. Book review: Distance Sampling: Estimating Abundance of Biological Populations, by S. T. Buckland, D. R. Anderson, K. P. Burnham, J. L. Laake. Biometrics, 50(3):891–892, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532812.

Kemp:1994:BRH

[948] A. W. Kemp. Book review: Handbook of the Logistic Distribution, by N. Balakrishnan. Biometrics, 50(3):892–893, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532813.

Kocherlakota:1994:BRI

[949] S. Kocherlakota. Book review: Identifiability in Stochastic Models: Characterization of Probability Distributions, by B. L. S. P. Rao. Biometrics, 50(3):893, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532814.

Reyment:1994:BRP

[950] R. A. Reyment. Book review: Problems of Relative Growth, by J. S. Huxley. Biometrics, 50(3):893–894, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532815.

Borchers:1994:BRR

[951] D. Borchers. Book review: Resource Selection, by Animals, by B. Manly, L. McDonald, D. Thomas. Biometrics, 50(3):894–895, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532816.

Goulding:1994:BRD

[952] K. W. T. Goulding. Book review: Dynamics of Nutrient Cycling and Food Webs, by D. L. DeAngelis. Biometrics, 50(3):895, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532817.

Anonymous:1994:BRSg

[953] Anonymous. Book review: Statistique Théorique et Appliquée, Tome 1: Les Bases Théoriques, by P. Dagnelie. Biometrics, 50(3):895–896, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532818.

Anonymous:1994:BRC

[954] Anonymous. Book reviews: The Chronological Annotated Bibliography of Order Statistics, Volume VII: 1968–1969, by H. L. Harter; The Chronological Annotated Bibliography of Order Statistics, Volume VIII: Indices with a Supplement on 1970–1992, by H. L. Harter. Biometrics, 50(3):896, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532820.

Anonymous:1994:BRE

[955] Anonymous. Book review: Evolution Equations, Control Theory, and Biomathematics, by P. Clément, G. Lumer. Biometrics, 50(3):896, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532823.

Anonymous:1994:BRFb

[956] Anonymous. Book review: The Frontiers of Modern Statistical Inference Procedures, II, by E. Bofinger, E. J. Dudewicz, G. J. Lewis, K. Mengersen. Biometrics, 50(3):896, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532819.

Anonymous:1994:BRIa

[957] Anonymous. Book review: Item Response Theory: Parameter Estimation Techniques, by F. B. Baker. Biometrics, 50(3):896, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532822.

Anonymous:1994:BRSh

[958] Anonymous. Book review: Statistical Methods in Analytical Chemistry, by P. C. Meier, R. E. Zünd. Biometrics, 50(3):896, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532821.

Anonymous:1994:BRG

[959] Anonymous. Book review: Guidelines for Laboratory Quality Auditing, by D. C. Singer, R. P. Upton. Biometrics, 50(3):897, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532827.

Anonymous:1994:BRIb

[960] Anonymous. Book review: Immunoassay Automation: a Practical Guide, by D. W. Chan. Biometrics, 50(3):897, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532826.

Anonymous:1994:BRMa

[961] Anonymous. Book review: Multidimensional Palaeobiology, by R. A. Reyment. Biometrics, 50(3):897, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532824.

Anonymous:1994:BRPb

[962] Anonymous. Book review: Practical Risk Theory for Actuaries, by C. D. Daykin, T. Pentikäinen, M. Pesonen. Biometrics, 50(3):897, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532828.

Anonymous:1994:BRSi

[963] Anonymous. Book review: The Shape of Powder-Particle Outlines, by A. E. Hawkins. Biometrics, 50(3):897, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532825.

Anonymous:1994:PPBc

[964] Anonymous. Papers to be published in biometrics. Biometrics, 50(3):898, September 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532829.

Anonymous:1994:FMd

[965] Anonymous. Front matter. Biometrics, 50(4):i–iv, December 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533430.

Anonymous:1994:VI

[966] Anonymous. Volume information. Biometrics, 50(4):vii–xxiv, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533429.

Billard:1994:WB

[967] L. Billard. The world of biometry. Biometrics, 50(4):899–916, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533431.

Aksland:1994:GCA

[968] Magnar Aksland. A general cohort analysis method. Biometrics, 50 (4):917–932, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533432.

Hedeker:1994:REO

[969] Donald Hedeker and Robert D. Gibbons. A random-effects ordinal regression model for multilevel analysis. Biometrics, 50(4):933–944, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533433.

Kenward:1994:AML

[970] M. G. Kenward, E. Lesaffre, and G. Molenberghs. An application of maximum likelihood and generalized estimating equations to the analysis of ordinal data from a longitudinal study with cases missing at random. Biometrics, 50(4):945–953, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533434.

Meester:1994:PMC

[971] Steven G. Meester and Jock MacKay. A parametric model for cluster correlated categorical data. Biometrics, 50(4):954–963, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533435.

Hirji:1994:EAP

[972] Karim F. Hirji. Exact analysis for paired binary data. Biometrics, 50(4):964–974, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533436.

Paik:1994:MSA

[973] Myunghee Cho Paik, Wei-Yann Tsai, and Ruth Ottman. Multivariate survival analysis using piecewise gamma frailty. Biometrics, 50(4):975– 988, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533437.

Dear:1994:IGL

[974] Keith B. G. Dear. Iterative generalized least squares for meta-analysis of survival data at multiple times. Biometrics, 50(4):989–1002, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533438.

DeGruttola:1994:MPC

[975] Victor De Gruttola and Xin Ming Tu. Modelling progression of CD4lymphocyte count and its relationship to survival time. Biometrics, 50 (4):1003–1014, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533439.

Tsai:1994:NAT

[976] Wei-Yann Tsai, James J. Goedert, John Orazem, Sheldon H. Landesman, Arye Rubinstein, Anne Willoughby, and Mitchell H. Gail. A nonparametric analysis of the transmission rate of human immunodeficiency virus from mother to infant. Biometrics, 50(4):1015–1028, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533440.

Bauer:1994:EEA

[977] P. Bauer and K. Köhne. Evaluation of experiments with adaptive interim analyses. Biometrics, 50(4):1029–1041, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533441. See correction [1306].

Mehta:1994:EPT

[978] Cyrus R. Mehta, Nitin Patel, Pralay Senchaudhuri, and Anastasios Tsiatis. Exact permutational tests for group sequential clinical trials. Biometrics, 50(4):1042–1053, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533442.

Amfoh:1994:ULM

[979] Kweku K. Amfoh, Richard F. Shaw, and George E. Bonney. The use of logistic models for the analysis of codon frequencies of DNA sequences in terms of explanatory variables. Biometrics, 50(4):1054–1063, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533443.

Barlow:1994:RVE

[980] William E. Barlow. Robust variance estimation for the case-cohort design. Biometrics, 50(4):1064–1072, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533444.

Whittemore:1994:SAC

[981] Alice S. Whittemore and Gail Gong. Segregation analysis of case-control data using generalized estimating equations. Biometrics, 50(4):1073– 1087, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533445.

Begg:1994:OCR

[982] Colin B. Begg and Madhuchhanda Mazumdar. Operating characteristics of a rank correlation test for publication bias. Biometrics, 50(4):1088– 1101, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533446.

Conaway:1994:CNM

[983] Mark R. Conaway. Causal nonresponse models for repeated categorical measurements. Biometrics, 50(4):1102–1116, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533447.

Zhao:1994:IDG

[984] Yuejen Zhao, Andy H. Lee, and Yer Van Hui. Influence diagnostics for generalized linear measurement error models. Biometrics, 50(4):1117– 1128, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533448.

Hargrove:1994:PPP

[985] J. W. Hargrove and C. H. Borland. Pooled population parameter estimates from mark-recapture data. Biometrics, 50(4):1129–1141, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533449.

Burr:1994:IBE

[986] Deborah Burr. On inconsistency of Breslow’s estimator as an estimator of the hazard rate in the Cox model. Biometrics, 50(4):1142–1145, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533450.

Cook:1994:GME

[987] Richard J. Cook and Vern T. Farewell. Guidelines for monitoring efficacy and toxicity responses in clinical trials. Biometrics, 50(4):1146–1152, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533451.

Edelman:1994:TFR

[988] David Edelman. On a truncation-flexible repeated significance test. Biometrics, 50(4):1153–1157, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533452.

Feng:1994:LRT

[989] Z. D. Feng and C. E. McCulloch. On the likelihood ratio test statistic for the number of components in a normal mixture with unequal variances. Biometrics, 50(4):1158–1162, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533453.

Rotnitzky:1994:NBE

[990] Andrea Rotnitzky and David Wypij. A note on the bias of estimators with missing data. Biometrics, 50(4):1163–1170, December 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533454.

Stram:1994:VCT

[991] Daniel O. Stram and Jae Won Lee. Variance components testing in the longitudinal mixed effects model. Biometrics, 50(4):1171–1177, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533455. See correction

[1200].

Hougaard:1994:HMD

[992] Philip Hougaard, Per Myglegaard, and Knut Borch-Johnsen. Heterogeneity models of disease susceptibility, with application to diabetic nephropathy. Biometrics, 50(4):1178–1188, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533456.

Jin:1994:TSD

[993] Kun Jin, T. P. Speed, W. Klitz, and G. Thomson. Testing for segregation distortion in the HLA complex. Biometrics, 50(4):1189–1198, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533457.

Murtaugh:1994:SAF

[994] Paul A. Murtaugh. Statistical analysis of food webs. Biometrics, 50 (4):1199–1202, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533458.

Pagano:1994:RAC

[995] Marcello Pagano, Xin Ming Tu, Victor De Gruttola, and Samantha MaWhinney. Regression analysis of censored and truncated data: Estimating reporting-delay distributions and AIDS incidence from surveillance data. Biometrics, 50(4):1203–1214, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533459.

Podgor:1994:CNA

[996] Marvin J. Podgor and Joseph L. Gastwirth. A cautionary note on applying scores in stratified data. Biometrics, 50(4):1215–1218, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533460.

Mallet:1994:CRE

[997] M. Mallet, L. P. Rivest, and William R. Bell. Capture–recapture estimation with known sex ratio. Biometrics, 50(4):1219–1223, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533461.

Loynes:1994:BRV

[998] R. M. Loynes. Book review: Variance Components, by S. R. Searle, G. Casella, C. E. McCulloch. Biometrics, 50(4):1224–1225, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533462.

Smith:1994:BRSa

[999] J. T. Smith. Book review: Sense and Nonsense of Statistical Inference, by C. Wang. Biometrics, 50(4):1225–1226, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533463.

Martin:1994:BRR

[1000] M. A. Martin. Book review: Resampling-Based Multiple Testing: Examples and Methods for p-Value Adjustment, by P. H. Westfall, S. S. Young. Biometrics, 50(4):1226–1227, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533464.

Do:1994:BRE

[1001] K.-A. Do. Book review: Efficient and Adaptive Estimation for Semiparametric Models, by P. J. Bickel, C. A. J. Klaassen, Y. Ritov, J. A. Wellner. Biometrics, 50(4):1227, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533465.

Jones:1994:BRM

[1002] M. C. Jones. Book review: Model-Free Curve Estimation, by M. E. Tarter, M. D. Lock. Biometrics, 50(4):1227–1228, December 1994. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533466.

Best:1994:BRN

[1003] D. J. Best. Book review: Nonparametric Regression and Generalized Linear Models: a Roughness Penalty Approach, by P. J. Green, B. W. Silverman. Biometrics, 50(4):1228, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533467.

Jordan:1994:BRL

[1004] B. Jordan. Book review: L’Analyse Statistique Bayesienne, by C. Robert. Biometrics, 50(4):1228–1229, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533468.

Papageorgiou:1994:BRF

[1005] H. Papageorgiou. Book review: A First Course in Order Statistics, by B. C. Arnold, N. Balakrishnan, H. N. Nagaraja. Biometrics, 50(4):1229– 1230, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533469.

Mellor:1994:BRS

[1006] D. Mellor. Book review: Statistical Analysis of Circular Data, by N. I. Fisher. Biometrics, 50(4):1230, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533470.

McLachlan:1994:BRM

[1007] G. J. McLachlan. Book review: Modelling Survival Data in Medical Research, by D. Collett. Biometrics, 50(4):1230–1231, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533471.

Howel:1994:BRS

[1008] D. Howel. Book review: Statistical First Aid: Interpretation of Health Research Data, by R. P. Hirsch, R. K. Riegelman. Biometrics, 50(4):1231, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533472.

Day:1994:BRD

[1009] S. Day. Book review: Drug Safety Assessment in Clinical Trials, by G. S. Gilbert. Biometrics, 50(4):1231–1232, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533473.

Fearn:1994:BRDb

[1010] T. Fearn. Book review: Design and Analysis of Bioavailability and Bioequivalence Studies, by S.-C. Chow, J.-P. Liu. Biometrics, 50(4):1232, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533474.

Solomon:1994:BRA

[1011] P. J. Solomon. Book review: AIDS Epidemiology: a Quantitative Approach, by R. Brookmeyer, M. H. Gail. Biometrics, 50(4):1232–1233, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533475.

Edwards:1994:BRCb

[1012] A. W. F. Edwards. Book review: The Calculation of Genetic Risks: Worked Examples in DNA Diagnostics, by P. J. Bridge. Biometrics, 50(4):1233–1234, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533476.

Abakuks:1994:BRM

[1013] A. Abakuks. Book review: Matrices and Graphs: Stability Problems in Mathematical Ecology, by D. O. Logofet. Biometrics, 50(4):1234–1235, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533477.

Smith:1994:BRSb

[1014] R. I. Smith. Book review: Sampling Methods for Multiresource Forest Inventory, by H. T. Schreuder, T. G. Gregoire, G. B. Wood. Biometrics, 50(4):1235, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533478.

Anonymous:1994:BRAd

[1015] Anonymous. Book review: Analysis of Quantal Response Data, by B. J. T. Morgan. Biometrics, 50(4):1235–1236, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533479.

Anonymous:1994:BRDc

[1016] Anonymous. Book review: Developments in Time Series Analysis: In Honour of Maurice B. Priestley, by T. Subba Rao. Biometrics, 50(4): 1236, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533483.

Anonymous:1994:BRIc

[1017] Anonymous. Book review: Introduction to Probability and Statistics, by N. Giri. Biometrics, 50(4):1236, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533482.

Anonymous:1994:BRId

[1018] Anonymous. Book review: Introductory Statistics for Biology Students, by T. A. Watt. Biometrics, 50(4):1236, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533481.

Anonymous:1994:BRMb

[1019] Anonymous. Book review: Mathematik für Biologen, by A. Riede. Biometrics, 50(4):1236, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533480.

Anonymous:1994:BRW

[1020] Anonymous. Book review: World Health Statistics Quarterly 1993, by World Health Organization. Biometrics, 50(4):1236–1237, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533484.

Anonymous:1994:PPBd

[1021] Anonymous. Papers to be published in biometrics. Biometrics, 50(4): 1238, December 1994. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533485.

Anonymous:1995:FMa

[1022] Anonymous. Front matter. Biometrics, 51(1):i–iv, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533308.

Billard:1995:RT

[1023] Lynne Billard. The roads travelled. Biometrics, 51(1):1–11, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533309.

Song:1995:EDP

[1024] K. S. Song, H. G. Müller, A. J. Clifford, H. C. Furr, and J. A. Olson. Estimating derivatives of pharmacokinetic response curves with varying bandwidths. Biometrics, 51(1):12–20, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533310.

Lee:1995:GSC

[1025] Jae Won Lee and David L. DeMets. Group sequential comparison of changes: Ad-hoc versus more exact method. Biometrics, 51(1):21–30, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533311.

Etzioni:1995:CTM

[1026] Ruth Etzioni and Steven G. Self. On the catch-up time method for analyzing cancer screening trials. Biometrics, 51(1):31–43, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533312. See corrections [1255].

Self:1995:LRT

[1027] Steven G. Self and Ruth Etzioni. A likelihood ratio test for cancer screening trials. Biometrics, 51(1):44–50, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533313.

Heisey:1995:MAS

[1028] Dennis M. Heisey and Erik V. Nordheim. Modelling age-specific survival in nesting studies, using a general approach for doubly-censored and truncated data. Biometrics, 51(1):51–60, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533314.

Li:1995:FAS

[1029] Jufang Li, Mei-Ching Chen, Hans T. Schreuder, and Tim G. Gregoire. Forestry applications of saddle-point approximations to construct confidence intervals for population means. Biometrics, 51(1):61–72, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533315.

Lu:1995:AAC

[1030] Ying Lu and Hina Mehta Malani. Analysis of animal carcinogenicity experiments with multiple tumor types. Biometrics, 51(1):73–86, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533316.

Rebai:1995:CPD

[1031] Ahmed Rebai, Bruno Goffinet, and Brigitte Mangin. Comparing power of different methods for QTL detection. Biometrics, 51(1):87–99, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533317.

Thorne:1995:ERM

[1032] Jeffrey L. Thorne and Gary A. Churchill. Estimation and reliability of molecular sequence alignments. Biometrics, 51(1):100–113, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533318.

Royston:1995:CNN

[1033] Patrick Royston and Simon G. Thompson. Comparing non-nested regression models. Biometrics, 51(1):114–127, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533319.

Hwang:1995:QEU

[1034] W.-D. Hwang and Anne Chao. Quantifying the effects of unequal catchabilities on Jolly–Seber estimators via sample coverage. Biometrics, 51(1): 128–141, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533320.

Cox:1995:ACB

[1035] Christopher Cox and Guangqin Ma. Asymptotic confidence bands for generalized nonlinear regression models. Biometrics, 51(1):142–150, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533321.

Follmann:1995:AGL

[1036] Dean Follmann and Margaret Wu. An approximate generalized linear model with random effects for informative missing data. Biometrics, 51(1):151–168, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533322.

Landin:1995:ATT

[1037] Richard Landin, Laurence S. Freedman, and Raymond J. Carroll. Adjusting for time trends when estimating the relationship between dietary intake obtained from a food frequency questionnaire and true average intake. Biometrics, 51(1):169–181, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2533323. See corrections [1199].

Benichou:1995:MIE

[1038] Jacques Benichou and Mitchell H. Gail. Methods of inference for estimates of absolute risk derived from population-based case-control studies. Biometrics, 51(1):182–194, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533324.

Barnhart:1995:MPM

[1039] Huiman X. Barnhart and Allan R. Sampson. Multiple population models for multivariate random length data — with applications in clinical trials. Biometrics, 51(1):195–204, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533325.

Chinchilli:1995:PLA

[1040] Vernon M. Chinchilli, James D. Esinhart, and W. Greg Miller. Partial likelihood analysis of within-unit variances in repeated measurement experiments. Biometrics, 51(1):205–216, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533326.

Dunnett:1995:SMT

[1041] Charles W. Dunnett and Ajit C. Tamhane. Step-up multiple testing of parameters with unequally correlated estimates. Biometrics, 51(1):217– 227, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533327.

Farrington:1995:RIE

[1042] C. P. Farrington. Relative incidence estimation from case series for vaccine safety evaluation. Biometrics, 51(1):228–235, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533328.

Kiuchi:1995:CPS

[1043] Amy S. Kiuchi, J. A. Hartigan, Theodore R. Holford, Pablo Rubinstein, and Cladd E. Stevens. Change points in the series of T4 counts prior to AIDS. Biometrics, 51(1):236–248, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533329.

Piantadosi:1995:IDP

[1044] Steven Piantadosi and John Crowley. An implicitly defined parametric model for censored survival data and covariates. Biometrics, 51(1):249– 258, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533330.

Preisler:1995:ATM

[1045] Haiganoush K. Preisler and R. Patrick Akers. Autoregressive-type models for the analysis of bark beetle tracks. Biometrics, 51(1):259–267, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533331.

Qu:1995:LVM

[1046] Yinsheng Qu, Marion R. Piedmonte, and Sharon V. Medendorp. Latent variable models for clustered ordinal data. Biometrics, 51(1):268–275, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533332.

Vogler:1995:PAE

[1047] George P. Vogler, Reimut Wette, Matthew K. McGue, and D. C. Rao. Properties of alternative estimators of familial correlations under variable sibship size. Biometrics, 51(1):276–283, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533333.

Dayananda:1995:ERP

[1048] P. W. A. Dayananda, L. Billard, and S. Chakraborty. Estimation of rate parameter and its relationship with latent and infectious periods in plant disease epidemics. Biometrics, 51(1):284–292, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533334.

Kendall:1995:LBA

[1049] William L. Kendall, Kenneth H. Pollock, and Cavell Brownie. A likelihood-based approach to capture–recapture estimation of demographic parameters under the robust design. Biometrics, 51(1):293–308, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533335.

Fitzmaurice:1995:CCI

[1050] Garrett M. Fitzmaurice. A caveat concerning independence estimating equations with multivariate binary data. Biometrics, 51(1):309–317, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533336.

Lin:1995:SCI

[1051] Shili Lin. A scheme for constructing an irreducible Markov chain for pedigree data. Biometrics, 51(1):318–322, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533337.

Andersen:1995:NTC

[1052] Per Kragh Andersen and Birgitte B. Rønn. A nonparametric test for comparing two samples where all observations are either left- or rightcensored. Biometrics, 51(1):323–329, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533338.

Baker:1995:EMD

[1053] Stuart G. Baker. Evaluating multiple diagnostic tests with partial verification. Biometrics, 51(1):330–337, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533339.

Wallach:1995:ROF

[1054] D. Wallach. Regional optimization of fertilization using a hierarchical linear model. Biometrics, 51(1):338–346, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533340.

Nagelkerke:1995:SIS

[1055] Nico J. D. Nagelkerke. Statistical inference in stationary populations. Biometrics, 51(1):347–350, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533341.

Tritchler:1995:ISD

[1056] David Tritchler. Interpreting the standardized difference. Biometrics, 51(1):351–353, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533342.

Nam:1995:STH

[1057] Jun mo Nam. Simple test for the Hardy–Weinberg law for HLA data with no observed double blanks. Biometrics, 51(1):354–357, March 1995.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533343.

James:1995:NAC

[1058] Ian R. James and Martin A. Tanner. A note on the analysis of censored regression data by multiple imputation. Biometrics, 51(1):358– 362, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533344.

Tyurin:1995:TMA

[1059] Yu. N. Tyurin, Andrej Yu. Yakovlev, J. Shi, and L. Bass. Testing a model of aging in animal experiments. Biometrics, 51(1):363–372, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533345.

Ripley:1995:SMP

[1060] B. D. Ripley, P. J. Solomon, M.-C. Wang, R. Brookmeyer, and N. P. Jewell. Statistical models for prevalent cohort data. Biometrics, 51(1): 373–375, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533346.

Glasbey:1995:USL

[1061] Chris Glasbey, Vicente Núñez-Antón, and George G. Woodworth. Unequally spaced longitudinal data. Biometrics, 51(1):375–377, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533347.

Rasch:1995:BRD

[1062] D. Rasch. Book review: Design and Analysis of Experiments. Design, by K. Hinkelmann, O. Kempthorne. Biometrics, 51(1):378–379, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533348.

Laycock:1995:BRO

[1063] P. J. Laycock. Book review: Optimum Experimental Designs, by A. C. Atkinson, A. N. Donev. Biometrics, 51(1):379, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533349.

Read:1995:BRM

[1064] C. B. Read. Book review: Multiple Comparisons, Selection, and Applications in Biometry: a Festschrift in Honor of Charles W. Dunnett, by F.

M. Hoppe. Biometrics, 51(1):380, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533350.

Martin:1995:BRC

[1065] M. A. Martin. Book review: The Collected Works of John W. Tukey: Volume VIII, Multiple Comparisons 1948–1983, by John W. Tukey, H. I. Braun. Biometrics, 51(1):380–381, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533351.

Balakrishnan:1995:BRO

[1066] N. Balakrishnan. Book review: Outliers in Statistical Data, by V. Barnett, T. Lewis. Biometrics, 51(1):381, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533352.

MacGillivray:1995:BRL

[1067] H. MacGillivray. Book review: Large Sample Methods in Statistics, by P. K. Sen, J. M. Singer. Biometrics, 51(1):381–382, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533353.

Papaioannou:1995:BRP

[1068] T. Papaioannou. Book review: Poisson Approximation, by A. D. Barbour, L. Holst, S. Janson. Biometrics, 51(1):382–383, March 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533354.

Elton:1995:BRS

[1069] R. A. Elton. Book review: Statistical Methods for Survival Data Analysis, by E. T. Lee. Biometrics, 51(1):383, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533355.

Panaretos:1995:BRS

[1070] J. Panaretos. Book review: Statistical Inference for Branching Processes, by P. Guttorp. Biometrics, 51(1):383–384, March 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533356.

Elston:1995:BRW

[1071] D. A. Elston. Book review: Wildlife Habitat Relationships in Forested Ecosystems, by D. R. Patton. Biometrics, 51(1):384–385, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533357.

Bjornsson:1995:BRA

[1072] H. Bjørnsson. Book review: Angler Survey Methods and Their Applications in Fisheries Management, by K. H. Pollock, C. M. Jones, T. L. Brown. Biometrics, 51(1):385, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533358.

Anderson:1995:BRA

[1073] A. J. B. Anderson. Book review: Algorithms, Routines, and S Functions for Robust Statistics, by A. Marazzi. Biometrics, 51(1):385–386, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533359.

Anonymous:1995:BRSa

[1074] Anonymous. Book review: Statistical Methods in Biology, by N. T. J. Bailey. Biometrics, 51(1):386, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533361.

Shotter:1995:BRU

[1075] M. V. Shotter. Book review: Understanding Probability and Statistics: a Book of Problems, by R. Falk. Biometrics, 51(1):386, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533360.

Anonymous:1995:BRBa

[1076] Anonymous. Book review: Biostatistics: a Foundation for Analysis in the Health Sciences, by W. W. Daniel. Biometrics, 51(1):386–387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533362.

Anonymous:1995:BRBb

[1077] Anonymous. Book review: Biostatistics: Perspectives in Health Care Research and Practice, by B. L. Verma, G. D. Shukla, R. N. Srivastava. Biometrics, 51(1):387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533363.

Anonymous:1995:BRBc

[1078] Anonymous. Book review: Biopharmaceutical Sequential Statistical Applications, by K. E. Peace. Biometrics, 51(1):387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533365.

Anonymous:1995:BREa

[1079] Anonymous. Book review: Essential Statistics, by D. G. Rees. Biometrics, 51(1):387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533366.

Anonymous:1995:BRFa

[1080] Anonymous. Book review: Foundations of Epidemiology, by D. E. Lilienfeld, P. D. Stolley. Biometrics, 51(1):387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533364.

Anonymous:1995:BRSb

[1081] Anonymous. Book review: Stochastic Processes, by J. Medhi. Biometrics, 51(1):387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533368.

Anonymous:1995:BRSc

[1082] Anonymous. Book review: Statistical Theory, by B. W. Lindgren. Biometrics, 51(1):387, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533367.

Anonymous:1995:BRG

[1083] Anonymous. Book review: The GLIM System: Release 4 Manual, by B. Francis, M. Green, C. Payne. Biometrics, 51(1):388, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533370.

Anonymous:1995:BRSd

[1084] Anonymous. Book review: SoftStat ’93: Advances in Statistical Software 4, by F. Faulbaum. Biometrics, 51(1):388, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533369.

Anonymous:1995:BRWa

[1085] Anonymous. Book review: World Health Statistics Annual 1993, by World Health Organization. Biometrics, 51(1):388, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533371.

Healy:1995:OFY

[1086] M. J. R. Healy. Obituary: Frank Yates 1902–1994. Biometrics, 51(1): 389–391, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533372.

Anonymous:1995:CMH

[1087] Anonymous. Correction: Mantel–Haenszel-type tests for testing equivalence or more than equivalence in comparative clinical trials. Biometrics, 51(1):392, March 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533373. See [935].

Anonymous:1995:FMb

[1088] Anonymous. Front matter. Biometrics, 51(2):i–iv, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532927.

Hogmander:1995:EDM

[1089] Harri Högmander and Jesper Møller. Estimating distribution maps from atlas data using methods of statistical image analysis. Biometrics, 51(2): 393–404, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532928.

Weinberg:1995:MEP

[1090] Clarice R. Weinberg and Allen J. Wilcox. A model for estimating the potency and survival of human gametes in vivo. Biometrics, 51(2):405– 412, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532929.

Zucker:1995:IAB

[1091] David M. Zucker, Gary O. Zerbe, and Margaret C. Wu. Inference for the association between coefficients in a multivariate growth curve model. Biometrics, 51(2):413–424, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532930.

Smith:1995:EMM

[1092] Fraser B. Smith and Ronald W. Helms. EM mixed model analysis of data from informatively censored normal distributions. Biometrics, 51(2):425– 436, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532931.

Tsodikov:1995:DSC

[1093] Alexander D. Tsodikov, Bernard Asselain, Alain Fourque, Thu Hoang, and Andrej Yu. Yakovlev. Discrete strategies of cancer post-treatment surveillance. estimation and optimization problems. Biometrics, 51(2): 437–447, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532932.

Raz:1995:FDE

[1094] Jonathan Raz, Valerie Cardenas, and Daniel Fletcher. Frequency domain estimation of covariate effects in multichannel brain evoked potential data. Biometrics, 51(2):448–460, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532933.

Rivest:1995:SMA

[1095] Louis-Paul Rivest, François Potvin, Hélène Crépeau, and Gaétan Daigle. Statistical methods for aerial surveys using the double-count technique to correct visibility bias. Biometrics, 51(2):461–470, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532934.

Udevitz:1995:UEI

[1096] Mark S. Udevitz and Kenneth H. Pollock. Using effort information with change-in-ratio data for population estimation. Biometrics, 51(2):471– 481, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532935.

Yamamura:1995:EPP

[1097] Kohji Yamamura and Tamio Sugimoto. Estimation of the pest prevention ability of the import plant quarantine in Japan. Biometrics, 51(2):482– 490, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532936.

Alho:1995:IEI

[1098] Juha M. Alho and Esko Valtonen. Interval estimation of inverse doseresponse. Biometrics, 51(2):491–501, June 1995. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532937. See corrections [1307].

Frydman:1995:SET

[1099] Halina Frydman. Semiparametric estimation in a three-state durationdependent Markov model from interval-censored observations with application to AIDS data. Biometrics, 51(2):502–511, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532938.

George:1995:FLP

[1100] E. Olusegun George and Dale Bowman. A full likelihood procedure for analysing exchangeable binary data. Biometrics, 51(2):512–523, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532939.

Lunn:1995:ACR

[1101] Mary Lunn and Don McNeil. Applying Cox regression to competing risks. Biometrics, 51(2):524–532, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532940.

ONeill:1995:TLR

[1102] Terence J. O’Neill and Simon C. Barry. Truncated logistic regression. Biometrics, 51(2):533–541, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532941.

Commenges:1995:STF

[1103] Daniel Commenges, Hélène Jacqmin, Luc Letenneur, and Cornelia M. Van Duijn. Score test for familial aggregation in probands studies: Application to Alzheimer’s disease. Biometrics, 51(2):542–551, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532942.

Yang:1995:MMA

[1104] Ziheng Yang and Tianlin Wang. Mixed model analysis of DNA sequence evolution. Biometrics, 51(2):552–561, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532943.

Lipsitz:1995:EMJ

[1105] Stuart R. Lipsitz, Garret M. Fitzmaurice, Lynn Sleeper, and L. P. Zhao. Estimation methods for the joint distribution of repeated binary observations. Biometrics, 51(2):562–570, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532944.

deBEdwardes:1995:CIEa

[1106] Michael D. deB. Edwardes. A confidence interval for Pr(X < Y) − Pr(X > Y) estimated from simple cluster samples. Biometrics, 51(2): 571–578, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deBEdwardes:1995:CIEb

[1107] Michael D. deB. Edwardes. A confidence interval for Pr(X < Y) − Pr(X > Y) estimated from simple cluster samples. Biometrics, 51(2): 571–578, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532945.

Liao:1995:EBA

[1108] Jiangang Liao and Ron Brookmeyer. An empirical Bayes approach to smoothing in backcalculation of HIV infection rates. Biometrics, 51(2): 579–588, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532946.

Weerahandi:1995:AUU

[1109] Samaradasa Weerahandi. ANOVA under unequal error variances. Biometrics, 51(2):589–599, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532947.

leCessie:1995:TFR

[1110] Saskia le Cessie and Hans C. van Houwelingen. Testing the fit of a regression model via score tests in random effects models. Biometrics, 51(2): 600–614, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532948.

Schall:1995:AIP

[1111] Robert Schall. Assessment of individual and population bioequivalence using the probability that bioavailabilities are similar. Biometrics, 51(2): 615–626, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532949.

Albert:1995:GEE

[1112] Paul S. Albert and Lisa M. McShane. A generalized estimating equations approach for spatially correlated binary data: Applications to the analysis of neuroimaging data. Biometrics, 51(2):627–638, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532950.

Barry:1995:BMG

[1113] Daniel Barry. A Bayesian model for growth curve analysis. Biometrics, 51(2):639–655, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532951.

Conaway:1995:BSD

[1114] Mark R. Conaway and Gina R. Petroni. Bivariate sequential designs for Phase II trials. Biometrics, 51(2):656–664, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532952.

Ferrandiz:1995:SIB

[1115] Juan Ferrándiz, Antonio López, Agustín Llopis, María Morales, and María Luisa Tejerizo. Spatial interaction between neighbouring counties: Cancer mortality data in Valencia (Spain). Biometrics, 51(2):665– 678, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532953.

Pierrat:1995:EMV

[1116] Jean Claude Pierrat, François Houllier, Jean Christophe Hervé, and Raúl Salas González. Estimation de la moyenne des valeurs les plus Élevées d’une population finie: Application aux inventaires forestiers. (French) [Estimation of the average of the highest values of a finite population: Application to forest inventories]. Biometrics, 51(2):679–686, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532954.

Vounatsou:1995:BAR

[1117] P. Vounatsou and A. F. M. Smith. Bayesian analysis of ring-recovery data via Markov chain Monte Carlo simulation. Biometrics, 51(2):687– 708, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532955.

Manly:1995:MSM

[1118] Bryan F. J. Manly. Measuring selectivity from multiple choice feedingpreference experiments. Biometrics, 51(2):709–715, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532956.

Drescher:1995:ACE

[1119] Karsten Drescher and Wolfgang Boedeker. Assessment of the combined effects of substances: The relationship between concentration addition and independent action. Biometrics, 51(2):716–730, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532957.

Larholt:1995:EHU

[1120] Kay M. Larholt and Allan R. Sampson. Effects of heteroscedasticity upon certain analyses when regression lines are not parallel. Biometrics, 51(2): 731–737, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532958.

vandenBroek:1995:STZ

[1121] Jan van den Broek. A score test for zero inflation in a Poisson distribution. Biometrics, 51(2):738–743, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532959.

Gogel:1995:REM

[1122] Beverley J. Gogel, Brian R. Cullis, and Arūnus P. Verbyla. REML estimation of multiplicative effects in multienvironment variety trails. Biometrics, 51(2):744–749, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532960.

Carothers:1995:MDR

[1123] Andrew D. Carothers and Wendy A. Bickmore. Models of DNA replication timing in interphase nuclei: an exercise in inferring process from state. Biometrics, 51(2):750–755, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532961.

Lee:1995:GSM

[1124] Jae Won Lee and Harland N. Sather. Group sequential methods for comparison of cure rates in clinical trials. Biometrics, 51(2):756–763, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532962.

Bieler:1995:CST

[1125] Gayle S. Bieler and Rick L. Williams. Cluster sampling techniques in quantal response teratology and developmental toxicity studies. Biometrics, 51(2):764–776, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532963.

Smith:1995:EAC

[1126] David R. Smith, Michael J. Conroy, and David H. Brakhage. Efficiency of adaptive cluster sampling for estimating density of wintering waterfowl. Biometrics, 51(2):777–788, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532964.

Finney:1995:E

[1127] David J. Finney, Ross H. Taplin, and Adrian E. Raftery. [Editorial]. Biometrics, 51(2):789–790, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532965.

Bohning:1995:BRT

[1128] D. Böhning. Book review: The Theory of Linear Models, by B. Jørgensen. Biometrics, 51(2):791–792, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532966.

Kemp:1995:BRM

[1129] A. W. Kemp. Book reviews: Models in Biology: Mathematics, Statistics and Computing, by D. Brown, P. Rothery; Computing Examples Supplement to Models in Biology: Mathematics, Statistics and Computing, by D. Brown, P. Rothery. Biometrics, 51(2):792–793, June 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532967.

Gordon:1995:BRS

[1130] A. D. Gordon. Book review: Statistical Applications Using Fuzzy Sets, by K. G. Manton, M. A. Woodbury, H. D. Tolley. Biometrics, 51(2):793– 794, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532968.

Lesaffre:1995:BRA

[1131] E. Lesaffre. Book review: Analysis of Systematic and Random Differences between Paired Ordinal Categorical Data, by E. Svensson. Biometrics, 51(2):794, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532969.

Howroyd:1995:BRF

[1132] J. D. Howroyd and M. M. Bayer. Book review: Fractal Modelling: Growth and Form in Biology, by J. A. Kaandorp. Biometrics, 51(2): 794–795, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532970.

Kemp:1995:BRS

[1133] A. W. Kemp. Book review: Statistics for the Environment 2: Water Related Issues, by V. Barnett, K. F. Turkman. Biometrics, 51(2):795– 796, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532971.

Anonymous:1995:BRAa

[1134] Anonymous. Book review: Aspects of Uncertainty: a Tribute to D. V. Lindley, by P. R. Freeman, A. F. M. Smith. Biometrics, 51(2):796, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532972.

Anonymous:1995:BRBd

[1135] Anonymous. Book review: Biostatistics: Experimental Design and Statistical Inference, by J. F. Zolman. Biometrics, 51(2):796–797, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532973.

Anonymous:1995:BREb

[1136] Anonymous. Book review: Elementary Applications of Probability Theory, by H. C. Tuckwell. Biometrics, 51(2):797, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532975.

Anonymous:1995:BRI

[1137] Anonymous. Book review: Introductory Statistics, by P. S. Mann. Biometrics, 51(2):797, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532974.

Anonymous:1995:BRSe

[1138] Anonymous. Book review: The Statistical Analysis of Time Series, by T. W. Anderson. Biometrics, 51(2):797, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532976.

Anonymous:1995:BRSf

[1139] Anonymous. Book review: Statistics for Long-Term Processes, by J. Beran. Biometrics, 51(2):797, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532977.

Anonymous:1995:BRCa

[1140] Anonymous. Book review: Counting the Dead: The Epidemiology of Skeletal Populations, by T. Waldron. Biometrics, 51(2):797–798, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532978.

Anonymous:1995:BRWb

[1141] Anonymous. Book review: World Health Statistics Quarterly 1994, Vol. 47 (1), (2), (3/4), by World Health Organization. Biometrics, 51(2): 798, June 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532979.

Anonymous:1995:FMc

[1142] Anonymous. Front matter. Biometrics, 51(3):i–iv, September 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532981.

Edwards:1995:XFM

[1143] A. W. F. Edwards. XVIIIth Fisher Memorial Lecture delivered at the Natural History Museum, London, on Thursday, 20th October, 1994: Fiducial inference and the fundamental theorem of natural selection. Biometrics, 51(3):799–809, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532982.

Link:1995:ECI

[1144] William A. Link and John R. Sauer. Estimation and confidence intervals for empirical mixing distributions. Biometrics, 51(3):810–821, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532983.

Orzack:1995:SSD

[1145] Steven Hecht Orzack and Rick Chappell. Sample-size determination for parent-offspring regression of a composite trait. Biometrics, 51(3):822– 830, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532984.

Brookmeyer:1995:SAP

[1146] Ron Brookmeyer and Yutaka Yasui. Statistical analysis of passive surveillance disease registry data. Biometrics, 51(3):831–842, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532985.

Gelfand:1995:BAP

[1147] Alan E. Gelfand and Bani K. Mallick. Bayesian analysis of proportional hazards models built from monotone functions. Biometrics, 51(3):843– 852, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532986.

Sheppard:1995:RPW

[1148] Lianne Sheppard and Ross L. Prentice. On the reliability and precision of within- and between-population estimates of relative rate parameters. Biometrics, 51(3):853–863, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532987.

Li:1995:MRE

[1149] Zhaohai Li. A multiplicative random effects model for meta-analysis with application to estimation of admixture component. Biometrics, 51(3):864–873, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532988.

Rosenberg:1995:HFE

[1150] Philip S. Rosenberg. Hazard function estimation using B-splines. Biometrics, 51(3):874–887, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532989.

Burnham:1995:MSS

[1151] Kenneth P. Burnham, Gary C. White, and David R. Anderson. Model selection strategy in the analysis of capture–recapture data. Biometrics, 51(3):888–898, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532990.

Taylor:1995:SPE

[1152] Jeremy M. G. Taylor. Semi-parametric estimation in failure time mixture models. Biometrics, 51(3):899–907, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532991.

Villarroya:1995:DAA

[1153] Angel Villarroya, Martín Ríos, and Josep M. Oller. Discriminant analysis algorithm based on a distance function and on a Bayesian decision. Biometrics, 51(3):908–919, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532992.

Young:1995:NPA

[1154] Stuart G. Young and Adrian W. Bowman. Non-parametric analysis of covariance. Biometrics, 51(3):920–931, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532993.

Cook:1995:IAC

[1155] Richard J. Cook. Interim analyses in 2 × 2 crossover trials. Biometrics, 51(3):932–945, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532994.

Liu:1995:GSP

[1156] Wei Liu. A group sequential procedure for all-pairwise comparisons of k treatments based on the range statistic. Biometrics, 51(3):946–955, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532995.

Baker:1995:REM

[1157] Rose D. Baker and Robert H. Stevens. A random-effects model for analysis of infectious disease final-state data. Biometrics, 51(3):956–968, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532996.

Flandre:1995:TSP

[1158] Philippe Flandre and John O’Quigley. A two-stage procedure for survival studies with surrogate endpoints. Biometrics, 51(3):969–976, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532997.

Wijesinha:1995:DRM

[1159] Manel Cooray Wijesinha and Steven Piantadosi. Dose-response models with covariates. Biometrics, 51(3):977–987, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532998.

Kim:1995:DAG

[1160] KyungMann Kim, Hélène Boucher, and Anastasios A. Tsiatis. Design and analysis of group sequential logrank tests in maximum duration versus information trials. Biometrics, 51(3):988–1000, September 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532999.

TenHave:1995:MCT

[1161] Thomas R. Ten Have and Mark P. Becker. Multivariate contingency tables and the analysis of exchangeability. Biometrics, 51(3):1001–1016, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533000.

vanEeuwijk:1995:MIG

[1162] Fred A. van Eeuwijk. Multiplicative interaction in generalized linear models. Biometrics, 51(3):1017–1032, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533001.

Lemdani:1995:TGL

[1163] Mohamed Lemdani and Odile Pons. Tests for genetic linkage and homogeneity. Biometrics, 51(3):1033–1041, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533002.

Baker:1995:MRR

[1164] Stuart G. Baker. Marginal regression for repeated binary data with outcome subject to non-ignorable non-response. Biometrics, 51(3):1042– 1052, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533003.

Lin:1995:UHM

[1165] Hung-Mo Lin and Michael D. Hughes. Use of historical marker data for assessing treatment effects in Phase I/II trials when subject selection is determined by baseline marker level. Biometrics, 51(3):1053–1063, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533004.

Jin:1995:TRM

[1166] Kun Jin, T. P. Speed, and Glenys Thomson. Tests of random mating for a highly polymorphic locus: Application to HLA data. Biometrics, 51(3):1064–1076, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533005.

Hurvich:1995:MSE

[1167] Clifford M. Hurvich and Chih-Ling Tsai. Model selection for extended quasi-likelihood models in small samples. Biometrics, 51(3):1077–1084, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533006.

Dellaportas:1995:BAE

[1168] Petros Dellaportas and David A. Stephens. Bayesian analysis of errorsin-variables regression models. Biometrics, 51(3):1085–1095, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533007.

Sun:1995:EED

[1169] Jianguo Sun. Empirical estimation of a distribution function with truncated and doubly interval-censored data and its application to AIDS studies. Biometrics, 51(3):1096–1104, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533008.

Ducharme:1995:RVO

[1170] Gilles R. Ducharme, Ali Gannoun, Marie-Claude Guertin, and Jean-Claude Jéquier. Reference values obtained by kernel-based estimation of quantile regressions. Biometrics, 51(3):1105–1116, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533009.

Scollnik:1995:BAT

[1171] David P. M. Scollnik. Bayesian analysis of two overdispersed Poisson models. Biometrics, 51(3):1117–1126, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533010. See corrections [1451].

Haukka:1995:CCM

[1172] Jari K. Haukka. Correction for covariate measurement error in generalized linear models — a bootstrap approach. Biometrics, 51(3):1127–1132, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533011.

Hanusz:1995:RPT

[1173] Zofia Hanusz. Relative potency of two preparations in two-way elimination of heterogeneity designs with multivariate responses. Biometrics, 51(3):1133–1139, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533012.

Zelterman:1995:SCS

[1174] Daniel Zelterman and James W. Curtsinger. Survival curves subjected to occasional insults. Biometrics, 51(3):1140–1146, September 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533013.

Smith:1995:PRS

[1175] Peter J. Smith. On plotting renovated samples. Biometrics, 51(3):1147– 1151, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533014.

Emerson:1995:SCT

[1176] Scott S. Emerson. Stopping a clinical trial very early based on unplanned interim analyses: a group sequential approach. Biometrics, 51(3):1152– 1162, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533015.

Davis:1995:RMM

[1177] Charles S. Davis and Younghae Chung. Randomization model methods for evaluating treatment efficacy in multicenter clinical trials. Biometrics, 51(3):1163–1174, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533016.

Lan:1995:SMS

[1178] K. K. Gordon Lan, William F. Rosenberger, and John M. Lachin. Sequential monitoring of survival data with the Wilcoxon statistic. Biometrics, 51(3):1175–1183, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533017.

deFalguerolles:1995:E

[1179] Antoine de Falguerolles, Stuart R. Lipsitz, and Garrett Fitzmaurice. [Editorial]. Biometrics, 51(3):1184–1185, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533018. See corrections [1256].

Kemp:1995:BRE

[1180] A. W. Kemp. Book review: Elsevier’s Dictionary of Biometry, by D. Rasch, M. L. Tiku, D. Sumpf. Biometrics, 51(3):1186–1187, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533019. See correction [1254].

Haigh:1995:BRS

[1181] J. Haigh. Book review: Studies in Applied Probability: Papers in Honour of Lajos Takács, by J. Galambos, J. Gani. Biometrics, 51(3):1187, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533020.

Hand:1995:BRA

[1182] D. J. Hand. Book review: Applied Discriminant Analysis, by C. J. Huberty. Biometrics, 51(3):1187–1188, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533021.

Edmonson:1995:BRA

[1183] R. N. Edmonson. Book review: Analysis of Two-Way Layouts, by J. Mandel. Biometrics, 51(3):1188–1189, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533022.

Seber:1995:BRM

[1184] G. A. F. Seber. Book review: Multivariate Environmental Statistics, by G. P. Patil, C. R. Rao. Biometrics, 51(3):1189–1190, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533023.

Smith:1995:BRE

[1185] R. I. Smith. Book review: Experimental Design and Analysis for Use in Tree Improvement, by E. R. Williams, A. C. Matheson. Biometrics, 51(3):1190, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533024.

Day:1995:BRS

[1186] S. Day. Book review: Statistical Design and Analysis in Pharmaceutical Science: Validation, Process Controls, and Stability, by S.-C. Chow, J.-P. Liu. Biometrics, 51(3):1190–1191, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533025.

McLachlan:1995:BRA

[1187] G. J. McLachlan. Book review: Analysing Survival Data from Clinical Trials and Observational Studies, by E. Marubini, M. G. Valsecchi. Biometrics, 51(3):1191, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533026.

Gani:1995:BRC

[1188] J. Gani. Book review: Cooperation and Conflict in General Evolutionary Processes, by J. L. Casti, A. Karlqvist. Biometrics, 51(3):1191–1193, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533027.

Lance:1995:BRL

[1189] N. Lance. Book review: Longitudinal Data with Serial Correlation: a State-Space Approach, by R. H. Jones. Biometrics, 51(3):1193–1194, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533028.

Anonymous:1995:BRAb

[1190] Anonymous. Book review: Aquatic Ecology: Scale, Pattern and Process, by P. S. Giller, A. G. Hildrew, D. G. Raffaelli. Biometrics, 51(3):1194, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533032.

Anonymous:1995:BRAc

[1191] Anonymous. Book review: Aquatic Microbiology: an Ecological Approach, by T. E. Ford. Biometrics, 51(3):1194, September 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533031.

Anonymous:1995:BRFb

[1192] Anonymous. Book review: A First Course in Probability Models and Statistical Inference, by J. H. C. Creighton. Biometrics, 51(3):1194, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533030.

Anonymous:1995:BRJ

[1193] Anonymous. Book review: Le Jeu de la Science et du Hasard, by D. Schwartz. Biometrics, 51(3):1194, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533029.

Anonymous:1995:BRCb

[1194] Anonymous. Book review: Choquet-Deny Type Functional Equations with Applications to Stochastic Models, by C. R. Rao, D. N. Shanbhag. Biometrics, 51(3):1195, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533034.

Anonymous:1995:BRCc

[1195] Anonymous. Book review: Conditional Measures and Applications, by M. M. Rao. Biometrics, 51(3):1195, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533033.

Anonymous:1995:BRMa

[1196] Anonymous. Book review: Monotone Structure in Discrete Event Systems, by P. Glasserman, D. D. Yao. Biometrics, 51(3):1195, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533035.

Anonymous:1995:BRO

[1197] Anonymous. Book review: Optimal Sequentially Planned Decision Procedures, by N. Schmitz, G. Duscha, J. Lübbert, T. Meyerthole. Biometrics, 51(3):1195, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533036.

Anonymous:1995:BRSg

[1198] Anonymous. Book review: Statistical Methods of Quality Assurance, by H.-J. Mittag, H. Rinne. Biometrics, 51(3):1195, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533037.

Anonymous:1995:CAT

[1199] Anonymous. Corrections: Adjusting for time trends when estimating the relationship between dietary intake obtained from a food frequency questionnaire and true average intake. Biometrics, 51(3):1196, September 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533039. See [1037].

Anonymous:1995:CVC

[1200] Anonymous. Corrections: Variance components testing in the longitudinal mixed effects model. Biometrics, 51(3):1196, September 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533038. See [991].

Anonymous:1995:FMd

[1201] Anonymous. Front matter. Biometrics, 51(4):i–iv, December 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533252.

Anonymous:1995:VI

[1202] Anonymous. Volume information. Biometrics, 51(4):v–xxvi, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533251.

Maller:1995:TPI

[1203] R. A. Maller and Sean Zhou. Testing for the presence of immune or cured individuals in censored survival data. Biometrics, 51(4):1197–1205, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533253.

Worton:1995:CHB

[1204] Bruce J. Worton. A convex hull-based estimator of home-range size. Biometrics, 51(4):1206–1215, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533254.

Sanchez:1995:STE

[1205] Alex Sanchez, Jordi Ocaña, and Frederic Utzet. Sampling theory, estimation, and significance testing for Prevosti’s estimate of genetic distance. Biometrics, 51(4):1216–1235, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533255.

Fattinger:1995:NME

[1206] Karin E. Fattinger, Lewis B. Sheiner, and Davide Verotta. A new method to explore the distribution of interindividual random effects in non-linear mixed effects models. Biometrics, 51(4):1236–1251, December 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533256.

Hackett:1995:GMQ

[1207] Christine A. Hackett and Joel I. Weller. Genetic mapping of quantitative trait loci for traits with ordinal distributions. Biometrics, 51(4):1252– 1263, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533257.

Gennings:1995:UCR

[1208] Chris Gennings and W. Hans Carter, Jr. Utilizing concentration-response data from individual components to detect statistically significant departures from additivity in chemical mixtures. Biometrics, 51(4):1264–1277, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533258.

Little:1995:TMS

[1209] M. P. Little. Are two mutations sufficient to cause cancer? Some generalizations of the two-mutation model of carcinogenesis of Moolgavkar, Venzon, and Knudson, and of the multistage model of Armitage and Doll. Biometrics, 51(4):1278–1291, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533259.

Houwing-Duistermaat:1995:TFA

[1210] Jeanine J. Houwing-Duistermaat, Bert H. F. Derkx, Frits R. Rosendaal, and Hans C. van Houwelingen. Testing familial aggregation. Biometrics, 51(4):1292–1301, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533260.

Singh:1995:PDC

[1211] M. Singh and K. Hinkelmann. Partial diallel crosses in incomplete blocks. Biometrics, 51(4):1302–1314, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533261.

Proschan:1995:DES

[1212] Michael A. Proschan and Sally A. Hunsberger. Designed extension of studies based on conditional power. Biometrics, 51(4):1315–1324, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533262.

Karunamuni:1995:BEA

[1213] Rohana J. Karunamuni and Terrance J. Quinn, II. Bayesian estimation of animal abundance for line transect sampling. Biometrics, 51(4):1325– 1337, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533263.

Dror:1995:DTA

[1214] Shuki Dror, David Faraggi, and Benjamin Reiser. Dynamic treatment allocation adjusting for prognostic factors for more than two treatments. Biometrics, 51(4):1338–1343, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533264.

Mehrotra:1995:REE

[1215] Devan V. Mehrotra. Robust elementwise estimation of a dispersion matrix. Biometrics, 51(4):1344–1351, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533265. See corrections [1404].

vanderLinde:1995:STA

[1216] A. van der Linde, K.-H. Witzko, and K.-H. Jockel. Spatial-temporal analysis of mortality using splines. Biometrics, 51(4):1352–1360, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533266.

Huh:1995:SAP

[1217] Myung-Hoe Huh, Douglas E. Critchlow, Joseph S. Verducci, Janice Kiecolt-Glaser, Ronald Glaser, and William B. Malarkey. A symmetric analysis of paired rankings with application to temporal patterns of hormonal concentration. Biometrics, 51(4):1361–1371, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533267.

Bryant:1995:ITC

[1218] John Bryant and Roger Day. Incorporating toxicity considerations into the design of two-stage Phase II clinical trials. Biometrics, 51(4):1372– 1383, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533268.

Shih:1995:IAP

[1219] Joanna H. Shih and Thomas A. Louis. Inferences on the association parameter in copula models for bivariate survival data. Biometrics, 51 (4):1384–1399, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533269.

Paul:1995:APP

[1220] Sudhir R. Paul and Ali S. Islam. Analysis of proportions in the presence of over-/under-dispersion. Biometrics, 51(4):1400–1410, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533270.

Nam:1995:IES

[1221] Jun mo Nam. Interval estimation and significance testing for cyclic trends in seasonality studies. Biometrics, 51(4):1411–1417, December 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533271.

Lebreton:1995:SSR

[1222] Jean-Dominique Lebreton, Byron J. T. Morgan, Roger Pradel, and Stephen N. Freeman. A simultaneous survival rate analysis of dead recovery and live recapture data. Biometrics, 51(4):1418–1428, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533272.

Richardson:1995:RRM

[1223] Alice M. Richardson and Alan H. Welsh. Robust restricted maximum likelihood in mixed linear models. Biometrics, 51(4):1429–1439, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533273.

Gilmour:1995:AIR

[1224] Arthur R. Gilmour, Robin Thompson, and Brian R. Cullis. Average information REML: an efficient algorithm for variance parameter estimation in linear mixed models. Biometrics, 51(4):1440–1450, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533274.

OGorman:1995:UKC

[1225] Thomas W. O’Gorman and Robert F. Woolson. Using Kendall’s τ_{b} correlations to improve variable selection methods in case-control studies. Biometrics, 51(4):1451–1460, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533275.

Gutierrez:1995:ATR

[1226] Roberto G. Gutierrez, Raymond J. Carroll, Naisyin Wang, Gyung-Hee Lee, and Brian H. Taylor. Analysis of tomato root initiation using a normal mixture distribution. Biometrics, 51(4):1461–1468, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533276.

Grambsch:1995:DPR

[1227] Patricia M. Grambsch, Terry M. Therneau, and Thomas R. Fleming. Diagnostic plots to reveal functional form for covariates in multiplicative intensity models. Biometrics, 51(4):1469–1482, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533277.

Hung:1995:EEM

[1228] H. M. James Hung, George Y. H. Chi, and Robert T. O’Neill. Efficacy evaluation for monotherapies in two-by-two factorial trials. Biometrics, 51(4):1483–1493, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533278.

Paula:1995:OST

[1229] Gilberto A. Paula and Pranab K. Sen. One-sided tests in generalized linear models with parallel regression lines. Biometrics, 51(4):1494–1501, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533279. See cor-

rections [1403].

Baade:1995:TLL

[1230] Ingrid A. Baade and Anthony N. Pettitt. Tests of loglinear and linear relative risks for Cox’s model. Biometrics, 51(4):1502–1513, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533280.

Murtaugh:1995:RCM

[1231] Paul A. Murtaugh. ROC curves with multiple marker measurements. Biometrics, 51(4):1514–1522, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533281.

Bowman:1995:EVF

[1232] Dale Bowman, James J. Chen, and E. Olusegun George. Estimating variance functions in developmental toxicity studies. Biometrics, 51(4):1523– 1528, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533282.

Schouten:1995:ABU

[1233] Hubert J. A. Schouten. Adaptive biased urn randomization in small strata when blinding is impossible. Biometrics, 51(4):1529–1535, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533283.

Etchison:1995:PTS

[1234] Tonya Etchison, Cavell Brownie, and Sastry G. Pantula. A portmanteau test for spatial ARMA models. Biometrics, 51(4):1536–1542, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533284.

Mehrabi:1995:LBM

[1235] Yadollah Mehrabi and J. N. S. Matthews. Likelihood-based methods for bias reduction in limiting dilution assays. Biometrics, 51(4):1543–1549, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533285.

Verweij:1995:TDE

[1236] Pierre J. M. Verweij and Hans C. van Houwelingen. Time-dependent effects of fixed covariates in Cox regression. Biometrics, 51(4):1550–1556, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533286.

Hanley:1995:PDM

[1237] James A. Hanley and Brenda MacGibbon. The Poisson distribution and “Multiple events” cell occupancy problems. Biometrics, 51(4):1557–1561, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533287.

Buonaccorsi:1995:PPM

[1238] John P. Buonaccorsi. Prediction in the presence of measurement error: General discussion and an example predicting defoliation. Biometrics, 51(4):1562–1569, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533288.

Moulton:1995:MMD

[1239] Lawrence H. Moulton and Neal A. Halsey. A mixture model with detection limits for regression analyses of antibody response to vaccine.

Biometrics, 51(4):1570–1578, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533289.

Bolfarine:1995:CFC

[1240] Heleno Bolfarine and Manuel Galea-Rojas. Comments on “Functional Comparative Calibration Using an EM Algorithm” by D. Kimura. Biometrics, 51(4):1579–1580, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533290. See [585].

Sankoh:1995:NOO

[1241] A. J. Sankoh and M. F. Huque. A note on O’Brien’s OLS and GLS tests. Biometrics, 51(4):1580–1581, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533291.

Kocherlakota:1995:BRC

[1242] S. Kocherlakota. Book reviews: Continuous Univariate Distributions, Vol. 1, by N. L. Johnson, S. Kotz, N. Balakrishnan; Continuous Univariate Distributions, Vol. 2, by N. L. Johnson, S. Kotz, N. Balakrishnan. Biometrics, 51(4):1582–1583, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533292.

Kemp:1995:BRK

[1243] C. D. Kemp. Book review: Kendall’s Advanced Theory of Statistics, Vol. 1: Distribution Theory, by A. Stuart, J. K. Ord. Biometrics, 51 (4):1583–1584, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533293.

Burnham:1995:BRH

[1244] K. P. Burnham. Book review: Handbook of Statistics, Vol. 12: Environmental Statistics, by G. P. Patil, C. R. Rao. Biometrics, 51(4):1584–1585, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533294.

Howel:1995:BRH

[1245] D. Howel. Book review: The Handbook of Research Synthesis, by H. Cooper, L. V. Hedges. Biometrics, 51(4):1585–1586, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533295.

Derr:1995:BRH

[1246] J. Derr. Book review: A Handbook of Small Data Sets, by D. J. Hand, F. Daly, A. D. Lunn, K. J. McConway, E. Ostrowski. Biometrics, 51(4): 1586, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533296.

Kemp:1995:BRH

[1247] G. C. R. Kemp. Book review: Handbook of Statistical Modeling for the Social and Behavioral Sciences, by G. Arminger, C. C. Clogg, M. E. Sobel. Biometrics, 51(4):1586–1587, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533297.

Aitken:1995:BRF

[1248] C. G. G. Aitken. Book review: Fifty Years of Folly and Fraud “in the Name of Science”: From Crossroads to the Health Care Crisis, by I. D. Bross. Biometrics, 51(4):1587, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533298.

Anonymous:1995:BRCd

[1249] Anonymous. Book review: Cyclic and Computer Generated Designs, by J. A. John, E. R. Williams. Biometrics, 51(4):1587–1588, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533299.

Anonymous:1995:BRD

[1250] Anonymous. Book review: Differential Geometry and Statistics, by M. K. Murray, J. W. Rice. Biometrics, 51(4):1588, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533302.

Anonymous:1995:BRMb

[1251] Anonymous. Book review: Mathematical Computation with Maple V: Ideas and Applications, by T. Lee. Biometrics, 51(4):1588, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533303.

Anonymous:1995:BRR

[1252] Anonymous. Book review: Runs and Patterns in Probability: Selected Papers, by A. P. Godbole, S. G. Papastavridis. Biometrics, 51(4):1588, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533301.

Anonymous:1995:BRSh

[1253] Anonymous. Book review: Subjective Probability, by G. Wright, P. Ayton. Biometrics, 51(4):1588, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533300.

Anonymous:1995:CBR

[1254] Anonymous. Corrections: Book reviews. Biometrics, 51(4):1589, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533304. See [1180].

Anonymous:1995:CCT

[1255] Anonymous. Corrections: On the catch-up time method for analyzing cancer screening trials. Biometrics, 51(4):1589, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533306. See [1026].

Anonymous:1995:CEF

[1256] Anonymous. Corrections: [Editorial] Falguerolles. Biometrics, 51(4): 1589, December 1995. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533305. See

[1179].

Anonymous:1996:FMa

[1257] Anonymous. Front matter. Biometrics, 52(1):i–iv, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533138.

Bedrick:1996:AFL

[1258] Edward J. Bedrick and Joe R. Hill. Assessing the fit of the logistic regression model to individual matched sets of case-control data. Biometrics, 52(1):1–9, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533139.

Bilker:1996:SEM

[1259] Warren B. Bilker and Mei-Cheng Wang. A semiparametric extension of the Mann–Whitney test for randomly truncated data. Biometrics, 52(1): 10–20, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533140.

Tamhane:1996:MTP

[1260] Ajit C. Tamhane, Yosef Hochberg, and Charles W. Dunnett. Multiple test procedures for dose finding. Biometrics, 52(1):21–37, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533141.

Stewart-Oaten:1996:SEL

[1261] Allan Stewart-Oaten. Sequential estimation of Log(Abundance). Biometrics, 52(1):38–49, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533142.

Lambert:1996:MNG

[1262] Philippe Lambert. Modeling of nonlinear growth curve on series of correlated count data measured at unequally spaced times: a full likelihood based approach. Biometrics, 52(1):50–55, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533143.

Shen:1996:PUD

[1263] Pao-Sheng Shen, P. L. Cornelius, and R. L. Anderson. Planned unbalanced designs for estimation of quantitative genetic parameters. I: Twoway matings. Biometrics, 52(1):56–70, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533144.

Gordon:1996:MCC

[1264] Ian Gordon and Ray Watson. The myth of continuity-corrected sample size formulae. Biometrics, 52(1):71–76, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533145.

Jordan:1996:ECI

[1265] Scott M. Jordan and K. Krishnamoorthy. Exact confidence intervals for the common mean of several normal populations. Biometrics, 52(1):77– 86, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533146.

Sun:1996:SPT

[1266] Yanqing Sun and Michael Sherman. Some permutation tests for survival data. Biometrics, 52(1):87–97, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533147.

Little:1996:PMM

[1267] Roderick J. A. Little and Yongxiao Wang. Pattern-mixture models for multivariate incomplete data with covariates. Biometrics, 52(1):98–111, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533148.

Zhang:1996:MRM

[1268] Ping Zhang and Barbara Medoff-Cooper. A Markov regression model for nutritive sucking data. Biometrics, 52(1):112–124, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533149.

Munholland:1996:SLS

[1269] Patricia L. Munholland and John J. Borkowski. Simple latin square sampling + 1: a spatial design using quadrats. Biometrics, 52(1):125– 136, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533150.

Murray:1996:NSE

[1270] Susan Murray and Anastasios A. Tsiatis. Nonparametric survival estimation using prognostic longitudinal covariates. Biometrics, 52(1):137–151, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533151.

Heffernan:1996:ISP

[1271] Peter M. Heffernan. Improved sequential probability ratio tests for negative binomial populations. Biometrics, 52(1):152–157, March 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533152.

Belanger:1996:EVF

[1272] Bruce A. Belanger, Marie Davidian, and David M. Giltinan. The effect of variance function estimation on nonlinear calibration inference in immunoassay data. Biometrics, 52(1):158–175, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533153.

Donner:1996:THK

[1273] Allan Donner, Michael Eliasziw, and Neil Klar. Testing the homogeneity of kappa statistics. Biometrics, 52(1):176–183, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533154.

McCulloch:1996:CRE

[1274] Charles E. McCulloch, Michael D. Boudreau, and Sara Via. Confidence regions for evolutionary trajectories. Biometrics, 52(1):184–192, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533155.

Sinsheimer:1996:BHT

[1275] Janet S. Sinsheimer, James A. Lake, and Roderick J. A. Little. Bayesian hypothesis testing of four-taxon topologies using molecular sequence data. Biometrics, 52(1):193–210, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533156.

Fusaro:1996:CRA

[1276] Robert E. Fusaro, Peter Bacchetti, and Nicholas P. Jewell. A competing risks analysis of presenting AIDS diagnoses trends. Biometrics, 52(1): 211–225, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533157.

Rathbun:1996:EPI

[1277] Stephen L. Rathbun. Estimation of Poisson intensity using partially observed concomitant variables. Biometrics, 52(1):226–242, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533158.

Edland:1996:BSE

[1278] Steven D. Edland. Bias in slope estimates for the linear errors in variables model by the variance ratio method. Biometrics, 52(1):243–248, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533159.

Rubin:1996:MUE

[1279] Donald B. Rubin and Neal Thomas. Matching using estimated propensity scores: Relating theory to practice. Biometrics, 52(1):249–264, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533160.

Tempelman:1996:MEM

[1280] Robert J. Tempelman and Daniel Gianola. A mixed effects model for overdispersed count data in animal breeding. Biometrics, 52(1):265–279, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533161.

Gromping:1996:NFM

[1281] Ulrike Grömping. A note on fitting a marginal model to mixed effects loglinear regression data via GEE. Biometrics, 52(1):280–285, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533162.

Chi:1996:AED

[1282] Vu Xuan Chi and Nguyen Phong Chau. Antihypertensive effect of drugs: Statistical distribution of the trough-to-peak ratio. Biometrics, 52(1): 286–290, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533163.

Lipsitz:1996:JEV

[1283] Stuart R. Lipsitz and Michael Parzen. A jackknife estimator of variance for Cox regression for correlated survival data. Biometrics, 52(1):291– 298, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533164.

Zhou:1996:NML

[1284] Xiao H. Zhou. A nonparametric maximum likelihood estimator for the receiver operating characteristic curve area in the presence of verification bias. Biometrics, 52(1):299–305, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533165.

Fayyad:1996:NET

[1285] Rana Fayyad, Franklin A. Graybill, and Richard K. Burdick. A note on exact tests for variance components in unbalanced random and mixed linear models. Biometrics, 52(1):306–308, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533166.

Christensen:1996:ETV

[1286] Ronald Christensen. Exact tests for variance components. Biometrics, 52(1):309–314, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533167.

Piepho:1996:SPC

[1287] Hans-Peter Piepho. A simplified procedure for comparing the stability of cropping systems. Biometrics, 52(1):315–320, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533168.

Zhan:1996:ISP

[1288] M. Zhan, C. B. Dean, R. Routledge, P. Gallaugher, A. P. Farrell, and H. Thorarensen. Inference on segmented polynomial models. Biometrics, 52(1):321–327, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533169.

Le:1996:DVT

[1289] Chap T. Le and Bruce R. Lindgren. Duration of ventilating tubes: a test for comparing two clustered samples of censored data. Biometrics, 52(1): 328–334, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533170.

Solow:1996:BEG

[1290] Andrew R. Solow and Debra Palka. On Bayesian estimation of group size. Biometrics, 52(1):335–340, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533171.

Chinchilli:1996:WCC

[1291] Vernon M. Chinchilli, Juliann K. Martel, Shiriki Kumanyika, and Tom Lloyd. A weighted concordance correlation coefficient for repeated measurement designs. Biometrics, 52(1):341–353, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533172.

Lumley:1996:GEE

[1292] Thomas Lumley. Generalized estimating equations for ordinal data: a note on working correlation structures. Biometrics, 52(1):354–361, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533173.

Baker:1996:ACC

[1293] Stuart G. Baker. The analysis of categorical case-control data subject to nonignorable nonresponse. Biometrics, 52(1):362–369, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533174.

Lawson:1996:SIB

[1294] Andrew Lawson, Juan Ferrándiz, Antonio López, Augustin Llopis, and Maria Luisa Tejerizo. Spatial interaction between neighbouring counties.

Biometrics, 52(1):370–371, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533175.

Conniffe:1996:IPR

[1295] Denis Conniffe, Bas Engel, and Pieter Walstra. Increasing precision or reducing expense in regression experiments by using information from concomitant variable. Biometrics, 52(1):371–372, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533176.

Kemp:1996:BRC

[1296] A. W. Kemp. Book reviews: Case Studies in Biometry, by N. Lange, L. Ryan, L. Billard, D. Brillinger, L. Conquest, J. Greenhouse; A Casebook for a First Course in Statistics and Data Analysis, by S. Chatterjee, M. S. Handcock, J. S. Simonoff. Biometrics, 52(1):373–376, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533177.

Bjornsson:1996:BRS

[1297] H. Bjørnsson. Book review: Statistics in Ecology and Environmental Monitoring, by D. J. Fletcher, B. F. J. Manly. Biometrics, 52(1):376, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533178.

Derr:1996:BRP

[1298] J. Derr. Book review: Problem Solving: a Statistician’s Guide, by C. Chatfield. Biometrics, 52(1):377, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533179.

Howroyd:1996:BRF

[1299] J. D. Howroyd. Book review: Fractals, Random Shapes and Point Fields: Methods of Geometrical Statistics, by D. Stoyan, H. Stoyan. Biometrics, 52(1):377–378, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533180.

Anonymous:1996:BRI

[1300] Anonymous. Book review: Introduction to Biostatistics: a Guide to Design, Analysis, and Discovery, by R. N. Forthofer, E. S. Lee. Biometrics, 52(1):378, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533181.

Anonymous:1996:BRD

[1301] Anonymous. Book review: Distribution-Free Statistical Methods, by J. S. Maritz. Biometrics, 52(1):378–379, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533182.

Anonymous:1996:BRAa

[1302] Anonymous. Book review: Agricultural Sustainability: Economic, Environmental and Statistical Considerations, by V. Barnett, R. Payne, R. Steiner. Biometrics, 52(1):379, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533184.

Anonymous:1996:BRC

[1303] Anonymous. Book review: Computational Statistics, by P. Dirschedl, R. Ostermann. Biometrics, 52(1):379, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533183.

Anonymous:1996:BRMa

[1304] Anonymous. Book review: Measures of Information and Their Applications, by J. N. Kapur. Biometrics, 52(1):379, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533186.

Anonymous:1996:BRNa

[1305] Anonymous. Book review: New Cambridge Statistical Tables, by D. V. Lindley, W. F. Scott. Biometrics, 52(1):379, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533185.

Anonymous:1996:CEE

[1306] Anonymous. Correction to “Evaluation of experiments with adaptive interim analysis”, by P. Bauer and K. Kohne; 50, 1029–1041, December 1994. Biometrics, 52(1):380, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533187;   https://www.jstor.org/stable/2533189. See

Anonymous:1996:CIE

[1307] Anonymous. Corrections: Interval estimation of inverse dose-response. Biometrics, 52(1):380, March 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533188. See [1098].

Anonymous:1996:FMb

[1308] Anonymous. Front matter. Biometrics, 52(2):i–iv, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532880.

Wang:1996:MPR

[1309] Peiming Wang, Martin L. Puterman, Iain Cockburn, and Nhu Le. Mixed Poisson regression models with covariate dependent rates. Biometrics, 52(2):381–400, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532881.

Wang:1996:QEM

[1310] Naisyin Wang, R. J. Carroll, and Kung-Yee Liang. Quasilikelihood estimation in measurement error models with correlated replicates. Biometrics, 52(2):401–411, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532882.

Lui:1996:IEI

[1311] Kung-Jong Lui, William G. Cumberland, and Lynn Kuo. An interval estimate for the intraclass correlation in beta-binomial sampling. Biometrics, 52(2):412–425, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532883.

Korol:1996:LBQ

[1312] Abraham B. Korol, Yefim I. Ronin, and Valery M. Kirzhner. Linkage between quantitative trait loci and marker loci: Resolution power of three statistical approaches in single marker analysis. Biometrics, 52(2): 426–441, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532884.

Cook:1996:CES

[1313] Richard J. Cook. Coupled error spending functions for parallel bivariate sequential tests. Biometrics, 52(2):442–450, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532885.

Lee:1996:ILG

[1314] Jae Kyun Lee, Erik V. Nordheim, and Hyun Kang. Inference for lethal gene estimation with application in plants. Biometrics, 52(2):451–462, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532886.

Cheung:1996:SCI

[1315] Siu Hung Cheung and Wai Sum Chan. Simultaneous confidence intervals for pairwise multiple comparisons in a two-way unbalanced design. Biometrics, 52(2):463–472, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532887.

TenHave:1996:MEMa

[1316] Thomas R. Ten Have. A mixed effects model for multivariate ordinal response data including correlated discrete failure times with ordinal responses. Biometrics, 52(2):473–491, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532888.

TenHave:1996:MEMb

[1317] Thomas R. Ten Have. A mixed effects model for multivariate ordinal response data including correlated discrete failure times with ordinal responses. Biometrics, 52(2):473–491, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marschner:1996:FMI

[1318] Ian C. Marschner. Fitting a multiplicative incidence model to age- and time-specific prevalence data. Biometrics, 52(2):492–499, June 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532889.

Mancl:1996:ERE

[1319] Lloyd A. Mancl and Brian G. Leroux. Efficiency of regression estimates for clustered data. Biometrics, 52(2):500–511, June 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532890.

Alioum:1996:PHM

[1320] Ahmadou Alioum and Daniel Commenges. A proportional hazards model for arbitrarily censored and truncated data. Biometrics, 52(2):512–524, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532891.

Wei:1996:EFN

[1321] Run-Peng Wei and Dag Lindgren. Effective family number following selection with restrictions. Biometrics, 52(2):525–535, June 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532892.

Stram:1996:MAP

[1322] Daniel O. Stram. Meta-analysis of published data using a linear mixedeffects model. Biometrics, 52(2):536–544, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532893.

Swallow:1996:URS

[1323] William H. Swallow and Farid Kianifard. Using robust scale estimates in detecting multiple outliers in linear regression. Biometrics, 52(2):545– 556, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532894.

Cook:1996:RTT

[1324] Richard J. Cook, Jerald F. Lawless, and Claude Nadeau. Robust tests for treatment comparisons based on recurrent event responses. Biometrics, 52(2):557–571, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532895.

Vonesh:1996:GFG

[1325] Edward F. Vonesh, Vernon M. Chinchilli, and Kewei Pu. Goodness-of-fit in generalized nonlinear mixed-effects models. Biometrics, 52(2):572– 587, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532896.

Wang:1996:FMH

[1326] Yuedong Wang and Morton B. Brown. A flexible model for human circadian rhythms. Biometrics, 52(2):588–596, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532897.

David:1996:DI

[1327] Olivier David and Rob A. Kempton. Designs for interference. Biometrics, 52(2):597–606, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532898.

Bellavance:1996:TAV

[1328] F. Bellavance, S. Tardif, and M. A. Stephens. Tests for the analysis of variance of crossover designs with correlated errors. Biometrics, 52(2): 607–612, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532899.

Heise:1996:ODB

[1329] Mark A. Heise and Raymond H. Myers. Optimal designs for bivariate logistic regression. Biometrics, 52(2):613–624, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532900.

Hawkins:1996:ETP

[1330] D. L. Hawkins, C. P. Han, and J. Eisenfeld. Estimating transition probabilities from aggregate samples augmented by haphazard recaptures. Biometrics, 52(2):625–638, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532901.

Norris:1996:NMU

[1331] James L. Norris, III and Kenneth H. Pollock. Nonparametric MLE under two closed capture–recapture models with heterogeneity. Biometrics, 52(2):639–649, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532902.

Sammel:1996:LVM

[1332] Mary Dupuis Sammel and Louise M. Ryan. Latent variable models with fixed effects. Biometrics, 52(2):650–663, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532903.

Catchpole:1996:MSR

[1333] E. A. Catchpole and B. J. T. Morgan. Model selection in ring-recovery models using score tests. Biometrics, 52(2):664–672, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532904.

OQuigley:1996:CRM

[1334] John O’Quigley and Larry Z. Shen. Continual reassessment method: a likelihood approach. Biometrics, 52(2):673–684, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532905.

Bhoj:1996:EPG

[1335] Dinesh S. Bhoj and M. Ahsanullah. Estimation of parameters of the generalized geometric distribution using ranked set sampling. Biometrics, 52(2):685–694, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532906.

Barlow:1996:MIA

[1336] William Barlow. Measurement of interrater agreement with adjustment for covariates. Biometrics, 52(2):695–702, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532907.

Pradel:1996:UCM

[1337] R. Pradel. Utilization of capture-mark-recapture for the study of recruitment and population growth rate. Biometrics, 52(2):703–709, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532908.

Cadima:1996:SSR

[1338] Jorge F. C. L. Cadima and Ian T. Jolliffe. Size- and shape-related principal component analysis. Biometrics, 52(2):710–716, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532909.

Singh:1996:UES

[1339] M. K. Singh and Ram A. Kumar. On the unique equilibrium state of tetraploids under selection. Biometrics, 52(2):717–720, June 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532910.

Lee:1996:SVT

[1340] Jae Won Lee. Some versatile tests based on the simultaneous use of weighted log-rank statistics. Biometrics, 52(2):721–725, June 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532911.

Sun:1996:NTT

[1341] Jianguo Sun and John D. Kalbfleisch. Nonparametric tests of tumor prevalence data. Biometrics, 52(2):726–731, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532912.

Cook:1996:DCC

[1342] Richard J. Cook and Andrew R. Willan. Design considerations in crossover trials with a single interim analysis and serial patient entry. Biometrics, 52(2):732–739, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532913.

Rochon:1996:ABR

[1343] James Rochon. Analyzing bivariate repeated measures for discrete and continuous outcome variables. Biometrics, 52(2):740–750, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532914.

Lipsitz:1996:STI

[1344] Stuart R. Lipsitz and Garrett M. Fitzmaurice. The score test for independence in R × C contingency tables with missing data. Biometrics, 52(2): 751–762, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532915.

Grondona:1996:AVY

[1345] M. O. Grondona, J. Crossa, P. N. Fox, and W. H. Pfeiffer. Analysis of variety yield trials using two-dimensional separable ARIMA processes. Biometrics, 52(2):763–770, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532916.

Hannah:1996:POR

[1346] Murray Hannah and Paul Quigley. Presentation of ordinal regression analysis on the original scale. Biometrics, 52(2):771–775, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532917.

Kemp:1996:BRI

[1347] C. D. Kemp. Book review: An Introduction to Regression Graphics, by R. D. Cook, S. Weisberg. Biometrics, 52(2):776–777, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532918.

Kemp:1996:BRA

[1348] A. W. Kemp. Book review: Applied Stochastic Processes: a Biostatistical and Population Oriented Approach, by S. Biswas. Biometrics, 52(2): 777, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532919.

Anonymous:1996:BRB

[1349] Anonymous. Book review: Biometrics in Plant Breeding: Applications of Molecular Markers, by J. W. van Ooijen, J. Jansen. Biometrics, 52(2): 778, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532921.

Anonymous:1996:BREa

[1350] Anonymous. Book review: Epidemic Models: Their Structure and Relation to Data, by D. Mollison. Biometrics, 52(2):778, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532920.

Anonymous:1996:BRW

[1351] Anonymous. Book review: Workshop Statistics: Discovery with Data, by A. J. Rossman. Biometrics, 52(2):778–779, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532922.

Anonymous:1996:EN

[1352] Anonymous. Editors’ note. Biometrics, 52(2):779, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532925.

Anonymous:1996:BREb

[1353] Anonymous. Book review: Elementary Probability, by D. Stirzaker. Biometrics, 52(2):779, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532924.

Anonymous:1996:BRF

[1354] Anonymous. Book review: Fundamentals of Applied Statistics and Surveys, by D. B. Orr. Biometrics, 52(2):779, June 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532923.

Anonymous:1996:FMc

[1355] Anonymous. Front matter. Biometrics, 52(3):i–iv, September 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533041.

Brookmeyer:1996:AES

[1356] Ron Brookmeyer. AIDS, epidemics, and statistics. Biometrics, 52 (3):781–796, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533042.

Qu:1996:REM

[1357] Yinsheng Qu, Ming Tan, and Michael H. Kutner. Random effects models in latent class analysis for evaluating accuracy of diagnostic tests. Biometrics, 52(3):797–810, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533043.

Fay:1996:RIT

[1358] Michael P. Fay. Rank invariant tests for interval censored data under the grouped continuous model. Biometrics, 52(3):811–822, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533044.

Keen:1996:LES

[1359] K. J. Keen. Limiting the effects of single-member families in the estimation of the intraclass correlation. Biometrics, 52(3):823–832, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533045.

Mahdi:1996:CCS

[1360] Smail Mahdi and Sabin Lessard. Convergence of covariance structures in additive Gaussian polygenic models. Biometrics, 52(3):833–845, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533046.

Troendle:1996:PSM

[1361] James F. Troendle. A permutational step-up method of testing multiple outcomes. Biometrics, 52(3):846–859, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533047.

Schwarz:1996:GMA

[1362] Carl James Schwarz and A. Neil Arnason. A general methodology for the analysis of capture–recapture experiments in open populations. Biometrics, 52(3):860–873, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533048.

Quantin:1996:RSM

[1363] Catherine Quantin, Thierry Moreau, Bernard Asselain, Jean Maccario, and Joseph Lellouch. A regression survival model for testing the proportional hazards hypothesis. Biometrics, 52(3):874–885, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533049.

Cade:1996:PTL

[1364] Brian S. Cade and Jon D. Richards. Permutation tests for least absolute deviation regression. Biometrics, 52(3):886–902, September 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533050.

Lipsitz:1996:EEM

[1365] Stuart R. Lipsitz and Garrett M. Fitzmaurice. Estimating equations for measures of association between repeated binary responses. Biometrics, 52(3):903–912, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533051.

Akritas:1996:NIF

[1366] Michael G. Akritas and Michael P. LaValley. Nonparametric inference in factorial designs with censored data. Biometrics, 52(3):913–924, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533052.

Ejigou:1996:PSS

[1367] Ayenew Ejigou. Power and sample size for matched case-control studies. Biometrics, 52(3):925–933, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533053.

Walker:1996:EAN

[1368] Stephen Walker. An EM algorithm for nonlinear random effects models. Biometrics, 52(3):934–944, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533054.

Zhou:1996:MCA

[1369] Haibo Zhou and Clarice R. Weinberg. Modeling conception as an aggregated Bernoulli outcome with latent variables via the EM algorithm. Biometrics, 52(3):945–954, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533055.

Paul:1996:STI

[1370] Sudhir R. Paul. Score tests for interclass correlation in familial data. Biometrics, 52(3):955–963, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533056.

Lauter:1996:ETAa

[1371] Jürgen Läuter. Exact t and F tests for analyzing studies with multiple endpoints. Biometrics, 52(3):964–970, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533057.

Lauter:1996:ETAb

[1372] Jürgen Läuter. Exact t and F tests for analyzing studies with multiple endpoints. Biometrics, 52(3):964–970, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Olson:1996:THH

[1373] Jane M. Olson and Marianne Foley. Testing for homogeneity of Hardy– Weinberg disequilibrium using data sampled from several populations. Biometrics, 52(3):971–979, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533058.

Matis:1996:AME

[1374] James H. Matis and Thomas R. Kiffe. On approximating the moments of the equilibrium distribution of a stochastic logistic model. Biometrics, 52(3):980–991, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533059.

Yao:1996:OSS

[1375] Tzy-Jyun Yao, Colin B. Begg, and Philip O. Livingston. Optimal sample size for a series of pilot trials of new agents. Biometrics, 52(3):992–1001, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533060.

Hogan:1996:ITA

[1376] Joseph W. Hogan and Nan M. Laird. Intention-to-treat analyses for incomplete repeated measures data. Biometrics, 52(3):1002–1017, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533061.

Hirji:1996:NEA

[1377] Karim F. Hirji. A note on exact analysis of several 2 × 2 tables. Biometrics, 52(3):1018–1025, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533062.

Umphrey:1996:MLE

[1378] Gary J. Umphrey and John N. Haddad. Maximum likelihood estimation of allele frequencies in the eusocial hymenoptera using offspring genotypes. Biometrics, 52(3):1026–1032, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533063.

Lin:1996:GSD

[1379] D. Y. Lin, L. Shen, Z. Ying, and N. E. Breslow. Group sequential designs for monitoring survival probabilities. Biometrics, 52(3):1033–1041, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533064.

Datta:1996:DST

[1380] Susmita Datta and Jonathan Arnold. Diagnostics and a statistical test of neutrality hypotheses using the dynamics of cytonuclear disequilibria. Biometrics, 52(3):1042–1054, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533065.

Chick:1996:BML

[1381] Stephen E. Chick. Bayesian models for limiting dilution assay and group test data. Biometrics, 52(3):1055–1062, September 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533066.

Smith:1996:EMT

[1382] Woollcott Smith and Andrew R. Solow. An exact McNemar test for paired binary Markov chains. Biometrics, 52(3):1063–1070, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533067.

Ibrahim:1996:PEI

[1383] Joseph G. Ibrahim and Stuart R. Lipsitz. Parameter estimation from incomplete data in binomial regression when the missing data mechanism is nonignorable. Biometrics, 52(3):1071–1078, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533068.

Stram:1996:ATD

[1384] Daniel O. Stram, Harland Sather, and Letao Wang. Analysis of time to death after transplantation when transplants are only given to patients in remission. Biometrics, 52(3):1079–1086, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533069.

McKeown:1996:TAE

[1385] Sean P. McKeown and William D. Johnson. Testing for autocorrelation and equality of covariance matrices. Biometrics, 52(3):1087–1095, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533070.

Gatto:1996:SAM

[1386] Riccardo Gatto. Saddlepoint approximations of marginal densities and confidence intervals in the logistic regression measurement error model. Biometrics, 52(3):1096–1102, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533071.

Agresti:1996:ORT

[1387] Alan Agresti and Brent A. Coull. Order-restricted tests for stratified comparisons of binomial proportions. Biometrics, 52(3):1103–1111, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533072.

Hamilton:1996:QAP

[1388] Martin A. Hamilton and Todd A. DeVries. Quantitative analysis of a presence/absence microbiological assay: The hard surface carrier test of disinfectant efficacy. Biometrics, 52(3):1112–1120, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533073.

Molenberghs:1996:MAM

[1389] Geert Molenberghs and Linda L. Ritter. Methods for analyzing multivariate binary data, with association between outcomes of interest. Biometrics, 52(3):1121–1133, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533074.

Hepworth:1996:ECI

[1390] Graham Hepworth. Exact confidence intervals for proportions estimated by group testing. Biometrics, 52(3):1134–1146, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533075.

Koziol:1996:CSA

[1391] James A. Koziol and Shu-Chen H. Wu. Changepoint statistics for assessing a treatment-covariate interaction. Biometrics, 52(3):1147–1152, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533076.

Best:1996:NAD

[1392] D. J. Best and J. C. W. Rayner. Nonparametric analysis for doubly ordered two-way contingency tables. Biometrics, 52(3):1153–1156, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533077.

Marimuthu:1996:SAA

[1393] P. Marimuthu, Laurence Freedman, Douglas Midthune, and Charles Brown. Statistical analysis of animal cancer chemoprevention experiments. Biometrics, 52(3):1157, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533078.

Ringrose:1996:BRK

[1394] T. J. Ringrose. Book review: Kendall’s Library of Statistics, Multivariate Analysis. Part 1: Distributions, Ordination and Inference, by W. J.

Krzanowski, F. H. C. Marriott. Biometrics, 52(3):1158–1159, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533079.

Gordon:1996:BRK

[1395] A. D. Gordon. Book review: Kendall’s Library of Statistics, Multivariate Analysis. Part 2: Classification, Covariance Structures and Repeated Measurements, by W. J. Krzanowski, F. H. C. Marriott. Biometrics, 52(3):1159–1160, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533080.

Fearn:1996:BRB

[1396] T. Fearn. Book review: Bayesian Data Analysis, by A. Gelman, J. B. Carlin, H. S. Stern, D. B. Rubin. Biometrics, 52(3):1160–1161, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533081.

Edwards:1996:BRH

[1397] A. W. F. Edwards. Book review: Handbook of Human Genetic Linkage, by J. D. Terwilliger, J. Ott. Biometrics, 52(3):1161, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533082.

Kemp:1996:BRM

[1398] A. W. Kemp. Book review: Mathematics Handbook for Science and Engineering (BETA), by L. Rade, B. Westergren. Biometrics, 52(3):1161– 1162, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533083.

Anonymous:1996:BRAb

[1399] Anonymous. Book reviews: The Analysis of Time Series: an Introduction, by C. Chatfield; Introduction to Statistical Time Series, by W. A. Fuller; Applied Econometric Time Series, by W. Enders; Time Series Models in Econometrics, Finance and Other Fields, by D. R. Cox, D. V. Hinkley, O. E. Barndoff-Nielson. Biometrics, 52(3):1162–1163, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533084.

Anonymous:1996:BREc

[1400] Anonymous. Book review: Essays on Probability and Statistics: Festschrift in Honour of Professor Anil Kumar Bhattacharyya, by S. P.

Mukherjee, A. Chaudhuri, S. K. Basu. Biometrics, 52(3):1163, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533086.

Anonymous:1996:BRSa

[1401] Anonymous. Book review: Statistical Methods for Groundwater Monitoring, by R. D. Gibbons. Biometrics, 52(3):1163, September 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533085.

Anonymous:1996:CES

[1402] Anonymous. Correction: Estimation of the size distribution of fibrillar centres in nucleoli — an example of the ‘Swiss Cheese’ problem in stereology. Biometrics, 52(3):1164, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533089. See [483].

Anonymous:1996:COS

[1403] Anonymous. Corrections: One-sided tests in generalized linear models with parallel regression lines. Biometrics, 52(3):1164, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533088. See [1229].

Anonymous:1996:CRE

[1404] Anonymous. Corrections: Robust elementwise estimation of a dispersion matrix. Biometrics, 52(3):1164, September 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533087. See [1215].

Anonymous:1996:FMd

[1405] Anonymous. Front matter. Biometrics, 52(4):i–iv, December 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532832.

Anonymous:1996:VI

[1406] Anonymous. Volume information. Biometrics, 52(4):i–xvi, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532831.

Zhao:1996:RAM

[1407] Lue Ping Zhao, Stuart Lipsitz, and Danika Lew. Regression analysis with missing covariate data using estimating equations. Biometrics, 52 (4):1165–1182, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532833.

Bock:1996:HDM

[1408] R. Darrell Bock and Robert D. Gibbons. High-dimensional multivariate probit analysis. Biometrics, 52(4):1183–1194, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532834.

Quan:1996:ARW

[1409] Hui Quan and Weichung J. Shih. Assessing reproducibility by the withinsubject coefficient of variation with random effects models. Biometrics, 52(4):1195–1203, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532835.

Heller:1996:RPC

[1410] Glenn Heller and E. S. Venkatraman. Resampling procedures to compare two survival distributions in the presence of right-censored data. Biometrics, 52(4):1204–1213, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532836.

Staudte:1996:RCL

[1411] R. G. Staudte, J. Zhang, and R. M. Huggins. A reexamination of the cell-lineage data of E. P. Powell. Biometrics, 52(4):1214–1222, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532837.

Liu:1996:MHT

[1412] I-Ming Liu and Alan Agresti. Mantel–Haenszel-type inference for cumulative odds ratios with a stratified ordinal response. Biometrics, 52 (4):1223–1234, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532838.

Fung:1996:DIO

[1413] Wing K. Fung. Diagnosing influential observations in quadratic discriminant analysis. Biometrics, 52(4):1235–1241, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532839.

Coakley:1996:VST

[1414] Clint W. Coakley and Mark A. Heise. Versions of the sign test in the presence of ties. Biometrics, 52(4):1242–1251, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532840.

Wang:1996:QLA

[1415] You-Gan Wang. A quasi-likelihood approach for ordered categorical data with overdispersion. Biometrics, 52(4):1252–1258, December 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532841.

Qu:1996:UPT

[1416] Roger P. Qu and Mari Palta. Using projection for testing goodness-of-fit in regression models for repeated measures. Biometrics, 52(4):1259–1267, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532842.

Cierco:1996:CCI

[1417] Christine Cierco and Brigitte Mangin. Construction of confidence intervals for quantitative trait loci location in F_{2} populations. Biometrics, 52(4):1268–1282, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532843.

Chen:1996:SSS

[1418] Song Xi Chen. Studying school size effects in line transect sampling using the kernel method. Biometrics, 52(4):1283–1294, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532844.

Steele:1996:MEA

[1419] Brian M. Steele. A modified EM algorithm for estimation in generalized mixed models. Biometrics, 52(4):1295–1310, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532845.

Cook:1996:IML

[1420] Richard J. Cook and Jerald F. Lawless. Interim monitoring of longitudinal comparative studies with recurrent event responses. Biometrics, 52(4):1311–1323, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532846.

Little:1996:ITA

[1421] Roderick Little and Linda Yau. Intent-to-treat analysis for longitudinal studies with drop-outs. Biometrics, 52(4):1324–1333, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532847.

Cao:1996:PBI

[1422] Guoliang Cao and Mike West. Practical Bayesian inference using mixtures of mixtures. Biometrics, 52(4):1334–1341, December 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532848.

Akritas:1996:UNR

[1423] Michael G. Akritas. On the use of nonparametric regression techniques for fitting parametric regression models. Biometrics, 52(4):1342–1362, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532849.

Pan:1996:SSA

[1424] Guohua Pan. Subset selection with additional order information. Biometrics, 52(4):1363–1374, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532850.

Conaway:1996:DPI

[1425] Mark R. Conaway and Gina R. Petroni. Designs for Phase II trials allowing for a trade-off between response and toxicity. Biometrics, 52 (4):1375–1386, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532851.

Garthwaite:1996:CIR

[1426] Paul H. Garthwaite. Confidence intervals from randomization tests. Biometrics, 52(4):1387–1393, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532852.

Lewy:1996:GDD

[1427] Peter Lewy. A generalized Dirichlet distribution accounting for singularities of the variables. Biometrics, 52(4):1394–1409, December 1996. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532853.

Barrowman:1996:ETS

[1428] N. J. Barrowman and R. A. Myers. Estimating tag-shedding rates for experiments with multiple tag types. Biometrics, 52(4):1410–1416, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532854.

Lin:1996:MLA

[1429] Shili Lin. Multipoint linkage analysis via Metropolis jumping kernels. Biometrics, 52(4):1417–1427, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532855.

Ronin:1996:SEL

[1430] Yefim I. Ronin, Abraham B. Korol, Tzion Fahima, Valery M. Kirzhner, and Eviatar Nevo. Sequential estimation of linkage between PCRgenerated markers and a target gene employing stepwise bulked analysis. Biometrics, 52(4):1428–1439, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2532856.

Shih:1996:TIB

[1431] Joanna H. Shih and Thomas A. Louis. Tests of independence for bivariate survival data. Biometrics, 52(4):1440–1449, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532857.

Tritchler:1996:EAR

[1432] David Tritchler. Explanatory analyses of randomized studies. Biometrics, 52(4):1450–1456, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532858.

Bishir:1996:CEM

[1433] John Bishir and Richard A. Lancia. On catch-effort methods of estimating animal abundance. Biometrics, 52(4):1457–1466, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532859.

Krzanowski:1996:RDB

[1434] W. J. Krzanowski. Rao’s distance between normal populations that have common principal components. Biometrics, 52(4):1467–1471, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532860.

Smith:1996:ESO

[1435] Woollcott Smith, Andrew R. Solow, and Peter E. Preston. An estimator of species overlap using a modified beta-binomial model. Biometrics, 52(4):1472–1477, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532861.

Ibrahim:1996:UHC

[1436] Joseph G. Ibrahim and Louise M. Ryan. Use of historical controls in timeadjusted trend tests for carcinogenicity. Biometrics, 52(4):1478–1485, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532862.

Lee:1996:CWI

[1437] Youngjo Lee and Justus Seely. Computing the Wald interval for a variance ratio. Biometrics, 52(4):1486–1491, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532863.

Giraudeau:1996:CII

[1438] Bruno Giraudeau, Alain Mallet, and Claude Chastang. Case influence on the intraclass correlation coefficient estimate. Biometrics, 52(4):1492– 1497, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532864.

Williams:1996:DMB

[1439] Paige Williams and Louise Ryan. Design of multiple binary outcome studies with intentionally missing data. Biometrics, 52(4):1498–1514, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532865.

Grieve:1996:CIA

[1440] Andrew P. Grieve, Stephen Senn, and R. J. Cook. A comment on interim analyses in crossover trials. Biometrics, 52(4):1515–1520, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2532866.

Li:1996:TH

[1441] Billy Y. G. Li and George E. Booney. Test of homogeneity. Biometrics, 52(4):1521–1523, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532867.

Faraway:1996:TGL

[1442] Julian Faraway, Mohamed Lemdani, and Odile Pons. Tests for genetic linkage and homogeneity. Biometrics, 52(4):1523–1524, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532868. See corrections [1499].

Dunsmore:1996:BRK

[1443] I. R. Dunsmore. Book review: Kendall’s Advanced Theory of Statistics, Vol. 2b: Bayesian Inference, by A. O’Hagan. Biometrics, 52(4):1525– 1526, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2532869.

Kemp:1996:BRL

[1444] C. D. Kemp. Book reviews: LaTEX: a Document Preparation System: User’s Guide and Reference Manual, by L. Lamport; The LaTEX Companion, by M. Goosens, F. Mittelbach, A. Samarin; A Guide to LaTEX_{2ε}, by H. Kopka, P. W. Daly. Biometrics, 52(4):1526–1527, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532870.

Anonymous:1996:BRAc

[1445] Anonymous. Book review: Activity-Based Statistics: Student Guide, by R. L. Scheaffer, M. Gnanadesikan, A. Watkins, J. A. Witmer. Biometrics, 52(4):1528, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2532872.

Anonymous:1996:BRMb

[1446] Anonymous. Book review: Modern Applied Statistics with S-Plus, by W. N. Venables, B. D. Ripley. Biometrics, 52(4):1528, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2532871.

Anonymous:1996:BRSb

[1447] Anonymous. Book review: Survival Analysis: a Self-Learning Text, by D. G. Kleinbaum. Biometrics, 52(4):1528, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532873.

Anonymous:1996:BRSc

[1448] Anonymous. Book review: Statistics in Medical Research: Developments in Clinical Trials, by E. A. Gehan, N. A. Lemak. Biometrics, 52(4):1528, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532874.

Anonymous:1996:BRAd

[1449] Anonymous. Book review: Astrostatistics, by G. J. Babu, E. D. Feigelson. Biometrics, 52(4):1528–1529, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532875.

Anonymous:1996:BRNb

[1450] Anonymous. Book review: Nonlinear Mathematics and Its Applications, by P. J. Aston. Biometrics, 52(4):1529, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2532876.

Anonymous:1996:CBA

[1451] Anonymous. Corrections: Bayesian analysis of two overdispersed Poisson models. Biometrics, 52(4):1530, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2532878. See [1171].

Anonymous:1996:CUR

[1452] Anonymous. Corrections: Using replicate observations in observer agreement studies with binary assessments. Biometrics, 52(4):1530, December 1996. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2532877. See [392].

Anonymous:1997:FMa

[1453] Anonymous. Front matter. Biometrics, 53(1):i–iv, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533091.

Spears:1997:EIK

[1454] Floyd M. Spears, Barry W. Brown, and E. Neely Atkinson. The effect of incomplete knowledge of parameter values on single- and multiplestage designs for logistic regression. Biometrics, 53(1):1–10, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533092.

Hung:1997:BVWa

[1455] H. M. James Hung, Robert T. O’Neill, Peter Bauer, and Karl Köhne. The behavior of the P -value when the alternative hypothesis is true. Biometrics, 53(1):11–22, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533093.

Hung:1997:BVWb

[1456] H. M. James Hung, Robert T. O’Neill, Peter Bauer, and Karl Köhne. The behavior of the P -value when the alternative hypothesis is true. Biometrics, 53(1):11–22, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinsheimer:1997:IPU

[1457] Janet S. Sinsheimer, James A. Lake, and Roderick J. A. Little. Inference for phylogenies under a hybrid parsimony method: Evolutionarysymmetric transversion parsimony. Biometrics, 53(1):23–38, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533094.

Bar-Hen:1997:TSC

[1458] Avner Bar-Hen and J.-J. Daudin. A test of a special case of typicality in linear discriminant analysis. Biometrics, 53(1):39–48, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533095.

George:1997:TTE

[1459] Varghese George, William D. Johnson, Aditi Shahane, and Todd G. Nick. Testing for treatment effect in the presence of regression toward the mean. Biometrics, 53(1):49–59, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533096.

Pradel:1997:CRS

[1460] Roger Pradel, James E. Hines, Jean-Dominique Lebreton, and James D. Nichols. Capture–recapture survival models taking account of transients. Biometrics, 53(1):60–72, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533097.

Bloch:1997:CTD

[1461] Daniel A. Bloch. Comparing two diagnostic tests against the same “Gold Standard” in the same sample. Biometrics, 53(1):73–85, March 1997.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533098. See corrections [1694].

Chan:1997:MLE

[1462] Jennifer S. K. Chan and Anthony Y. C. Kuk. Maximum likelihood estimation for probit-linear mixed models with correlated random effects. Biometrics, 53(1):86–97, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533099.

Britton:1997:TDC

[1463] Tom Britton. Tests to detect clustering of infected individuals within families. Biometrics, 53(1):98–109, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533100.

Fitzmaurice:1997:RMM

[1464] Garrett M. Fitzmaurice and Nan M. Laird. Regression models for mixed discrete and continuous responses with potentially missing values. Biometrics, 53(1):110–122, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533101.

Kaur:1997:UAM

[1465] Amarjot Kaur, G. P. Patil, and Charles Taillie. Unequal allocation models for ranked set sampling with skew distributions. Biometrics, 53(1): 123–130, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533102.

Wang:1997:RCF

[1466] C. Y. Wang, Li Hsu, Z. D. Feng, and Ross L. Prentice. Regression calibration in failure time regression. Biometrics, 53(1):131–145, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533103.

Candy:1997:EFY

[1467] Steven G. Candy. Estimation in forest yield models using composite link functions with random effects. Biometrics, 53(1):146–160, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533104. See correction [1581].

Heuer:1997:MTT

[1468] Carsten Heuer. Modeling of time trends and interactions in vital rates using restricted regression splines. Biometrics, 53(1):161–177, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533105.

Schwarz:1997:ETM

[1469] Carl James Schwarz and Wayne T. Stobo. Estimating temporary migration using the robust design. Biometrics, 53(1):178–194, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533106.

Cardwell:1997:RTD

[1470] Kitty F. Cardwell and Thomas E. Wehrly. A rank test for distinguishing environmentally and genetically induced disease resistance in plant varieties. Biometrics, 53(1):195–206, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533107.

Gould:1997:CEE

[1471] William R. Gould and Kenneth H. Pollock. Catch-effort estimation of population parameters under the robust design. Biometrics, 53(1):207– 216, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533108.

Straatman:1997:ELT

[1472] Huub Straatman, Petronella G. M. Peer, and André L. M. Verbeek. Estimating lead time and sensitivity in a screening program without estimating the incidence in the screened group. Biometrics, 53(1):217– 229, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533109.

Gustafson:1997:LHB

[1473] Paul Gustafson. Large hierarchical Bayesian analysis of multivariate survival data. Biometrics, 53(1):230–242, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533110.

Rubin:1997:MSB

[1474] Donald B. Rubin and Ying Nian Wu. Modeling schizophrenic behavior using general mixture components. Biometrics, 53(1):243–261, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533111.

Eckert:1997:TAM

[1475] R. Stephen Eckert, Raymond J. Carroll, and Naisyin Wang. Transformations to additivity in measurement error models. Biometrics, 53(1): 262–272, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533112.

Garcia-Soidan:1997:SRM

[1476] Pilar H. Garcia-Soidan and Peter Hall. On sample reuse methods for spatial data. Biometrics, 53(1):273–281, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533113.

Li:1997:SEM

[1477] Hongzhe Li and Elizabeth Thompson. Semiparametric estimation of major gene and family-specific random effects for age of onset. Biometrics, 53(1):282–293, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533114.

Stephane:1997:FDA

[1478] Champely Stéphane, Guinand Bruno, Thioulouse Jean, and Clermidy Annabelle. Functional data analysis of curve asymmetry with application to the color pattern of hydropsyche contubernalis head capsule. Biometrics, 53(1):294–305, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533115.

Hoenig:1997:CCR

[1479] John M. Hoenig, Cynthia M. Jones, Kenneth H. Pollock, Douglas S. Robson, and David L. Wade. Calculation of catch rate and total catch in roving surveys of anglers. Biometrics, 53(1):306–317, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533116.

Scheike:1997:DSM

[1480] Thomas H. Scheike and Tina Kold Jensen. A discrete survival model with random effects: an application to time to pregnancy. Biometrics, 53(1): 318–329, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533117.

Wulfsohn:1997:JMS

[1481] Michael S. Wulfsohn and Anastasios A. Tsiatis. A joint model for survival and longitudinal data measured with error. Biometrics, 53(1):330–339, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533118.

Piepho:1997:DFT

[1482] Hans-Peter Piepho. Distribution-free tests for one-way homoscedasticity in a two-way classification. Biometrics, 53(1):340–351, March 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533119.

Cook:1997:LMT

[1483] Richard J. Cook and Stephen D. Walter. A logistic model for trend in 2 × 2 × K tables with applications to meta-analyses. Biometrics, 53(1): 352–357, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533120.

Cook:1997:LBN

[1484] Richard J. Cook and Edmund T. M. Ng. A logistic-bivariate normal model for overdispersed two-state Markov processes. Biometrics, 53(1): 358–364, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533121.

Emerson:1997:CSA

[1485] Scott S. Emerson and John M. Kittelson. A computationally simpler algorithm for the UMVUE of a normal mean following a group sequential trial. Biometrics, 53(1):365–369, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533122.

Song:1997:ACR

[1486] Hae Hiang Song. Analysis of correlated ROC areas in diagnostic testing. Biometrics, 53(1):370–382, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533123.

Roy:1997:AIP

[1487] Tapon Roy and Robert Schall. Assessment of individual and population bioequivalence using the probability that bioavailabilities are similar. Biometrics, 53(1):383–384, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533124.

Follmann:1997:AGL

[1488] Dean A. Follmann. An appropriate generalized linear model with random effects for informative missing data. Biometrics, 53(1):384, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533125.

Leonard:1997:BRB

[1489] T. Leonard. Book review: Bayes and Empirical Bayes Methods for Data Analysis, by B. P. Carlin, T. A. Louis. Biometrics, 53(1):385–386, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533126.

Brown:1997:BRS

[1490] R. A. Brown. Book review: Statistical Methods: a Geometric Primer, by D. J. Saville, G. R. Wood. Biometrics, 53(1):386–387, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533127.

Hand:1997:BRA

[1491] D. J. Hand. Book review: Analyse Discriminante sur Variables Qualitatives, by G. Celeux, J.-P. Nakache. Biometrics, 53(1):387–388, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533128.

Jolliffe:1997:BRS

[1492] I. T. Jolliffe. Book review: Statistical Factor Analysis and Related Methods: Theory and Applications, by A. Basilevsky. Biometrics, 53(1):388, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533129.

Wood:1997:BRA

[1493] S. N. Wood. Book review: Applied Population Ecology: a Supply-Demand Approach, by A. P. Gutierrez. Biometrics, 53(1):388–389, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533130.

Kemp:1997:BRPa

[1494] A. W. Kemp. Book reviews: Probability with a View toward Statistics, Volumes I and II, by J. Hoffmann-Jorgensen; Elementary Applications of Probability Theory, by H. C. Tuckwell; Problems and Snapshots from the World of Probability, by G. Blom, L. Holst, D. Sandell; Algorithmic Probability: a Collection of Problems, by M. F. Neuts. Biometrics, 53(1): 389–391, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533131.

Anonymous:1997:BRS

[1495] Anonymous. Book review: Statistical Methods in Medical Research, by P. Armitage, G. Berry. Biometrics, 53(1):391, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533132.

Anonymous:1997:BRN

[1496] Anonymous. Book review: Nonparametric Methods for Quantitative Analysis, by J. D. Gibbons. Biometrics, 53(1):391–392, March 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533133.

Anonymous:1997:BRG

[1497] Anonymous. Book review: Genetic Data Analysis II, by B. S. Weir. Biometrics, 53(1):392, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533134.

Anonymous:1997:BRRa

[1498] Anonymous. Book reviews: Readings from The Environmental Professional: Risk Assessment, by J. Lemons; Readings from The Environmental Professional: Natural Resources, by J. Lemons; Readings from The Environmental Professional: National Environmental Policy Act, by J. Lemons. Biometrics, 53(1):392, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533135.

Anonymous:1997:CTG

[1499] Anonymous. Corrections: Tests for genetic linkage and homogeneity. Biometrics, 53(1):393, March 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533136. See [1442].

Anonymous:1997:FMb

[1500] Anonymous. Front matter. Biometrics, 53(2):i–iv, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533944.

Spiegelman:1997:FPS

[1501] Donna Spiegelman and Mario Casella. Fully parametric and semiparametric regression models for common events with covariate measurement error in main study/validation study designs. Biometrics, 53(2): 395–409, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533945.

Lim:1997:NTC

[1502] Dong Hoon Lim and Douglas A. Wolfe. Nonparametric tests for comparing umbrella pattern treatment effects with a control in a randomized block design. Biometrics, 53(2):410–418, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533946.

Lin:1997:EMC

[1503] D. Y. Lin, E. J. Feuer, R. Etzioni, and Y. Wax. Estimating medical costs from incomplete follow-up data. Biometrics, 53(2):419–434, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533947.

Ahn:1997:TSL

[1504] Hongshik Ahn and James J. Chen. Tree-structured logistic model for over-dispersed binomial data with application to modeling developmental effects. Biometrics, 53(2):435–455, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533948.

Simon:1997:BDA

[1505] Richard Simon and Laurence S. Freedman. Bayesian design and analysis of two × two factorial clinical trials. Biometrics, 53(2):456–464, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533949.

Ibrahim:1997:PVS

[1506] Joseph G. Ibrahim and Ming-Hui Chen. Predictive variable selection for the multivariate linear model. Biometrics, 53(2):465–478, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533950.

George:1997:TTR

[1507] E. Olusegun George and Hulin Wu. Testing for treatment related trend with partially exchangeable clustered data. Biometrics, 53(2):479–487, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533951.

Link:1997:EPT

[1508] William A. Link and John R. Sauer. Estimation of population trajectories from count data. Biometrics, 53(2):488–497, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533952.

Wang:1997:LST

[1509] Sue-Jane Wang and H. M. James Hung. Large sample tests for binary outcomes in fixed-dose combination drug studies. Biometrics, 53(2):498– 503, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533953.

Gillet:1997:MLE

[1510] Elizabeth M. Gillet. Maximum likelihood estimators of the gametic contributions to single-plant progenies. Biometrics, 53(2):504–523, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533954.

Craig:1997:AAS

[1511] Bruce A. Craig, Michael A. Newton, Robert A. Garrott, John E. Reynolds III, and J. Ross Wilcox. Analysis of aerial survey data on Florida manatee using Markov chain Monte Carlo. Biometrics, 53(2): 524–541, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533955.

Sy:1997:SMA

[1512] J. P. Sy, J. M. G. Taylor, and W. G. Cumberland. A stochastic model for the analysis of bivariate longitudinal AIDS data. Biometrics, 53(2):542– 555, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533956.

Rosenbaum:1997:SRS

[1513] Paul R. Rosenbaum. Signed rank statistics for coherent predictions. Biometrics, 53(2):556–566, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533957.

Obuchowski:1997:NAC

[1514] Nancy A. Obuchowski. Nonparametric analysis of clustered ROC curve data. Biometrics, 53(2):567–578, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533958.

Xu:1997:NES

[1515] Jian-Lun Xu and Philip C. Prorok. Nonparametric estimation of solid cancer size at metastasis and probability of presenting with metastasis at detection. Biometrics, 53(2):579–591, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533959.

Weiss:1997:PMS

[1516] Robert E. Weiss, Yan Wang, and Joseph G. Ibrahim. Predictive model selection for repeated measures random effects models using Bayes factors. Biometrics, 53(2):592–602, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533960.

Buckland:1997:MSI

[1517] S. T. Buckland, K. P. Burnham, and N. H. Augustin. Model selection: an integral part of inference. Biometrics, 53(2):603–618, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533961.

Cerioli:1997:MTI

[1518] Andrea Cerioli. Modified tests of independence in 2×2 tables with spatial data. Biometrics, 53(2):619–628, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533962.

Muranty:1997:SGL

[1519] Hélène Muranty and Bruno Goffinet. Selective genotyping for location and estimation of the effect of a quantitative trait locus. Biometrics, 53(2):629–643, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533963.

Reiser:1997:CIG

[1520] Benjamin Reiser and David Faraggi. Confidence intervals for the generalized ROC criterion. Biometrics, 53(2):644–652, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533964.

Kao:1997:GFO

[1521] Chen-Hung Kao and Zhao-Bang Zeng. General formulas for obtaining the MLEs and the asymptotic variance–covariance matrix in mapping quantitative trait loci when using the EM algorithm. Biometrics, 53(2): 653–665, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533965.

Barker:1997:JML

[1522] Richard J. Barker. Joint modeling of live-recapture, tag-resight, and tagrecovery data. Biometrics, 53(2):666–677, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533966.

Chao:1997:EOC

[1523] Wei-Hsiung Chao, Mari Palta, and Terry Young. Effect of omitted confounders on the analysis of correlated binary data. Biometrics, 53(2): 678–689, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533967.

Borgan:1997:EER

[1524] Ørnulf Borgan and Bryan Langholz. Estimation of excess risk from casecontrol data using Aalen’s linear regression model. Biometrics, 53(2): 690–697, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533968.

McShane:1997:LPR

[1525] Lisa M. McShane, Paul S. Albert, and Meg A. Palmatier. A latent process regression model for spatially correlated count data. Biometrics, 53(2): 698–706, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533969.

Williamson:1997:AIA

[1526] John M. Williamson and Amita K. Manatunga. Assessing interrater agreement from dependent data. Biometrics, 53(2):707–714, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533970.

Finney:1997:RR

[1527] David J. Finney. The responsible referee. Biometrics, 53(2):715–719, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533971.

Stukel:1997:TSM

[1528] Thérèse A. Stukel and Eugene Demidenko. Two-stage method of estimation for general linear growth curve models. Biometrics, 53(2):720–728, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533972.

Singer:1997:RMA

[1529] Julio M. Singer and Dalton F. Andrade. Regression models for the analysis of pretest/posttest data. Biometrics, 53(2):729–735, June 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533973.

Wallenstein:1997:WLT

[1530] Sylvan Wallenstein and Agnes Berger. Weighted logrank tests to detect a transient improvement in survivorship. Biometrics, 53(2):736–744, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533974.

Durham:1997:RWR

[1531] Stephen D. Durham, Nancy Flournoy, and William F. Rosenberger. A random walk rule for Phase I clinical trials. Biometrics, 53(2):745–760, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533975.

Piepho:1997:AGE

[1532] H.-P. Piepho. Analyzing genotype-environment data by mixed models with multiplicative terms. Biometrics, 53(2):761–766, June 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533976.

Langholz:1997:EAR

[1533] Bryan Langholz and Ørnulf Borgan. Estimation of absolute risk from nested case-control data. Biometrics, 53(2):767–774, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533977. See correction [2115].

Atkinson:1997:CUC

[1534] Greg Atkinson and Alan Nevill. Comment on the use of concordance correlation to assess the agreement between two variables. Biometrics, 53(2): 775–777, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533978. See

rejoinder [1535].

I-Kuei:1997:RLE

[1535] Lawrence I-Kuei and Vernon Chinchilli. Rejoinder to the Letter to the Editor from Atkinson and Nevill. Biometrics, 53(2):777–778, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533979. See [1534].

Buckland:1997:BRA

[1536] S. T. Buckland. Book review: Advances in Biometry, by P. Armitage, H. A. David. Biometrics, 53(2):779–780, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533980.

Leonard:1997:BRO

[1537] T. Leonard. Book review: Operational Subjective Statistical Methods: a Mathematical, Philosophical, and Historical Introduction, by F. Lad. Biometrics, 53(2):780–781, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533981.

Kemp:1997:BRPb

[1538] A. W. Kemp. Book review: Probability and Information, by D. Applebaum. Biometrics, 53(2):781–782, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533982.

Anonymous:1997:BRAa

[1539] Anonymous. Book review: Analysis of Longitudinal Data, by P. J. Diggle, K.-Y. Liang, S. L. Zeger. Biometrics, 53(2):782, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533983.

Anonymous:1997:BRCa

[1540] Anonymous. Book review: Computer-Aided Multivariate Analysis, by A. A. Afifi, V. Clark. Biometrics, 53(2):782, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533984.

Anonymous:1997:BRMa

[1541] Anonymous. Book review: MSI-2000: Multivariate Statistical Analysis in Honor of Professor Minoru Siotani on his 70th Birthday, Vols. 1 and 2, by T. Hayakawa, M. Aoshima, K. Shimizu. Biometrics, 53(2):782– 783, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533985.

Anonymous:1997:BRAb

[1542] Anonymous. Book review: Applied Factor Analysis in the Natural Sciences, by R. Reyment, K. G. Jöreskog. Biometrics, 53(2):783, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533986.

Anonymous:1997:BRIa

[1543] Anonymous. Book review: Image Analysis for the Biological Sciences, by C. A. Glasbey, G. W. Horgan. Biometrics, 53(2):783, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533987.

Anonymous:1997:BRIb

[1544] Anonymous. Book review: Introduction to Stochastic Processes, by G. F. Lawler. Biometrics, 53(2):783, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533988.

Anonymous:1997:FMc

[1545] Anonymous. Front matter. Biometrics, 53(3):i–iv, September 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533541.

Yue:1997:DFM

[1546] Huibin Yue and K. S. Chan. A dynamic frailty model for multivariate survival data. Biometrics, 53(3):785–793, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533542.

Betensky:1997:ESA

[1547] Rebecca A. Betensky. Early stopping to accept H_{0} based on conditional power: approximations and comparisons. Biometrics, 53(3):794–806, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533543.

Betensky:1997:SAC

[1548] Rebecca A. Betensky. Sequential analysis of censored survival data from three treatment groups. Biometrics, 53(3):807–822, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533544.

Coffin:1997:ROC

[1549] Marie Coffin and Shashikala Sukhatme. Receiver operating characteristic studies and measurement errors. Biometrics, 53(3):823–837, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533545.

Young:1997:FTS

[1550] David A. Young, Gary O. Zerbe, and William W. Hay, Jr. Fieller’s theorem, Scheffé simultaneous confidence intervals, and ratios of parameters of linear and nonlinear mixed-effects models. Biometrics, 53(3):838–847, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533546.

Carothers:1997:LBA

[1551] Andrew D. Carothers. A likelihood-based approach to the estimation of relative DNA copy number by comparative genomic hybridization. Biometrics, 53(3):848–856, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533547.

Troxel:1997:WEE

[1552] Andrea B. Troxel, Stuart R. Lipsitz, and Troyen A. Brennan. Weighted estimating equations with nonignorably missing response data. Biometrics, 53(3):857–869, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533548.

Klein:1997:AMP

[1553] Stefan Klein, Jens Timmer, and Josef Honerkamp. Analysis of multichannel patch clamp recordings by hidden Markov models. Biometrics, 53(3):870–884, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533549.

Kang:1997:MGF

[1554] Shin-Soo Kang and Kenneth J. Koehler. Modification of the Greenwood formula for correlated response times. Biometrics, 53(3):885–899, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533550. See cor-

rections [1642, 1643].

Tu:1997:GCA

[1555] X. M. Tu, J. Kowalski, J. Randall, J. Mendoza-Blanco, M. K. Shear, T. H. Monk, E. Frank, and D. J. Kupfer. Generalized covariance-adjusted discriminants: Perspective and application. Biometrics, 53(3):900–909, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533551.

Lin:1997:LMM

[1556] Xihong Lin, Jonathan Raz, and Sioḃán D. Harlow. Linear mixed models with heterogeneous within-cluster variances. Biometrics, 53(3):910–923, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533552.

Lin:1997:AEI

[1557] Hung-Mo Lin and Michael D. Hughes. Assessing the effects of interventions using longitudinal data with samples subject to selection. Biometrics, 53(3):924–936, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533553.

Liu:1997:SSC

[1558] Guanghan Liu and Kung-Yee Liang. Sample size calculations for studies with correlated observations. Biometrics, 53(3):937–947, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533554.

Yang:1997:LVM

[1559] Ilsoon Yang and Mark P. Becker. Latent variable modeling of diagnostic accuracy. Biometrics, 53(3):948–958, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533555.

M:1997:TSA

[1560] Mohammad Salehi M. and George A. F. Seber. Two-stage adaptive cluster sampling. Biometrics, 53(3):959–970, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533556.

Waller:1997:LLM

[1561] Lance A. Waller and Daniel Zelterman. Log-linear modeling with the negative multinomial distribution. Biometrics, 53(3):971–982, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533557.

Kenward:1997:SSI

[1562] Michael G. Kenward and James H. Roger. Small sample inference for fixed effects from restricted maximum likelihood. Biometrics, 53(3):983– 997, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533558.

Le:1997:ECE

[1563] Chap T. Le. Evaluation of confounding effects in ROC studies. Biometrics, 53(3):998–1007, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533559.

Lyles:1997:DEA

[1564] Robert H. Lyles and Lawrence L. Kupper. A detailed evaluation of adjustment methods for multiplicative measurement error in linear regression with applications in occupational epidemiology. Biometrics, 53 (3):1008–1025, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533560.

Goodnight:1997:BCG

[1565] Charles J. Goodnight and James M. Schwartz. A bootstrap comparison of genetic covariance matrices. Biometrics, 53(3):1026–1039, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533561.

Hitchcock:1997:BIR

[1566] D. Hitchcock and C. A. Glasbey. Binary image restoration at subpixel resolution. Biometrics, 53(3):1040–1053, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533562.

Borkowf:1997:ABC

[1567] Craig B. Borkowf, Mitchell H. Gail, Raymond J. Carroll, and Richard D. Gill. Analyzing bivariate continuous data grouped into categories defined by empirical quantiles of marginal distributions. Biometrics, 53(3):1054– 1069, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533563.

Yu:1997:RER

[1568] Philip L. H. Yu and K. Lam. Regression estimator in ranked set sampling. Biometrics, 53(3):1070–1080, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2533564. See correction [1797].

Wang:1997:UTS

[1569] Sue-Jane Wang and H. M. James Hung. Use of two-stage test statistic in the two-period crossover trials. Biometrics, 53(3):1081–1091, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533565.

Pontius:1997:SAC

[1570] Jeffrey S. Pontius. Strip adaptive cluster sampling: Probability proportional to size selection of primary units. Biometrics, 53(3):1092–1096, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533566.

Brooks:1997:FMM

[1571] S. P. Brooks, B. J. T. Morgan, M. S. Ridout, and S. E. Pack. Finite mixture models for proportions. Biometrics, 53(3):1097–1115, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533567. See correction [1642].

Albert:1997:GEE

[1572] Paul S. Albert, Dean A. Follmann, and Huiman X. Barnhart. A generalized estimating equation approach for modeling random length binary vector data. Biometrics, 53(3):1116–1124, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533568.

Bauer:1997:NMT

[1573] P. Bauer. A note on multiple testing procedures in dose finding. Biometrics, 53(3):1125–1128, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533569.

Zhou:1997:MCM

[1574] Xiao-Hua Zhou, Sujuan Gao, and Siu L. Hui. Methods for comparing the means of two independent log-normal samples. Biometrics, 53(3):1129– 1135, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533570.

Jolliffe:1997:MMC

[1575] Ian T. Jolliffe and Anna R. Jolliffe. Modelling memory in coal tits: an illustration of the EM algorithm. Biometrics, 53(3):1136–1142, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533571.

Dette:1997:OAT

[1576] Holger Dette and Axel Munk. Optimum allocation of treatments for Welch’s test in equivalence assessment. Biometrics, 53(3):1143–1150, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533572.

Hertz-Picciotto:1997:VEA

[1577] Irva Hertz-Picciotto and Beverly Rockhill. Validity and efficiency of approximation methods for tied survival times in Cox regression. Biometrics, 53(3):1151–1156, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533573.

Barlow:1997:GML

[1578] William E. Barlow. Global measures of local influence for proportional hazards regression models. Biometrics, 53(3):1157–1162, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533574.

Gautam:1997:TLT

[1579] Shiva Gautam. Test for linear trend in 2 × K ordered tables with openended categories. Biometrics, 53(3):1163–1169, September 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533575.

Drescher:1997:EGI

[1580] Karsten Drescher and Heiko Becher. Estimating the generalized impact fraction from case-control data. Biometrics, 53(3):1170–1176, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533576.

Anonymous:1997:CEF

[1581] Anonymous. Correction: Estimation in forest yield models using composite link functions with random effects. Biometrics, 53(3):1177, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533577. See [1467].

Kemp:1997:BRS

[1582] A. W. Kemp. Book review: Stochastic Modeling of Scientific Data, by P. Guttorp. Biometrics, 53(3):1178–1179, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533578.

Crowder:1997:BRN

[1583] M. J. Crowder. Book review: Nonlinear Models for Repeated Measurement Data, by M. Davidian, D. M. Giltinan. Biometrics, 53(3):1179– 1180, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533579.

Hutton:1997:BRM

[1584] J. L. Hutton. Book review: Measurement Error in Nonlinear Models, by R. J. Carroll, D. Ruppert, L. A. Stefanski. Biometrics, 53(3):1180–1181, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533580.

Ringrose:1997:BRC

[1585] T. J. Ringrose. Book review: Construction and Assessment of Classification Rules, by D. J. Hand. Biometrics, 53(3):1181–1182, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533581.

Anonymous:1997:BRBa

[1586] Anonymous. Book review: Beyond Anova: Basics of Applied Statistics, by W. R. G. Miller. Biometrics, 53(3):1182, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533582.

Anonymous:1997:BRBb

[1587] Anonymous. Book review: Biostatistics: a Methodology for the Health Sciences, by L. D. Fisher, G. Van Belle. Biometrics, 53(3):1182, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533583.

Anonymous:1997:BRCb

[1588] Anonymous. Book review: Clustering and Classification, by P. Arabie, L. J. Hubert, G. de Soete. Biometrics, 53(3):1182, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533584.

Anonymous:1997:BRF

[1589] Anonymous. Book review: Fractal Geometry, by K. Falconer. Biometrics, 53(3):1183, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533585.

Anonymous:1997:BRTa

[1590] Anonymous. Book review: A Tour of the Calculus, by D. Berlinsky. Biometrics, 53(3):1183, September 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533586.

Anonymous:1997:FMd

[1591] Anonymous. Front matter. Biometrics, 53(4):i–iv, December 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533488.

Anonymous:1997:VI

[1592] Anonymous. Volume information. Biometrics, 53(4):i–xvi, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533487.

Zhang:1997:MHT

[1593] Jie Zhang and Dennis D. Boos. Mantel–Haenszel test statistics for correlated binary data. Biometrics, 53(4):1185–1198, December 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533489.

Younes:1997:LBM

[1594] Naji Younes and John Lachin. Link-based models for survival data with interval and continuous time censoring. Biometrics, 53(4):1199–1211, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533490.

Banneheka:1997:STS

[1595] Sarath G. Banneheka, Richard D. Routledge, and Carl J. Schwarz. Stratified two-sample tag-recovery census of closed populations. Biometrics, 53(4):1212–1224, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533491.

Hougaard:1997:AOC

[1596] Philip Hougaard, Mei-Ling Ting Lee, and G. A. Whitmore. Analysis of overdispersed count data by mixtures of Poisson variables and Poisson processes. Biometrics, 53(4):1225–1238, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533492.

Dey:1997:BAN

[1597] Dipak K. Dey, Ming-Hui Chen, and Hong Chang. Bayesian approach for nonlinear random effects models. Biometrics, 53(4):1239–1252, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533493.

Sasieni:1997:GGD

[1598] Peter D. Sasieni. From genotypes to genes: Doubling the sample size. Biometrics, 53(4):1253–1261, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533494.

Sun:1997:BDD

[1599] Dongchu Sun and Robert K. Tsutakawa. Bayesian design for doseresponse curves with penalized risk. Biometrics, 53(4):1262–1273, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533495.

Kim:1997:RAI

[1600] Dong K. Kim. Regression analysis of interval-censored survival data with covariates using log-linear models. Biometrics, 53(4):1274–1283, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533496.

Lachenbruch:1997:DD

[1601] Peter A. Lachenbruch. Discriminant diagnostics. Biometrics, 53(4):1284– 1292, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533497.

Salter:1997:TRE

[1602] A. Salter and P. J. Solomon. Truncated recurrent event survival models for methadone data. Biometrics, 53(4):1293–1303, December 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533498.

Zeng:1997:CIB

[1603] Qi Zeng and Marie Davidian. Calibration inference based on multiple runs of an immunoassay. Biometrics, 53(4):1304–1317, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533499.

Burch:1997:ECI

[1604] Brent D. Burch and Hari K. Iyer. Exact confidence intervals for a variance ratio (or heritability) in a mixed linear model. Biometrics, 53(4):1318– 1333, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533500.

Staniswalis:1997:SRA

[1605] Joan G. Staniswalis, Peter F. Thall, and John Salch. Semiparametric regression analysis for recurrent event interval counts. Biometrics, 53 (4):1334–1353, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533501.

Wang:1997:EES

[1606] Jinliang Wang. Effect of excluding sib matings on inbreeding coefficient and effective size of finite diploid populations. Biometrics, 53(4):1354– 1365, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533502.

Ng:1997:NRA

[1607] Meei Pyng Ng and Gary K. Grunwald. Nonlinear regression analysis of the joint-regression model. Biometrics, 53(4):1366–1372, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533503.

Frensham:1997:GEV

[1608] Alison Frensham, Brian Cullis, and Arūnas Verbyla. Genotype by environment variance heterogeneity in a two-stage analysis. Biometrics, 53(4):1373–1383, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533504.

Marschner:1997:MAA

[1609] Ian C. Marschner. A method for assessing age-time disease incidence using serial prevalence data. Biometrics, 53(4):1384–1398, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533505.

Strawderman:1997:ACL

[1610] Robert L. Strawderman, Michael I. Parzen, and Martin T. Wells. Accurate confidence limits for quantiles under random censoring. Biometrics, 53(4):1399–1415, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533506.

Rayner:1997:HOA

[1611] J. C. W. Rayner and D. J. Best. How order affects the sign test. Biometrics, 53(4):1416–1421, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533507.

Nam:1997:EET

[1612] Jun mo Nam. Establishing equivalence of two treatments and sample size requirements in matched-pairs design. Biometrics, 53(4):1422–1430, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533508.

Wu:1997:MDS

[1613] Tiee-Jian Wu, John P. Burke, and Daniel B. Davison. A measure of DNA sequence dissimilarity based on Mahalanobis distance between frequencies of words. Biometrics, 53(4):1431–1439, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533509.

Carroll:1997:DAC

[1614] Raymond J. Carroll, Laurence Freedman, and David Pee. Design aspects of calibration studies in nutrition, with analysis of missing data in linear measurement error models. Biometrics, 53(4):1440–1457, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533510.

Xie:1997:GEE

[1615] Fang Xie and Myunghee Cho Paik. Generalized estimating equation model for binary outcomes with missing covariates. Biometrics, 53 (4):1458–1466, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533511.

Kodell:1997:AAT

[1616] Ralph L. Kodell and Hongshik Ahn. An age-adjusted trend test for the tumor incidence rate for multiple-sacrifice experiments. Biometrics, 53 (4):1467–1474, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533512.

Andersen:1997:EVC

[1617] Per Kragh Andersen, John P. Klein, Kim M. Knudsen, and René Tabanera y Palacios. Estimation of variance in Cox’s regression model with shared gamma frailties. Biometrics, 53(4):1475–1484, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533513.

Kooperberg:1997:HRI

[1618] Charles Kooperberg and Douglas B. Clarkson. Hazard regression with interval-censored data. Biometrics, 53(4):1485–1494, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533514.

Tsodikov:1997:DTS

[1619] Alexander D. Tsodikov, Bernard Asselain, and Andrej Y. Yakovlev. A distribution of tumor size at detection: an application to breast cancer data. Biometrics, 53(4):1495–1502, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533515.

Nickerson:1997:NCC

[1620] Carol A. E. Nickerson. A note on “A concordance correlation coefficient to evaluate reproducibility”. Biometrics, 53(4):1503–1507, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533516.

Liu:1997:SRS

[1621] Wei Liu. Some results on step-up tests for comparing treatments with a control in unbalanced one-way layouts. Biometrics, 53(4):1508–1512, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533517.

Qu:1997:MCP

[1622] Yinsheng Qu and Kirk A. Easley. A method for comparing positive rates of two blood culture systems. Biometrics, 53(4):1513–1519, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533518.

Padmanabhan:1997:RAW

[1623] A. R. Padmanabhan, V. M. Chinchilli, and Gutti Jogesh Babu. Robust analysis of within-unit variances in repeated measurement experiments. Biometrics, 53(4):1520–1526, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533519.

Gibbons:1997:REP

[1624] Robert D. Gibbons and Donald Hedeker. Random effects probit and logistic regression models for three-level data. Biometrics, 53(4):1527– 1537, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533520.

Xie:1997:MIM

[1625] Fang Xie and Myunghee Cho Paik. Multiple imputation methods for the missing covariates in generalized estimating equation. Biometrics, 53(4):1538–1546, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533521.

Douglas:1997:LAO

[1626] J. B. Douglas. Likelihood analyses of overdispersed Poisson models. Biometrics, 53(4):1547–1551, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533522.

Hines:1997:ACP

[1627] R. J. O’Hara Hines. Analysis of clustered polytomous data using generalized estimating equations and working covariance structures. Biometrics, 53(4):1552–1556, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533523.

Edwards:1997:FML

[1628] A. W. F. Edwards. Fisher memorial lectures. Biometrics, 53(4):1557– 1558, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533524.

Brown:1997:BRI

[1629] R. A. Brown. Book review: Intuitive Biostatistics, by H. Motulsky. Biometrics, 53(4):1559–1560, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533525.

Fearn:1997:BRB

[1630] T. Fearn. Book review: Bayesian Biostatistics, by D. A. Berry, D. K. Stangl. Biometrics, 53(4):1560, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533526.

Kemp:1997:BRR

[1631] A. W. Kemp. Book review: Randomization, Bootstrap and Monte Carlo Methods in Biology, by B. F. J. Manly. Biometrics, 53(4):1560–1561, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533527.

Papaioannou:1997:BRM

[1632] T. Papaioannou. Book review: Multiple Comparisons: Theory and Methods, by J. C. Hsu. Biometrics, 53(4):1561–1562, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533528.

Kemp:1997:BRM

[1633] C. D. Kemp. Book review: Math into LaTEX: an Introduction to LaTEX and AMS-LaTEX, by G. Grätzer. Biometrics, 53(4):1562–1563, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533529.

Anonymous:1997:BRLa

[1634] Anonymous. Book review: Leading Personalities in Statistical Sciences, by N. L. Johnson, S. Kotz. Biometrics, 53(4):1563, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533530.

Anonymous:1997:BRTb

[1635] Anonymous. Book review: Testing Statistical Hypotheses, by E. L. Lehmann. Biometrics, 53(4):1563, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533531.

Anonymous:1997:BRLb

[1636] Anonymous. Book review: Linear Models, by S. R. Searle. Biometrics, 53(4):1563–1564, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533532.

Anonymous:1997:BRD

[1637] Anonymous. Book review: The Design and Analysis of Sequential Clinical Trials, by J. Whitehead. Biometrics, 53(4):1564, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533535.

Anonymous:1997:BRMb

[1638] Anonymous. Book review: Modern Digital Simulation Methodology, II: Univariate and Bivariate Distribution Fitting, Bootstrap Methods, & Applications, by E. J. Dudewiczo. Biometrics, 53(4):1564, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533536.

Anonymous:1997:BRP

[1639] Anonymous. Book review: Plane Answers to Complex Questions: The Theory of Linear Models, by R. Christensen. Biometrics, 53(4):1564, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533533.

Anonymous:1997:BRRb

[1640] Anonymous. Book review: Recent Advances in Clinical Trial Design and Analysis, by P. H. Thall. Biometrics, 53(4):1564, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533534.

Anonymous:1997:BRW

[1641] Anonymous. Book reviews: World Health Statistics Annual, 1995, by World Health Organization; World Health Statistics Quarterly, Vol. 48, by World Health Organization; World Health Statistics Quarterly, Vol. 49, by World Health Organization. Biometrics, 53(4):1564–1565, December 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533537.

Anonymous:1997:C

[1642] Anonymous. Corrections. Biometrics, 53(4):1566, December 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533538. See [1571, 1554].

Anonymous:1997:CMG

[1643] Anonymous. Corrections: Modification of the Greenwood formula for correlated response times. Biometrics, 53(4):1566, December 1997. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533539. See [1554].

Anonymous:1998:FMa

[1644] Anonymous. Front matter. Biometrics, 54(1):i–iv, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533990.

Cullis:1998:SAM

[1645] Brian Cullis, Bev Gogel, Arūnas Verbyla, and Robin Thompson. Spatial analysis of multi-environment early generation variety trials. Biometrics, 54(1):1–18, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533991.

Higgins:1998:ESD

[1646] Karen M. Higgins, Marie Davidian, Ginger Chew, and Harriet Burge. The effect of serial dilution error on calibration inference in immunoassay. Biometrics, 54(1):19–32, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533992.

Catchpole:1998:IRR

[1647] E. A. Catchpole, S. N. Freeman, B. J. T. Morgan, and M. P. Harris. Integrated recovery/recapture data analysis. Biometrics, 54(1):33–46, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533993.

Plante:1998:SCR

[1648] Nathalie Plante, Louis-Paul Rivest, and Gilles Tremblay. Stratified capture–recapture estimation of the size of a closed population. Biometrics, 54(1):47–60, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533994.

Pledger:1998:TMN

[1649] Shirley Pledger and Leigh Bullen. Tests for mate and nest fidelity in birds with application to little blue penguins (eudyptula minor). Biometrics, 54(1):61–66, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533995.

See:1998:ARI

[1650] Kyoungah See and A. John Bailer. Added risk and inverse estimation for count responses in reproductive aquatic toxicology studies. Biometrics, 54(1):67–73, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533996.

Nettleton:1998:IMQ

[1651] Dan Nettleton and Jens Praestgaard. Interval mapping of quantitative trait loci through order-restricted inference. Biometrics, 54(1):74–87, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533997.

Mangin:1998:PQA

[1652] B. Mangin, P. Thoquet, and N. Grimsley. Pleiotropic QTL analysis. Biometrics, 54(1):88–99, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533998.

deGunst:1998:MPE

[1653] Mathisca C. M. de Gunst and E. Georg Luebeck. A method for parametric estimation of the number and size distribution of cell clusters from observations in a section plane. Biometrics, 54(1):100–112, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533999.

Belisle:1998:CPA

[1654] Patrick Bélisle, Lawrence Joseph, Brenda MacGibbon, David B. Wolfson, and Roxane du Berger. Change-point analysis of neuron spike train data. Biometrics, 54(1):113–123, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2534000.

Pepe:1998:TAR

[1655] Margaret Sullivan Pepe. Three approaches to regression analysis of receiver operating characteristic curves for continuous test results. Biometrics, 54(1):124–135, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534001.

Hanfelt:1998:IOR

[1656] John J. Hanfelt and Kung-Yee Liang. Inference for odds ratio regression models with sparse dependent data. Biometrics, 54(1):136–147, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2534002.

Lipsitz:1998:THR

[1657] Stuart R. Lipsitz, Keith B. G. Dear, Nan M. Laird, and Geert Molenberghs. Tests for homogeneity of the risk difference when data are sparse. Biometrics, 54(1):148–160, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2534003.

Crouchley:1998:IBG

[1658] R. Crouchley and A. Dassios. Interpreting the beta geometric in comparative fecundability studies. Biometrics, 54(1):161–167, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2534004.

Heimann:1998:PDL

[1659] G. Heimann and G. Neuhaus. Permutational distribution of the log-rank statistic under random censorship with applications to carcinogenicity assays. Biometrics, 54(1):168–184, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2534005.

Joly:1998:PLA

[1660] Pierre Joly, Daniel Commenges, and Luc Letenneur. A penalized likelihood approach for arbitrarily censored and truncated data: Application to age-specific incidence of dementia. Biometrics, 54(1):185–194, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2534006.

Fay:1998:CLR

[1661] Michael P. Fay, Barry I. Graubard, Laurence S. Freedman, and Douglas N. Midthune. Conditional logistic regression with sandwich estimators: Application to a meta-analysis. Biometrics, 54(1):195–208, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2534007.

Snapinn:1998:SAU

[1662] Steven M. Snapinn. Survival analysis with uncertain endpoints. Biometrics, 54(1):209–218, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534008.

Cheng:1998:PCI

[1663] S. C. Cheng, Jason P. Fine, and L. J. Wei. Prediction of cumulative incidence function under the proportional hazards model. Biometrics, 54(1): 219–228, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2534009.

Betensky:1998:MIE

[1664] Rebecca A. Betensky. Multiple imputation for early stopping of a complex clinical trial. Biometrics, 54(1):229–242, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2534010.

Wang:1998:ODS

[1665] You-Gan Wang and Denis Heng-Yan Leung. An optimal design for screening trials. Biometrics, 54(1):243–250, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2534011.

Thall:1998:SDF

[1666] Peter F. Thall and Kathy E. Russell. A strategy for dose-finding and safety monitoring based on efficacy and adverse outcomes in Phase I/II clinical trials. Biometrics, 54(1):251–264, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2534012.

Kleinman:1998:BFI

[1667] Ken P. Kleinman, Joseph G. Ibrahim, and Nan M. Laird. A Bayesian framework for intent-to-treat analysis with missing data. Biometrics, 54(1):265–278, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534013.

Stallard:1998:SSD

[1668] Nigel Stallard. Sample size determination for Phase II clinical trials based on Bayesian decision theory. Biometrics, 54(1):279–294, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2534014.

Lipsitz:1998:IUC

[1669] Stuart R. Lipsitz, Michael Parzen, and Marian Ewell. Inference using conditional logistic regression with missing covariates. Biometrics, 54(1): 295–303, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2534015.

Viennet:1998:PLE

[1670] Gabrielle Viennet, Frédéric Ménard, and Guy Thomas. Partial likelihood estimation in categorical time series with stochastic covariates. Biometrics, 54(1):304–311, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534016.

Hines:1998:CTC

[1671] R. J. O’Hara Hines. Comparison of two covariance structures in the analysis of clustered polytomous data using generalized estimating equations. Biometrics, 54(1):312–316, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2534017.

Olkin:1998:CMA

[1672] Ingram Olkin and Allan Sampson. Comparison of meta-analysis versus analysis of variance of individual patient data. Biometrics, 54(1):317– 322, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2534018.

Dawson:1998:SSC

[1673] Jeffrey D. Dawson. Sample size calculations based on slopes and other summary statistics. Biometrics, 54(1):323–330, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2534019.

Nam:1998:PSS

[1674] Jun mo Nam. Power and sample size for stratified prospective studies using the score method for testing relative risk. Biometrics, 54(1):331– 336, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2534020.

Bracci:1998:ACB

[1675] Paige M. Bracci, Shelley B. Bull, and Marc D. Grynpas. Analysis of compositional bone density data using log ratio transformations. Biometrics, 54(1):337–349, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534021.

Chao:1998:GSL

[1676] Edward C. Chao. Gibbs sampling for long-term survival data with competing risks. Biometrics, 54(1):350–366, March 1998. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2534022.

TenHave:1998:MEL

[1677] Thomas R. Ten Have, Allen R. Kunselman, Erik P. Pulkstenis, and J. Richard Landis. Mixed effects logistic regression models for longitudinal binary response data with informative drop-out. Biometrics, 54(1): 367–383, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2534023.

Zheng:1998:UUB

[1678] Qi Zheng. To use or not to use? Backward equations in stochastic carcinogenesis models. Biometrics, 54(1):384–388, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2534024.

Guiard:1998:RDE

[1679] Volker Guiard, Edward J. Bedrick, and Chih-Ling Tsai. Robust designs for the estimation of the ED50 using Fieller’s confidence interval. Biometrics, 54(1):389–390, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534025.

Sato:1998:MSM

[1680] Tosiya Sato, Edward J. Bedrick, and Chih-Ling Tsai. Model selection for multivariate regression in small samples. Biometrics, 54(1):391, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2534026.

Best:1998:BRG

[1681] J. Best. Book review: A Guide to Chi-Squared Testing, by P. E. Greenwood, M. S. Nikulin. Biometrics, 54(1):392–393, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2534027.

Murphy:1998:BRM

[1682] S. A. Murphy. Book review: Multivariate Dependencies: Models, Analysis and Interpretation, by D. R. Cox, N. Wermouth. Biometrics, 54(1): 393, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2534028.

Bowman:1998:BRK

[1683] A. W. Bowman. Book review: Kernel Smoothing, by M. P. Wand, M. C. Jones. Biometrics, 54(1):393–394, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2534029.

Campbell:1998:BRH

[1684] M. J. Campbell. Book review: Hidden Markov and other Models for Discrete-Valued Time Series, by I. L. Macdonald, W. Zucchini. Biometrics, 54(1):394–395, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2534030.

Xekelaki:1998:BRB

[1685] E. Xekelaki. Book review: Bonferroni-Type Inequalities with Applications, by J. Galambos, G. Simonelli. Biometrics, 54(1):395, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2534031.

Gentle:1998:BRE

[1686] J. E. Gentle. Book review: The EM Algorithm and Extensions, by G. J. McLachlan, T. Krishnan. Biometrics, 54(1):395–396, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2534032.

McLachlan:1998:BRP

[1687] G. J. McLachlan. Book review: Pattern Classification: a Unified View of Statistical and Neural Approaches, by J. Schürmann. Biometrics, 54(1): 396–397, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2534033.

Anonymous:1998:BRIa

[1688] Anonymous. Book review: Introductory Statistics for Biology Students, by T. A. Watt. Biometrics, 54(1):397, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2534034.

Anonymous:1998:BRMa

[1689] Anonymous. Book review: Modelling Frequency and Count Data, by J. K. Lindsey. Biometrics, 54(1):397, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2534035.

Anonymous:1998:BRC

[1690] Anonymous. Book review: A Course in Large Sample Theory, by T. S. Ferguson. Biometrics, 54(1):398, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2534036.

Anonymous:1998:BRMb

[1691] Anonymous. Book review: Morphometric Tools for Landmark Data, by F. Bookstein. Biometrics, 54(1):398, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2534038.

Anonymous:1998:BRSa

[1692] Anonymous. Book review: Statistics for the Environment 3: Polution Assessment and Control, by V. Barnett, K. F. Turkman. Biometrics, 54 (1):398, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2534037.

Anonymous:1998:BRIb

[1693] Anonymous. Book review: Introduction to Computational Biology: Maps, Sequences and Genomes, by M. S. Waterman. Biometrics, 54(1): 398–399, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2534039.

Anonymous:1998:CCT

[1694] Anonymous. Corrections: Comparing two diagnostic tests against the same “Gold Standard” in the same sample. Biometrics, 54(1):399, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2534042. See [1461].

Anonymous:1998:BRPa

[1695] Anonymous. Book review: Protect Your Privacy on the Internet, by B. Pfaffenberger. Biometrics, 54(1):399, March 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2534041.

Anonymous:1998:BRRa

[1696] Anonymous. Book review: Reliability: Probabilistic Models and Statistical Methods, by L. M. Leemis. Biometrics, 54(1):399, March 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2534040.

Anonymous:1998:FMb

[1697] Anonymous. Front matter. Biometrics, 54(2):i–iv, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109750.

Diggle:1998:NEC

[1698] Peter J. Diggle and Arūnas P. Verbyla. Nonparametric estimation of covariance structure in longitudinal data. Biometrics, 54(2):401–415, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3109751.

Jensen:1998:PDN

[1699] C. S. Jensen and N. Sheehan. Problems with determination of noncommunicating classes for Monte Carlo Markov chain applications in pedigree analysis. Biometrics, 54(2):416–425, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3109752.

McKnight:1998:LBI

[1700] Barbara McKnight, Camlin Tierney, Susan P. McGorray, and Nicholas E. Day. Likelihood-based inference for the genetic relative risk based on affected-sibling-pair marker data. Biometrics, 54(2):426–443, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109753.

Leisenring:1998:RMD

[1701] Wendy Leisenring and Margaret Sullivan Pepe. Regression modelling of diagnostic likelihood ratios for the evaluation of medical diagnostic tests. Biometrics, 54(2):444–452, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3109754.

Zhou:1998:CCA

[1702] Xiao-Hua Zhou. Comparing correlated areas under the ROC curves of two diagnostic tests in the presence of verification bias. Biometrics, 54(2): 453–470, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3109755.

Federer:1998:RII

[1703] Walter T. Federer. Recovery of interblock, intergradient, and intervariety information in incomplete block and lattice rectangle designed experiments. Biometrics, 54(2):471–481, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3109756.

Bloch:1998:MMR

[1704] Juliette Bloch and Michel Chavance. A mixed model for repeated dilution assays. Biometrics, 54(2):482–492, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3109757.

DiRienzo:1998:TSA

[1705] A. Gregory DiRienzo, Igor G. Zurbenko, and David O. Carpenter. Time series analysis of aplysia total motion activity. Biometrics, 54(2):493– 508, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3109758.

Huffer:1998:MCM

[1706] Fred W. Huffer and Hulin Wu. Markov chain Monte Carlo for autologistic regression models with application to the distribution of plant species. Biometrics, 54(2):509–524, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3109759.

Bohning:1998:RDC

[1707] Dankmar Böhning, Ekkehart Dietz, and Peter Schlattmann. Recent developments in computer-assisted analysis of mixtures. Biometrics, 54(2): 525–536, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3109760.

StLaurent:1998:EAGa

[1708] Roy T. St. Laurent. Evaluating agreement with a gold standard in method comparison studies. Biometrics, 54(2):537–545, June 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109761.

StLaurent:1998:EAGb

[1709] Roy T. St. Laurent. Evaluating agreement with a gold standard in method comparison studies. Biometrics, 54(2):537–545, June 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:1998:DDU

[1710] Chang Yu, Lance A. Waller, and Daniel Zelterman. Discrete distributions for use in twin studies. Biometrics, 54(2):546–557, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109762.

Dunson:1998:DDN

[1711] David B. Dunson. Dose-dependent number of implants and implications in developmental toxicity. Biometrics, 54(2):558–569, June 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109763.

Lesaffre:1998:LIL

[1712] Emmanuel Lesaffre and Geert Verbeke. Local influence in linear mixed models. Biometrics, 54(2):570–582, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3109764.

Huggins:1998:AML

[1713] R. M. Huggins and D. Z. Loesch. On the analysis of mixed longitudinal growth data. Biometrics, 54(2):583–595, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3109765.

Navidi:1998:BCC

[1714] William Navidi. Bidirectional case-crossover designs for exposures with time trends. Biometrics, 54(2):596–605, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3109766.

Mack:1998:KML

[1715] Y. P. Mack and Pham X. Quang. Kernel methods in line and point transect sampling. Biometrics, 54(2):606–619, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3109767.

Evans:1998:TVS

[1716] Geoffrey T. Evans and John M. Hoenig. Testing and viewing symmetry in contingency tables, with application to readers of fish ages. Biometrics, 54(2):620–629, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3109768.

Loughin:1998:TAC

[1717] Thomas M. Loughin and Peter N. Scherer. Testing for association in contingency tables with multiple column responses. Biometrics, 54(2):

630–637, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3109769.

Neuhaus:1998:BWC

[1718] J. M. Neuhaus and J. D. Kalbfleisch. Between- and within-cluster covariate effects in the analysis of clustered data. Biometrics, 54(2):638–645, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3109770.

Petersen:1998:AFM

[1719] Jørgen Holm Petersen. An additive frailty model for correlated life times. Biometrics, 54(2):646–661, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3109771.

Walker:1998:NAS

[1720] Stephen Walker. A nonparametric approach to a survival study with surrogate endpoints. Biometrics, 54(2):662–672, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109772.

Hommel:1998:IMT

[1721] Gerhard Hommel and Frank Krummenauer. Improvements and modifications of Tarone’s multiple test procedure for discrete data. Biometrics, 54(2):673–681, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3109773.

Tan:1998:CTD

[1722] Ming Tan, Xiaoping Xiong, and Michael H. Kutner. Clinical trial designs based on sequential conditional probability ratio tests and reverse stochastic curtailing. Biometrics, 54(2):682–695, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109774.

Wassmer:1998:CTM

[1723] Gernot Wassmer. A comparison of two methods for adaptive interim analyses in clinical trials. Biometrics, 54(2):696–705, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109775.

Lui:1998:IER

[1724] Kung-Jong Lui. Interval estimation of the risk ratio between a secondary infection, given a primary infection, and the primary infection. Biometrics, 54(2):706–711, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3109776.

Heisey:1998:MTD

[1725] Dennis M. Heisey and Annie P. Foong. Modelling time-dependent interaction in a time-varying covariate and its application to rejection episodes and kidney transplant failure. Biometrics, 54(2):712–719, June 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109777.

Barnhart:1998:GFT

[1726] Huiman X. Barnhart and John M. Williamson. Goodness-of-fit tests for GEE modeling with binary responses. Biometrics, 54(2):720–729, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3109778.

Becker:1998:ETR

[1727] Niels G. Becker and A. M. Hasofer. Estimating the transmission rate for a highly infectious disease. Biometrics, 54(2):730–738, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109779.

Giesbrecht:1998:IPA

[1728] F. G. Giesbrecht and T. B. Whitaker. Investigations of the problems of assessing aflatoxin levels in peanuts. Biometrics, 54(2):739–753, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3109780.

Murtaugh:1998:MCF

[1729] Paul A. Murtaugh and DeWayne R. Derryberry. Models of connectance in food webs. Biometrics, 54(2):754–761, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3109781.

Weller:1998:TTC

[1730] Edie A. Weller and Louise M. Ryan. Testing for trend with count data. Biometrics, 54(2):762–773, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3109782.

Stewart:1998:ACI

[1731] Susan L. Stewart, Karen C. Swallen, Sally L. Glaser, Pamela L. Horn-Ross, and Dee W. West. Adjustment of cancer incidence rates for ethnic misclassification. Biometrics, 54(2):774–781, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109783.

Shih:1998:STT

[1732] Weichung Joseph Shih and Hui Quan. Stratified testing for treatment effects with missing data. Biometrics, 54(2):782–787, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109784.

Hutchinson:1998:DRN

[1733] T. P. Hutchinson and D. Cairns. Difference or ratio? A note on twotreatment, four-sequence analysis. Biometrics, 54(2):788–789, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109785.

Wittkowski:1998:VST

[1734] Knut M. Wittkowski, Clint W. Coakley, and Mark A. Heise. Versions of the sign test in the presence of ties. Biometrics, 54(2):789–791, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3109786.

Jolliffe:1998:MMC

[1735] Ian T. Jolliffe and Anna Jolliffe. Modelling memory in coal tits: an illustration of the EM algorithm. Biometrics, 54(2):792–793, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109787.

Balding:1998:BRE

[1736] D. J. Balding. Book review: Encyclopedia of Statistical Sciences: Update Volume 1, by S. Kotz, C. B. Read, D. L. Banks. Biometrics, 54(2):794– 795, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3109788.

Buckland:1998:BRB

[1737] S. T. Buckland. Book review: Bootstrap Methods and Their Application, by A. C. Davison, D. V. Hinkley. Biometrics, 54(2):795, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109789.

Papageorgiou:1998:BRD

[1738] H. Papageorgiou. Book review: Discrete Multivariate Distributions, by N. L. Johnson, S. Kotz, N. Balakrishnan. Biometrics, 54(2):795–796, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3109790.

Kemp:1998:BRP

[1739] A. W. Kemp. Book review: Probability and Statistical Inference, by R. Bartoszyński, M. Niewiadomska-Bugaj. Biometrics, 54(2):796, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3109791.

Anonymous:1998:BRAa

[1740] Anonymous. Book review: Advances in the Theory and Practice of Statistics: a Volume in Honor of Samuel Kotz, by N. L. Johnson, N. Balakrishnan. Biometrics, 54(2):797, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3109793.

Anonymous:1998:BRBa

[1741] Anonymous. Book review: Breakthroughs in Statistics, Vol. III, by S. Kotz, N. L. Johnson. Biometrics, 54(2):797, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109792.

Anonymous:1998:BRSb

[1742] Anonymous. Book review: Series Approximation in Statistics, 2nd Edition, by J. E. Kolassa. Biometrics, 54(2):797, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109794.

Anonymous:1998:BRT

[1743] Anonymous. Book review: Tools for Statistical Inference: Methods for the Exploration of Posterior Distributions and Likelihood Functions, 3rd Edition, by M. A. Tanner. Biometrics, 54(2):797, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3109795.

Anonymous:1998:BRAb

[1744] Anonymous. Book review: Applied Survival Analysis, by C. T. Le. Biometrics, 54(2):798, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3109796.

Anonymous:1998:BRE

[1745] Anonymous. Book review: Elements of Multivariate Time Series Analysis, 2nd Edition, by G. C. Reinsel. Biometrics, 54(2):798, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109798.

Anonymous:1998:BRMc

[1746] Anonymous. Book review: MSI-2000: Multivariate Statistical Analysis in Honor of Professor Minoru Siotani on His 70th Birthday, Vol 3, by T. Hayakawa, M. Aoshima, K. Shimizu. Biometrics, 54(2):798, June 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3109797.

Anonymous:1998:FMc

[1747] Anonymous. Front matter. Biometrics, 54(3):i–iv, September 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533834.

Gilbert:1998:SMA

[1748] Peter B. Gilbert, Steven G. Self, and Mark A. Ashby. Statistical methods for assessing differential vaccine protection against human immunodeficiency virus types. Biometrics, 54(3):799–814, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533835.

Chen:1998:CCR

[1749] Chiu-Lan Chen, Kenneth H. Pollock, and John M. Hoenig. Combining change-in-ratio, index-removal, and removal models for estimating population size. Biometrics, 54(3):815–827, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533836.

Cowling:1998:SML

[1750] Ann Cowling. Spatial methods for line transect surveys. Biometrics, 54(3):828–839, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533837.

Engen:1998:DES

[1751] Steinar Engen, Øyvind Bakke, and Aminul Islam. Demographic and environmental stochasticity-concepts and definitions. Biometrics, 54(3):840– 846, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533838.

Cannon:1998:SCM

[1752] R. M. Cannon. Sampling to comply with a maximum pest limit. Biometrics, 54(3):847–858, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533839.

Xu:1998:EDF

[1753] Jian-Lun Xu and Philip C. Prorok. Estimating a distribution function of the tumor size at metastasis. Biometrics, 54(3):859–864, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533840.

Rivest:1998:SME

[1754] Louis-Paul Rivest, Serge Couturier, and Hélène Crépeau. Statistical methods for estimating caribou abundance using postcalving aggregations detected by radio telemetry. Biometrics, 54(3):865–876, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533841.

Billard:1998:APS

[1755] L. Billard, P. W. A. Dayananda, and S. Elmes. Assessment of plant stress due to the presence of disease. Biometrics, 54(3):877–887, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533842.

Pledger:1998:CBD

[1756] Shirley Pledger and Murray Efford. Correction of bias due to heterogeneous capture probability in capture–recapture studies of open populations. Biometrics, 54(3):888–898, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533843.

Chen:1998:MEL

[1757] Song Xi Chen. Measurement errors in line transect surveys. Biometrics, 54(3):899–908, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533844.

FitzGerald:1998:IRL

[1758] Patrick E. B. FitzGerald and Matthew W. Knuiman. Interpretation of regressive logistic regression coefficients in analyses of familial data. Biometrics, 54(3):909–920, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533845.

Kleinman:1998:SBA

[1759] Ken P. Kleinman and Joseph G. Ibrahim. A semiparametric Bayesian approach to the random effects model. Biometrics, 54(3):921–938, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533846.

Chen:1998:MFR

[1760] Shande Chen, Christopher Cox, and Lu Cui. A more flexible regressionto-the-mean model with possible stratification. Biometrics, 54(3):939– 947, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533847.

Lin:1998:ASR

[1761] D. Y. Lin, B. M. Psaty, and R. A. Kronmal. Assessing the sensitivity of regression results to unmeasured confounders in observational studies. Biometrics, 54(3):948–963, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533848.

Carlin:1998:AOS

[1762] Bradley P. Carlin, Joseph B. Kadane, and Alan E. Gelfand. Approaches for optimal sequential decision analysis in clinical trials. Biometrics, 54(3):964–975, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533849.

Gray:1998:ETE

[1763] Sarah M. Gray and Ron Brookmeyer. Estimating a treatment effect from multidimensional longitudinal data. Biometrics, 54(3):976–988, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533850.

Martinussen:1998:ADU

[1764] Torben Martinussen and Thorkild I. A. Sørensen. Age-dependent Ushaped risk functions and Aalen’s additive risk model. Biometrics, 54 (3):989–1001, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533851.

Lipsitz:1998:EEI

[1765] Stuart R. Lipsitz and Joseph G. Ibrahim. Estimating equations with incomplete categorical covariates in the Cox model. Biometrics, 54 (3):1002–1013, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533852.

Buyse:1998:CVS

[1766] Marc Buyse and Geert Molenberghs. Criteria for the validation of surrogate endpoints in randomized experiments. Biometrics, 54(3):1014–1029, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533853.

Li:1998:AAO

[1767] Hongzhe Li, Ping Yang, and Ann G. Schwartz. Analysis of age of onset data from case-control family studies. Biometrics, 54(3):1030–1039, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533854.

Yang:1998:SSE

[1768] Song Yang. Some scale estimators and lack-of-fit tests for the censored two-sample accelerated life model. Biometrics, 54(3):1040–1052, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533855.

Pan:1998:ESC

[1769] Wei Pan and R. Chappell. Estimating survival curves with left-truncated and interval-censored data under monotone hazards. Biometrics, 54 (3):1053–1060, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533856.

Betensky:1998:CCS

[1770] Rebecca A. Betensky. Construction of a continuous stopping boundary from an alpha spending function. Biometrics, 54(3):1061–1071, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533857.

Sooriyarachchi:1998:MSA

[1771] M. R. Sooriyarachchi and John Whitehead. A method for sequential analysis of survival data with nonproportional hazards. Biometrics, 54 (3):1072–1084, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533858.

Elston:1998:EDD

[1772] D. A. Elston. Estimation of denominator degrees of freedom of fdistributions for assessing Wald statistics for fixed-effect factors in unbalanced mixed models. Biometrics, 54(3):1085–1096, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533859.

Garcia-Dorado:1998:MDE

[1773] A. García-Dorado and J. M. Marín. Minimum distance estimation of mutational parameters for quantitative traits. Biometrics, 54(3):1097– 1114, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533860.

Shih:1998:MMD

[1774] Joanna H. Shih. Modeling multivariate discrete failure time data. Biometrics, 54(3):1115–1128, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533861.

Baumgartner:1998:NTG

[1775] W. Baumgartner, P. Weiß, and H. Schindler. A nonparametric test for the general two-sample problem. Biometrics, 54(3):1129–1135, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533862.

Minami:1998:ECC

[1776] Mihoko Minami and Kunio Shimizu. Estimation for a common correlation coefficient in bivariate normal distributions with missing observations. Biometrics, 54(3):1136–1146, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533863.

Liu:1998:ODS

[1777] Qing Liu. An order-directed score test for trend in ordered 2 × K tables. Biometrics, 54(3):1147–1154, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533864.

Engel:1998:BRA

[1778] Bas Engel and Willem Buist. Bias reduction of approximate maximum likelihood estimates for heritability in threshold models. Biometrics, 54 (3):1155–1164, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533865.

Leon:1998:AMA

[1779] Larry F. Léon and Chih-Ling Tsai. Assessment of model adequacy for Markov regression time series models. Biometrics, 54(3):1165–1175, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533866.

Lee:1998:STI

[1780] Seung-Yeoun Lee and Robert A. Wolfe. A simple test for independent censoring under the proportional hazards model. Biometrics, 54(3):1176– 1182, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533867.

Yao:1998:OTS

[1781] Tzy-Jyun Yao and E. S. Venkatraman. Optimal two-stage design for a series of pilot trials of new agents. Biometrics, 54(3):1183–1189, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533868.

Bross:1998:LE

[1782] Irwin D. Bross, Catherine Hill, and Agnès Laplanche. Letter to the editor. Biometrics, 54(3):1190–1193, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533869.

Midthune:1998:TLR

[1783] Douglas N. Midthune, Edward L. Korn, Barry I. Graubard, Simon Barry, and Terry O’Neill. Truncated logistic regression and residual intracluster correlation. Biometrics, 54(3):1193–1196, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533870.

Kemp:1998:BRE

[1784] C. D. Kemp. Book review: Encyclopedia of Biostatistics, by P. Armitage, T. Coulton. Biometrics, 54(3):1197–1199, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533871. See correction [1858].

Gordon:1998:BRE

[1785] A. D. Gordon. Book review: Encyclopedia of Statistical Sciences: Update Volume 2, by S. Kotz, C. B. Read, D. L. Banks. Biometrics, 54(3):1199, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533872.

Edwards:1998:BRE

[1786] A. W. F. Edwards. Book review: Error and the Growth of Experimental Knowledge, by D. G. Mayo. Biometrics, 54(3):1199–1200, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533873.

Kemp:1998:BRH

[1787] A. W. Kemp. Book review: A History of Mathematical Statistics from 1750 to 1930, by A. Hald. Biometrics, 54(3):1200–1201, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533874.

Ringrose:1998:BRM

[1788] T. Ringrose. Book review: Multivariate Models and Dependence Concepts, by H. Joe. Biometrics, 54(3):1201–1202, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533875.

Leonard:1998:BRC

[1789] T. Leonard. Book review: Case Studies in Bayesian Statistics, by C. Gatsonis, J. S. Hodges, R. E. Kass, R. McCulloch, P. Rossi, N. D. Singpurwalla. Biometrics, 54(3):1202, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533876.

Anonymous:1998:BRSc

[1790] Anonymous. Book review: Statistics and Public Policy, by B. D. Spencer. Biometrics, 54(3):1202–1203, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533877.

Anonymous:1998:BRAc

[1791] Anonymous. Book review: Applied Multivariate Statistical Analysis, by R. A. Johnson, D. W. Wichern. Biometrics, 54(3):1203, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533879.

Anonymous:1998:BRIc

[1792] Anonymous. Book review: Intelligence, Genes, and Success: Scientists Respond to The Bell Curve, by B. Devlin, S. E. Fienberg, D. P. Resnick, K. Roeder. Biometrics, 54(3):1203, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533878.

Anonymous:1998:BRPb

[1793] Anonymous. Book review: Probability without Equations: Concepts for Clinicians, by B. K. Holland. Biometrics, 54(3):1203–1204, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533880.

Anonymous:1998:BRAd

[1794] Anonymous. Book review: Angewandte Statistik: Anwendung statistischer Methoden, 8. Auflage, by L. Sachs. Biometrics, 54(3):1204, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533881.

Anonymous:1998:BRId

[1795] Anonymous. Book review: Introduction to Time Series and Forecasting, by P. J. Brockwell, R. A. Davis. Biometrics, 54(3):1204, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533882.

Anonymous:1998:BRRb

[1796] Anonymous. Book reviews: Random Walks and Random Environments, Volume 1: Random Walks, by B. D. Hughes; Random Walks and Random Environments, Volume 2: Random Environments, by B. D. Hughes. Biometrics, 54(3):1204–1205, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533883.

Anonymous:1998:CRE

[1797] Anonymous. Correction: Regression estimator in ranked set sampling. Biometrics, 54(3):1205, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533886. See [1568].

Anonymous:1998:BRMd

[1798] Anonymous. Book review: Matrix Algebra From a Statistician’s Perspective, by D. A. Harville. Biometrics, 54(3):1205, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533884.

Anonymous:1998:BRSd

[1799] Anonymous. Book review: Statistical Thinking for Managers, by D. H. Hildebrand, R. L. Ott. Biometrics, 54(3):1205, September 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533885.

Anonymous:1998:FMd

[1800] Anonymous. Front matter. Biometrics, 54(4):i–iv, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533650.

Anonymous:1998:VI

[1801] Anonymous. Volume information. Biometrics, 54(4):i–xvi, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533649.

Borchers:1998:MRM

[1802] David L. Borchers, Walter Zucchini, and Rachel M. Fewster. Markrecapture models for line transect surveys. Biometrics, 54(4):1207–1220, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533651.

Borchers:1998:HTE

[1803] David L. Borchers, Stephen T. Buckland, Paul W. Goedhart, Elizabeth D. Clarke, and Sharon L. Hedley. Horvitz-Thompson estimators for double-platform line transect surveys. Biometrics, 54(4):1221–1237, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533652.

Hauser:1998:GLA

[1804] Elizabeth R. Hauser and Michael Boehnke. Genetic linkage analysis of complex genetic traits by using affected sibling pairs. Biometrics, 54 (4):1238–1246, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533653.

Teng:1998:MLAa

[1805] Jun Teng and David Siegmund. Multipoint linkage analysis using affected relative pairs and partially informative markers. Biometrics, 54(4):1247– 1265, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533654. See

discussion [1806] and rejoinder [1808].

Olson:1998:DTS

[1806] Jane M. Olson. Discussion of Teng and Siegmund article. Biometrics, 54(4):1266–1270, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533655. See [1805] and rejoinder [1808].

Nicolae:1998:MLA

[1807] Dan L. Nicolae, Michael L. Frigge, Nancy J. Cox, and Augustine Kong. [Multipoint linkage analysis using affected relative pairs and partially informative markers]: Discussion. Biometrics, 54(4):1271–1274, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533656.

Teng:1998:MLAb

[1808] Jun Teng and David Siegmund. [Multipoint linkage analysis using affected relative pairs and partially informative markers]: Rejoinder. Biometrics, 54(4):1275–1279, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533657. See [1805, 1806].

Hiby:1998:UAT

[1809] Lex Hiby and Phil Lovell. Using aircraft in tandem formation to estimate abundance of harbour porpoise. Biometrics, 54(4):1280–1289, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533658.

Newman:1998:SSM

[1810] Ken B. Newman. State-space modeling of animal movement and mortality with application to salmon. Biometrics, 54(4):1290–1314, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533659.

vanEeuwijk:1998:MMI

[1811] Fred A. van Eeuwijk and Pieter M. Kroonenberg. Multiplicative models for interaction in three-way ANOVA, with applications to plant breeding. Biometrics, 54(4):1315–1333, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533660.

Stephens:1998:BAQ

[1812] D. A. Stephens and R. D. Fisch. Bayesian analysis of quantitative trait locus data using reversible jump Markov chain Monte Carlo. Biometrics, 54(4):1334–1347, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533661.

Chen:1998:HEH

[1813] Ming-Hui Chen, Amita K. Manatunga, and Christopher J. Williams. Heritability estimates from human twin data by incorporating historical prior information. Biometrics, 54(4):1348–1362, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533662.

Young:1998:TEI

[1814] David James Young and Madhusudan Bhandary. Test for equality of intraclass correlation coefficients under unequal family sizes. Biometrics, 54(4):1363–1373, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533663.

Azais:1998:IDV

[1815] J.-M. Azaïs, H. Monod, and R. A. Bailey. The influence of design on validity and efficiency of neighbour methods. Biometrics, 54(4):1374– 1387, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533664.

Huang:1998:SCM

[1816] Yu-Chuang Huang and Weng-Kee Wong. Sequential construction of multiple-objective optimal designs. Biometrics, 54(4):1388–1397, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533665.

Mehrabi:1998:IBD

[1817] Yadollah Mehrabi and J. N. S. Matthews. Implementable Bayesian designs for limiting dilution assays. Biometrics, 54(4):1398–1406, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533666.

Hu:1998:EPC

[1818] Ping Hu, Anastasios A. Tsiatis, and Marie Davidian. Estimating the parameters in the Cox model when covariate variables are measured with error. Biometrics, 54(4):1407–1419, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533667.

Huang:1998:PES

[1819] Xin Huang, Shande Chen, and Seng jaw Soong. Piecewise exponential survival trees with time-dependent covariates. Biometrics, 54(4):1420– 1433, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533668.

Kim:1998:NRE

[1820] Haesook T. Kim and Young K. Truong. Nonparametric regression estimates with censored data: Local linear smoothers and their applications. Biometrics, 54(4):1434–1444, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533669.

Dafni:1998:ESM

[1821] Urania G. Dafni and Anastasios A. Tsiatis. Evaluating surrogate markers of clinical outcome when measured with error. Biometrics, 54(4):1445– 1462, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533670.

Sinha:1998:PLM

[1822] Debajyoti Sinha. Posterior likelihood methods for multivariate survival data. Biometrics, 54(4):1463–1474, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533671.

Faraggi:1998:BVS

[1823] David Faraggi and Richard Simon. Bayesian variable selection method for censored survival data. Biometrics, 54(4):1475–1485, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533672.

Sargent:1998:GFR

[1824] Daniel J. Sargent. A general framework for random effects survival analysis in the Cox proportional hazards setting. Biometrics, 54(4):1486–1497, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533673.

Goggins:1998:MCM

[1825] William B. Goggins, Dianne M. Finkelstein, David A. Schoenfeld, and Alan M. Zaslavsky. A Markov chain Monte Carlo EM algorithm for analyzing interval-censored data under the Cox proportional hazards model. Biometrics, 54(4):1498–1507, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533674.

Tsodikov:1998:PHM

[1826] Alexander Tsodikov. A proportional hazards model taking account of long-term survivors. Biometrics, 54(4):1508–1516, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533675.

Verweij:1998:GFT

[1827] Pierre J. M. Verweij, Hans C. van Houwelingen, and Theo Stijnen. A goodness-of-fit test for Cox’s proportional hazards model based on martingale residuals. Biometrics, 54(4):1517–1526, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533676.

Anderson:1998:ADR

[1828] Aparna B. Anderson and Ross L. Prentice. On the accommodation of disease rate correlations in aggregate data studies of disease risk factors. Biometrics, 54(4):1527–1540, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533677.

Berger:1998:CHT

[1829] Vance W. Berger, Thomas Permutt, and Anastasia Ivanova. Convex hull test for ordered categorical data. Biometrics, 54(4):1541–1550, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533678.

Flyer:1998:CCU

[1830] Paul A. Flyer. A comparison of conditional and unconditional randomization tests for highly stratified designs. Biometrics, 54(4):1551–1559, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533679.

Morrell:1998:LRT

[1831] Christopher H. Morrell. Likelihood ratio testing of variance components in the linear mixed-effects model using restricted maximum likelihood.

Biometrics, 54(4):1560–1568, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533680.

Baker:1998:EAB

[1832] Stuart G. Baker. Evaluating the age to begin periodic breast cancer screening using data from a few regularly scheduled screenings. Biometrics, 54(4):1569–1578, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533681.

Park:1998:ACD

[1833] Taesung Park. An approach to categorical data with nonignorable nonresponse. Biometrics, 54(4):1579–1590, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533682.

Causeur:1998:FSP

[1834] David J. Causeur and Thierry J. Dhorne. Finite sample properties of a multivariate extension of double regression. Biometrics, 54(4):1591–1601, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533683.

Falk:1998:TSD

[1835] Raymond W. Falk and Gary G. Koch. Testing a specified difference between proportions. Biometrics, 54(4):1602–1614, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533684.

Mehta:1998:EPS

[1836] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Exact power and sample-size computations for the cochran-armitage trend test. Biometrics, 54(4):1615–1621, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533685.

Shults:1998:ASC

[1837] Justine Shults and N. Rao Chaganty. Analysis of serially correlated data using quasi-least squares. Biometrics, 54(4):1622–1630, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533686.

Xue:1998:MSD

[1838] Xiaonan Xue. Multivariate survival data under bivariate frailty: an estimating equation approach. Biometrics, 54(4):1631–1637, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533687.

Lee:1998:ACP

[1839] Eric W. Lee and Mimi Y. Kim. The analysis of correlated panel data using a continuous-time Markov model. Biometrics, 54(4):1638–1644, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533688.

Park:1998:GEE

[1840] Taesung Park, Dong Wan Shin, and Chul Gyu Park. A generalized estimating equations approach for testing ordered group effects with repeated measurements. Biometrics, 54(4):1645–1653, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533689.

Variyam:1998:RRM

[1841] Jayachandran N. Variyam, James Blaylock, and David Smallwood. Reduced-rank models for nutrition knowledge assessment. Biometrics, 54(4):1654–1661, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533690.

Lunn:1998:AKS

[1842] Mary Lunn. Applying k-sample tests to conditional probabilities for competing risks in a clinical trial. Biometrics, 54(4):1662–1672, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533691.

McShane:1998:SSR

[1843] Phil McShane, Sylvan Wallenstein, and Agnes Berger. Small sample results with weighted logrank tests. Biometrics, 54(4):1673–1674, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533692.

Tomassone:1998:BRL

[1844] R. Tomassone. Book review: L’expérience et le modele: Un discours sur la methode, by J.-M. Legay. Biometrics, 54(4):1675–1676, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533693.

Leonard:1998:BRB

[1845] T. Leonard. Book review: Bayesian Statistics 5, by J. M. Bernado, J. O. Berger, A. P. Dawid, A. F. M. Smith. Biometrics, 54(4):1676–1677, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533694.

Talbot:1998:BRP

[1846] M. Talbot. Book review: Practical Data Analysis for Designed Experiments, by B. S. Yandell. Biometrics, 54(4):1678, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533695.

Wood:1998:BRD

[1847] S. N. Wood. Book review: Data Driven Statistical Methods, by P. Sprent. Biometrics, 54(4):1678–1679, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533696.

Hinkelmann:1998:BRK

[1848] K. Hinkelmann. Book review: Kurzgefasste Statistik für die klinische Forschung: Ein Praktischer Leitfaden für die Analyse kleiner Stichproben, by J. Bortz, G. A. Lienert. Biometrics, 54(4):1679–1680, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533697.

Hand:1998:BRH

[1849] D. J. Hand. Book review: Handbook of Modern Item Response Theory, by W. J. van der Linden, R. K. Hambleton. Biometrics, 54(4):1680, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533698.

Kemp:1998:BRV

[1850] A. W. Kemp. Book review: Visual Explanations: Images and Quantities, Evidence and Narrative, by E. R. Tufte. Biometrics, 54(4):1680–1681, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533699.

Rawlings:1998:BRA

[1851] D. P. Rawlings. Book review: Advances in Combinatorial Methods and Applications to Probability and Statistics, by N. Balakrishnan. Biometrics, 54(4):1681–1682, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533700.

Anonymous:1998:BRBb

[1852] Anonymous. Book review: Biostatistics: a Manual of Statistical Methods for Use in Health, Nutrition and Anthropology, by K. V. Rao. Biometrics, 54(4):1682, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533701.

Anonymous:1998:BRRc

[1853] Anonymous. Book review: Robust Statistical Procedures: Asymptotics and Interrelations, by J. Jurečková, P. K. Sen. Biometrics, 54(4):1682, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533702.

Anonymous:1998:BRMe

[1854] Anonymous. Book review: Multiple Tests und Auswahlverfahren, by M. Horn, R. Vollandt. Biometrics, 54(4):1682–1683, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533703.

Anonymous:1998:BRH

[1855] Anonymous. Book review: Handbook of the Normal Distribution, 2nd Edition, by J. K. Patel, C. B. Read. Biometrics, 54(4):1683, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533704.

Anonymous:1998:BRPc

[1856] Anonymous. Book review: Population Genetics: a Concise Guide, by J. Gillespie. Biometrics, 54(4):1683, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533705.

Anonymous:1998:BRMf

[1857] Anonymous. Book review: Modern Digital Simulation Methodology, III, by E. J. Dudewicz. Biometrics, 54(4):1683–1684, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533706.

Anonymous:1998:CEB

[1858] Anonymous. Correction: Encyclopedia of Biostatistics. Biometrics, 54(4):1684, December 1998. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533708. See [1784].

Anonymous:1998:BRMg

[1859] Anonymous. Book review: Mathematical Models in the Applied Sciences, by A. C. Fowler. Biometrics, 54(4):1684, December 1998. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533707.

Anonymous:1999:FMa

[1860] Anonymous. Front matter. Biometrics, 55(1):i–iv, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533888.

Mau:1999:BPI

[1861] Bob Mau, Michael A. Newton, and Bret Larget. Bayesian phylogenetic inference via Markov chain Monte Carlo methods. Biometrics, 55(1):1– 12, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533889.

Hougaard:1999:FSD

[1862] Philip Hougaard. Fundamentals of survival data. Biometrics, 55(1):13– 22, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533890.

Keiding:1999:GVL

[1863] Niels Keiding, Marusca Filiberti, Sille Esbjerg, James M. Robins, and Niels Jacobsen. The graft versus leukemia effect after bone marrow transplantation: a case study using structural nested failure time models. Biometrics, 55(1):23–28, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533891.

Skaug:1999:HML

[1864] Hans J. Skaug and Tore Schweder. Hazard models for line transect surveys with independent observers. Biometrics, 55(1):29–36, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533892.

Robin:1999:MT

[1865] Stéphane Robin. A model for thermograms. Biometrics, 55(1):37–43, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533893.

Carroll:1999:FPM

[1866] Raymond J. Carroll, Kathryn Roeder, and Larry Wasserman. Flexible parametric measurement error models. Biometrics, 55(1):44–54, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533894.

Darling:1999:GMR

[1867] R. W. R. Darling and Tim Holt. Genetic models with reduced penetrance related to the Y chromosome. Biometrics, 55(1):55–64, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533895.

Liang:1999:HTU

[1868] Kung-Yee Liang and Paul J. Rathouz. Hypothesis testing under mixture models: Application to genetic linkage analysis. Biometrics, 55(1):65– 74, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533896.

Wu:1999:USM

[1869] Margaret C. Wu and Dean A. Follmann. Use of summary measures to adjust for informative missingness in repeated measures data with random effects. Biometrics, 55(1):75–84, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533897.

TenHave:1999:MEMa

[1870] Thomas R. Ten Have and Alfredo Morabia. Mixed effects models with bivariate and univariate association parameters for longitudinal bivariate binary response data. Biometrics, 55(1):85–93, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533898.

TenHave:1999:MEMb

[1871] Thomas R. Ten Have and Alfredo Morabia. Mixed effects models with bivariate and univariate association parameters for longitudinal bivariate binary response data. Biometrics, 55(1):85–93, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Golm:1999:SMM

[1872] Gregory T. Golm, M. Elizabeth Halloran, and Ira M. Longini, Jr. Semiparametric methods for multiple exposure mismeasurement and a bivariate outcome in HIV vaccine trials. Biometrics, 55(1):94–101, March 1999.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533899.

Tao:1999:EMS

[1873] Huageng Tao, Mari Palta, Brian S. Yandell, and Michael A. Newton. An estimation method for the semiparametric mixed effects model. Biometrics, 55(1):102–110, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533900.

Wang:1999:BCC

[1874] Jinping Wang and Sabyasachi Basu. Bias-corrected confidence intervals for the concentration parameter in a dilution assay. Biometrics, 55(1): 111–116, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533901.

Aitkin:1999:GML

[1875] Murray Aitkin. A general maximum likelihood analysis of variance components in generalized linear models. Biometrics, 55(1):117–128, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533902.

Vangel:1999:MLA

[1876] Mark G. Vangel and Andrew L. Rukhin. Maximum likelihood analysis for heteroscedastic one-way random effects ANOVA in interlaboratory studies. Biometrics, 55(1):129–136, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533903.

Ridout:1999:EIC

[1877] Martin S. Ridout, Clarice G. B. Demétrio, and David Firth. Estimating intraclass correlation for binary data. Biometrics, 55(1):137–148, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533904. See correction [2114].

Lindsey:1999:RSO

[1878] J. K. Lindsey. Response surfaces for overdispersion in the study of the conditions for fish eggs hatching. Biometrics, 55(1):149–155, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533905.

Gignoux:1999:CPD

[1879] Jacques Gignoux, Camille Duby, and Sébastien Barot. Comparing the performances of Diggle’s tests of spatial randomness for small samples with and without edge-effect correction: Application to ecological data. Biometrics, 55(1):156–164, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533906.

Greenhalgh:1999:PRM

[1880] David Greenhalgh, Murray T. Doyle, and Janet Mortimer. A partial ranking method for identifying repeated inclusion of individuals in anonymized HIV infection reports. Biometrics, 55(1):165–173, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533907.

Do:1999:DAE

[1881] Kim-Anh Do and Katherine Kirk. Discriminant analysis of eventrelated potential curves using smoothed principal components. Biometrics, 55(1):174–181, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533908.

Cai:1999:REM

[1882] Jianwen Cai, Pranab K. Sen, and Haibo Zhou. A random effects model for multivariate failure time data from multicenter clinical trials. Biometrics, 55(1):182–189, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533909.

Shen:1999:SIS

[1883] Yu Shen and Lloyd Fisher. Statistical inference for self-designing clinical trials with a one-sided hypothesis. Biometrics, 55(1):190–197, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533910.

Huzurbazar:1999:SHF

[1884] S. Huzurbazar and Aparna V. Huzurbazar. Survival and hazard functions for progressive diseases using saddlepoint approximations. Biometrics, 55(1):198–203, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533911.

LeBlanc:1999:ARS

[1885] Michael LeBlanc and John Crowley. Adaptive regression splines in the Cox model. Biometrics, 55(1):204–213, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533912.

Lipsitz:1999:LMI

[1886] Stuart R. Lipsitz, Joseph G. Ibrahim, and Garrett M. Fitzmaurice. Likelihood methods for incomplete longitudinal binary responses with incomplete categorical covariates. Biometrics, 55(1):214–223, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533913.

Routledge:1999:EST

[1887] R. D. Routledge, G. E. J. Smith, L. Sun, N. Dawe, E. Nygren, and J. S. Sedinger. Estimating the size of a transient population. Biometrics, 55(1):224–230, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533914.

Hung:1999:RGT

[1888] M. Hung and William H. Swallow. Robustness of group testing in the estimation of proportions. Biometrics, 55(1):231–237, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533915.

Betensky:1999:LEE

[1889] Rebecca A. Betensky, Jane C. Lindsey, Louise M. Ryan, and M. P. Wand. Local EM estimation of the hazard function for interval-censored data. Biometrics, 55(1):238–245, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533916.

Kolassa:1999:AMC

[1890] John E. Kolassa and Martin A. Tanner. Approximate Monte Carlo conditional inference in exponential families. Biometrics, 55(1):246–251, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533917.

Lin:1999:PRM

[1891] D. Y. Lin, P. Arbogast, D. S. Siscovick, and R. N. Lemaitre. Poisson regression with missing durations of exposure. Biometrics, 55(1):252– 257, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533918.

Ryan:1999:PRM

[1892] Louise M. Ryan, D. Y. Lin, P. Arbogast, D. S. Siscovick, and R. N. Lemaitre. Poisson regression with missing durations of exposure. Biometrics, 55(1):252–257, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:1999:RLV

[1893] Ming Tan, Yinsheng Qu, and J. Sunil Rao. Robustness of the latent variable model for correlated binary data. Biometrics, 55(1):258–263, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533919.

Freidlin:1999:UVS

[1894] Boris Freidlin and Joseph L. Gastwirth. Unconditional versions of several tests commonly used in the analysis of contingency tables. Biometrics, 55(1):264–267, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533920.

Liao:1999:HBM

[1895] J. G. Liao. A hierarchical Bayesian model for combining multiple 2 × 2 tables using conditional likelihoods. Biometrics, 55(1):268–272, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533921.

Shukla:1999:NET

[1896] G. K. Shukla and G. S. V. Subrahmanyam. A note on an exact test and confidence interval for competition and overlap effects. Biometrics, 55(1): 273–276, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533922.

Li:1999:GST

[1897] Zhengqing Li. A group sequential test for survival trials: an alternative to rank-based procedures. Biometrics, 55(1):277–283, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533923.

Cook:1999:UMF

[1898] Richard J. Cook and Vernon T. Farewell. The utility of mixed-form likelihoods. Biometrics, 55(1):284–288, March 1999. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533924.

Nam:1999:PSS

[1899] Jun-Mo Nam. Power and sample size for testing homogeneity of relative risks in prospective studies. Biometrics, 55(1):289–293, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533925.

Coull:1999:UML

[1900] Brent A. Coull and Alan Agresti. The use of mixed logit models to reflect heterogeneity in capture–recapture studies. Biometrics, 55(1):294–301, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533926.

Begg:1999:ATUa

[1901] Melissa Dowd Begg. Analyzing k (2 × 2) tables under cluster sampling. Biometrics, 55(1):302–307, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533927.

Begg:1999:ATUb

[1902] Melissa Dowd Begg. Analyzing k(2 × 2) tables under cluster sampling. Biometrics, 55(1):302–307, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:1999:RMO

[1903] Minge Xie and Douglas G. Simpson. Regression modeling of ordinal data with nonzero baselines. Biometrics, 55(1):308–316, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533928.

Betensky:1999:MSSa

[1904] Rebecca A. Betensky and Daniel Rabinowitz. Maximally selected χ^{2} statistics for k × 2 tables. Biometrics, 55(1):317–320, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533929.

Betensky:1999:MSSb

[1905] Rebecca A. Betensky and Daniel Rabinowitz. Maximally selected x^{2} statistics for k × 2 tables. Biometrics, 55(1):317–320, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bakke:1999:EOP

[1906] Øyvind Bakke and Svein-Håkon Lorentsen. Estimation of offspring production from a limited number of stage-structured censuses. Biometrics, 55(1):321–325, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533930.

Sykes:1999:BRM

[1907] A. M. Sykes. Book review: Methods and Applications of Linear Models: Regression and the Analysis of Variance, by R. R. Hocking. Biometrics, 55(1):326–327, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533931.

Kemp:1999:BRa

[1908] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(1):326– 332, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lachenbruch:1999:BRL

[1909] P. A. Lachenbruch. Book review: Linear Mixed Models in Practice: a SAS-Oriented Approach, by G. Verbeke, G. Molenberghs. Biometrics, 55 (1):327, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533932.

Nevill:1999:BRA

[1910] A. Nevill and R. Holder. Book review: Analysis of Variance, Design and Regression: Applied Statistical Methods, by R. Christensen. Biometrics, 55(1):327–328, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533933.

Kemp:1999:BRNa

[1911] A. W. Kemp. Book reviews: Neural Networks for Pattern Recognition, by C. M. Bishop; Pattern Recognition and Neural Networks, by B. D. Ripley; Elements of Pattern Theory, by U. Grenander. Biometrics, 55(1):328– 329, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533934.

Xekelaki:1999:BRP

[1912] E. Xekelaki. Book reviews: Prove It with Figures: Empirical Methods in Law and Litigation, by H. Zeisel; Statistics and the Evaluation of Evidence for Forensic Scientists, by C. G. G. Aitken. Biometrics, 55(1):

329–330, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533935.

Anonymous:1999:BRDa

[1913] Anonymous. Book review: Dictionary of Statistics, by B. S. Everitt. Biometrics, 55(1):330, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533936.

Anonymous:1999:BRSa

[1914] Anonymous. Book review: Statistics in Ecology and Environmental Monitoring 2, by D. J. Fletcher, L. Kavalieris, B. F. J. Manly. Biometrics, 55(1):330–331, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533937.

Anonymous:1999:BRAa

[1915] Anonymous. Book review: Applied Survival Analysis, by C. T. Le. Biometrics, 55(1):331, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533939.

Anonymous:1999:BRAb

[1916] Anonymous. Book review: Advances in Statistical Decision Theory and Applications, by S. Panchapakesan, N. Balakrishnan. Biometrics, 55(1): 331, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533940.

Anonymous:1999:BRCa

[1917] Anonymous. Book review: Case Studies in Environmental Statistics, by D. Nychka, W. W. Piegorsch, L. H. Cox. Biometrics, 55(1):331, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533938.

Anonymous:1999:BRI

[1918] Anonymous. Book review: An Introduction to the Mathematics of Neurons: Modeling in the Frequency Domain, by F. C. Hoppensteadt. Biometrics, 55(1):331, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533941.

Anonymous:1999:BRW

[1919] Anonymous. Book reviews: World Health Statistics Annual, 1996, by World Health Organization; World Health Statistics Quarterly, Vol. 50, by World Health Organization. Biometrics, 55(1):331–332, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533942.

Anonymous:1999:FMb

[1920] Anonymous. Front matter. Biometrics, 55(2):i–iv, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533775.

Wilson:1999:EB

[1921] Susan R. Wilson. Evolution and biometry. Biometrics, 55(2):333–337, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533776.

Morrissey:1999:MME

[1922] Mary J. Morrissey and Donna Spiegelman. Matrix methods for estimating odds ratios with misclassified exposure data: Extensions and comparisons. Biometrics, 55(2):338–344, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533777.

Cui:1999:NED

[1923] Jisheng Cui. Nonparametric estimation of a delay distribution based on left-censored and right-truncated data. Biometrics, 55(2):345–349, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533778.

Wei:1999:ODD

[1924] Run-Peng Wei and Francis C. Yeh. Optimal diversity-dependent contributions of genotypes to mixtures. Biometrics, 55(2):350–354, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533779.

Rongling:1999:MEM

[1925] Wu Rongling and Bailian Li. A multiplicative–epistatic model for analyzing interspecific differences in outcrossing species. Biometrics, 55 (2):355–365, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1999:MEM

[1926] Rongling Wu and Bailian Li. A multiplicative-epistatic model for analyzing interspecific differences in outcrossing species. Biometrics, 55(2):

355–365, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533780.

Stanley:1999:GFT

[1927] Thomas R. Stanley and Kenneth P. Burnham. A goodness-of-fit test for capture–recapture model M_{t} under closure. Biometrics, 55(2):366– 375, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533781.

Lee:1999:SDG

[1928] Jae Kyun Lee, Martin Lascoux, Michael A. Newton, and Erik V. Nordheim. A study of deleterious gene structure in plants using Markov chain Monte Carlo. Biometrics, 55(2):376–386, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533782.

Huggins:1999:ESO

[1929] Richard M. Huggins and Paul S. F. Yip. Estimation of the size of an open population from capture–recapture data using weighted martingale methods. Biometrics, 55(2):387–395, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533783.

Sitter:1999:TSD

[1930] R. R. Sitter and C. F. J. Wu. Two-stage design of quantal response studies. Biometrics, 55(2):396–402, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533784.

Follmann:1999:RPR

[1931] Dean A. Follmann, Sally A. Hunsberger, and Paul S. Albert. Repeated probit regression when covariates are measured with error. Biometrics, 55(2):403–409, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533785.

Wu:1999:PHD

[1932] Hulin Wu and A. Adam Ding. Population HIV-1 dynamics in vivo: Applicable models and inferential tools for virological data from AIDS clinical trials. Biometrics, 55(2):410–418, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533786.

George:1999:BAO

[1933] A. W. George, K. L. Mengersen, and G. P. Davis. A Bayesian approach to ordering gene markers. Biometrics, 55(2):419–429, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533787.

Barber:1999:STC

[1934] Stuart Barber and Christopher Jennison. Symmetric tests and confidence intervals for survival probabilities and quantiles of censored survival data. Biometrics, 55(2):430–436, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533788.

Zocchi:1999:OED

[1935] Silvio S. Zocchi and Anthony C. Atkinson. Optimum experimental designs for multinomial logistic models. Biometrics, 55(2):437–444, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533789.

Goggins:1999:ACP

[1936] William B. Goggins, Dianne M. Finkelstein, and Alan M. Zaslavsky. Applying the Cox proportional hazards model when the change time of a binary time-varying covariate is interval censored. Biometrics, 55(2): 445–451, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533790.

Zhang:1999:AIG

[1937] Heping Zhang. Analysis of infant growth curves using multivariate adaptive splines. Biometrics, 55(2):452–459, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533791.

Quan:1999:ART

[1938] Hui Quan and Thomas Capizzi. Adjusted regression trend test for a multicenter clinical trial. Biometrics, 55(2):460–462, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533792.

Muthen:1999:FMM

[1939] Bengt Muthén and Kerby Shedden. Finite mixture modeling with mixture outcomes using the EM algorithm. Biometrics, 55(2):463–469, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533793.

Berlin:1999:ECS

[1940] Jesse A. Berlin, Stephen E. Kimmel, Thomas R. Ten Have, and Mary D. Sammel. An empirical comparison of several clustered data approaches under confounding due to cluster effects in the analysis of complications of coronary angioplasty. Biometrics, 55(2):470–476, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533794.

Walker:1999:BSA

[1941] Stephen Walker and Bani K. Mallick. A Bayesian semiparametric accelerated failure time model. Biometrics, 55(2):477–483, June 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533795.

Simon:1999:BDA

[1942] Richard Simon. Bayesian design and analysis of active control clinical trials. Biometrics, 55(2):484–487, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533796.

Toledano:1999:GEE

[1943] Alicia Y. Toledano and Constantine Gatsonis. Generalized estimating equations for ordinal categorical data: Arbitrary patterns of missing responses and missingness in a key covariate. Biometrics, 55(2):488–496, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533797.

Klein:1999:MRE

[1944] John P. Klein, Corey Pelz, and Mei-Jie Zhang. Modeling random effects for censored data by a multivariate normal regression model. Biometrics, 55(2):497–506, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533798.

Satten:1999:FSM

[1945] Glen A. Satten and Maya R. Sternberg. Fitting semi-Markov models to interval-censored data with unknown initiation times. Biometrics, 55(2): 507–513, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533799.

Sternberg:1999:DTN

[1946] Maya R. Sternberg and Glen A. Satten. Discrete-time nonparametric estimation for semi-Markov models of chain-of-events data subject to interval censoring and truncation. Biometrics, 55(2):514–522, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533800.

Nardi:1999:NRC

[1947] Alessandra Nardi and Michael Schemper. New residuals for Cox regression and their application to outlier screening. Biometrics, 55(2):523– 529, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533801.

vanderLaan:1999:LEE

[1948] Mark J. van der Laan and Alan Hubbard. Locally efficient estimation of the quality-adjusted lifetime distribution with right-censored data and covariates. Biometrics, 55(2):530–536, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533802.

Dunson:1999:SMS

[1949] D. B. Dunson, C. R. Weinberg, S. D. Perreault, and R. E. Chapin. Summarizing the motion of self-propelled cells: Applications to sperm motility. Biometrics, 55(2):537–543, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533803.

Kulldorff:1999:KMO

[1950] Martin Kulldorff and Ulf Hjalmars. The Knox method and other tests for space–time interaction. Biometrics, 55(2):544–552, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533804.

Natarajan:1999:MHN

[1951] Ranjini Natarajan and Charles E. McCulloch. Modeling heterogeneity in nest survival data. Biometrics, 55(2):553–559, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533805.

Rosenbaum:1999:RSH

[1952] Paul R. Rosenbaum. Reduced sensitivity to hidden bias at upper quantiles in observational studies with dilated treatment effects. Biometrics, 55(2):560–564, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533806.

Deltour:1999:SAM

[1953] Isabelle Deltour, Sylvia Richardson, and Jean-Yves Le Hesran. Stochastic algorithms for Markov models estimation with intermittent missing data. Biometrics, 55(2):565–573, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533807.

Preisser:1999:RRC

[1954] John S. Preisser and Bahjat F. Qaqish. Robust regression for clustered data with application to binary responses. Biometrics, 55(2):574–579, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533808.

Parzen:1999:GGF

[1955] Michael Parzen and Stuart R. Lipsitz. A global goodness-of-fit statistic for Cox regression models. Biometrics, 55(2):580–584, June 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533809.

Sinha:1999:BAM

[1956] Debajyoti Sinha, Ming-Hui Chen, and Sujit K. Ghosh. Bayesian analysis and model selection for interval-censored survival data. Biometrics, 55(2):585–590, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533810.

Ibrahim:1999:MCE

[1957] Joseph G. Ibrahim, Ming-Hui Chen, and Stuart R. Lipsitz. Monte Carlo EM for missing covariates in parametric regression models. Biometrics, 55(2):591–596, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533811.

Agresti:1999:LCI

[1958] Alan Agresti. On logit confidence intervals for the odds ratio with small samples. Biometrics, 55(2):597–602, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533812.

Follmann:1999:BME

[1959] Dean A. Follmann and Paul S. Albert. Bayesian monitoring of event rates with censored data. Biometrics, 55(2):603–607, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533813. See [2116].

Brookmeyer:1999:AMP

[1960] Ron Brookmeyer. Analysis of multistage pooling studies of biological specimens for estimating disease incidence and prevalence. Biometrics, 55(2):608–612, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533814.

Lin:1999:SVC

[1961] Xihong Lin and Raymond J. Carroll. SIMEX variance component tests in generalized linear mixed measurement error models. Biometrics, 55(2): 613–619, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533815.

McDonald:1999:ETG

[1962] John W. McDonald, Peter W. F. Smith, and Jonathan J. Forster. Exact tests of goodness of fit of log–linear models for rates. Biometrics, 55(2): 620–624, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533816.

Hughes:1999:MEM

[1963] James P. Hughes. Mixed effects models with censored data with application to HIV RNA levels. Biometrics, 55(2):625–629, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533817.

Greenwood:1999:SIM

[1964] Celia M. T. Greenwood, Julian P. Midgley, Andrew G. Matthew, and Alexander G. Logan. Statistical issues in a metaregression analysis of randomized trials: Impact on the dietary sodium intake and blood pressure relationship. Biometrics, 55(2):630–636, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533818.

Qiou:1999:MSA

[1965] Zuqiang Qiou, Nalini Ravishanker, and Dipak K. Dey. Multivariate survival analysis with positive stable frailties. Biometrics, 55(2):637–644, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2533819.

Xiao-Hua:1999:CSI

[1966] Zhou Xiao-Hua and Wanzhu Tu. Comparison of several independent population means when their samples contain log–normal and possibly zero observations. Biometrics, 55(2):645–651, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:1999:CSI

[1967] Xiao-Hua Zhou and Wanzhu Tu. Comparison of several independent population means when their samples contain log-normal and possibly zero observations. Biometrics, 55(2):645–651, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533820.

Graubard:1999:PMS

[1968] Barry I. Graubard and Edward L. Korn. Predictive margins with survey data. Biometrics, 55(2):652–659, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533821.

Ridout:1999:MCT

[1969] Martin S. Ridout. Memory in coal tits: an alternative model. Biometrics, 55(2):660–662, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533822.

Rayner:1999:MTS

[1970] J. C. W. Rayner and D. J. Best. Modelling ties in the sign test. Biometrics, 55(2):663–665, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533823.

Kemp:1999:BRNb

[1971] C. D. Kemp. Book reviews: Numerical Linear Algebra for Applications in Statistics, by J. E. Gentle; Random Number Generation and Monte Carlo Methods, by J. E. Gentle. Biometrics, 55(2):666–668, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533824.

Kemp:1999:BRb

[1972] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(2):666– 670, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1999:BRAa

[1973] A. W. Kemp. Book review: Analytical Theory of Biological Populations, by A. J. Lotka, D. P. Smith, H. Rossert. Biometrics, 55(2):668, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533825.

Anonymous:1999:BRDb

[1974] Anonymous. Book review: Design and Analysis of Experiments, by A. Dean, D. Voss. Biometrics, 55(2):669, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533827.

Anonymous:1999:BRMa

[1975] Anonymous. Book review: MSI-2000: Multivariate Statistical Analysis in Honor of Professor Minoru Siotani on his 70th Birthday, Vol. IV, by T. Hayakawa, M. Aoshima, K. Shimizu, V. S. Taneja. Biometrics, 55(2): 669, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533829.

Anonymous:1999:BRN

[1976] Anonymous. Book reviews: The New Statistical Analysis of Data, by T. W. Anderson, J. D. Finn; The SPSS Guide to the New Statistical Analysis of Data, by S. B. Gerber, K. E. Voelkl. Biometrics, 55(2): 669, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533826.

Anonymous:1999:BRPa

[1977] Anonymous. Book review: The Practice of Time Series Analysis, by H. Akaike, G. Kitagawa. Biometrics, 55(2):669, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533828.

Anonymous:1999:BRAc

[1978] Anonymous. Book review: Analyzing and Modeling Rank Data, by J. I. Marden. Biometrics, 55(2):669–670, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533830.

Anonymous:1999:BRMb

[1979] Anonymous. Book review: Matrix Differential Calculus with Applications in Statistics and Econometrics, by J. R. Magnus, H. Neudecker. Biometrics, 55(2):670, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533832.

Anonymous:1999:BRT

[1980] Anonymous. Book review: The Theory of Canonical Moments with Applications in Statistics, Probability, and Analysis, by H. Dette, W. J. Studden. Biometrics, 55(2):670, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533831.

Anonymous:1999:FMc

[1981] Anonymous. Front matter. Biometrics, 55(3):i–iv, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533588.

Stockmarr:1999:LRE

[1982] Anders Stockmarr. Likelihood ratios for evaluating DNA evidence when the suspect is found through a database search. Biometrics, 55(3):671– 677, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533589.

Schweder:1999:SLM

[1983] Tore Schweder, Hans J. Skaug, Mette Langaas, and Xeni K. Dimakos. Simulated likelihood methods for complex double-platform line transect surveys. Biometrics, 55(3):678–687, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533590.

Heagerty:1999:MSL

[1984] Patrick J. Heagerty. Marginally specified logistic-normal models for longitudinal binary data. Biometrics, 55(3):688–698, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533591.

Durban:1999:ASE

[1985] Maria Durban, Christine A. Hackett, and Iain D. Currie. Approximate standard errors in semiparametric models. Biometrics, 55(3):699–703, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533592.

Opsomer:1999:KNV

[1986] J. D. Opsomer, D. Ruppert, M. P. Wand, U. Holst, and O. Hössjer. Kriging with nonparametric variance function estimation. Biometrics, 55(3):704–710, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533593.

Wang:1999:ECB

[1987] C. Y. Wang, Garnet L. Anderson, and Ross L. Prentice. Estimation of the correlation between nutrient intake measures under restricted sampling. Biometrics, 55(3):711–717, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533594.

Weinberg:1999:UPE

[1988] Clarice R. Weinberg and David M. Umbach. Using pooled exposure assessment to improve efficiency in case-control studies. Biometrics, 55(3):718–726, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2533595. See correction [2117].

Chavez-Demoulin:1999:BIS

[1989] V. Chavez-Demoulin. Bayesian inference for small-sample capture– recapture data. Biometrics, 55(3):727–731, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533596.

Follmann:1999:VIR

[1990] Dean A. Follmann and Michael A. Proschan. Valid inference in random effects meta-analysis. Biometrics, 55(3):732–737, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533597.

Heikkinen:1999:MPF

[1991] Juha Heikkinen and Elja Arjas. Modeling a Poisson forest in variable elevations: a nonparametric Bayesian approach. Biometrics, 55(3):738– 745, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533598.

Thall:1999:TCB

[1992] Peter F. Thall and Su-Chun Chengt. Treatment comparisons based on two-dimensional safety and efficacy alternatives in oncology trials. Biometrics, 55(3):746–753, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533599.

Chen:1999:EIO

[1993] Song Xi Chen. Estimation in independent observer line transect surveys for clustered populations. Biometrics, 55(3):754–759, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533600.

Regan:1999:LMC

[1994] Meredith M. Regan and Paul J. Catalano. Likelihood models for clustered binary and continuous out comes: Application to developmental toxicology. Biometrics, 55(3):760–768, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533601.

Gerard:1999:LBS

[1995] Patrick D. Gerard and William R. Schucany. Local bandwidth selection for kernel estimation of population densities with line transect sampling. Biometrics, 55(3):769–773, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533602.

Pennello:1999:UME

[1996] Gene A. Pennello, Susan S. Devesa, and Mitchell H. Gail. Using a mixed effects model to estimate geographic variation in cancer rates. Biometrics, 55(3):774–781, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533603.

Follmann:1999:SPT

[1997] Dean A. Follmann and Michael A. Proschan. A simple permutation-type method for testing circular uniformity with correlated angular measurements. Biometrics, 55(3):782–791, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533604.

Datta:1999:EEV

[1998] Susmita Datta, M. Elizabeth Halloran, and Ira M. Longini, Jr. Efficiency of estimating vaccine efficacy for susceptibility and infectiousness: Randomization by individual versus household. Biometrics, 55(3):792–798, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533605.

Chen:1999:CDB

[1999] Song Xi Chen and Jodie L. Woolcock. A condition for designing bus-route type access site surveys to estimate recreational fishing effort. Biometrics, 55(3):799–804, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533606.

Clegg:1999:MMB

[2000] Limin X. Clegg, Jianwen Cai, and Pranab K. Sen. A marginal mixed baseline hazards model for multivariate failure time data. Biometrics, 55(3):805–812, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533607.

Ross:1999:MCD

[2001] Eric A. Ross and Dirk Moore. Modeling clustered, discrete, or grouped time survival data with covariates. Biometrics, 55(3):813–819, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533608.

Dryden:1999:HRR

[2002] Ian L. Dryden and Gary Walker. Highly resistant regression and object matching. Biometrics, 55(3):820–825, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533609.

Cho:1999:FLAa

[2003] Meehyung Cho and Nathaniel Schenker. Fitting the log-F accelerated failure time model with incomplete covariate data. Biometrics, 55(3): 826–833, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cho:1999:FLAb

[2004] Meehyung Cho and Nathaniel Schenker. Fitting the log-F accelerated failure time model with incomplete covariate data. Biometrics, 55(3):826–833, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533610.

Rabinowitz:1999:TAO

[2005] Daniel Rabinowitz and Qiong Yang. Testing for age-at-onset anticipation with affected parent–child pairs. Biometrics, 55(3):834–838, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533611.

Reboussin:1999:EEL

[2006] Beth A. Reboussin, Kung-Yee Liang, and David M. Reboussin. Estimating equations for a latent transit ion model with multiple discrete indicators. Biometrics, 55(3):839–845, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533612.

Ofversten:1999:SQC

[2007] Jukka Öfversten. Statistical quality control of variety trial data. Biometrics, 55(3):846–852, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533613.

Cui:1999:MSS

[2008] Lu Cui, H. M. James Hung, and Sue-Jane Wang. Modification of sample size in group sequential clinical trials. Biometrics, 55(3):853–857, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533614.

Muller:1999:BHA

[2009] Peter Müller, Giovanni Parmigiani, Joellen Schildkraut, and Luca Tardella. A Bayesian hierarchical approach for combining case-control and prospective studies. Biometrics, 55(3):858–866, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533615.

Johnson:1999:DS

[2010] Wesley O. Johnson and Larry M. Pearson. Dual screening. Biometrics, 55(3):867–873, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533616.

Kittelson:1999:UFG

[2011] John M. Kittelson and Scott S. Emerson. A unifying family of group sequential test designs. Biometrics, 55(3):874–882, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533617.

Freidlin:1999:ERT

[2012] Boris Freidlin, Marvin J. Podgor, and Joseph L. Gastwirth. Efficiency robust tests for survival or ordered categorical data. Biometrics, 55(3):883– 886, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533618.

Ryan:1999:ERT

[2013] Louise M. Ryan, Boris Freidlin, Marvin J. Podgor, and Joseph L. Gastwirth. Efficiency robust tests for survival or ordered categorical data. Biometrics, 55(3):883–886, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joly:1999:PLA

[2014] Pierre Joly and Daniel Commenges. A penalized likelihood approach for a progressive three-state model with censored and truncated data: Application to AIDS. Biometrics, 55(3):887–890, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533619.

Real:1999:SEH

[2015] Daniel Real and Ian L. Gordon. Standard errors of heritabilities for forage breeding nurseries. Biometrics, 55(3):891–895, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533620.

Sun:1999:ADE

[2016] Yanqing Sun, Shein-Chung Chow, Gang Li, and Keh-Wei Chen. Assessing distributions of estimated drug shelf lives in stability analysis. Biometrics, 55(3):896–899, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533621.

Wang:1999:EEP

[2017] You-Gan Wang. Estimating equations for parameters in stochastic growth models from tag-recapture data. Biometrics, 55(3):900–903, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533622.

Yip:1999:EPS

[2018] Paul S. F. Yip, Yong Zhou, D. Y. Lin, and Xiang-Zhong Fang. Estimation of population size based on additive hazards models for continuous-time recapture experiments. Biometrics, 55(3):904–908, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533623.

Sun:1999:RAD

[2019] Jianguo Sun, Qiming Liao, and Marcello Pagano. Regression analysis of doubly censored failure time data with applications to AIDS studies. Biometrics, 55(3):909–914, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533624.

Cook:1999:MMT

[2020] Richard J. Cook. A mixed model for two-state Markov processes under panel observation. Biometrics, 55(3):915–920, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533625.

Weerahandi:1999:EIG

[2021] Samaradasa Weerahandi and Vance W. Berger. Exact inference for growth curves with intraclass correlation structure. Biometrics, 55 (3):921–924, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533626.

Dette:1999:ODW

[2022] Holger Dette and Weng Kee Wong. Optimal designs when the variance is a function of the mean. Biometrics, 55(3):925–929, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533627.

Qiu:1999:MDS

[2023] Peihua Qiu, Rick Chappell, William Obermeyer, and Ruth Benca. Modeling daily and subdaily cycles in rat sleep data. Biometrics, 55(3):930– 935, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533628.

Agresti:1999:MCV

[2024] Alan Agresti and I-Ming Liu. Modeling a categorical variable allowing arbitrarily many category choices. Biometrics, 55(3):936–943, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533629.

Pepe:1999:PUP

[2025] Margaret Sullivan Pepe, Patrick Heagerty, and Robert Whitaker. Prediction using partly conditional time-varying coefficients regression models. Biometrics, 55(3):944–950, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533630.

Fahrmeir:1999:AIL

[2026] Ludwig Fahrmeir, Christian Gieger, and Christian Heumann. An application of isotonic longitudinal marginal regression to monitoring the healing process. Biometrics, 55(3):951–956, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533631.

Mikulich:1999:RCC

[2027] Susan K. Mikulich, Gary O. Zerbe, Richard H. Jones, and Thomas J. Crowley. Relating the classical covariance adjustment techniques of multivariate growth curve models to modern univariate mixed effects models. Biometrics, 55(3):957–964, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533632.

Dunson:1999:MTO

[2028] D. B. Dunson and J. K. Haseman. Modeling tumor onset and multiplicity using transition models with latent variables. Biometrics, 55(3):965–970, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533633.

Stallard:1999:DTD

[2029] Nigel Stallard, Peter F. Thall, and John Whitehead. Decision theoretic designs for Phase II clinical trials with multiple outcomes. Biometrics, 55(3):971–977, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533634.

Michiels:1999:SMP

[2030] Bart Michiels, Geert Molenberghs, and Stuart R. Lipsitz. Selection models and pattern-mixture models for incomplete data with covariates. Biometrics, 55(3):978–983, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533635.

Wang:1999:EEN

[2031] You-Gan Wang. Estimating equations with nonignorably missing response data. Biometrics, 55(3):984–989, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533636.

Greenland:1999:LEB

[2032] Sander Greenland, D. Y. Lin, R. A. Kronmal, and B. M. Psaty. Letter to the Editor of Biometrics. Biometrics, 55(3):990–991, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533637.

Kemp:1999:BRE

[2033] C. D. Kemp. Book review: Encyclopedia of Statistical Sciences: Update Volume 3, by S. Kotz, C. B. Read, D. L. Banks. Biometrics, 55(3):992– 993, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533638.

Kemp:1999:BRc

[2034] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(3):992–996, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cormack:1999:BRS

[2035] R. M. Cormack. Book review: Sampling and Statistical Methods for Behavioral Ecologists, by J. Bart, M. A. Fligner, W. I. Notz. Biometrics, 55(3):993, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533639.

Kemp:1999:BRCa

[2036] C. D. Kemp. Book review: CRC Concise Encyclopedia of Mathematics, by E. W. Weisstein. Biometrics, 55(3):993–994, September 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533640.

Anonymous:1999:BRCb

[2037] Anonymous. Book review: Comparative Statistical Inference, by V. Barnett. Biometrics, 55(3):994, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533641.

Anonymous:1999:BRPb

[2038] Anonymous. Book review: Practical Nonparametric Statistics, by W. J. Conover. Biometrics, 55(3):994–995, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533642.

Anonymous:1999:BRE

[2039] Anonymous. Book review: Environmental Statistics: Analysing Data for Environmental Policy, by G. R. Bock, J. A. Goode. Biometrics, 55(3): 995, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533645.

Anonymous:1999:BRSb

[2040] Anonymous. Book review: Statistical Methods for Plant Variety Evaluation, by R. A. Kempton, P. N. Fox. Biometrics, 55(3):995, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533644.

Anonymous:1999:BRSc

[2041] Anonymous. Book review: Sampling of Populations: Methods and Applications, by P. S. Levy, S. Lemeshow. Biometrics, 55(3):995, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533643.

Anonymous:1999:BRDc

[2042] Anonymous. Book review: Design and Analysis of Animal Studies in Pharmaceutical Development, by S. C. Chow, J. P. Liu. Biometrics, 55(3):996, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533647.

Anonymous:1999:BRSd

[2043] Anonymous. Book review: Survival Models and Data Analysis, by R. C. Elandt-Johnson, N. J. Johnson. Biometrics, 55(3):996, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533646.

Anonymous:1999:FMd

[2044] Anonymous. Front matter. Biometrics, 55(4):i–iv, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533711.

Anonymous:1999:VI

[2045] Anonymous. Volume information. Biometrics, 55(4):v–xix, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533710.

Devlin:1999:GCA

[2046] B. Devlin and Kathryn Roeder. Genomic control for association studies. Biometrics, 55(4):997–1004, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533712.

Royston:1999:NAM

[2047] Patrick Royston and Alberto Ferreira. A new approach to modeling daily probabilities of conception. Biometrics, 55(4):1005–1013, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533713.

Schwarz:1999:SPE

[2048] Carl J. Schwarz, Myra Andrews, and Michael R. Link. The stratified Petersen estimator with a known number of unread tags. Biometrics, 55(4):1014–1021, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533714.

TenHave:1999:EBEa

[2049] Thomas R. Ten Have and A. Russell Localio. Empirical Bayes estimation of random effects parameters in mixed effects logistic regression models. Biometrics, 55(4):1022–1029, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533715.

TenHave:1999:EBEb

[2050] Thomas R. Ten Have and A. Russell Localio. Empirical Bayes estimation of random effects parameters in mixed effects logistic regression models. Biometrics, 55(4):1022–1029, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Minkin:1999:LBE

[2051] Salomon Minkin and Kiran Kundhal. Likelihood-based experimental design for estimation of Ed 50. Biometrics, 55(4):1030–1037, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533716.

Chen:1999:EBE

[2052] D. G. Chen, E. M. Carter, J. J. Hubert, and P. T. Kim. Empirical Bayes estimation for combinations of multivariate bioassays. Biometrics, 55(4):1038–1043, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533717.

Halpern:1999:MSO

[2053] Aaron L. Halpern. Minimally selected P and other tests for a single abrupt changepoint in a binary sequence. Biometrics, 55(4):1044–1050, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Halpern:1999:MSP

[2054] Aaron L. Halpern. Minimally selected p and other tests for a single abrupt changepoint in a binary sequence. Biometrics, 55(4):1044–1050, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533718.

Hogmander:1999:MSP

[2055] Harri Högmander and Aila Särkkä. Multitype spatial point patterns with hierarchical interactions. Biometrics, 55(4):1051–1058, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533719.

Reiczigel:1999:AED

[2056] Jenö Reiczigel. Analysis of experimental data with repeated measurements. Biometrics, 55(4):1059–1063, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533720.

McGuire:1999:IEB

[2057] Gráinne McGuire, Michael J. Prentice, and Frank Wright. Improved error bounds for genetic distances from DNA sequences. Biometrics, 55(4):1064–1070, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533721.

Mallick:1999:BSA

[2058] Bani K. Mallick, David G. T. Denison, and Adrian F. M. Smith. Bayesian survival analysis using a Mars model. Biometrics, 55(4):1071–1077, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533722.

Mahe:1999:ERP

[2059] Cédric Mahé and Sylvie Chevret. Estimating regression parameters and degree of dependence for multivariate failure time data. Biometrics, 55 (4):1078–1084, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533723.

Murray:1999:SMC

[2060] Susan Murray and Anastasios A. Tsiatis. Sequential methods for comparing years of life saved in the two-sample censored data problem. Biometrics, 55(4):1085–1092, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533724.

Shen:1999:CBC

[2061] Yu Shen and S. C. Cheng. Confidence bands for cumulative incidence curves under the additive risk model. Biometrics, 55(4):1093–1100, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533725.

Zhao:1999:EED

[2062] Hongwei Zhao and Anastasios A. Tsiatis. Efficient estimation of the distribution of quality-adjusted survival time. Biometrics, 55(4):1101– 1107, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533726.

Huang:1999:TSP

[2063] Yijian Huang. The two-sample problem with induced dependent censorship. Biometrics, 55(4):1108–1113, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533727.

Chang:1999:NAD

[2064] Chung-Chou H. Chang and Lisa A. Weissfeld. Normal approximation diagnostics for the Cox model. Biometrics, 55(4):1114–1119, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533728.

Piepho:1999:FRM

[2065] Hans-Peter Piepho. Fitting a regression model for genotype-by-environment data on heading dates in grasses by methods for nonlinear mixed models. Biometrics, 55(4):1120–1128, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533729.

Langholz:1999:ABR

[2066] Bryan Langholz, Argyrios Ziogas, Duncan C. Thomas, Cheryl Faucett, Mark Huberman, and Larry Goldstein. Ascertainment bias in rate ratio estimation from case-sibling control studies of variable age-at-onset diseases. Biometrics, 55(4):1129–1136, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533730.

Gail:1999:TAE

[2067] Mitchell H. Gail, Larry Kessler, Douglas Midthune, and Steven Scoppa. Two approaches for estimating disease prevalence from population-based registries of incidence and total mortality. Biometrics, 55(4):1137–1144, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533731.

Erkanli:1999:BIP

[2068] Alaattin Erkanli, Refik Soyer, and Elizabeth J. Costello. Bayesian inference for prevalence in longitudinal two-phase studies. Biometrics, 55 (4):1145–1150, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533732.

Follmann:1999:MTI

[2069] Dean A. Follmann and Michael A. Proschan. A multivariate test of interaction for use in clinical trials. Biometrics, 55(4):1151–1155, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533733.

Shih:1999:CPT

[2070] Joanna H. Shih and Michael P. Fay. A class of permutation tests for stratified survival data. Biometrics, 55(4):1156–1161, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533734.

Carlin:1999:HPH

[2071] Bradley P. Carlin and James S. Hodges. Hierarchical proportional hazards regression models for highly stratified data. Biometrics, 55(4):1162– 1170, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533735.

Venkatraman:1999:PNT

[2072] E. S. Venkatraman and Colin B. Begg. Properties of a nonparametric test for early comparison of treatments in clinical trials in the presence of surrogate endpoints. Biometrics, 55(4):1171–1176, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533736.

Grigoletto:1999:ACI

[2073] Matteo Grigoletto and Michael G. Akritas. Analysis of covariance with incomplete data via semiparametric model transformations. Biometrics, 55(4):1177–1187, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533737.

Tang:1999:CTP

[2074] Dei-In Tang and Nancy L. Geller. Closed testing procedures for group sequential clinical trials with multiple endpoints. Biometrics, 55(4):1188– 1192, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533738.

Holcroft:1999:DVS

[2075] Christina A. Holcroft and Donna Spiegelman. Design of validation studies for estimating the odds ratio of exposure-disease relationships when exposure is misclassified. Biometrics, 55(4):1193–1201, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533739.

Chan:1999:TBE

[2076] Ivan S. F. Chan and Zhongxin Zhang. Test-based exact confidence intervals for the difference of two binomial proportions. Biometrics, 55 (4):1202–1209, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533740.

Bartoletti:1999:AE

[2077] Stefania Bartoletti, Bernard D. Flury, and Daan G. Nel. Allometric extension. Biometrics, 55(4):1210–1214, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533741.

Lee:1999:SIS

[2078] Mei-Ling Ting Lee and G. A. Whitmore. Statistical inference for serial dilution assay data. Biometrics, 55(4):1215–1220, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533742.

Mathew:1999:EMA

[2079] Thomas Mathew and Kenneth Nordström. On the equivalence of metaanalysis using literature and using individual patient data. Biometrics, 55(4):1221–1223, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533743.

Satten:1999:VMB

[2080] Glen A. Satten, Robert Janssen, Michael P. Busch, and Somnath Datta. Validating marker-based incidence estimates in repeatedly screened populations. Biometrics, 55(4):1224–1227, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533744.

Satten:1999:EET

[2081] Glen A. Satten. Estimating the extent of tracking in interval-censored chain-of-events data. Biometrics, 55(4):1228–1231, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533745.

Shih:1999:LMC

[2082] Joanna H. Shih and Paul S. Albert. Latent model for correlated binary data with diagnostic error. Biometrics, 55(4):1232–1235, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533746.

Chen:1999:NIM

[2083] Yuh-Ing Chen. Nonparametric identification of the minimum effective dose. Biometrics, 55(4):1236–1240, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2533747.

Regal:1999:ETA

[2084] Ronald R. Regal and Ernest B. Hook. An exact test for all-way interaction in a 2^{M} contingency table: Application to interval capture– recapture estimation of population size. Biometrics, 55(4):1241–1246, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533748.

Zhang:1999:BRR

[2085] Heping Zhang and Daniel Zelterman. Binary regression for risks in excess of subject-specific thresholds. Biometrics, 55(4):1247–1251, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533749.

Albert:1999:MSM

[2086] Paul S. Albert. A mover–stayer model for longitudinal marker data. Biometrics, 55(4):1252–1257, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533750.

Chen:1999:RBT

[2087] Yuh-Ing Chen. Rank-based tests for dose finding in nonmonotonic doseresponse settings. Biometrics, 55(4):1258–1262, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533751.

Wang:1999:EER

[2088] You-Gan Wang. Estimating equations for removal data analysis. Biometrics, 55(4):1263–1268, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533752.

Chen:1999:VDC

[2089] John J. Chen and GlenysM Thomson. The variance for the disequilibrium coefficient in the individual Hardy–Weinberg test. Biometrics, 55(4):1269–1272, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2533753.

Sun:1999:UHC

[2090] Jianguo Sun. On the use of historical control data for trend test in carcinogenicity studies. Biometrics, 55(4):1273–1276, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533754.

Lindsey:1999:MEC

[2091] J. K. Lindsey. Multivariate elliptically contoured distributions for repeated measurements. Biometrics, 55(4):1277–1280, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533755.

Louzada-Neto:1999:PML

[2092] Francisco Louzada-Neto. Polyhazard models for lifetime data. Biometrics, 55(4):1281–1285, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2533756.

Lehmacher:1999:ASS

[2093] Walter Lehmacher and Gernot Wassmer. Adaptive sample size calculations in group sequential trials. Biometrics, 55(4):1286–1290, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2533757.

Kolassa:1999:SSC

[2094] John E. Kolassa and Martin A. Tanner. Small-sample confidence regions in exponential families. Biometrics, 55(4):1291–1294, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533758.

Wei:1999:MCI

[2095] Wen Hsiang Wei and Ju Shiang Su. Model choice and influential cases for survival studies. Biometrics, 55(4):1295–1299, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533759.

Fritsch:1999:MCE

[2096] Kathleen S. Fritsch and Jason C. Hsu. Multiple comparison of entropies with application to dinosaur biodiversity. Biometrics, 55(4):1300–1305, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2533760.

Auranen:1999:HBM

[2097] Kari Auranen, Martin Eichner, Helena Käyhty, Aino K. Takala, and Elja Arjas. A hierarchical Bayesian model to predict the duration of immunity to Haemophilus influenzae Type b. Biometrics, 55(4):1306– 1313, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2533761.

Senn:1999:COA

[2098] Stephen Senn and Andrew P. Grieve. A comment on optimal allocations for bioequivalence studies. Biometrics, 55(4):1314–1315, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533762.

Anonymous:1999:EB

[2099] Anonymous. To the Editor of Biometrics. Biometrics, 55(4):1316–1317, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hernan:1999:ASR

[2100] Miguel A. Hernán and James M. Robins. [Assessing the sensitivity of regression results to unmeasured confounders in observational studies]. Biometrics, 55(4):1316–1317, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533763.

Anonymous:1999:ARF

[2101] Anonymous. The authors replied as follows. Biometrics, 55(4):1316– 1370, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1999:BRNc

[2102] A. W. Kemp. Book review: The Nature of Mathematical Modeling, by N. Gershenfeld. Biometrics, 55(4):1318–1319, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533764.

Kemp:1999:BRd

[2103] C. D. Kemp and A. W. Kemp. Book reviews. Biometrics, 55(4):1318– 1322, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thomas:1999:BRJ

[2104] L. Thomas. Book review: JMP Start Statistics: a Guide to Statistics and Data Analysis Using JMP and JMP IN Software, by J. Sall, A. Lehman. Biometrics, 55(4):1319–1320, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533765.

Buckland:1999:BRQ

[2105] S. T. Buckland. Book review: Quantitative Fish Dynamics, by T. J. Quinn, R. B. Deriso. Biometrics, 55(4):1320, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533766.

Anonymous:1999:BRPc

[2106] Anonymous. Book review: Practical Longitudinal Data Analysis, by D. Hand, M. Crowder. Biometrics, 55(4):1320–1321, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533767.

Anonymous:1999:BRB

[2107] Anonymous. Book review: Bayesian Forecasting and Dynamic Models, by M. West, J. Harrison. Biometrics, 55(4):1321, December 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533771.

Anonymous:1999:BRG

[2108] Anonymous. Book review: The Genetical Analysis of Quantitative Traits, by M. J. Kearsey, H. S. Pooni. Biometrics, 55(4):1321, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533769.

Anonymous:1999:BRL

[2109] Anonymous. Book review: Local Polynomial Modelling and Its Applications, by J. Fan, I. Gijbels. Biometrics, 55(4):1321, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533768.

Anonymous:1999:BRR

[2110] Anonymous. Book review: Random and Quasi-Random Point Sets, by P. Hellekalek, G. Larcher. Biometrics, 55(4):1321, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2533772.

Anonymous:1999:BRSe

[2111] Anonymous. Book review: The Statistical Theory of Shape, by C. G. Small. Biometrics, 55(4):1321, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2533770.

Kemp:1999:BRNd

[2112] C. D. Kemp and A. W. Kemp. Book review: The Northeast Shelf Ecosystem: Assessment, Sustainability, and Management, by K. Sherman, N. A. Jaworski, T. J. Smayda. Biometrics, 55(4):1321–1322, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2533773.

Anonymous:1999:A

[2113] Anonymous. Acknowledgements. Biometrics, 55(4):1323, December 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:CEI

[2114] Anonymous. Correction to “Estimating Intraclass Correlation for Binary Data” by M. S. Ridout, C. G. B. Demétrio, and D. Firth; 55, 137–148, March 1999. Biometrics, 59(1):199, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1877].

Anonymous:2003:CEA

[2115] Anonymous. Correction to “Estimation of absolute risk from nested casecontrol data” by B. Langholz and Ø. Borgan; 53, 767–774, June 1997. Biometrics, 59(2):451, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1533].

Anonymous:2013:CBM

[2116] Anonymous. Correction to “Bayesian Monitoring of Event Rates with Censored Data,” by D. A. Follmann and P. S. Albert; 55, 603–607, DOI: 10.1111/j.0006-341X.1999.00603.x June 1999. Biometrics, 69(2): 556, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1959].

Anonymous:2014:CUP

[2117] Anonymous. Correction to “Using Pooled Exposure Assessment to Improve Efficiency in Case-Control Studies,” by Clarice R. Weinberg and David M. Umbach; 55, 718–726, September 1999. Biometrics, 70(4): 1061, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1988].